A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.03

Title word cross-reference

−0 [TBCK91]. ^{2} [JCT97]. _{2} [FPD90]. b [TMWD97]. F_{1} [KJ95, TMD96]. × [WA91]. W_{r} [BM91, LPW^{+}95].

-Induced [FPD90].

1-m [JCT97]. 12-Month [CB95]. 1970s [Kra93]. 1988 [Har98]. 1989 [EMC90].

222 [MZ97]. 2nd [KOH96, McG91].

3-Trifluoromethyl-4-nitrophenol [KKR^{+}94].

8-Tetrachlorodibenzo-p-dioxin [JHT96].

Abalone [DHR96]. Ability [HSES98, Mes94, WK91]. Ablation [VE99]. Abnormal [PGG90]. Abnormalities [FSW96, NW95]. Absaroka [KHR97]. Absence [LRW97, LRW98, SO96, You98]. Absolute [KC98]. Absorptive [FF92]. Abundance [BS93a, BR91a, BV97, BMN91, BG99, BTA97, Bra99, BC94, CH95, CR97b, CST96, DE92, Dib91, EKH93, EO91, Fle94a, GE91b, GHG98, HSH91, HTS92, HW96, HWW^{+}97, Jag95, KC98, Lue90, LW93, MSC90, MSP95, MP98, MMH94, Mic97a, Mic97b, Mor92, PT94, RN92, SAG95, SR95c, SV95, UMT96, WPW94, ZHP99]. Abundances [CZG90, GE91a, HM92, MP97, PGG90]. AC [Gri91]. Acanthurids [MKT92]. Acceleration [HONCS98]. Acceptable [HJDI98, Hou97]. Access [BBA98]. Acclimation [KN90, KS97]. according [WH95]. Accumulation [HBMB98, MC93a]. Accumulations [CST96]. Accuracy [PVE94, RB94]. Achieving [Sca94]. Acid [ESM^{+}92, GHR91, GNM92, HIG^{+}99, SI91, TR90, VRF94]. Acidic [GSC93, Gra93, WFL^{+}91]. Acidification [BRWB92, HSCK97, LHS95]. Acidified [FC93, ML99]. Acoustic [ABM98, ACB^{+}99, PRR94]. across [DDS96, HK93, SAG95]. Act [ADHB98]. Activated [SBRH94]. Activation [BWT96]. Activities [CH93]. Activity [ABM98, BS93b, BFM97, CDAM90, CCD90, DMWM96, Hoc90, LBCB96, Sta92, YRB97]. Actograph [ABM98]. Acute [HIG^{+}99, Mes94, NWGA99]. Adaptation [Bur99]. Adaptive [KJ95]. Adding [BLW98]. Additive [RH91]. Adfluvial [KOB99]. Adhesive [PJ91]. Adult [BDZ^{+}99, CC90, CM97, DGHP99, DBCM92, DHP99, GWF99, HSH91, HDL^{+}96, HD97, KL96b, LD98, LP94, MK91, MT91, Nak94, ORT^{+}98, PE94, SMHP92, TF99, UJ98, WK90, You98]. Affect [BB98, HC96b]. Affecting [BJ92, CR97a, CRRH93, LO95, Mic97a, PQF95, SHK93, SI91]. Affects [HSES98, WGFS99]. AFLP [LLK^{+}99]. Africa [Kel99]. African [BW99]. after [BMN91, BSS98, DHRK94, DFO94, HBMB98, KW99, LHS95, Lue90, MMH90, MSC90, MSP91, MHS^{+}99, OW96, PD98, SC92b, Swi90, Toh93, TBM95, UHG94, WBI97, WE92]. Age [BW99, BMWB97, BSRW90, BFOS95, BM94, BLC98, BCB96, CBB^{+}96, CAM95, CDK97, DSRS93, DGHP99, DSL98, DB95, DBCM92, Dyg90, FP99, GS98b, GM93, Goo95, GM94, HR95, Hal91, HB93, HW96, HDS91, Hug98b, Jen94, JP96, JE96, JKM96, JMMC93, KJ93, MFWP96, MSP95, Mic97a, Mic97b, MH94, MVW99, MIPGV97, MT94b, Mur94, OJS97, PRR94, PVE94, PJN95, PRS95, PPJ98, PMV98, PM99, PvdH99, PMF90, RB94, RB92, RSV95, RNB95, SO94, SSP96, SM96a, SB94, TBCK91, TB95, TMD96, TH91, UMT96, VHRR98, WH95, War92, WGFS99, ZMH97]. Age-0 [BLC98, CDK97, DSRS93, DGHP99, DSL98, DB95, FP99, GS98b, GM93, HB93, HW96, JP96, JE96, JKM96, JMMC93, MFWP96, MSP95, Mic97a, Mic97b, MH94, PJN95, PRS95, PPJ98, PvdH99, PMF90, SO94, SM96a, SB94, TMD96, TH91, VHRR98, WGFS99]. Age-1 [JKM96]. Age-at-Maturity [BFOS95]. Age-Class [Dyg90]. Age-Related [GM94]. Age-Specific [PJN95]. Ages [MT91]. Aggression [MW98, Mes91]. Aging [Hal91]. Agricultural [SJW92, SO95]. Aided [CBB^{+}96]. Air [Bur99, Hen90, MK90a]. Air-Breathing [Bur99]. Alabama [HD97, BM97, LD98, SMD94, TMD96, TMWD97]. Alaska [BKM91, BRQ93, BSWW98, BSS98, BBM91, CB92, ENB92, FDGG92, FGB^{+}95, FGW^{+}95, FGMG96, FP99, GGGD92, HBMB98, HBB^{+}98, KBM^{+}98, KWE99, OW95, SC99a, SC99b, UMT96, WSB^{+}92, Whi92, WGG92]. Alaskan [BSC97, Hug98a, Hug98b]. Albemarle [OR92, HCH96]. Alberta [AT99, BM95, CHI94]. Alevins [BM90]. Alewives [Jes93, Kra93, OJS97, TN91]. Alimentary [RN98]. Alizarin [BS96a]. Alkaline [MWW^{+}94]. Alkalinity [BMWB97]. All-Female [SSS^{+}99]. Allocation [JN97]. Allochthonous [HBMB98]. Allometry [McG93]. Allozyme [GRKM93, GKM^{+}94, MFL99, Ret98, SA92, WAW96]. Allozymes [EKNP96, MSC94, SC99a, VW98]. Allozymic [KP92]. along [GSS90, HBMB98, LLP^{+}94, SK91, SC99a, WDRM90, XC95]. Alpine [WPSD99]. Alterations [BKM91]. Altered [BFB98]. Alternating [BWH93]. Alternative [BC97, MS96b, VPPH91]. Aluminum [WFL^{+}91]. alvei [TKOS92]. Amazon [Bay98, Lov98]. Ambient [MS99]. America [Bra99, Gre99, HB95c, MK90a, SM91]. American [Van93, BSEK99, CDAM90, GJ95, HONCS98, MHR94, Mor92, NAC99, Oli97, RB92, RBL92, RRBA92, SC99a, SAG95, VMM99, WK90]. Amisk [AT99]. Ammonia [LRW97, LRW98, LRW99]. Ammonium [BG94]. among [BHSH96, BEC98, CHSM93, CW96, DS96, DLF97, EHBP99, FBHB97, GRKM93, HD94, HC96b, Lyo97, MCC^{+}91, PKM93, SO96, You99]. Amount [KR94a]. Amplification [WTC92]. Amplified [LLK^{+}99, WKW98]. Amprolium [FABH^{+}98]. Anadromous [BHJ^{+}95, DHP99, FGB^{+}95, Gar92, GWL98, IB96, Jen94, Jes93, JMMC93, RES93, SSGE94, SSG94, SH94, Zim99]. Anaerobic [KFTT96]. Analogy [Mur93]. Analyses [BR91b, DV97, HJWS99, WMC93]. Analysis [BM90, BS95b, BSEK99, DHP99, EA92, FBHB97, GSR97, Ham91, HW93, JRLK95, LJ92, MFL99, MVW99, PPF99, PH92, Ret98, RBM94, Sai92, SSEO99, SM90a, Smi91, SMHP92, TFS92, TMWD97, VSSS91, WW94, WHW96, WTC92, XC95]. Analyzing [HT91]. Ancestral [QGWF98]. Anchor [WDM90]. Anchovy [LM91]. Anesthetics [SMDC96]. Angled [TIT96]. Angler [CB95, Fis96a, JRLK95]. Angling [GWW91, JROG90, KKP^{+}95, Li99, WDB^{+}96]. anguillarum [ACH^{+}98]. Anguillicola [FWJ96]. Anna [SO94]. Annaquatucket [Oli97]. Annual [JS95, KK93, LBCB96, NW93, ROI^{+}97, RJK^{+}99, RKBC95, SSEO99, WDK90]. Annuli [HB92]. Anoxia [NÜr95, SS90a]. Anoxic [NÜr95]. Antibodies [BFBC92, GNZO97]. Antigen [BFBC92]. Apache [CHSM93]. Apalachicola [ZHP99]. Apicomplexa [LF95]. Appalachian [BIBV98, CR97a, DFO94, ESM90, Fle94a, LMC95, LW92, PM95]. Appearance [GS98a]. Applicability [BBA^{+}99, RTA95]. Application [DC90, FGB^{+}95, HH93, LSC93, PC98, PGS96, Sai92, SMHP92, VHRR98]. Applications [BH93, Cow97, HBB^{+}93, IMMP90]. Applied [EA92, Lis98, WRS^{+}97]. Approach [CP94, CV99, KR96, MC93b, Men93a, VRC93]. Approaches [BH93, Iwa96, LJS^{+}97, Wea90]. Aquaculture [Lov98]. Aquatic [BB92, BMN91, CZG90, Eis95, Gre99, JHT96, Koh97, Lem96, RHP90, Ton93, WD90]. Arctic [BM98, DHRK94, DGHP99, FGMG96, GGGD92, GN99, HGK^{+}94, HDS91, Hug98a, Hug98b, Jen94, KJ95, KWE99, LB94, LB92, LBKD96, Nil94, OS93, SSGE94, SSG94, SH94, UHG94, UMT96, WSB^{+}92]. Area [CW95, DRF90]. Areas [DHMS94, FF92, HDL^{+}96, MHS^{+}99, SH98]. Arizona [CHSM93, CCR98, CW95, MD97]. Arkansas [CR97b, JHT96, SF97, WW96]. Arripis [DSD99]. Art [DMM98]. Artificial [BGB90, HJDI98, HW96, KT94, Mes91, PE94]. Artificially [GVMP^{+}95]. Asian [SC99a]. Assay [BZBE98, HJWS99]. Assemblage [BC94, LO91, LC96, MCB90, Pal94, Ret98, TBM95, WG94, YL93]. Assemblages [GS98b, LB92, MRM97, PLL92, RH91, RES93, REM92, SH97a, WMC93]. Assess [ABM98, HDL^{+}96, PRD96, PBH92]. Assessed [SV95]. Assessing [BFB98, BR91b, HD94]. Assessment [ABG93, BS97b, Car92, Con98, Cow97, JKM96, LPW^{+}95, LCR98, Lyo97, RTA95, TSM98]. Assessments [SMPS98]. Associated [CB95, Dyg90, MM92b, PCGW90, ZWLB90]. Associations [Dol90, HN90]. Assortative [EHBP99]. Assumptions [BS93b]. Asymmetries [WP93]. Asymmetry [MIPGV97]. Atchafalaya [SKR99]. Athabasca [CHI94]. Atherinid [BMN91]. Atlantic [BM94, RT99, BFOS95, DH94, DCC99, DLF97, DWRM97, FFCS96, FNAR97, FPD90, FABH^{+}98, GHR91, GVMP^{+}95, Gar92, GSS90, HONCS98, HBS90, JDW^{+}97, KK93, LHS95, LP94, MT94a, MIPGV97, MR95b, Nor91, PMSM97, PP94, RE97a, RE97b, RD95, SH90, Sau91, SD90, SMPS98, SSG94, Tri98, VVF92, WDRM90, WF94, WHW96, War92, WPM99, WDB^{+}96, WGG92, WMWC93, ZMH97]. Atnsjø [HDS91]. ATPase [SBRH94]. Attached [CF99]. Attachment [PMSM97]. Attack [Swi90]. Attributable [KOB98]. Attributes [TK94]. Australia [AW94, Li99, SSG93]. Australian [PG96]. Austrian [WS99]. Automated [BFM97]. Autumn [WSB^{+}92]. Autumnal [You98]. Availability [PM92]. Available [DSB98, Kel96]. Average [HH93]. Avoid [MKT92]. Avoidance [JPL93, Mes94, ODS92, WGFB97].

Back [KBF99, PRL96]. Back-Calculation [KBF99, PRL96]. Backwaters [ZJ93]. Bacterial [BE92, NW95, WE90]. Bacterium [TKOS92]. Bag [PF90]. Bait [RO93]. Balance [Sca94]. Baltic [Nak99]. Banding [FBHB97, MB91a, ZT98]. Bank [CBB^{+}96, HPH93, RN92]. Bar [Ken95, Mor95, Sig94, Sig95, VSGA94]. Bar-Coded [Ken95, Mor95, Sig95]. Barium [PG96]. Barramundi [PG96, SSG93]. Barriers [WP98]. Based [BPL99, BRTV95, BA90, Bre93b, BHSH96, CBB^{+}96, Cha93, CP94, CR97a, CRRH93, DD98, EA92, GNM92, GKM^{+}94, PG96, RC93, TMWD97, VRC93, VRW^{+}93, WHW96, XC95]. Baseline [RC93]. Basin [BFH99, LLP^{+}94, MMT96, PPF99, RSL94, SKR99, Tyu90, TH91, Hol90, MRM97, Zim99, ZW99]. Basins [CH99, You99, RES93]. Basis [PS91, RO93].

Bass [BB92, BWH93, BR91a, BFOS95, Bev96, BK93, BZBE98, BM91, BEC98, BCSD97, BK96, CHH98, CBYH96, CV99, CM97, CB90, Cou90, CRRH93, DMWM96, DDS96, DSL98, DLF97, DB95, DIT97, DR96, DD98, DHMS94, DWRM97, ESM^{+}92, FSO91, FK90, GS98b, GGM95, GM93, GNM92, Gri93, GM94, GC99, GKF94, GWW91, HCH96, HAHD94, Ham91, HNW^{+}98, HB93, Har98, HW96, HB94, HIG^{+}99, HC91, HD94, HC92, HLH92, JP96, KKP^{+}95, KFTT96, KBF99, KSH^{+}94, LRC93, LT96, LPFA97, LO95, Lyo97, MM92a, MHW^{+}91, MSP91, MK90a, MK90b, MO91, Men93a, Men93b, MH94, NBGF93, NP94a, NWGA99, OFM91, OR92, PW91, PJN95, PV97, PB98, ROI^{+}97, RE97a, RE97b, RBVP91, RLCB98, RO93, RC93, RD95, RT99, SO94, SO96, SJW92, SD90, SWD91, SI91, SS90b, SS91]. Bass [SR95c, SE98, SPR94, SR95d, SC92b, TIT96, TWC91, VM94, VHRR98, VPPH91, WDM90, WDRM90, WW94, WF94, WW95, Wal95, WRS^{+}97, WV95, WW96, WZ99, WKPK93, WH97b, WKW98, WL94, WGGM91, WMWC93, WA91, WKSS95, WGFS99, ZWLB90, Zim99]. Basses [HJWS99, MCC^{+}91]. Batch [GA93]. Bay [BRQ93, DS90, FDGG92, FGB^{+}95, FGW^{+}95, FGMG96, FTB90, GGGD92, HB99, Li99, LM91, RS94, BK93, CB90, DHMS94, GE91b, GT96, HR95, LT96, LBCB96, LM91, MH94, WMWC93]. Bayou [KR94b, JHT96]. Beam [ABM98, SW95]. Bear [BL97, NW92, RW99]. Beaufort [KWE99, LB92]. Beaver [ME94]. Become [MKT92]. Bed [BLB98, KCSF91]. Beds [KC98]. before [CDL^{+}99, DHRK94, DFO94, KW99]. Behavior [AREK98, CP97, Fle94b, HONCS98, Hug98a, JROG90, KJ95, KMNF94, LAD97, LT92, Nie92, OW96, PC98, RB92, SHK93, WEH95]. Behavioral [Bir98, BLP^{+}94, BKN90, GFD95, LB94, SG95]. Behaviour [Lav92]. Bellied [BCWB97]. Benefit [GN99]. Benthic [HTS92, LAD97, Lue90, MBS^{+}92, WBD^{+}94]. Bertalanffy [VB94]. Better [CG97, SSS^{+}99]. between [Abr96, ADHB98, BS97a, BTS^{+}99, BG99, BM98, Con90, DR94, DW92, Don97, FGB^{+}95, GS98a, GFD95, GSS90, HNW^{+}98, HK93, JCP96, JKO93, KBSM90, KFTT96, KSZ99, MPG93, Mes91, Pal94, PV97, PB98, PD94, PMF90, PPD95, RCR93, SK91, SK96, Sca94, Sco93, SC99b, SMD94, TL96, TMWD97, VWBM91, VRF94, WSCL96, WE90]. Bias [Fis96a, HB93, Pie97, RSBN98]. Biases [CB95]. Big [Bre93a, BM98, LBKD96]. Bight [CC90, WHW96, ZMH97, Phe92]. Bioaccumulation [MS97]. Biochemical [BGB90, DNR^{+}94, Eis95, LLD96, MT94a, TSM98]. Biodiversity [Kel99]. Bioenergetic [BT90, RLC93, WS91]. Bioenergetics [BS93b, BH93, HBB^{+}93, HB93, Lab93, LB93, Lyo97, MO99, MC93b, Ney93, PW99, RR99, RBM94, WH97b, WHNW98]. Biological [KW99, PZW95, Rab97, Rei92, SI91]. Biology [Far90, Lem96, LM91, NW93, RW92, RS94, RSV95, See93, Wal95, WD90]. Biomanipulation [DDS96]. Biomarkers [SSGI99]. Biomass [CW95, CH99, DH94, ESM90, LPW^{+}95, OFM91, WDK90]. Biomes [Gre99]. Biotelemetry [DMWM96, GWJ94]. Biotic [BF91, CWBC90, OKB^{+}92, RH91, SHC99]. Biphenyls [FSO91]. Birth [FNAR97]. Bivalve [Lei90]. Black [BSRW90, DNR^{+}94, ESM^{+}92, HD94, NW93, NW95, PJ91, SMD94, Ste91, TMD96, TMWD97]. Blackfish [SC92a]. Blackfoot [Swa97]. Blacknose [MZ97]. Blackwater [Mef91]. Bladder [FWJ96]. Bladders [MHC95]. Bloater [BR94b, MCB90, RBM94]. Bloaters [TF99]. Blood [FG98, MM92b]. Blue [BWD98, CSM92, Fis99, SBSN95]. Blueback [DF91, Jes93]. Bluefin [SMPS98]. Bluefish [BC96, BLC98, CC90, JMMC93, MSP95]. Bluegill [GS98b, HIG^{+}99, SCJ93]. Bluegills [Bre93b, FK90, HB92, HW96, Joh93, Kel96, KL96b, Mef91, MS96b, PD99]. Board [Sig94]. Boat [LD97]. Body [BRTV95, FGW^{+}95, FGMG96, Ham91, HJLM99, KFTT96, KBF99]. Bonneville [BL97, KOB99]. Book [Gro99, Nev99, PT99]. Bore [RT99]. Boreal [ACB^{+}99]. Boron [HW90]. Bottom [RN92]. Bottom-Water [RN92]. Boulder [KCSF91]. Boulder-Bed [KCSF91]. Bounded [EA92]. Box [WM97]. Brain [Sta92]. Breathing [Bur99]. Bristol [BRQ93]. British [BPL99, Hea91, LM93, PPF99, SMC90, ENB92, PGG90, QGWF98, WE90]. Broad [FDGG92, FGW^{+}95, GGGD92, OSBS98]. Broad-Scale [OSBS98]. Brood [Hub99]. Brook [Abr96, BC97, Bir98, BR96, BM98, CR97a, CS99, CBND97, DR94, DW92, DHP99, ESM90, FC93, GSC93, Hal91, KFTT96, KMNF94, LHS95, LW92, MW98, MPG93, MGM93, ME94, Mei90, PKM93, SH98]. Brown [AW94, BDL97, BC91, BIBV98, CCD90, CR97b, CNWT93, DW92, GFD95, GHW91, HW93, HDS91, HBS90, KT94, KT96, LCR98, MKR97, Pow94, RLC93, SSGI99, WS99]. Brunswick [DHP99, Jes93]. Brush [Cou96, QPS96]. Brushy [SPQ95]. Bubble [Lut95]. Budget [HC91]. Buffalo [WW96]. Bull [BFH99, BS96b, JMTC98, KMNF94, LA97, RM95, Swa97, WPSD99]. Bullhead [CNWT93]. Bullheads [SSGI99]. Burbot [Car95, EKH93, PPF99]. Bycatch [DCC99]. Bypasses [HONCS98].

calcarifer [PG96]. Calcein [BK96, GCM^{+}97, Mon93]. Calcified [Cas90]. Calcium [GM94, GC99]. Calculation [KBF99, PRL96]. California [GJ95, BGB90, BG90, CBB^{+}96, DHR96, DE92, HN96, KV96, KVR96, MBAL94, Mat96, MHSM92, Nak94, NLO94, NCFG97, RMSL92, SJW92, TSM98]. Can [PH92]. Canada [Hen90, KWE99, RD95, RT99]. Canadian [FS96, WPSD99]. Canal [RN98]. Cannibalism [DD98, JP96]. Canopy [HBMB98, HBB^{+}98]. Canyon [CCR98]. Capabilities [ODS92]. Capacity [Con90]. Cape [MR95b, Jag95]. Captive [BTS^{+}99]. Captive-Reared [BTS^{+}99]. Captivity [KSH^{+}94, NAC99]. Capture [Lin96, WK91]. Captured [GKM^{+}94, GHG98]. Capturing [HC92]. Carbon [KOB98]. Carcasses [Gar92]. Carolina [BS93a, BCSD97, CHH98, CS91, DF91, Fle94a, HCH96, MR95b, Pal94, PRD96, PD98, RSV95, War92]. Carp [BWTM90, CP97, FG93, FG98, GNZO97, LB99]. Cartilage [GCM^{+}97].

Case [MP94]. Catch [DCC99, EA92, JRLK95, KKP^{+}95, LSC93, LJ92, WMC93]. Catch-and-Release [KKP^{+}95]. Catch-Per-Unit-Effort [LJ92]. Catchability [Pie97]. Catfish [BW99, BWD98, BWT96, Bay98, BMB98, CSM92, LGS^{+}96, LLK^{+}99, Mar90, MSC94, RO93, Sie94, TR90, TGC94, VPPH91, ZT98]. Catfishes [Kel95, ZTC94]. Catostomidae [PNSF99]. Caught [Jag99]. Cause [MCWN93, WH97a]. Caused [Kit90]. Causes [HES^{+}96]. Causeway [GGGD92]. Causeway-Induced [GGGD92]. Causing [FFCS96]. Cautionary [DV97]. Cayuga [FABH^{+}98]. Cell [HMR^{+}91]. Cellular [Eis95, Iwa96, Kol91]. Central [BDL97, CW95, HB95c, CM97, CH99, CWBC90, GJ95, JDW^{+}97, MCC^{+}91, NÜr95, PW91, SA92]. Centrarchid [Kol91, MCC^{+}91]. Centromere [LGS^{+}96]. Centropomine [TBS99]. Chain [BZBE98, ZTC94]. Challenges [Rey94]. Champlain [KL96a, MP97]. Change [Aue96, FPD90, GN99, Hea90, Hen90, Hol90, Hug98a, KSZ99, Mur93, RH91, RMC90, Ton90, VRW^{+}93, ZMH97]. Changes [BVBM91, BGR92, BMN91, CWBC90, DC90, DE92, Dyg90, FG98, FGW^{+}95, FP99, GM94, HT94, HSCK97, HBMB98, JVH99, JKM96, Lue90, MMH90, MM92b, MSC90, McC90, MMH94, MCB90, PD98, RHP90, Sta92, SSG94, TR90, WG94, You98, YL93]. Changing [Cou90, KN90, OJS97, PGS96]. Channel [BWD98, BWT96, BS97a, BMB98, CSM92, DHR96, Kel95, Mar90, Sie94, TR90, TGC94, VPPH91, ZT98]. Channelized [BG99]. Char [DHRK94, GN99, HGK^{+}94, HDS91, Jen94, LB94, Nil94, SSGE94, SSG94, SH94, UHG94]. Char-Possibilities [HGK^{+}94]. Characteristic [WGFB97]. Characteristics [BB92, BS97a, BM97, BRQ93, Gri91, HNW^{+}98, JCT97, KV96, KW99, Kwa91, LD98, LLD96, LD96, MSP91, ME94, MKR97, Nak94, PZW95, RD95, RKBC95]. Characterization [SSG93]. Chars [Rey94]. Chatanika [CB92]. Chemical [AA90, HNW^{+}98, LT92, Mon93, SI91, SJMM92]. Chemistry [LRC93]. Chemometric [VSSS91]. Chemometry [GNM92]. Chequamegon [HB99]. Chesapeake [LT96, BK93, CB90, DHMS94, LBCB96, LM91, WMWC93]. Chile [BR94a]. China [XC95]. Chinook [AREK98, ACH^{+}98, BGB90, BG90, BE92, BTM92, BDZ^{+}99, Bro92, DJG99, GHG92, GFD95, GWF99, HW90, MSP97, MP98, Mes94, Mur94, OW95, PQF95, QNM91, RGF90, RSL94, SJW92, UJ98]. Chloride [GC99, MHW^{+}91]. Choice [Bev96, CDL^{+}92, Joh93, Mes91, MS96b]. Cholinesterase [Sta92]. Chondroitin [Gri91]. Chromosome [GVMP^{+}95]. Chronic [DV99a]. Chub [KBSM90, MD97, TL96]. Chum [BE92, Hea91, SC99a, SC99b]. Cichlid [FF92, Lav92]. Cichlids [LT92, SB94]. Ciguatoxic [MKT92]. cis [MCWN93]. Cisco [AT99, BL97, CB92, GGGD92, MSC90]. Ciscoes [FGMG96, LB92]. Clair [HN90]. Clark [WBD^{+}94]. clarkii [IE98]. Class [DBCM92, Don97, Dyg90, FTB90]. Classical [Hou97]. Classification [Dol90, Ton93]. Clear [CH99]. Clear-Cut [CH99]. Clearwater [SPS96].

Climate [BD90, Cou90, DC90, FPD90, Hea90, Hen90, HM90, Hol90, JVH99, JE90, KSZ99, MMH90, Mur93, RMC90, SP90, Ton90, Woo94]. Climatic [Mei90]. Clinicopathological [NWGA99]. Close [PJ91]. Clupeid [BG94]. Co [Spa90, FPD90]. Co-Operative [Spa90]. Coast [CH99, XC95, BFOS95, GJ95, GSS90, JDW^{+}97, KP92, MT94a, NP94b, SC99a, VP98, WDRM90, WMWC93]. Coastal [BSEK99, CH99, Cou90, Gar92, HN96, HCG93, Ken98, KWE99, LP94, LB92, MBS^{+}92, MKR97, Pal94, PRD96, PS91, RES93, SH97a, WF94, WH95, WSB^{+}92, You99]. Cod [Tri98]. Code [VSGA94]. Coded [HSES98, IE98, Ken95, Mor95, Sig95, WGC97]. Codes [Sig94]. Coding [SAMW90]. Coeur [FWB^{+}99, GWF99, WGFB97]. Coho [BGB90, BE92, BTM92, BTS^{+}99, BTA97, Bra99, Bro92, BSWW98, BB91, Cou96, DS97, GH98, HN96, HBB^{+}98, Nie92, ODS92, OW95, QPS96, SPQ95, TJ95, VW98, Whi92, WE90, You99]. Cohort [CP94, DSRS93, XC95]. Cohort-Based [CP94]. Cold [BCWB97]. Collected [KWE99]. Collection [ASOM92]. Colonization [PB93]. Colorado [CCR98, KBSM90, MD97, MK91, ORM97, ORT^{+}98, OB98, Tyu90, TH91, BW94, BB98, BRCR93, BR91b, CC96, CR97b, DR94, DMM98, MK91, MB91b, ORM97, ORT^{+}98, OB98, SF97, TBCK91, Tyu90, TH91, YRB97, You98]. Columbia [BPL99, BR91a, BRN95, DJTH95, MS97, PHV^{+}91, QNM91, RBVP91, VPPH91, WPL95, Zim99, ZW99, BPL99, DP94, ENB92, Hea91, KW99, LM93, LB99, PPF99, PZW95, PBM93, PQF95, Pet94, PW99, PGG90, QGWF98, RGF90, RRBA92, SMC90, WZ99, WE90]. columnaris [Gri91]. Combination [SM90a]. Combinations [SEDP94]. Combined [WBI97]. Commensalistic [SLH96]. Comment [Cou96, JP97, Mor95, QPS96, Sig95, SR95d, WW95, Wil97]. Commentary [SMPS98]. Comments [BCJ97]. Commercial [LB94, Mac97, UHG94, UJ98]. Common [FG93, FG98, GNZO97, RJK^{+}99, WDFY97]. Communities [BF91, DS92, HB99, Mef91, PRD96, Ton90]. Community [Dol90, HTS92, HK90, HN90, PD98, RH91]. Comparative [BM90, NWGA99, TFS92]. Compare [VRW^{+}93]. Comparing [Bor97]. Comparison [CW96, Dib91, FSSH96, FF92, HBS90, LS93, Mon93, MVW99, ODS92, PRR94, PRL96, RCR93, VRF94, WRS^{+}97]. Comparisons [HK93, KHTC94, LBKD96, Par93b]. Compatibility [TN91]. Compensation [Eva90, Kol91]. Compensatory [HNW97, WHNW98]. Competition [DD98, FGB^{+}95, GS98a, MBD91, SK96, TL96, WPW94]. Competitive [BTS^{+}99, BM98, CR97a, PMF90, WP93]. Complexity [PLL92]. Composition [BM91, DE92, Dyg90, FGW^{+}95, FGMG96, GLHH99, HT94, LLD96, MH94, PRR94, Ret98, SR95b, TFS92, TR90, WHW96]. Compounds [BS96a]. Computer [CBB^{+}96, JROG90]. Computer-Aided [CBB^{+}96]. Concentration [CDAM90, FG98, LRC93]. Concentrations [BD90, BHSH96, BG94, DP94, FSO91, GM94, GC99, HW90, HNW^{+}98, MHW^{+}91, SH97b, WMSG90]. Conceptual [Hoc90, Ton90]. Condition [ABG93, CG97, FGW^{+}95, FGMG96, FH95, GT96, GGGD92, HCH96, JKM96, MT94a, PCGW90, SM90b, TBCK91, WK91]. Conditions [LPW^{+}95, SPQ95, WGFB97, ZWLB90]. Coney [PM99]. Confidence [RSBN98, Wil96]. Confinement [SC92b]. Confirmation [LA97]. Confirms [BS95b]. Conflict [JKO93]. Confluences [BG99]. Congener [WE92]. Congeners [ADHB98]. Connecticut [HNW^{+}98, WK90]. Connection [Bay98]. Consequences [DSMM91, DD98, LRW97, LRW98, PPJ98, RLC93, SJ91]. Conservation [BFH99, Bay98, Kel99, Lov98, Mef91, Wal95, War93]. Conserving [Fle92]. Consideration [WAW96]. Consistency [CAM95]. Constant [BW94]. Constitutive [ZT98]. Constraints [JP96, JE96]. Consuming [MKT92]. Consumption [BS93b, CW96, DS92, FF90, HM92, HA96, HK93, HM90, LPFA97, MHC95, RR99, RBM94, SM96a, VPPH91, WS91, YL93]. Contact [SC99b]. Contaminant [JMS^{+}98]. Contaminants [SJMM92]. Contaminated [ACH^{+}98, FWB^{+}99, GWF99, JHT96, WBD^{+}94]. Contemporary [Ker90]. Content [Dyg90, FP99, TBCK91]. Contents [GHR91, HT91]. Continuous [LF95, LRCB96]. Contraction [Mar90]. Contributing [ML99]. Contributions [MBS^{+}92, WF94, WMWC93]. Control [BB92, Lab93, PC98, Rab97, WDFY97]. Controlled [KSH^{+}94]. Conversion [KP91, VVF92]. Convict [LT92]. Cool [Bir98, MBAL94]. Cooper [BCSD97]. Copepods [MO91]. Coping [SG95]. Coral [Par93a]. Correcting [And95]. Correlated [MSP95]. Correlates [FOS93, MRM97, RJK^{+}99, Ste91, You99]. Corridor [MFL99]. Corroboration [PW99]. Corticosteroid [ODS92]. Cost [HMPV91, Li99, SC92a]. Costs [LRW97, LRW98]. Countergradient [Con90]. Coupled [VE99]. Covariances [Gra99]. Cover [BKM91, BB91, BBM91, Joh93, KBM^{+}98, MKR97, SPQ95, TW91]. Coverage [RSBN98]. Crabs [Fis99, SJ91]. Crappie [BBM90, ESM^{+}92, GDT^{+}90, HA96, PD94, SP97, SMD94, TMWD97]. Crappies [DNR^{+}94, GWJ94, Par93b, PJ91, TMD96]. crassus [FWJ96]. Crayfish [DRWD94, IE98]. Creek [BS96b, DE92, RA92, Wat99]. Creel [BBA98, RSBN98]. Criteria [BC97, BFB98]. Critical [HAHD94, SS90b, VP98]. Croaker [DCC99]. Croakers [GHR91, Nor91]. Crop [BKM91]. Cross [GNZO97, Sie94]. Cross-Reactivity [GNZO97]. Crosses [MSC94]. Crossings [WP98]. Crude [VVF92]. Crustacea [BFF93]. Crustacean [DHRK94]. Cryopreservation [BWD98, TGC94, TFW^{+}98]. Cryopreserved [CDL^{+}99, TGC94]. Cryoprotectants [TGC94]. Cucumber [MCWN93]. Culture [HGK^{+}94]. Cultured [Koh97, MKT92, VRF94]. Cultures [LF95]. Culverts [TFH99]. Cumulative [LLP^{+}94]. Curie [VSSS91]. Curie-Point [VSSS91]. Current [OS93, VP98, WGFS99]. Cut [CH99]. Cutthroat [BR91b, BM95, CHSM93, CH99, DR94, DV99a, DV99b, DPR^{+}99, FWB^{+}99, JMTC98, KOB99, KHR97, Lue90, MMT96, Mes91, TK94, VM97, WFL^{+}91, WGFB97, YRB97, You98]. Cutting [Ney93]. Cyanoacrylate [PJ91]. Cycle [JS95, McC90, RJK^{+}99]. Cycleptus [PNSF99]. Cycles [DV99a, TV94]. Cycling [SKH^{+}93].

Cyprinid [PM95]. Cyprinidae [MGW90]. Cytochrome [TMWD97]. Cytochrome- [TMWD97]. Cytogenetic [IMMP90]. Cytophaga [Gri91].

D [MS97]. Dace [GHG98, MZ97]. Daily [BB98, BR94b, DB95, HMPV91, HH93, HA96, KV96, MS96a, PBH92, PVE94, PF90, SM96a, TV94]. Dakota [GWJ94]. d’Alene [WGFB97, FWB^{+}99, GWF99]. Dam [MHR94, TBM95, IB96, PBM93]. Damage [BSS98]. Damming [WTP91]. Daphnia [SM96a, SBJ99]. Daphnid [LB93]. Darkblotched [NP94b]. Dart [MG91]. Darter [BBFW98, KR94b, SF97, TFH99, VBKS97]. Data [BA90, Goo95, GKM^{+}94, LJ92, MVW99, RHP90, SHC99, SM90a, WMC93, WRC90]. Date [PJN95]. Dauphin [SD91]. Day [BR91a, LLP^{+}94, PHV^{+}91, RBVP91, VPPH91, VVF92, Wil90]. Daylight [LW93]. Dead [GHG92]. Debris [BS97a, CST96, OS93]. Decline [CB90, Esh92, WH97a]. Declines [Mor92]. Declining [HES^{+}96]. Decouple [DIT97]. Deep [MBAL94]. Deepwater [HPH93]. Defaunated [PB93]. Defense [HC91]. Deficiency [FFCS96]. Definition [WF94]. Degradation [SH97a]. Delaware [GE91b, GT96, RRBA92, WW94]. Delayed [BLC98]. Delineation [GCMR99]. Delta [MHSM92, SMDC96]. Demand [JKO93, Li99]. Demography [JN97]. Demonstration [Hug98b]. Densities [KT94]. Density [AW94, BS97b, DCE97, FF90, HB95a, HN96, HW96, HY97, JKM96, LB93, PRR94, RLO^{+}94, RM92, SW95, WP93, YDS96]. Density-Dependent [WP93]. Dependence [HA96, MP93]. Dependent [JE90, JKO93, LRCB96, MH94, WDM90, WP93]. Deposition [BB98, DB95]. Deprivation [JM96]. Depth [BCJ95, Ham91]. Derivation [HSH91]. Dermal [SJMM92]. Deschutes [BDZ^{+}99]. Describe [FSSH96]. Description [GHW91, RC93]. Desert [LLP^{+}94]. Design [Ang91, KOH96]. Designation [MSC94]. Designed [GHG98, WM97]. Detection [BGB90, MS96b, MVW99, WSCL96]. Determination [CBB^{+}96, Mur94]. Determinations [CAM95]. Determine [DBCM92, JROG90, PH92]. Determined [DMWM96, MT91]. Determining [CAM95, Dib91, LJS^{+}97]. Development [AA90, BC97, BM90, BW94, CBB^{+}96, FK90, GE91a, HLH92, JMS^{+}98, Ken98, LLK^{+}99, LT92, OW96, PW99, PS99, RTA95, SH90, Sca94, Ton93, VWBM91]. Developmental [EAEA95]. Device [Fle92]. Diaries [CB95]. Did [DIT97]. Diel [Bea90, BIBV98, CH93, DV99a, Mat96, SM96a, SKR99]. Diet [Abr96, EO91, GLHH99, GE91b, HBB^{+}98, HD97, HLH92, LP94, MDS98, Mic97b, Nie92, OSNH95, PNSF99, PJN95, PMF90, Toh93, WBD^{+}94]. Dietary [FWB^{+}99, FABH^{+}98, KR94a]. Diets [HB95b, Har98, Hea91, KH93, KVD^{+}92, Mic97b]. Differences [ADHB98, BRN95, BTS^{+}99, RRBA92]. Different [BCJ95, BEC98, DS90, GE91a, GM94, GHG98, LM93, LB94, RES93, VRW^{+}93]. Differentiating [WMC93]. Differentiation [BFHB95, BSC97, CW96, MPG93, NP94a]. Difficult [HDL^{+}96]. Digestion [HW93]. Dimension [BT90]. Dimensional [LBBC95]. Dimorphism [MVW99]. Dinoflagellates [MKT92]. dioxin [JHT96]. Diploid [LGS^{+}96, Par93b, SSS^{+}99]. Direct [BS93b, HK90, HC91, HD94, JW98, MT91]. Discharged [KR94a]. Discharges [KH93]. Discriminant [VSSS91]. Discriminate [SC99a]. Discriminating [RE97b]. Discrimination [BMWB97, BZBE98, BHSH96, DMM98, FBHB97, GNM92, SD90]. Disease [BE92, Lei90, MM92b, Ste91, WE90]. Disease-Associated [MM92b]. Diseases [McG91]. Dispersal [JHT96, ORT^{+}98, RCF98]. Dissolved [BD90, CMMP98, DV99b]. Distance [BGR92]. Distances [LGS^{+}96]. Distribution [BS93a, BFH99, BR91a, BCJ95, BS96b, Cam98, CP97, CM97, DPR^{+}99, EKH93, FWLC96, Fle94a, GA93, HCH96, HCG93, Jen94, KCSF91, KHR97, Lue90, Lyo97, MMH94, NPLJ92, OW95, PRS95, RLCB98, RSL94, SRD95, SAG95, SR95c, SPQ95, SH94, SH98, TF99, TH91, UJ98, War92, WK90]. Distributions [CZG90, KR96, Mur93, PM95, ROI^{+}97, TFS92]. Disturbance [ELAD97]. Disturbances [LLP^{+}94]. Disturbed [MRM97]. Diurnal [HC91, SV95]. Divergence [DLF97, QGWF98]. Divergent [BS95b]. Divers [SV95, SW95]. Diverse [BCB96]. Diversion [VM94]. Diversity [BMWB97, Bur99, Gra93, HSCK97, NCFG97, RES93, SPR94]. Diving [Dib91]. DNA [BPL99, BS95b, BZBE98, BSEK99, DLF97, EKNP96, GRKM93, GSR97, GKM^{+}94, GKF94, LT96, MFL99, MK96, NP94a, NCFG97, PPF99, RKF91, SC99a, SKG92, SPR94, SR95d, WW94, WW95, WHW96, WTC92, WKW98, WDFY97, WGGM91, WMWC93, ZTC94]. Do [ADHB98, HBB^{+}93, Kol91, MKT92, WGFS99]. Dogfish [BTM92]. Dolly [BHJ^{+}95, KWE99, LA97, UMT96]. Dominance [CR97a, WK99]. Dominant [DBCM92]. Dominated [CZG90]. Double [Fle94b, HNW97, HWW^{+}97]. Double-Entry [Fle94b]. Double-Pass [HWW^{+}97]. Downstream [HONCS98, Lut95, PBM93, TBM95]. Drainage [BM97, BM98, HIG^{+}99, KMNF94, LD98, NPLJ92, OKB^{+}92, REM92]. Drainage-Wide [REM92]. Drainwater [SJW92, SO95]. Drawdown [AKKH99]. Drift [OS93, YRB97]. Drinking [GC99]. Drum [BSRW90, NW93, NW95, VRF94]. Drums [Bum91, GRKM93, MT91, RSV95]. Dual [SW95]. Dual-Beam [SW95]. Due [KR96, McC90, Mei90, WTP91, Hug98a]. Dungeness [SJ91]. during [ACB^{+}99, Bir98, BG94, Don97, DWRM97, FBHB97, GSC93, GWF99, Har98, Ken91, MSP91, MK91, SH90, SG94, TR90, VVF92, Zim99]. Dwelling [Jen94, RO93]. Dynamic [BT90, BSEK99, PF90]. Dynamics [BRN95, BCSD97, CS99, CRRH93, DJTH95, Fle90, Fro97, GLHH99, Kel99, KW97, Mac97, PP94, Qui90, RC93, RN92, Smi91, VRC93, WDRM90, WKPK93, Wat99]. Dysprosium [GA93].

Early [Arg95, BW94, BBFW98, CCR98, CZG90, DH94, GM93, Hol90, LT92, Nak99, SC98, TMD96]. Earth [GA93]. East [CW95]. East-Central [CW95]. Eastern [OWJH98, KCSF91, PS91]. Ebb [LMC95]. Ecological [Ang91, Gal92, Kwa91, LS93, Ort96]. Ecologically [ADHB98]. Ecology [AT99, Bay98, BH93, Dev93, Fis96b, HPH93, HBB^{+}93, Ker90, Lav92, Lov98, Par93a, Pow94, RGF90, Sau91, Sny96, SM96b, Wis95]. Economic [CB91, LS97, Sca94]. Ecosystem [JHT96]. Ecosystems [Gre99, Koh97, RHP90]. Edge [Ney93]. edition [KOH96, McG91]. Eels [CDAM90, FWJ96, Oli97, SAG95]. Effect [BB92, BMEV95, BWT96, BV97, DSD99, Fis96a, Fis99, FF90, FTB90, FABH^{+}98, GVMP^{+}95, HONCS98, Hub99, JH99, KT96, KS97, KVD^{+}92, Lab93, MW98, MG91, MK90a, MB91b, SW95, SB94, SB90, Swi90, Swi93, WEH95, WK99]. Effects [Abr96, AREK98, AT99, ABM98, AW94, BTM92, BW94, BRCR93, BD90, BBFW98, BS99, BLC98, Cam98, CV99, CC96, CDL^{+}99, CR97b, CMMP98, Cou96, DC90, DS92, DW92, DCC99, DV99a, DV99b, ESM^{+}92, EWE97, Eva90, FWB^{+}99, FNAR97, GH98, GA93, GGGD92, GC99, GDT^{+}90, HM92, HN96, HW96, HK90, HBB^{+}98, HM90, HY97, HC92, JMS^{+}98, JW98, JCT97, KKP^{+}95, KL96b, Koh97, LJS^{+}97, LRCB96, LLP^{+}94, LD96, LRW99, LS97, LT92, LW93, MZ97, MC93a, MT94a, MS99, MR93, MSP97, MP98, Mes94, MCB90, MS96b, OS93, PM92, PT94, PH92, QPS96, Rab97, RR99, RMC90, RB92, REM92, SP97, SSGE94, SSG94, VM94, WBI97, WV95, WS99, WL94, WA91, WBD^{+}94]. Efficiencies [KP91]. Efficiency [ASOM92]. Effluent [KH93]. Effluents [Ken91, KR94a]. Effort [HH93, LSC93, LJ92]. Egg [AA90, BM90, BBFW98, BL97, FWLC96, FTB90, NAC99, OFM91, Tri98]. Eggs [BWD98, BL97, CWK99, MB96, VSSS91, WL94]. Eigenmann [BR94a]. Elasmobranch [GCM^{+}97]. Electrofishing [And95, Hug98a, LD97, SB99]. Electromyogram [DMWM96, HDL^{+}96]. Electronic [Sig94]. Electrophoretic [CSM92, SSG93]. Element [BHSH96]. Elemental [VE99]. Elements [GA93]. Elevated [WFL^{+}91]. Elevation [RK96]. Elicit [WHNW98]. ELISA [HJWS99]. Embryonic [OW96]. Embryos [BM90, FC93]. Emerge [CG97]. Emerged [BTS^{+}99]. Emphasis [SE98]. Empirical [BTA97, HW93, KSZ99, MS96a, PS91, PRL96, RHP90, WRS^{+}97]. Encroachment [LMC95]. Endangered [BLW98, MD97, MS99, OB98, TFW^{+}98]. Endemic [RK96]. Endocrine [RJK^{+}99]. Energetic [DS97, SC92a, SHDB99]. Energetics [CG97, LS93, LRW99, Men93a, RSS^{+}93, RLO^{+}94, WGFS99]. Energy [DS90, Dyg90, FP99, FS96, HB95a, HC91, JKM96, LB93, RLO^{+}94]. Engines [KOB98]. England [MSP95]. English [ACC94, Dyg90]. Enhance [BLW98]. Entrainment [TZF93]. Entry [Fle94b]. Environment [CBYH96, EN90, LJP93, SG95]. Environmental [BB98, DGHP99, DFT94, Don97, HC96b, LPW^{+}95, MRM97, ML99, PRD96, PPD95, SH97a, Ste91, VRW^{+}93, WPM99, YC96, You99]. Environmentally [Gri93]. Environments [MK90b]. Enzyme [ZT98, HJWS99]. Enzyme-Linked [HJWS99]. Epilepsy [SB99]. Epilimnetic [JO91]. Episodes [GSC93]. Episodic [HSCK97]. Episodically [FC93]. Epithelialization [Mar90]. Epoch [PH92]. Equal [KT94]. Equilibrium [DRF90]. Equivalent [ADHB98]. Erie [HN90, MFL99, BD90, HM92, Har98, MTK^{+}96, PVE94]. Ernest [EMC90]. Errata [Ano94]. Esocids [WS91]. Essential [GCMR99]. Establishment [WE92]. Estimate [SEDP94, WMWC93]. Estimated [RBVP91]. Estimates [BFOS95, BBA98, BM94, DHMS94, Gra99, HSH91, LW93, PVE94, Pie97, RB94, SW95, TB95, XC95]. Estimating [CP94, Fis96a, HB95a, HWW^{+}97, LJP93, LB93, Men93a, PRR94, Pet94, WF94]. Estimation [BS93b, CJ95, FNAR97, FF92, HMPV91, HH93, JMBK98, RNB95, VB94]. Estimator [JROG90]. Estimators [RSBN98]. Estuaries [Ayv99, HC96a, HCG93]. Estuarine [ASOM92, BFF93, FSW96, HB95b, HJDI98, LP94, Nor91, RTA95, WS90, WM97]. Estuary [ACH^{+}98, LPFA97, MHSM92, OSNH95, RS94, RLCB98, RT99, TSM98, DH94, DWRM97, JMMC93, MM92b, MSP95, MM95, SBH97, VM94]. Eurasian [MVW99]. Europium [GA93]. Euryhaline [KS97]. Evacuation [BC96, HB94, HW93, RN98]. Evaluated [RSBN98]. Evaluating [BL97]. Evaluation [ABG93, BC97, BA90, BFB98, CBB^{+}96, DBCM92, Goo95, HKC^{+}93, KVR96, MO99, SR95c, WH97b, WM97]. Event [PH92]. Evidence [BS95a, BSS98, BFHB95, Con90, GB96, Hug98a, JCP96, Kel96, Mur94, RK96, SO95, TKOS92]. Evolution [Bur99, Lav92, Win94]. Examination [CDK97, Plu90, SJ99, VE99]. Examine [CB95]. Examining [ADHB98]. Example [MVW99, VB94]. Examples [DV97, HB93]. Excavated [SK91]. Excretion [MWW^{+}94]. Execution [Ang91]. Exercise [CBYH96]. Existence [BS95b]. Exotic [BW91, FWJ96]. Experience [WEH95]. Experiment [HK90]. Experimental [BS97b, CV99, Ham91, MC93b, RCR93, SPQ95, TKOS92]. Experimentally [PB93]. Experiments [And95, Ang91, Bev96, Fle90, WP93]. Expert [HAHD94]. Explanation [WDM90]. Explicit [BK93]. Exploitation [KW99, Wal95]. Exploited [HDS91]. Exposed [FG98, LRW99]. Exposure [OW96, VVF92]. Expression [LGS^{+}96]. Expulsion [BW99]. Extenders [CDL^{+}99]. Extension [CSK99]. Extent [WK99]. External [ABM98, MMSW92]. Externally [CF99]. Extirpation [WTP91]. Extreme [HY98]. Extremes [DFT94]. Exxon [MHS^{+}99, BMEV95, BSS98].

F [EN90]. F. [WRC90]. Facilities [KOH96]. Facility [Aue96, TZF93]. Factor [NÜr95]. Factors [ASOM92, BJ92, BB98, BG99, Bro92, CR97a, ELAD97, HAHD94, HSH91, LAD97, LO95, ML99, Mic97a, PQF95, SHK93, SI91, WPM99, WMSG90, CRRH93]. Failed [Mef91]. Failure [FFCS96, Fle92, ML99]. Fall [BM95, DJG99, JMTC98, PQF95, QNM91, TWC91]. Family [SS93]. Fare [WGFS99]. Fate [BL97, Gar92]. Fathead [Abr96, Sta92]. Fatty [GHR91, GNM92, TR90, VRF94]. Fauna [FK90]. Fear [MR95b]. Features [BC91]. Fecundity [CB92, EJ95, Jes93, KGJ92, LBCB96, OR92, RKR92]. Feed [EWE97]. Feeding [AREK98, AT99, Bro92, CR97b, CSK99, DW92, FP99, GT96, HPH93, Hea91, HD94, KR94b, LRCB96, MP93, Mes91, NW92, OS93, OSNH95, PW99, PHV^{+}91, RGF90, RB92, SC92a, SHK93, SM96a, SPS96, SLH96, SM96b, WH97b, WHNW98, Wis95, YDS96]. Female [BFOS95, Dyg90, Fis99, OFM91, SSS^{+}99, SMC90, WKSS95]. Fertility [BFBC92, CS99, NAC99]. Fertilization [BWD98, CDL^{+}99, TGC94]. Field [ABG93, BMEV95, ESM^{+}92, HMPV91, LPW^{+}95, LCR98, MZ97, MG91, SHDB99, WS91, WP93, WRC90]. Fifteen [SSEO99]. Filaments [GNZO97]. Fillets [TR90]. Fin [RB94, VE99]. Fine [CST96]. Fines [YHW90]. Finger [FFCS96]. Fingerlings [WV95]. Fingerprinting [WGGM91]. Fingerprints [GKF94, RKF91]. Fire [Gre99]. First [Car92, CC90, Don97, MZ97, MTK^{+}96, ML99, MB96, SG94]. First-Year [CC90, MTK^{+}96, ML99]. Fish [ABG93, Ayv99, BMEV95, BS96a, BS97a, BS93b, Bor97, BF91, BH93, BLP^{+}94, Bre93a, BC94, BFM97, CG97, Cas90, Cha93, CP94, CDL^{+}92, Dab90, DSRS93, DE92, Dev93, Dib91, DIT97, DRWD94, DFO94, Dol90, DV97, ELAD97, EN90, FAC97, Fle90, Fle92, Fle94b, FF90, FPD90, FF92, Gar92, HBB^{+}93, HB95a, HT91, HK90, HW93, HSCK97, HWW^{+}97, HK93, HM90, HB99, Hou97, Iwa96, JMBK98, KKR^{+}94, Kel99, KOB98, Ken90, KSZ99, KV96, Kol91, KOH96, LJS^{+}97, LAD97, Lin96, LS97, LO91, LC96, Lov98, LJP93, LW93, Lut95, Mac97, MMH90, MRM97, MP93, MGW90, MJ93, McG91, MWW^{+}94, MVW99, MIPGV97, Mur93, NÜr95, Pal94, PRD96, PD98, PLL92, PRL96, Plu90, PC98, PRS95, Qui90, RH91, Rei92, RM92, REM92, SK91, SHC99, SH97a, SAMW90, SR95b]. Fish [Smi91, SW95, SLH96, TCCT96, TN91, Ton90, TBM95, VSSS91, VRC93, VRW^{+}93, VPPH91, VRF94, WP93, WP98, Wea90, WMC93, WG94, WS99, WTC92, WK99, Wis95, YDS96, YL93, RMC90]. Fish-Conserving [Fle92]. Fished [DRF90]. Fisheries [Ayv99, CJ95, CB91, Con98, Cow97, DC90, DCC99, Ens94, EN90, GJ95, HA96, Hea90, IMMP90, JKO93, LSC93, PF90, SC99a, Spa90, Woo94]. Fishery [MBS^{+}92, PS91, Sai92, UJ98, WF94, WHW96, WH97a, WMWC93]. Fishes [AKKH99, ASOM92, Arg95, Arm92, BFB98, BG99, BG94, BKN90, Bur99, CCR98, CZG90, Dan95, DSMM91, DSB98, EO91, FGB^{+}95, FSW96, GS98a, Gra93, HH93, HSH98, IMMP90, JCT97, JMMC93, KN90, Koh97, Kol91, Kol95, Lav92, MD97, May98, ML99, MMH94, MB96, Nak99, Par93a, PB93, PHV^{+}91, PPD95, RLO^{+}94, RCF98, RKBC95, SSEO99, Sco93, SP90, SR95a, SJ99, SMHP92, SV95, SH97b, Ste91, TBS99, VSSS91, WPW94]. Fishing [Bor97, PD98, SJ91]. Fishless [Lue90]. Fishways [KOH96]. Five [BM90]. Fixed [SV95]. Fl [PW91]. Flaming [TL96, YL93]. Flathead [KMNF94, RO93, Toh93]. Flattery [Jag95]. Flexibility [DFT94]. Flooding [PLL92]. Floodplain [SK91]. Florida [BZBE98, CZG90, FC97, GGM95, LRC93, MM92a, MC93c, MT94b, PW91, SHC99, SC98, ZHP99]. Flounder [GHR91, HR95, HC96a, MM92b, MT94a, Mon93, Phe92, RS94, RN92, RA92]. Flow [BG90, BCJ97, BR96, BFB98, Fle90, GH98, JP97, Ken91, LMC95, LBBC95, PE94, See93, TBM95, VM94, Wil97]. Flowing [KBM^{+}98]. Flows [BCJ95, Fle94b]. Fluctuating [BW94, MIPGV97, SBSN95]. Fluctuations [PE94]. Fluorescent [GCM^{+}97]. Fluvial [KJ95, LBKD96, Swa97]. following [ELAD97, HT94, Lab93, WDB^{+}96]. Food [AT99, Bea90, Bev96, CH95, CW96, DF91, DS92, DD98, ESM90, FH95, FF90, Fro97, GH98, HH93, HKC^{+}93, JO91, JDW^{+}97, JM96, LM93, LB94, MGH92, MP94, Men93a, PM92, RR99, RBM94, SCJ93, SKH^{+}93, SM96a, VBKS97, VVF92, WS91, WPSD99, Zim99]. Foods [MC93c]. Forage [Bre93a, TN91]. Foraging [Bre93a, CH93, MW98, Nie92, RA92, SBJ99, WEH95]. Forecasting [DC90, Mur93]. Forest [Arm92]. Forested [BFH99, Gre99]. Forests [KBM^{+}98]. Fork [NPLJ92, WBD^{+}94, SG94]. Form [BHJ^{+}95]. Formation [MBAL94]. Formed [FBHB97]. Formulated [KVD^{+}92]. Fossil [VM97]. Found [GVMP^{+}95]. Fountain [BBFW98]. Four [DRWD94, HCG93, MWW^{+}94, Ret98, RKBC95, SKG92, WTP91]. Fourier [FBHB97]. Fox [AA90, KOB98]. Fragment [BZBE98, LLK^{+}99]. Fragmented [RM95]. Framework [BS99, Ton90]. Franklin [MS97]. Frederick [EMC90]. Free [CCD90]. Free-Ranging [CCD90]. Freezing [CDL^{+}99]. French [BDL97]. Frequencies [EA92]. Frequency [SM90a, TFS92, UMT96]. Freshwater [BG94, Car92, CBYH96, CB91, Cou90, Dan95, Kel99, KS97, LSC93, NCFG97, SM91, TIT96]. Fruitful [Lov98]. Fry [CST96, EMC90, EN90, FBHB97, GKM^{+}94, Hoc90, Ker90, WBD^{+}94, WRC90]. Fujian [XC95]. Function [Fis96b]. Functions [SMHP92]. Furunculosis [BE92]. Future [BH93]. Fuzzy [Sai92]. Fyke [HC92].

Gags [KC98]. Gain [BMB98]. Galaxias [AW94, BR94a]. Gambusia [Hub99]. Gamete [FOS93]. Gametogenesis [OWJH98]. Gametogenic [JS95]. Gape [DSB98]. Gar [JN97, SKR99]. Gas [CMMP98, Lut95]. Gastric [BC96, HB94, HW93, RN98]. Gastrically [AREK98]. Gear [PGS96]. Gears [WMC93]. Gene [BG90, GRKM93, LGS^{+}96, SAMW90, TMWD97, WTC92]. Genera [SS93]. Generalized [BFB98]. Generations [BSS98]. Genetic [BGB90, BG90, BE92, BM97, BSC97, CSM92, DLF97, JLV96, Ken90, LT96, LA97, LLK^{+}99, MPG93, MFL99, MK96, MSC94, OSBS98, PKM93, RMSL92, SC99b, SSG93, SE98, TBS99, WE92, WSCL96, WDFY97, WE90, ZTC94]. Genetically [BS95b]. Genetics [DNR^{+}94, Nil94]. Genome [GGM95, LB92]. Genotype [MR93]. Genotypes [GRKM93]. Genotypic [QGWF98]. Genus [HJWS99]. Geographic [ACB^{+}99, DPR^{+}99, EKNP96, KR96]. Geologically [BSC97]. Geology [NPLJ92]. Geomorphic [KHR97]. Geomorphology [BFH99, NPLJ92]. Georges [RN92]. Georgia [SMC90, BS93a, KP91, OWJH98, SM90a, WH95, WL94]. Gila [DMM98]. Gill [HC92, PS99, SBRH94]. GIS [KR96]. Gizzard [ADHB98, BS97b, BS99, DS92, DS96, DSL98, GS98b, GDT^{+}90, HH93, Mic97a, Mic97b, PD94, SM96a, TN91, YDS96]. Glacial [GWJ94]. Glacially [BFHB95]. Glacier [VM97]. Global [Car92, Hen90, HM90, KR96, MMH90, McC90, PD98]. globiceps [BR94a]. Glucuronyltransferase [KKR^{+}94]. Go [HBB^{+}93]. Gobies [RK96]. Goby [SM96b]. Golden [BW91, KV96, Mat96, Mat96]. Goldeyes [DBCM92, Don97]. Gonadal [HT94]. Good [WRC90]. Gorge [TL96, YL93]. Gradient [BS96b, KT96, See93]. Gradients [HC96b, RH91]. Grand [CCR98, FTB90]. Granite [BS96b]. Granulation [Mar90]. Graph [Sai92]. Graphical [CAM95]. Grass [BWTM90, CP97, KC98, LB99]. Gravel [BLB98]. Gravels [KCSF91]. Grayling [BM98, DGHP99, Hug98a, Hug98b, KJ95, LBKD96, OS93, WSB^{+}92]. Graysby [PM99]. Great [BF91, CBFS90, JE90, LS93, RH91, SS90a, BR94b, FFCS96, FAC97, HM90, Hol90, KT96, MMH90, MPG93, Nak99, RS94, SHK93]. Green [RN98, VMM99, FH95, Tyu90, TH91]. Greenback [WFL^{+}91]. Greenland [HB95c]. Gross [FSW96, KP91]. Ground [PGG90]. Groundwater [BR96]. Group [MSC94, RB92, RBL92]. Group-Size-Mediated [RBL92]. Grouper [BDCMSC99, EWE97, Goo95, VB94]. Growing [Con90, Ney93]. Growth [Abr96, AREK98, BE92, BSRW90, BDZ^{+}99, BB98, BV97, BS93b, BK93, BM94, Bre93b, BEC98, BLW98, Cas90, CH95, CC90, CV99, CW96, Con90, DSRS93, DGHP99, DH94, DW92, DS90, DV99a, DV99b, DIT97, Dyg90, Eds99, EWE97, FDGG92, FGB^{+}95, FF90, GS98b, GA93, GM93, GE91a, GB93, GGGD92, GI96, GN99, HAHD94, HM92, HB95b, HN96, HW96, HNW97, HC93, HK93, HM90, Hug98a, IE98, Jen94, JN97, KBM^{+}98, KSZ99, KL96a, KBF99, KOB99, KT94, KVD^{+}92, LPW^{+}95, LRW99, LCR98, MTK^{+}96, MDS98, MW98, MT94a, MG91, MSC90, MSP95, MK90a, MSP97, MP98, MK90b, Mic97a, Mic97b, MCB90, MVW99, MBD91, MT94b, Nie92, Nil94, OJS97, Oli97, ORM97, PM92, PBH92, PD99, PNSF99, PW91, PJN95, PMV98, PM99, PM97, PvdH99, PMF90]. Growth [PPD95, RR99, RKR92, RM92, RLCB98, RN92, RSV95, RA92, RKBC95, SH90, SSS^{+}99, SM90a, SMHP92, SPQ95, SH94, Swi93, Swi95, SD91, TR90, TGC94, Toh93, TMD96, TH91, UHG94, VB94, VVF92, WS91, WEH95, WP93, War92, Wea90, WPW94, WHNW98, WA91, ZMH97, ZJ93]. Gulf [BDCMSC99, GCMR99, Gra91, HC96a, KP92, KC98, NW93, RW92, RLCB98, BSRW90, CDK97, DLF97, FC97, GRKM93, GSR97, GI96, KS97, MC93c, SC98, VP98, ZHP99]. Gynogenesis [IMMP90].

Habitat [ACB^{+}99, AW94, BWTM90, BC97, BC91, BS97a, Bev96, BFB98, BR91b, BM95, BB91, BBM91, CCD90, CW95, CB90, Cou90, DCE97, DW92, DFO94, DV97, FWLC96, FS98, GCMR99, GN99, HC93, HBS90, HD97, JMTC98, JCP96, JCT97, KR96, KSZ99, KT96, KVR96, LAD97, LO91, MMT96, MMH90, MR95a, MGH92, MBAL94, Mat96, ME94, Mei90, MS96b, MR95b, NPLJ92, PBM93, PLL92, PG96, PE94, PM95, QTBG99, RH91, RM95, RSL94, RA92, RW99, SS90b, SKR99, SR95c, Swa97, TCCT96, Ton93, TH91, WH95, WW96, WK91, You98, ZMDE98]. Habitats [Arm92, GHG98, JCT97, LP94, LS97, NBGF93, NCFG97, RGF90]. Habits [Bea90, CR97b, JDW^{+}97, KR94b, LM93, RA92, WPSD99, Zim99]. Habituation [KSH^{+}94]. Haddock [RN92]. Hafnia [TKOS92]. Hake [EAEA95]. Hampshire [RS94]. Handbook [Ort96]. Handling [CBYH96, DS97, SMDC96, WV95]. Hanford [DP94]. Hardness [MSP91, MR93]. Harvest [BBA98, HSH91, RES93, RO93]. Harvesting [REM92]. Harvests [BSEK99]. Hatch [NAC99]. Hatcheries [BDZ^{+}99, PVE94, PQF95]. Hatchery [BTM92, CAB^{+}91, DW92, GKM^{+}94, HJLM99, JKO93, Ken91, KR94a, MMSW92, MPG93, MGM93, MSP97, Mes91, MIPGV97, NCFG97, PBH92, PMSM97, SWD91, SPS96, Sie94, WS99]. Hatchery-Dependent [JKO93]. Hatchery-Produced [BTM92, SWD91]. Hatchery-Reared [MMSW92, MGM93, MSP97, PBH92, PMSM97, Sie94, WS99]. Hatching [PJN95, TMD96]. Hawaii [HPH93]. Hawaiian [RK96]. Head [MHR94]. Headwater [HSCK97, JMTC98, Kel95, ME94, SR95a, SR95b, VM97]. Healing [Mar90]. Health [ABG93, Hou97, SSGI99]. Heart [GNM92]. Heat [BT90, BMB98]. Heavily [HDS91]. Heavy [CR97b]. Hecate [Hea91]. Held [NAC99]. Hematocrit [MZ97]. Hematological [Hou97, MM92b, MHW^{+}91]. Hematopoietic [LGRF90]. Henry [SG94]. Here [HBB^{+}93]. Heritable [JCP96]. Heritage [PKM93]. Herring [BHSH96, DF91, FP99, GSS90, Jes93, Lin96, McG93, PGG90]. Heterobranchus [BW99]. Heterochromatin [ZT98]. Heterogeneity [AD93, MS97, SC99b]. Heterologous [GNZO97]. High [DV99a, DV99b, LT96, LLP^{+}94, LBKD96, RK96, SC99b, WBI97, GN99]. High-Arctic [GN99]. High-Elevation [RK96]. High-Temperature [LBKD96]. Highlands [SE98, SJ99]. Highly [RKF91]. Histological [NW95]. Historical [Mur93, VM97, WG94]. Histories [BWH93, Win94]. History [BRQ93, MHSM92, PG96, RD95, SSP96, SC98, VRW^{+}93, You99]. Hog [JPL93, MR95a]. Hole [BM98, LBKD96]. Homing [BS95a, HSES98, HSH98, PQF95, QNM91]. Hooked [DR96]. Hormonal [FOS93]. Hormone [Whi92]. Hormones [CDAM90]. Host [Swi91]. Host-Size [Swi91]. Hudson [DWRM97, DH94, JMMC93, MM92b, NBGF93, WDRM90, WMWC93]. Hugo [DFO94]. Humboldt [NPLJ92]. Humpback [CB92, KBSM90, MD97]. Humper [BCB96]. Huron [Esh92, MFL99]. Hurricane [DFO94]. Hurry [Bre93b]. Hybrid [BM91, BW91, DDS96, DSL98, HAHD94, HNW97, LGS^{+}96, PW91, WE92, WA91, WKSS95]. Hybridization [BGB90, DNR^{+}94, KMNF94, MPG93, MCC^{+}91, PV97, PB98, SMD94, TMWD97, WSCL96]. Hybrids [EHBP99, TMD96]. Hydroacoustic [LW93]. Hydroacoustics [SW95]. Hydrodynamic [LBBC95]. Hydroelectric [Aue96, DJG99, MHR94, TZF93, TBM95]. Hydrologic [JVH99]. Hydrological [Ort96, ROI^{+}97]. Hyperthermia [SR95a, SR95b]. Hypervariable [LT96]. Hypolimnetic [AT99]. Hypolimnion [SS90a]. Hypomesus [TSM98]. Hypothesis [TWC91, BHJ^{+}95]. Hypoxia [SR95a, SR95b].

Ice [JMTC98]. Ichthyofauna [WM97]. ICP [VE99]. Ictalurus [MSC94]. Idaho [BS96b, FWB^{+}99, GWF99, GWL98, PPF99, RM92, SPS96, SG94, TK94, WAW96, WGFB97]. Idealized [Wil96]. Identification [BPL99, BFBC92, BLP^{+}94, CSM92, DHP99, Sie94, SR95d, TSM98, TBS99, WW95, WRS^{+}97, WKW98, WGGM91]. Identify [Ret98]. II [CRRH93, LD96]. III [GNZO97]. Illinois [PW91, PPD95, ROI^{+}97]. Immersion [SWD91]. Immunological [GNZO97, SD90]. Immunosorbent [HJWS99]. Impacts [Bor97, BFB98, BLB98, PRD96]. Impaired [SO95]. Implanted [AREK98, BW99, KL96b]. Implications [BFH99, CC96, CR97a, Con90, DS92, DDS96, HA96, Hea90, Hen90, ML99, ME94, Mef91, MCB90, PD98, PD99, Qui90, RLO^{+}94, SS90a]. Importance [Bev96, HC91, JW98, Pet94, TW91]. Impounded [BRN95]. Impoundment [AD93]. Impoundments [HC93]. Improved [LF95]. Incisions [PJ91]. Incorporate [DRF90]. Incorporating [DV97]. Increase [CST96]. Increased [ACH^{+}98, BG94, Lab93]. Increment [BT90, BB98, FNAR97, HLM97, SM90a]. Incremental [LBBC95]. Increments [PBH92, RKBC95]. Incubation [FBHB97]. Independence [BCJ97, JP97, Wil97]. Index [ABG93, BF91, OKB^{+}92, SHC99, WPL95, ZW99]. Indicate [SSGI99]. Indicators [GGGD92, Hou97, Rei92, SH97a]. Indices [SM90b]. Indirect [HK90, JW98]. Individual [Bre93b, Cha93, CR97a, CRRH93, DD98, FH95, HJLM99, RC93, VRC93, VRW^{+}93]. Individual-Based [Bre93b, Cha93, CR97a, CRRH93, DD98, RC93, VRC93, VRW^{+}93]. Individually [KBF99]. Individuals [CG97]. Induced [AKKH99, BGR92, FPD90, GGGD92, ODS92, VSGA94]. Induction [BBM90]. Inductively [VE99]. Industrialized [SSGI99]. Industry [HAHD94]. Infectious [LGRF90]. Infer [Hug98b]. Inferred [FGB^{+}95, GSR97]. Inflated [MHC95]. Influence [BC91, BS97b, CF99, DR94, DGHP99, GI96, JVH99, KH93, LRC93, LCR98, Lyo97, MHW^{+}91, MSP91, Mic97b, OKB^{+}92, PMSM97, PLL92, PJN95, RHP90, RM92, RKBC95, SH90, SSEO99, YRB97, ZMDE98]. Influenced [DSB98, YDS96]. Influences [Arm92, KHR97, RN92, SR95b]. Influencing [ASOM92, ELAD97, LAD97, WMSG90]. Information [ACB^{+}99, KR96]. Informative [RKF91]. Ingestion [MJ93]. Inhabiting [BSS98, KP92, SO95]. Inheritance [VW98]. Initiated [LB94]. Initiation [LF95, TBM95]. Injected [GA93, TJ95]. Inland [Con98, GB96, Mac97]. Inner [Phe92]. Innovative [BH93]. Input [HBMB98]. Insights [HC91]. Instability [EAEA95]. Instance [FGB^{+}95]. Instream [Cou96, KBM^{+}98, LBBC95, QPS96, SPQ95]. Intake [Fle90, SCJ93, SM96a]. Integrated [BLP^{+}94]. Integrating [ACB^{+}99]. Integration [Fro97]. Integrity [BF91, CWBC90, OKB^{+}92, SHC99]. Intense [PD98]. Intensity [Bro92, Hub99, Joh93]. Intensive [SJ91]. Interaction [Abr96, GFD95]. Interactions [BM98, DR94, DW92, HD94, PM95, PD94, PMF90, RCR93, Sco93]. Interactive [WV95]. Interannual [FGMG96, Hug98a, Hug98b]. Intergeneric [BW91]. Intergrades [WKW98]. Interior [Hug98a, Hug98b, SE98]. Interlake [LS93]. Intermediate [GNZO97]. Intermittent [LRCB96]. Internal [GA93]. Internally [Cue94]. Interpretation [OSBS98, RHP90]. Interseasonal [FGW^{+}95]. Interspecific [FGB^{+}95, MW98, MBD91]. Intertidal [BMEV95]. Interval [RSBN98]. Intervals [Wil96]. Intestinal [FF92]. Intra [FGW^{+}95, FGMG96]. Intra- [FGW^{+}95, FGMG96]. Intracohort [JP96]. Intraspecific [WW96]. Introduced [MD97, MO91, MMH94, MCC^{+}91, PV97, SS91, WK99, WE92, WSCL96]. Introduction [EN90, Lue90, MSC90, RMC90, SB90, Toh93]. Introductions [Sca94]. Introgression [CHSM93, GGM95]. Intrusion [VM94]. Invading [TSM98]. Invasion [JE90, Mef91]. Invasive [WTC92]. Invertebrate [FPD90, HTS92, JMBK98, RO93]. Invertebrates [FWB^{+}99, HBB^{+}98, WBD^{+}94]. Investigation [CW96, ESM^{+}92]. Involving [WE92]. Ionic [Iwa96]. Ions [LD96]. Island [BSWW98, SA92, Oli97]. Islands [DHR96]. Isochronal [RTA95]. Isolated [RK96]. Isolation [ZTC94]. Isotope [DHP99, VHRR98]. Isotopes [FS96]. Isozyme [LGS^{+}96, MIPGV97].

J [EN90, WRC90]. Japanese [MHC95]. Jersey [JDW^{+}97, Gra93, RA92]. Jet [TJ95]. Jet-Injected [TJ95]. Joaquin [MM95, MHSM92, SJW92, SBH97, TSM98]. Jocassee [DF91]. John [BR91a, LLP^{+}94, PHV^{+}91, RBVP91, VPPH91]. Johnny [SF97]. Joint [MSC94]. Joseph [EMC90]. Julian [EKH93, Wil90]. June [Har98]. Just [Bre93a]. Juvenile [AREK98, ACH^{+}98, BRTV95, BS96b, BS99, Bro92, BM91, BK96, CDAM90, CC96, CCR98, CF99, Cou96, DS97, DCE97, DR94, DH94, Eds99, EO91, FK90, FF90, FS96, GHG92, GCMR99, GH98, GT96, GL98, GM94, GN99, HONCS98, HB94, Hea91, HIG^{+}99, HBB^{+}98, HD97, HLH92, IB96, KBM^{+}98, KOB99, KT96, KC98, LHS95, LP94, LRW97, LRW98, LRW99, MC93b, MT94a, MG91, MHR94, MP98, Mes94, NW92, Nie92, Nil94, ODS92, OW95, OB94, PD99, Pet94, PW99, PHV^{+}91, QPS96, RES93, RBVP91, RW99, SJW92, SC92a, SH90, SHK93, SBRH94, SWD91, SPS96, SMHP92, SPQ95, TW91, TJ95, VWBM91, VPPH91, WPL95, Zim99, ZW99]. Juveniles [Bea90, GFD95, LRCB96, MGH92, Whi92].

Kahawai [DSD99]. Kansas [QTBG99]. Karluk [Whi92]. Karst [BSWW98]. Karyotype [VMM99, ZT98]. Kendall [GHG98]. Key [PH92]. Key-Event [PH92]. Keystone [ZWLB90]. Kidney [BE92, TKOS92, WE90]. Killifish [KP91, KS97]. Kills [KOB98]. Kinetics [KKR^{+}94]. King [GI96]. Klamath [MS99]. Kokanee [PRR94, RM92, SBJ99]. Kokanees [PBH92, TL96]. Kootenai [PPF99]. Kudoa [EAEA95]. Kvichak [BRQ93].

Laboratory [Bev96, BB91, CDK97, DR94, ESM^{+}92, HY97, MO99, MW98, MBD91, SK96, WH97b]. Lack [BS95a]. Lahontan [DV99a, DV99b, DPR^{+}99]. Lake [AT99, ABM98, ACB^{+}99, Aue96, Bir98, BBA^{+}99, BHSH96, BLW98, BFHB95, BCB96, CBFS90, CBND97, DCE97, DV99b, EKH93, Eds99, EO91, Esh92, FSSH96, FFCS96, FOS93, FS96, FTB90, GKM^{+}94, GWL98, GWJ94, HES^{+}96, HM92, HK90, HSH98, HB99, HJLM99, Jen94, JMBK98, Lab93, LM93, Lin96, LJ92, Lue90, MC93a, MTK^{+}96, MDS98, MO99, MS99, MSC90, MR93, MGM93, MWW^{+}94, MFL99, OSNH95, OB94, PM97, PMF90, RNB95, SHK93, SM96a, SKG92, SMD94, SG95, SBJ99, Swi90, Swi93, SD91, Toh93, TMD96, VM97, Whi92, WPSD99, WAW96, YL93, Bea90, BD90, BL97, BHSH96, BSC97, BCB96, Car95, CM97, CZG90, Esh92, GB93, GKM^{+}94, GKF94, HES^{+}96, Har98, HDS91, JO91, JKO93, KL96a, Kra93, Lin96, MDS98, MSC90, McC90, MFL99, MP94, MP97, MS97, NW92]. Lake [OJS97, PVE94, PRS95, RLO^{+}94, RBM94, RW99, See93, SLH96, SB90, TF99, WDK90]. Lake-Dwelling [Jen94]. Lake-Resident [Whi92]. Lakes [FS96, Gra93, HNW^{+}98, Hol90, HC96b, LRC93, MGW90, MS96a, NÜr95, Ret98, RM92, SKH^{+}93, SHC99, SS91, Ton93, WK99, WMSG90, BR94b, CBFS90, FFCS96, FAC97, HN90, HM90, JE90, KT96, LS93, MMH90, Nak99, RSS^{+}93, SHK93, SS90a]. Lamprey [BMWB97, CSK99, Hol90, Kit90, Lab93, MB91a, Swi90, Swi93]. Lampreys [BS95a, CBFS90, HY97, HY98, MBD91, Swi91, Swi95]. Landscape [BSWW98]. Large [AD93, BWTM90, CCD90, CSBH95, ELAD97, Fle94b, LO91, WD90, WH97a, YHW90]. Large-Scale [ELAD97]. Largemouth [BB92, BZBE98, DMWM96, ESM^{+}92, FK90, GS98b, GGM95, GM93, GWW91, Ham91, HNW^{+}98, HW96, HB94, HIG^{+}99, JP96, KFTT96, LRC93, MM92a, MK90a, MK90b, MH94, NBGF93, NP94a, PW91, PJN95, ROI^{+}97, SR95c, WH97b, WKW98, WGFS99]. Larger [CG97]. Larvae [BC94, CMMP98, GI96, JM96, KP91, LRCB96, MJ93, Men93a, Men93b, MB91b, PCGW90, PGG90, WL94]. Larval [AD93, ADHB98, BV97, BS99, BEC98, CV99, CC96, CCR98, CDL^{+}92, DS92, DSB98, DH94, FTB90, GS98a, GE91b, GE91a, GSS90, GM94, HH93, HD94, HY97, JPL93, KVD^{+}92, LPFA97, MC93b, MHC95, MO91, MP97, MBD91, Nor91, OW96, PM92, PD99, PS99, PD94, RK96, Ret98, RCF98, RB92, SO96, SWD91, War92, WPW94, ZJ93]. LAS-ICP-MS [VE99]. Laser [VE99]. Lates [PG96]. Latitude [Jen94]. Latitudes [BEC98]. Latitudinal [PM97]. Laurentian [JE90, LS93]. Lawrence [FWLC96, HC96a, RLCB98]. lchthyoplankton [TZF93]. Leachates [HJDI98]. Least [CB92, MRM97]. Least-Disturbed [MRM97]. Lemon [MG91]. Length [BZBE98, Con90, EA92, FH95, LLK^{+}99, MH94, MVW99, PRL96, SM90a, TFS92, UMT96, XC95]. Length-at-Age [MVW99]. Length-Based [EA92, XC95]. Length-Dependent [MH94]. Length-Frequency [SM90a, TFS92]. Length-Increment [SM90a]. Lengths [GHG98]. Lentic [TW91]. Leopard [TFH99]. Lethal [FG98]. Lethality [SBSN95]. Leukocytes [ACC94]. Level [BTA97, DV97, PPJ98]. Levels [DDS96, DLF97, DV99b, MK96, ODS92, PG96, RES93, Whi92, WH97b]. Life [Arg95, BWH93, BRQ93, BBFW98, CZG90, Don97, MHSM92, Nak99, RD95, SC98, VRW^{+}93, Win94, You99]. Light [Hub99, Joh93, YRB97]. Like [PGS96]. Limitation [DSB98, ESM90, MT94a]. Limitations [TCCT96]. Limits [KR96, PF90]. Limnological [SHC99, ZWLB90]. Limnology [Dol90, Gal92]. Line [BA90]. Lineages [SC99b]. Lines [HMR^{+}91]. Lingcod [Jag95, JLV96, Jag99, SM90a]. Lingeod [SMC90]. Linkage [LLK^{+}99, MSC94]. Linkages [DS96]. Linked [Eil95, HJWS99]. Linking [GS98b, VRW^{+}93]. Links [KSZ99]. Lipid [GHR91, MH94, TBCK91]. Lipids [HJLM99]. Little [BRCR93, CCR98, MD97]. Littoral [MS96a, WMC93]. Live [GHG92]. Livebearers [Hub99]. Living [DPR^{+}99, Sai97]. Loading [GWL98]. Local [DPR^{+}99, Spa90]. Localization [GNZO97, ZT98]. Location [OKB^{+}92, You98]. Locations [DS90, RK96]. Loci [SAMW90]. Locomotion [WD90]. Locomotor [CDAM90]. Logging [BFH99]. Long [FSSH96, HKC^{+}93, Wat99]. Long-Term [FSSH96, HKC^{+}93, Wat99]. longifilis [BW99]. Longnose [JN97]. Loss [BMB98, BL97, KR96, LRCB96, MSP95, Mei90, RBVP91, SH98, WDM90]. Lost [Wil96, HLM97, MS99]. Louis [OSNH95]. Louisiana [BSRW90, HCG93, RKR92, RTA95, SKR99]. Low [BJ92, CSM92, CS99, DLF97, GC99, HMPV91, HH93, Ken91, KT96, MT94a, MHR94, MK90b, Mor92, See93, SS91, SBSN95]. Low-Cost [HMPV91]. Low-Effort [HH93]. Low-Flow [Ken91]. Low-Gradient [KT96, See93]. Low-Head [MHR94]. Low-pH [SS91]. Low-Risk [CSM92]. Low-Salinity [MK90b]. Low-Temperature [BJ92, Mor92]. Lower [AA90, BRN95, BG99, BC94, BRWB92, DJG99, DJTH95, DWRM97, HD97, IB96, KW99, LT96, LB99, MR95b, PZW95, QNM91, RKBC95, SS90a, SKR99, WPL95, WZ99, Zim99, ZW99]. Lutjanid [HPH93]. Lyase [Gri91]. Lymphocystis [SJMM92].

m [JCT97, MG91]. M-Type [MG91]. Mackerel [GI96]. Macrohabitat [BR91b, MW98]. Macroinvertebrate [YRB97]. Macroinvertebrates [Lue90, SM91]. Macrophytes [CZG90]. Madtoms [MKR97]. Mail [CB95]. Maine [CM97, GSS90, SH97b]. Mainstream [BWTM90]. Maintenance [JM96]. Major [RSL94]. Malaŵi [SLH96]. Male [HC91, KKP^{+}95, SMC90, WKSS95, You99]. Males [SJ91]. Management [Arm92, BFH99, BH93, BS99, HBB^{+}93, HA96, Hea90, LSC93, MP94, Qui90, Rey94, Sai97, Sau91, Spa90, War93, WGGM91]. Manipulation [IMMP90]. Manitoba [SD91]. Mapping [LLK^{+}99]. Margin [Fle94a]. Marine [KOB98, Lei90, McG91, Mur93, OW95, RK96, Sai97, Sca94, WS90, WGG92, Woo94]. marinus [LF95]. Mark [And95]. Marked [IE98, MFWP96]. Marker [GCM^{+}97]. Markers [CSM92, DLF97, LT96, LLK^{+}99, MFL99, MK96, Mon93, SKG92]. Marking [BS96a, Bum91, SWD91]. Marks [GA93, MB91b, TJ95, VSGA94]. Marsh [FWLC96, KP91, MMH94, OWJH98, RA92]. Maryland [CRRH93, PM95, SH97a]. Mass [DHRK94, LRCB96, PGG90, VSSS91, VE99]. Massachusetts [KK93, KK96]. Materials [RN98]. Mathematical [Smi91]. Mating [EHBP99, SJ91]. Maturation [DS90, OR92, RJK^{+}99, SSGE94]. Maturity [BFOS95, Fis99, HDS91, KJ93, MTK^{+}96, UMT96]. Maximum [BR94b, BKN90, HA96]. May [Hug98a]. Mayan [SB94]. McNary [PBM93]. Mean [CJ95, Goo95]. Means [DHP99, Sie94]. Measure [BRTV95, BR96, NÜr95]. Measurements [MZ97]. Measuring [And95, Ken95, Mor95, Sig94, Sig95]. Mechanisms [BL97, Nor91, PD94]. Medaka [MHC95]. Mediated [RBL92]. Mediterranean [SV95]. Mendota [PRS95, MP94]. Menhaden [FNAR97, GI96, LP94, War92]. Merced [KVR96]. mercedis [BFF93]. Mercury [GLHH99, HNW^{+}98, LRC93, MS97, SH97b, WMSG90]. meridionalis [PNSF99]. Merrimack [KK93, KK96]. Mesohabitat [BDL97]. Metabolic [Eva90, KN90, LRW97, LRW98, RBL92]. Metabolism [CW96, KFTT96, LJP93]. Metal [CR97b]. Metals [CR97b, FWB^{+}99, GWF99, WBD^{+}94, WGFB97]. Metals-Contaminated [FWB^{+}99, GWF99, WBD^{+}94]. Metamorphosis [HY97, HY98]. Meters [BR96, HB95c]. Methanesulfonate [MZ97]. Method [BS96a, GNM92, LF95, LJP93, MS96a, MJ93, Mur94]. Methodology [LBBC95]. Methods [Bor97, CAM95, Cow97, Dib91, DBCM92, Hal91, HBS90, PRR94, PRL96, Plu90, RTA95]. Methylmercury [OW96]. Meto [JHT96]. Mexico [BSRW90, BDCMSC99, DLF97, GCMR99, GRKM93, GSR97, Gra91, KC98, NW93, RW92]. Michigan [RSS^{+}93, CCD90, DS90, EKH93, FTB90, Kra93, MDS98, McC90, Ret98, RBM94, See93, TF99]. Microalgae [MBS^{+}92]. Microassay [SBRH94]. Microbiological [Plu90]. Microdistribution [BB92]. Microgeographical [OWJH98]. Microhabitat [BVBM91, BCJ97, BR91b, EKH93, JP97, KV96, LHS95, LW92, Nie92, RLC93, SO94, Wil97]. Microhabitat-Specific [Nie92]. Micropterus [MGH92]. Microsatellite [BPL99, MK96, NCFG97, OSBS98]. Microstructure [FNAR97, GB96]. Mid [Kra93]. Mid-1970s [Kra93]. Middle [WPL95]. Midsummer [HB99]. Midwestern [Lut95]. Migrating [GL98, HDL^{+}96, PHV^{+}91]. Migration [Bay98, CHH98, DHMS94, FC97, GWF99, IB96, KOB99, LD98, TF99, WSB^{+}92, WPM99, BHJ^{+}95]. Migratory [BG94, CDAM90, HONCS98, Whi92]. Mimic [GLHH99]. Mine [HIG^{+}99]. Minimally [WTC92]. Minimize [SMDC96]. Minimum [TBM95, Wil96]. Minimum-Flow [TBM95]. Mining [BLB98, WGFB97]. Minnesota [KW97, SKG92, WKPK93, Wat99]. Minnow [BW91, Sta92, WTP91]. Minnows [Abr96]. Miramichi [RLCB98]. Mirror [Hoc90]. Mississippi [BRWB92, RKBC95, ZJ93, BS95b, MH94, PCGW90, SK91]. Missouri [BG99, BC94, KHTC94, SR95c]. Misuses [Wil90]. Mitochondrial [BS95b, BZBE98, BSEK99, EKNP96, GRKM93, GSR97, GKM^{+}94, MFL99, NP94a, NCFG97, PPF99, SC99a, SKG92, SPR94, SR95d, WW94, WW95, WHW96, WTC92, WDFY97, WMWC93]. Mitogens [ACC94]. Mixed [BSEK99, MFL99, SC99a, WMWC93]. Mixed-Stock [BSEK99, MFL99, SC99a]. Mixing [HIG^{+}99]. Mobile [BM97]. Model [AO96, Bre93b, CP94, CR97a, CSK99, CRRH93, Cue94, FGB^{+}95, HB93, HW93, HKC^{+}93, JROG90, Li99, MO99, MC93b, MP93, PS91, PW99, PGS96, RCR93, RC93, RBM94, SEDP94, WS91, WP93, WH97b, RC93]. Modeling [DSRS93, DD98, FDGG92, LBBC95, Lyo97, Ney93, PP94, RR99, RSS^{+}93, TCCT96, VRW^{+}93]. Models [BT90, BS93b, BH93, BK93, Cha93, DC90, DRF90, EA92, FSSH96, GB93, HBB^{+}93, KBF99, Lab93, LAD97, LB93, SHDB99, SAG95, WHNW98, WGFS99]. Moderate [PE94]. Modifications [KBM^{+}98]. Molecular [Eis95, Iwa96, May98]. Molluscs [Lei90]. Molybdenum [HW90]. Monitoring [BLP^{+}94, BFM97]. Monoclonal [BFBC92]. Monoxide [KOB98]. Montana [BM98, JMTC98, KMNF94, LBKD96, PPF99, Swa97, Toh93, VM97, WBD^{+}94]. Month [CB95, FNAR97, GC99]. Moonlight [LW93]. Morone [HJWS99, MGH92]. Morphological [JP96, WDM90]. Morphologically [TSM98]. Morphology [BRTV95, BRQ93, HLM97, Lis98, RE97a, RE97b]. Morphometric [KHTC94, Kel95]. Morphometry [SK91]. Mortality [CP94, DH94, DR96, EA92, GSC93, GB96, GI96, Jag99, JE90, Kit90, MHR94, McG93, MT94b, Nak99, OFM91, PD99, RKR92, RSV95, SO96, SMC90, SMDC96, SD91, TIT96, VB94, WBI97, ZWLB90, ZJ93]. Mosquitofish [SO95]. Motility [CDL^{+}99, LD96]. Mottled [GSC93]. Mountain [GHW91, RH91, KR96]. Mountains [CW95, Fle94a, KHR97, LMC95, MPG93, WPSD99]. Mouth [Ham91]. Movement [ACB^{+}99, BIBV98, CP97, Hug98a, IB96, Jag99, MR95a, Mat96, RN98, SJ99, SKR99, WP98, YRB97]. Movements [BWTM90, BM95, CM97, GH98, GWF99, Hug98b, JMTC98, KK93, MK91, MR95b, Oli97, Phe92, SMC90, Swa97]. MS [MZ97, VE99]. MS-222 [MZ97]. Much [WH97a]. Muddy [Sco93]. Mullet [GHR91, RTA95]. Multi [BS99]. Multi-System [BS99]. Multicohort [AO96]. MULTIFAN [TFS92]. Multiple [DDS96, DHP99, HT91, Mes94, SSG93, VSGA94, WMC93]. Multiple-Regression [HT91]. Multiple-Stable-Isotope [DHP99]. Multiplicative [Gra99]. Multispecies [MCC^{+}91, Sai92]. Multivariate [DMM98]. Mummichog [KP91]. Mummichogs [OW96, WK91]. Muscle [VRF94]. Muskellunge [CDL^{+}99, CW96, FWLC96, JKM96, LLD96, LD96, ZMDE98]. mykiss [NCFG97]. Mysidacea [BFF93]. Mysis [Toh93]. Myxosporeans [EAEA95].

Narragansett [HR95]. Nassau [EWE97]. National [MPG93, UMT96, VM97, WSB^{+}92]. Native [BW91, CCR98, DW92, GJ95, MM92a, MD97, MPG93, Mef91, MO91, MMH94, MCC^{+}91, SK96, Sco93, TSM98, WK99, WSCL96]. nattereri [BCWB97]. Natural [BGB90, CNWT93, LJP93, SMD94, ZMDE98]. Naturally [FFCS96, Gra93, GKM^{+}94, RM95, TMD96]. near [Ste91]. Nearshore [UMT96]. Necrosis [LGRF90]. Neglected [LBBC95].

Nematode [FWJ96]. Neomysis [BFF93]. nerka [WAW96]. Nest [GKF94, HC91]. Nest-Specific [GKF94]. Nesting [KKP^{+}95, LO95]. Net [CB91, SH98]. Nets [HC92]. Network [OKB^{+}92]. Nevada [MBAL94, DV99b, JVH99, KCSF91, NPLJ92, Sco93]. Newfoundland [CS99]. Newly [BTS^{+}99, Swi95]. Nighttime [TF99]. Nile [SMHP92]. Nine [MMSW92, VBKS97]. Nitrite [MHW^{+}91]. Nitrogen [MWW^{+}94]. nitrophenol [KKR^{+}94]. No [SH98]. Nocturnal [HC91]. Nomenclature [Ken90, SAMW90]. nonadienal [MCWN93]. Nonanadromous [DHP99]. Nonindigenous [SK96]. Nonlethal [BRTV95]. Nonlinear [FSSH96]. Nonnative [LMC95, Sco93]. Nonpolar [HJLM99]. Nonresponse [Fis96a]. Nonstressed [Gri93]. Normal [MM92b]. North [Bra99, CHH98, HCH96, MR95b, SM91, SA92, BSWW98, CWBC90, Fle94a, Gre99, HSH98, MK90a, NPLJ92, PD98, PG96, RSV95, SO94, SC99a, Van93, War92]. North-Central [CWBC90]. Northeastern [NPLJ92, BS95b, KC98, PKM93, SM91, WK99]. Northern [BR91a, BZBE98, DHRK94, FWLC96, GHG92, GGM95, HC93, IB96, JPL93, KW99, Kwa91, MR95a, MO93, MK96, PZW95, PW99, PW91, Pie97, Rey94, RBVP91, SC99b, SPS96, VPPH91, WH95, WPL95, WZ99, Zim99, ZW99, BGB90, LC96, NW93, NLO94, RMSL92, RW92, SH94, WMSG90]. Northwest [FPD90, WAW96]. Northwestern [BS97a, BM94, GCMR99]. Norway [DHRK94, HGK^{+}94, HDS91, Jen94, SH94]. Notes [BMWB97, BCJ95, BS95b, BRCR93, CG97, CSBH95, FWLC96, Fle92, FWJ96, Gri93, HT91, HLM97, JPL93, JE96, KJ95, KMNF94, KP91, LA97, LB93, MM92b, MHC95, MHR94, MS96a, MWW^{+}94, MK96, Mon93, MT91, ODS92, PVE94, PJ91, ROI^{+}97, RTA95, RBL92, RRBA92, SJ91, SA92, SBSN95, TIT96, WDM90, WK91, WK90]. Nova [Jes93]. Nuclear [DLF97, GRKM93, NCFG97]. Null [SAG95]. Number [CDL^{+}92]. Numbers [GVMP^{+}95]. Nurseries [MT94a, SK91]. Nursery [NBGF93]. Nutrient [GWL98]. Nutrients [BLW98]. Nutrition [Dab90]. Nutritional [PCGW90, SEDP94].

Obligate [Lin96]. Observations [HC91, PM99]. Observe [MJ93]. Occupied [ME94]. Occurrence [CNWT93, FWJ96, LB99, MGW90, RM95, SSEO99, SS90a, Ste91, UJ98]. Occurring [FFCS96]. Ocean [BM94]. October [Har98]. Odd [OSBS98]. Odd-Year [OSBS98]. Odor [MCWN93]. off [BS93a, CBB^{+}96, CS91, Jag99, JDW^{+}97, NP94b, War92]. Offspring [BTS^{+}99, LT92, Tri98, WKSS95]. Ohio [DSL98, Har98, HC93, PVE94]. Oil [BMEV95, BSS98, MHS^{+}99, VVF92]. Oklahoma [GGM95, REM92, ZWLB90]. Old [GC99]. Oligotrophic [BLW98, RM92]. Omnivorous [YDS96]. On-Site [BBA98]. Oncorhynchus [NCFG97, SS93, WAW96]. One [GC99]. One-Month-Old [GC99]. Oneida [PMF90, SM96a]. Onset [BLC98]. Ontario [Car95, CH93, EO91, GB93, GKM^{+}94, GKF94, JO91, JKO93, MSC90, MGW90, Mei90, NÜr95, OJS97, OB94, RSS^{+}93, RLO^{+}94, SJMM92, Ste91]. Ontogenetic [RW99]. Ontogenic [MDS98]. Ontogeny [MB91a]. Opeongo [Car95, GKF94, MSC90]. Operated [DCE97]. Operation [Aue96]. Operative [Spa90]. Operculum [DBCM92]. Opinions [HAHD94]. Opportunities [Rey94, SJ91]. Opportunity [PM97]. Optima [KN90]. Optimal [HY98]. Optimum [SEDP94]. Oral [FG93]. Order [Pal94]. Oregon [BDZ^{+}99, CH99, LLP^{+}94, MS99, NP94b, RES93, RMSL92, RSL94, You99]. Orientation [Joh93]. Origin [BBA^{+}99, WW94]. Origins [GKM^{+}94, KWE99]. Osmerid [MCWN93]. Osmolality [LD96]. Osmolarity [KS97]. Osmotic [BWT96]. Other [BC91, MGW90, DWRM97, KOH96, MD97, PC98, SM90b]. Otolith [BB98, DIT97, FBHB97, FNAR97, GB96, HB92, HLM97, MT91, VSGA94]. Otoliths [BS96a, BCB96, CBB^{+}96, CSBH95, DB95, Mur94, MB91b, PBH92, TB95]. Ouachita [SJ99]. Out-Migrating [PHV^{+}91]. Outmigration [Zim99]. Ovarian [JMS^{+}98, TV94]. Ovaries [NW95]. Overfishing [CBFS90]. Overlap [Mic97b]. Oversummering [Nak94]. Overview [VRC93]. Overwinter [JE96, QTBG99, TBCK91]. Overwintering [BCWB97, WSB^{+}92]. Oxygen [BD90, Cou90, MHC95]. Oxygenation [AT99]. Oxytetracycline [Bum91, BK96, SWD91]. Oyster [LF95, OWJH98]. Ozark [MR95a].

p [JHT96]. Pacific [BM90, Eil95, EAEA95, FP99, GL98, PGG90, SBRH94, Sny96, SS93, WAW96]. Paddlefish [EKNP96, HD97, KVD^{+}92, LD97, LD98, RKR92, SSP96, WGC97]. Pains [Ney93]. Pairs [MW98]. Pallid [KGJ92, KJ93, Kwa91]. Palmetto [NWGA99]. Paradigm [Ker90, SJ99]. Parameters [VB94]. Parasite [FK90]. Parasitic [FWJ96, Swi91, Swi95]. Parasitism [SD91]. Parasitized [EAEA95]. Parental [GVMP^{+}95, GM93, LT92]. Park [VM97, MPG93]. Parr [BCJ95, SSG94, VVF92]. Part [JP97, Wil97]. Particle [MJ93]. Particles [Lem96]. Particulate [SC92a]. Partitioning [HB95b, VBKS97]. Pass [HWW^{+}97]. Passage [HONCS98, HDL^{+}96, MHR94]. Passive [BLP^{+}94]. past [HSH91]. Patches [RM95]. Pathogen [LF95]. Pathological [FSW96]. Pathology [TKOS92]. Pattern [BFHB95, Pet94]. Patterns [BFM97, DMWM96, FBHB97, FNAR97, Fro97, GVMP^{+}95, GT96, Hug98b, JMBK98, Kel95, KOB99, MDS98, MR95a, MB91a, ORT^{+}98, PPJ98, QNM91, RH91, RA92, RBM94, SJ99, SKR99, SV95, TCCT96]. Paxton [LM93]. PCB [MC93a]. PCR [BZBE98]. Pecos [WE92]. Pectoral [RB94, Sie94]. Pelagia [OJS97]. Pelagic [DHRK94, FAC97, LW93, MP93, PRS95, RLO^{+}94]. Pen [WS90]. Pen-Reared [WS90]. Penguin [HPH93]. Peninsula [SC99a]. Pennsylvania [HSCK97]. Pentachlorophenol [SCJ93]. Perch [Cam98, DS90, ESM^{+}92, HTS92, HT94, HMPV91, HN90, JS95, JE90, JE96, LRCB96, LPFA97, MKR97, MVW99, PT94, PvdH99, PMF90, SK96, SHDB99, SSG93, SMHP92, SD91]. Performance [CC96, CF99, DSD99, Kel96, KS97, KVR96, MMSW92, MO93, PMSM97, SP97, SC92b, TFS92, TFH99, WS99]. Performances [DRWD94]. Period [TR90]. Periodicity [OSNH95]. Periphyton [HBMB98]. Perkinsus [LF95]. Permits [WTC92]. Persistence [JVH99, JHT96]. Perspective [KP92]. Perspectives [Eis95, Gal92]. pH [MGW90, PM95, SS91, WBI97, WFL^{+}91]. PHABSIM [Wil96]. Phase [Oli97, SMHP92]. Phenotypes [LM93]. Phenotypic [Cha93, QGWF98]. Philadelphia [SSGI99]. Phillip [Li99]. Phosphorus [HKC^{+}93, KH93, KR94a, Kra93, SKH^{+}93]. Photoperiod [MB91a, SH90, SP97, WA91]. Phylogeny [SS93]. Physical [Bev96, HNW^{+}98]. Physicochemical [BG99, RKBC95]. Physiological [CBYH96, GWW91, HC92, LRW97, LRW98, LJP93, MSP91, MR93, Sny96, TIT96, WPM99]. Physiology [Kol95, Mes94, WDB^{+}96]. Phytoplankton [DS96]. Pike [FWLC96, HC93, MK96, Pie97]. Pikeminnow [KW99, PW99, WZ99, Zim99, ZW99]. Pinfish [BJ92]. Pink [BSS98, FS98, HSES98, Hea91, MHS^{+}99, OSBS98]. Piranha [BCWB97]. Pirate [MKR97]. Piscivorous [Bre93a, BC96, HW93, HSH98]. Piscivory [BLC98, PT94, PPJ98, VWBM91, WEH95]. PIT [BLP^{+}94]. Placement [HSES98]. Plaice [Mor92, NAC99]. Plain [MKR97, Pal94, PRD96, SH97a]. Plains [SF97, BF91, RH91]. Planktivores [FAC97]. Planktivorous [Lin96]. Plankton [DHRK94]. Planning [KVR96]. Plant [BB92]. Plasma [HJWS99, KS97, VE99, Whi92]. Plateau [DMM98]. Ploidy [IMMP90]. Plume [PCGW90]. Plunge [BS96b]. Point [VSSS91]. Polar [HJLM99]. Policy [Hea90]. Polluted [WK91]. Pollution [Bor97, Far90, LJS^{+}97, LS97]. Polychlorinated [FSO91]. Polymerase [BZBE98, ZTC94]. Polymers [SMDC96]. Polymorphic [WKW98]. Polymorphism [BZBE98, LLK^{+}99, OSBS98]. Polymorphisms [WMWC93]. Pomoxis [EHBP99]. Pond [CH93]. Ponds [Bre93b, FF90, MC93b, PM92, PD94, SK91]. Pool [BS96b, MBAL94]. Pools [Nak94, NLO94]. Population [AW94, BB92, BPL99, BS95b, BM97, CHI94, CW96, CB90, CRRH93, Cue94, DCE97, DCC99, DD98, GSR97, HD97, JLV96, Jen94, KP92, Kwa91, LD98, LCR98, Mac97, MFL99, PT94, Pie97, PPJ98, Qui90, RSS^{+}93, RC93, RD95, See93, SKG92, SP90, SH94, TSM98, VRC93, WW94, WGGM91, XC95]. Populations [BDL97, BG90, BRN95, Bor97, BSWW98, BSS98, Cha93, CDL^{+}92, DLF97, DFO94, FBHB97, GGM95, Gri93, HK93, HC96b, JCP96, JO91, LJS^{+}97, LT96, MP93, PKM93, QGWF98, RCR93, SC99a, SKG92, SE98, WP93, WE92, WSCL96, WAW96, You99]. Porgy [CS91]. Port [Li99]. Position [Mes91, WGC97]. Possibilities [HGK^{+}94]. Possible [SJ91]. Postmortem [Sta92]. Postsettlement [EWE97]. Postspawning [Gar92]. Potamodromy [Tyu90]. Potassium [SBRH94]. Potassium-Activated [SBRH94]. Potential [AKKH99, BRN95, BTM92, BCWB97, BK93, DC90, DSL98, DJTH95, Eva90, Gar92, HM90, HD94, KR96, KCSF91, MMH90, McC90, Mei90, PMF90]. Potomac [CRRH93, TWC91]. Power [BKN90, Eva90, Kel96, PH92]. Practical [HLH92]. Prairie [WTP91]. Precision [PVE94, RB94]. Predaceous [PHV^{+}91]. Predation [BTM92, BRCR93, Cam98, CWK99, DHRK94, DDS96, GHG92, GL98, HW96, HK90, JO91, JMBK98, JH99, JMMC93, Lab93, MD97, MO91, MB96, OB94, Pet94, RBVP91, SD91, TW91, WPL95, ZW99]. Predator [Ham91, HM92, JPL93, JKO93, MFWP96, Mes94, Mic97b, ODS92, RCR93]. Predators [Bay98, BB91, BBM91, HB95b, MS96b]. Predict [Lab93]. Predicted [FPD90]. Predicting [BCJ95, SBSN95, WGFS99]. Prediction [BTA97, MS96a]. Preference [CDK97, KN90, SF97, SB94]. Preferences [BCJ95, BCJ97, JP97, JCP96, KV96, Wil97, YDS96]. Prehistoric [GJ95]. Preliminary [VSSS91]. Preparedness [PNSF99]. Preparing [Mur94]. Presence [AW94, LRW97, LRW98, VB94]. Present [WW94]. Preservatives [DRWD94]. Pressure [BWT96]. Prevalence [FSW96]. Previously [CH99, Lue90]. Prey [Cam98, CR97b, DSB98, DWRM97, EO91, GS98b, GE91b, GE91a, Ham91, HM92, HM90, JKO93, KL96a, LPW^{+}95, LB93, MS96b, PT94, RLO^{+}94, RCR93, VPPH91, WK91]. Prince [BSWW98, BSS98, FP99]. Principal [McG91]. Principle [BKN90, SB99]. Principles [Ort96]. Priority [Ken95]. Probabilities [Lin96]. Probe [RKF91]. Problems [SSGI99, WF94]. Procambarus [IE98]. Procedure [ABM98, CSBH95]. Procedures [PMSM97]. Proceedings [EN90, RMC90]. Processes [Lei90]. Processing [CSBH95]. Produce [VSGA94]. Produced [BTM92, GVMP^{+}95, GKM^{+}94, Gri91, SWD91, TGC94, TMD96]. Produces [WKSS95]. Production [AO96, BRN95, BBFW98, BTA97, CS99, DJTH95, DH94, DRF90, GDT^{+}90, HAHD94, Hub99, JMBK98, KW97, NAC99, OFM91, PP94, PGS96, RO93, Tri98, WDK90, WKPK93, Wat99]. Productivity [BS99, GWL98, RM92]. Profile [GNM92]. Profiles [BHSH96, Gri93, MHW^{+}91, VRF94]. Progeny [DHP99]. Project [KVR96]. Projects [DJG99]. Proliferation [NW95]. Proliferative [ACC94]. Prolonged [MO93]. Properties [JRLK95, SM90b]. Proportional [PRL96]. Proportionality [KBF99]. Proposal [SB99]. Prospect [WDFY97]. Protection [Sca94]. Protein [SAMW90]. Protein-Coding [SAMW90]. Provenance [FS96]. Provoke [WRC90]. Proximate [BM91, Dyg90, FGW^{+}95, FGMG96, TR90]. Prudhoe [FDGG92, FGB^{+}95, FGW^{+}95, FGMG96, GGGD92]. Puget [JMS^{+}98, LJS^{+}97]. Pulsed [LPFA97]. Pumpkinseeds [CH93]. punctatus [MSC94]. Pupfish [WE92]. Purpose [Ang91]. Putative [WSCL96]. Putting [WH97a]. Pygocentrus [BCWB97]. Pyrenées [BDL97]. Pyrolysis [VSSS91].

Quadratic [SEDP94]. Quahog [WH95]. Qualitative [DCC99]. Quality [Arg95, Car92, Ken91, KW97, MS99, Van93]. Quantifying [DS96, DDS96, PD94]. Quantitative [ABG93, Cow97, NÜr95, Pow94].

Queensland [SSG93]. Questionnaires [CB95]. Questions [WRC90]. Quiescent [Kol91]. Quotas [HSH91].

Radio [AREK98, CP97, Eil95]. Radio-Tagged [CP97]. Radionuclide [DP94]. Rainbow [BVBM91, BGR92, Bea90, BFBC92, BRCR93, CHI94, CHSM93, CR97a, CST96, DSD99, ESM90, Eva90, FH95, GA93, HC96b, KH93, KR94a, KL96a, Lab93, LS93, LMC95, LRW97, LRW98, LRW99, LW92, LB93, MW98, MMSW92, MSP97, MP98, PE94, PPJ98, SSS^{+}99, SG94, SA92, SH98, SB90, TW91, WBI97, WH97a, WSCL96, WBD^{+}94]. Ram [BM95]. Random [WKW98]. Range [CCD90, CH99, MM92a, WDRM90]. Rangeland [Arm92]. Ranging [CCD90]. RAPD [WKW98]. Rapid [ABG93, CSBH95, ZTC94]. Rare [GA93]. Raritan [MM92b]. Rate [BEC98, Eva90, LCR98, RBL92]. Rates [BMB98, BS93b, BC96, CP94, DCC99, GB93, GC99, HB94, HNW97, HW93, Hea91, JRLK95, MSP95, Oli97, PB93, PPD95, VPPH91]. Ratio [WH95]. Ration [BR94b, HMPV91, LRW99, MT94a, WA91]. Ratios [OWJH98, WDK90]. Ray [VE99]. Rays [RB94]. Razorback [JPL93, JH99, PM92, TFW^{+}98]. Reach [DP94]. Reaction [BZBE98, ZTC94]. Reactive [BGR92]. Reactivity [GNZO97]. Reanalysis [CBFS90]. Reared [BTS^{+}99, MMSW92, MGM93, MSP97, PBH92, PMSM97, Sie94, WS90, WS99]. Rearing [FF90, PBM93, UHG94]. Recall [CB95]. Recapture [And95]. Recaptured [KBF99]. Reciprocal [PW91]. Recognition [JPL93]. Recommendations [WP93]. Record [MB96]. Recovery [CBYH96, CC96, DHMS94, ELAD97, Fle90, MHS^{+}99, TBM95]. Recreational [CB91, JROG90, Li99]. Recruitment [BS97b, BCSD97, CRRH93, DH94, Gra91, MTK^{+}96, ML99, MCB90, Nor91, PD99, PH92]. Recycling [HKC^{+}93]. Red [BCWB97, BDCMSC99, Bum91, CS91, GCMR99, GRKM93, GSR97, Goo95, IE98, MT91, NW95, RSV95, VB94, VRF94]. Red-Bellied [BCWB97]. Redd [KV96]. Redds [BR96, GHW91, TK94, YHW90]. Redeye [PB98]. Redfish [GWL98, WAW96]. Rediscovery [BR94a]. Reduced [SCJ93, SC92b]. Reduces [ESM90, GL98]. Reduction [HTS92, Hug98a, RBL92, WK91]. Reef [EKH93, PD98]. Reefs [HJDI98, Par93a]. Reference [DBCM92, KMNF94]. Reflections [CB90]. Refuge [UMT96, WSB^{+}92]. Regeneration [Kra93]. Regimes [BFB98, ROI^{+}97, WHNW98]. Region [FGW^{+}95, FGMG96, GGGD92, KR96, WDFY97]. Regional [BTA97, Fle94a, SR95c, WK99]. Regression [CP94, HT91, HH93]. Regulated [LO95, PF90, RCF98]. Regulation [DSL98, Iwa96, PD99]. Rehabilitation [Fle92, KVR96]. Related [Bro92, GM94, LGRF90, PM97]. Relation [BKM91, CM97, Con90, Dol90, Eds99, EO91, FH95, GM93, GSS90, MGW90, MIPGV97, Mor92, NPLJ92, RO93, SSP96, SS90b, SR95c, TFH99]. Relations [BG99, FAC97, KBSM90]. Relationship [BM91, DIT97, Don97, HK93, KR94a, KFTT96, PvdH99, SK91].

Relationships [BS97a, CW95, DV97, FDGG92, Gra91, HNW^{+}98, HJLM99, LPW^{+}95, Pal94, PPD95, SLH96, VWBM91]. Relative [BS93a, Bev96, BM91, Cas90, Dib91, DRWD94, FGW^{+}95, FGMG96, GS98a, GHG98, IB96, JW98, LPW^{+}95, RM92, SM90b, UMT96, WF94, WGC97, WPM99, WMWC93]. Release [KKP^{+}95, SPS96, TBM95]. Released [DR96]. Releases [WS90]. relicta [Toh93]. Remote [Eil95]. Remotely [BA90, DCE97]. Removal [BMN91, DHRK94, HT94, HWW^{+}97, HBMB98, HBB^{+}98]. Removals [WZ99]. Representation [Cha93]. Representative [Dyg90, Swi95]. Reproduction [BDCMSC99, Dyg90, JS95, NP94b, SO95, Tyu90, ZMDE98]. Reproductive [CAB^{+}91, EHBP99, FFCS96, JN97, LM91, NW93, RW92, RTA95, RS94, RJK^{+}99, RSV95, VM94]. Required [BFBC92]. Requirement [KR94a]. Requirements [BR91b, JM96, Men93a, YC96]. Research [ML99]. Reservoir [BWTM90, BR91a, BS99, DF91, DS92, DDS96, Gal92, HB92, KOB99, Lut95, MGH92, MH94, PHV^{+}91, RBVP91, RGF90, SSP96, SRD95, TL96, TN91, TMWD97, VPPH91, ZWLB90, YL93]. Reservoirs [CP97, DS96, DSL98, Dol90, GS98a, GS98b, HD94, HC96b, Mac97, PV97, RM92]. Residence [BG94]. Resident [RLC93, Whi92]. Residual [MSP97]. Resilience [PLL92, SBH97]. Resistance [BE92, PLL92, WE90]. Resource [CCR98, HB95b, Sca94, VBKS97]. Resources [AT99, Ens94, Sai97]. Respect [BE92]. Respiratory [FG98]. Responder [Hol90]. Response [ACC94, Aue96, BCJ97, DS97, FPD90, GWW91, HTS92, JM96, KBM^{+}98, Kel96, LB94, LC96, MB91a, OB94, QPS96, SEDP94, SPS96, Sig95, SR95d, WBI97, WZ99]. Responses [BB91, BBM91, CBYH96, DDS96, DFT94, MR93, MGM93, NWGA99, TIT96, VRW^{+}93]. Resting [EKH93]. Restricted [HC93, LRW99, SJ99]. Restriction [BZBE98, ZT98]. Retention [GA93, IE98, KJ95, TJ95]. Retina [VWBM91]. Retropinnid [MCWN93]. Return [BDZ^{+}99]. Returns [WS90]. Reveal [BCB96, DLF97, LT96, MK96]. Review [BTA97, CBFS90, Hou97]. Reviews [Gro99, Nev99, PT99]. Revisited [DMM98]. RFLP [BZBE98]. Rheotactic [KJ95]. Rhode [Oli97]. Rich [MMT96]. Richness [NÜr95, WK99]. Ring [DB95]. Riparian [KBM^{+}98, LLP^{+}94]. Ripening [FOS93]. Rise [TWC91]. Risk [CSM92, PPJ98, TW91]. River [AA90, BFH99, BDZ^{+}99, BRQ93, BF91, BM95, CHI94, CHH98, CCR98, CB92, CR97b, CRRH93, CWBC90, DH94, DHP99, DWRM97, ENB92, FWB^{+}99, FH95, FC97, Gar92, GWF99, HDL^{+}96, IB96, Jen94, JMMC93, KOB98, KK93, KK96, KVR96, Lis98, LBKD96, MRM97, MD97, MC93c, MR95b, MS97, Nak94, NPLJ92, OSNH95, Oli97, OR92, PPF99, RLCB98, RCF98, RSL94, RT99, SO94, Sco93, SPS96, SJMM92, SG94, SKR99, SSGI99, SC98, Swa97, TK94, Tyu90, TH91, VM94, WW96, War93, Whi92, WL94, WBD^{+}94, WGFB97, ZHP99, Zim99, ZW99, BPL99, BR91a, BRN95, BR91b, BG99, BC94, BRWB92, BIBV98, BM98, DP94, DJG99, DR94, DJTH95, FWLC96, HLM97, HD97, KBSM90, KHTC94, LD98, MS99, MK91, NBGF93, ORM97, ORT^{+}98, OB98, PBM93, PQF95, PG96, Pet94]. River [PW99, PHV^{+}91, PCGW90, QNM91, QTBG99, ROI^{+}97, RBVP91, RGF90, RKBC95, SK91, SPR94, TWC91, Van93, VPPH91, WDRM90, WW94, WPL95, WSCL96, WMWC93, WK90, YRB97, You98, ZJ93]. River-Dwelling [Jen94]. River-Migrating [HDL^{+}96]. Riverine [ENB92, RGF90, WS90]. Rivers [Ort96, KW99, LB99, OFM91, PZW95, RRBA92, WZ99]. Road [TFH99, WP98]. Roads [BFH99]. Roanoke [CHH98, OR92, SPR94]. Robust [BBA98, CP94, CJ95]. Rock [ESM^{+}92, JMS^{+}98, RO93]. Rockfish [CBB^{+}96, EJ95, NP94b]. Rocky [WPSD99, KR96, RH91]. Role [GS98a, GWL98, Ham91, JMTC98, KL96a, SD91, TBCK91, TWC91]. Roles [JO91, WPW94]. Roosevelt [MS97]. Rostrums [WGC97]. Rotenone [Dib91, FG93, FG98]. Roundtail [KBSM90]. Roving [BBA98]. Roving-Access [BBA98]. Rudd [BW91]. Ruffe [OSNH95, SK96]. Running [Fis96b].

Sacramento [MM95, MHSM92, SC92a, SBH97, TSM98]. Safety [HSH91]. Sagehen [DE92]. Saginaw [DS90]. Sagitta [BBA^{+}99]. Sagittae [BHSH96]. Sakakawea [SSP96]. Salinities [HJDI98]. Salinity [Fis99, MK90b, SH90, SB94, WV95, WL94]. Salmon [AREK98, ACH^{+}98, BGB90, BG90, BM90, BE92, BTM92, BDZ^{+}99, BTS^{+}99, BRQ93, BTA97, Bra99, Bro92, BSWW98, BLW98, BSS98, BB91, BFHB95, BSC97, Cou96, DJG99, DS97, ENB92, Eil95, FBHB97, FFCS96, FS98, FABH^{+}98, GHG92, GVMP^{+}95, GH98, GFD95, GWF99, GL98, GWL98, HSES98, HW90, HONCS98, HN96, Hea91, HBS90, HBMB98, HBB^{+}98, HDL^{+}96, KVR96, LHS95, MSP97, MP98, Mes94, MIPGV97, MHS^{+}99, Mur94, Nie92, ODS92, OSBS98, OW95, PQF95, PMSM97, PP94, QNM91, QPS96, QGWF98, RGF90, RSL94, SJW92, SH90, Sau91, SC99a, SC99b, Sny96, SPQ95, SSG94, TJ95, UJ98, VW98, VVF92, WPM99, Whi92, WDB^{+}96, WGG92, WE90, You99]. Salmonid [Arm92, DFT94, Fle94a, KR96, Ken91, KCSF91, RES93]. Salmonidae [SS93]. Salmonids [BKM91, BBM91, Hug98a, Hug98b, IB96, KBM^{+}98, KL96a, LGRF90, Pet94, PW99, PHV^{+}91, RBVP91, SBRH94, SPS96, VPPH91, WPL95, YHW90, Zim99, ZW99]. Salmonine [GB93, JKO93, KT94, RLO^{+}94]. Salmons [SS93]. Salt [DR96, KP91, RA92]. Salt-Marsh [RA92]. Salts [MSP91, SMDC96]. Saltwater [VM94]. Samarium [GA93]. Same [Jen94]. Sample [CDL^{+}92]. Samples [MVW99]. Sampling [BA90, CDL^{+}92, Goo95, HWW^{+}97, JCT97, LD97, LSC93, VE99, WMC93]. San [MM95, MHSM92, SBH97, TSM98, SJW92]. Sand [BM94]. Santee [BCSD97]. Sarcoma [SJMM92]. Sardine [XC95]. Satellite [Eil95]. Satellite-Linked [Eil95]. Savannah [VM94, WL94]. Scale [Cow97, DBCM92, ELAD97, KBF99, OSBS98, RE97a, RE97b]. Scales [PG96, PRL96, SAG95]. Scenario [LRW97, LRW98]. Schuylkill [SSGI99]. Scientific [Rey94]. Scope [Hoc90, Kit90]. Scotia [Jes93]. Scrap [HJDI98]. Screening [Fle94b, ZTC94]. Scuba [SV95, SW95]. Sculpin [NW92, RW99].

Sculpins [GSC93, OB94]. Sea [BMWB97, BS95a, CBFS90, CSK99, Hol90, HY97, HY98, Kit90, KC98, Lab93, Swi90, Swi91, Swi93, Swi95, Wal95, KWE99, LB92, Nak99]. Season [CC90, Con90, Ken91, MK91]. Seasonal [BVBM91, Bea90, CH93, DE92, Dyg90, FC97, HC96a, JKM96, KV96, MM92b, SKR99, SSG94, Tri98, WPSD99]. Seasonally [LPFA97]. Seatrout [HCG93, KP92, MT94b]. Seawater [SSGE94]. Second [KBM^{+}98]. Second-Growth [KBM^{+}98]. Sections [MT91, Sie94]. Sedentary [CDAM90]. Sediment [JH99, MMT96, Rab97]. Sediment-Rich [MMT96]. Seeing [DMM98]. Seepage [BR96]. Segregation [FS98, Hug98b, MSC94, WW96]. Selected [Don97]. Selection [Abr96, And95, BVBM91, DSB98, Ham91, KN90, Kel96, LB94, LW92, Swi91]. Selective [DHRK94, GB96, GI96, MO91, SO96, VB94]. Selectivity [Jag99, LPFA97, PGS96]. Seleniferous [SO95]. Selenium [HW90]. Self [LB94]. Self-Initiated [LB94]. Semen [LLD96]. Sensed [BA90]. Sensitivity [WFL^{+}91]. Separation [VSSS91]. Series [LJ92, PH92]. Setting [HSH91, VRW^{+}93]. Severe [LC96]. Sewage [TWC91]. Sex [OWJH98, SH94, WH95]. Sexual [HDS91, KJ93, MVW99]. Shad [AD93, ADHB98, BV97, BS97b, BS99, BSEK99, DF91, DS92, DS96, DSL98, DSMM91, GS98b, GDT^{+}90, HONCS98, HH93, HD94, MHR94, Mic97a, Mic97b, PD94, RB92, RBL92, RRBA92, SRD95, SM96a, TN91, WK90, YDS96]. Shaped [TCCT96]. Shark [HB95c]. Sharks [MG91]. Shedding [FSSH96]. Sheepshead [RW92]. Shellfish [McG91, Plu90]. Shepherd [TFS92]. Shield [FS96]. Shifts [RW99]. Shiner [BW91, GLHH99, Kwa91]. Shock [BBM90]. Shoreline [BFHB95]. Short [DE92, JM96]. Short-Term [DE92, JM96]. Shortnose [HLM97, KK93, KK96, MR95b]. Shovelnose [QTBG99]. Shown [FS96, WW94]. Shrimp [Gra91]. Shubenacadie [RT99]. Sierra [JVH99, KCSF91, MBAL94]. Significance [Gar92, Ker90, PH92]. Silversides [GB96]. Similar [TSM98]. Similarities [ADHB98]. Simple [BS96a]. Simulated [AKKH99, CST96, GFD95, Goo95, LRW97, LRW98, LRW99, PGS96, TFS92, VB94]. Simulating [PF90]. Simulation [HKC^{+}93, JROG90, MC93a, RSBN98, SB90]. Simulations [Bre93b, RC93]. Single [Swi90]. Site [BBA98, DV97, WV95, WGFB97]. Sites [KWE99, OWJH98]. Situ [FDGG92, FGB^{+}95, FAC97, GNZO97, MC93b]. Sixty [SHC99]. Size [And95, BW99, BM90, BS97b, BLC98, Cam98, Cas90, CM97, CDL^{+}92, DIT97, EO91, Fis99, FAC97, GB96, GM93, Goo95, GI96, HDS91, Jag99, JE90, KFTT96, LGRF90, LRCB96, LB94, LB92, MC93a, MSP95, Mic97b, PT94, PM97, PPD95, ROI^{+}97, RLO^{+}94, RCR93, RM95, RB92, RBL92, SO96, Swi90, Swi91, TBCK91, Tri98, VB94, WDM90, WH95, WP93, War92, WA91, YDS96, ZMH97]. Size-Dependent [JE90, LRCB96, WDM90]. Size-Related [LGRF90]. Size-Selective [GB96, GI96, SO96, VB94]. Size-Specific [PPD95]. Size-Structured [RCR93, WP93]. Sizes [LB94]. Skate [CWK99]. Slimy [GSC93, OB94]. Slow [DIT97]. Small [BVBM91, BKM91, BBM91, Cam98, Cow97, CBND97, FS96, HN96, HBMB98, HBB^{+}98, KBM^{+}98, LMC95, MGW90, MS96b, MVW99, PD94, PF90, WP98, WKPK93, WS99, WPSD99, You98]. Small-Scale [Cow97]. Small-Stream [WP98]. Smallmouth [BWH93, BR91a, Bev96, CM97, DMWM96, DD98, GKF94, HC91, KKP^{+}95, KBF99, LO95, Lyo97, PV97, PB98, RBVP91, RO93, SO94, SO96, SI91, SS90b, SS91, SR95c, SE98, VHRR98, VPPH91, WW96, WZ99, WKPK93, Zim99]. Smelt [KL96a, Lab93, LS93, MHSM92, SMDC96, TSM98]. Smelts [MCWN93]. Smoky [MPG93]. Smolt [BDZ^{+}99, BRTV95, BTA97, PP94, SSG94]. Smolt-to-Adult [BDZ^{+}99]. Smoltification [BDZ^{+}99]. Smolts [Bra99, HONCS98, PMSM97, WPM99]. Smooth [RS94]. Snake [DJG99, IB96, MRM97, WPL95, WSCL96, KW99, LB99, PZW95, SG94, TK94, WZ99]. Snapper [CS91, GCMR99, GSR97, Li99, PMV98, ZMH97]. Snappers [HPH93]. Snook [RJK^{+}99, WDFY97]. Social [OW96]. Sockeye [BRQ93, BLW98, BFHB95, BSC97, ENB92, FBHB97, FS98, GWL98, Hea91, HDL^{+}96, QGWF98, TJ95]. Sodium [GM94, GC99, SBRH94]. Solar [HBMB98]. Sole [ACC94, Dyg90, JMS^{+}98]. Solids [DV99b]. Somatic [Dyg90, FP99, HT94]. Some [BHJ^{+}95, DV97, ML99]. Sonic [Car95]. Sound [BSS98, FP99, HCH96, JMS^{+}98, PC98, LJS^{+}97, OR92]. Sources [HB93, Rab97]. South [DF91, BS93a, BCSD97, CS91, GWJ94, Pal94, PRD96, RSL94, SC99a, ZMH97]. Southeast [BSWW98, BKM91, BBM91, HBMB98, HBB^{+}98, KBM^{+}98]. Southeastern [HB92, HDS91, Ken98, Mef91, SRD95, DNR^{+}94, FWJ96, HD94, OW95, PMV98, PM99, REM92, WGG92, WGGM91]. Southern [BHJ^{+}95, SC99b, BPL99, BR94a, BDCMSC99, BIBV98, CR97a, DFO94, EO91, ESM90, Fle94a, HC96a, HC93, LMC95, LW92, MSP95, Mei90, NCFG97, OB94, XC95]. Southwest [Ret98]. Southwestern [EKH93, MGH92, Gra91]. Space [MP93, Wil96]. Spain [LC96]. Spanish [XC95]. Spate [LC96]. Spatial [AD93, BC91, Bra99, CV99, EJ95, GB93, GT96, GKM^{+}94, KBSM90, MDS98, MMT96, MP97, MS97, Pet94, PRS95, PPJ98, SRD95, SAG95]. Spatially [BK93]. Spawned [CBND97]. Spawning [Aue96, BFH99, BC91, BHSH96, BFHB95, CHH98, CC90, DJG99, DHMS94, ENB92, FWLC96, FS98, GWF99, Gra91, JCP96, JN97, JMS^{+}98, KBSM90, KK96, KMNF94, KV96, KSH^{+}94, KCSF91, KVR96, LD98, LBCB96, MMT96, MK91, MHS^{+}99, NBGF93, PBM93, PNSF99, PGG90, RT99, SJMM92, TV94, WKSS95, ZMDE98]. Spawningin [ACB^{+}99]. Special [DBCM92, KMNF94, WAW96]. Specially [GHG98]. Species [BM90, BMN91, CC96, EHBP99, FF92, Gra93, MD97, MGH92, MGW90, MWW^{+}94, NÜr95, RTA95, Ret98, Sca94, SKG92, SR95b, SLH96, TSM98, VBKS97, VRW^{+}93, WDK90, WK99, WTP91]. Specific [GKF94, Nie92, PJN95, PPD95, RE97a]. Spectrometry [VSSS91, VE99]. Speeds [Men93b, VP98]. Sperm [BWT96, BFBC92, LD96, TGC94, TFW^{+}98]. Spermatozoa [BWD98, BFBC92, CDL^{+}99, LLD96, LD96]. Spill [BMEV95, BSS98, MHS^{+}99]. Spine [Sie94]. Spinedace [BRCR93]. Spiny [BTM92]. Splake [MGM93]. Splenic [ACC94]. Splittail [MM95, SBH97, YC96]. Splitter [WM97]. Sport [Jag99, SH97b]. Sport- [Jag99]. Spot [PCGW90, Ste91]. Spotfin [KP91]. Spotted [DB95, HCG93, KP92, MT94b, PV97, SKR99]. spp [DMM98, EAEA95, Hub99]. Spring [BDZ^{+}99, HY97, MSP97, MP98, SH90]. Springs [GHG98, MH94]. Squawfish [BR91a, BW94, BB98, CC96, GHG92, IB96, MK91, MO93, MB91b, ORM97, ORT^{+}98, OB98, PZW95, RBVP91, SPS96, TBCK91, Tyu90, TH91, VPPH91, WPL95]. SRLCA [TFS92]. SS [HB95c]. St [HC96a, HN90, RLCB98]. St. [FWLC96, OSNH95]. Stability [GRKM93, HB99, KCSF91, LHS95, SR95d, VE99, WW95]. Stable [DHP99, FS96, See93, VHRR98]. Stable-Flow [See93]. Stage [EWE97, Nak99, RK96]. Stages [Arg95, BBFW98, CZG90, LT92, SC98]. Standard [Eva90]. Standardization [ZT98]. Standardized [Ken90]. Standing [BKM91, TCCT96]. Starvation [JW98, LRCB96, Lyo97]. Starving [LRCB96]. States [BCWB97, FWJ96, PMV98, PM99, WGGM91]. Static [BT90]. Stationary [SW95]. Statistical [CAM95]. Statolith [BMWB97, MB91a]. Status [BRTV95, CS91, DHR96, MM95, MHSM92, OB98, VM97, WE92]. Staying [Bir98]. Steelhead [BPL99, BCJ95, BRTV95, BB91, CAB^{+}91, HN96, KT94, KT96, MFWP96, MSP97, Nak94, NLO94, RSS^{+}93, RMSL92, RSL94, See93, WS90]. Steep [KCSF91]. Stem [HW96]. Stickleback [LM93, RKF91]. Sticklebacks [Abr96]. Stimuli [PC98]. Stock [AO96, BPL99, BBA^{+}99, BSEK99, Con98, Cow97, Gra91, KWE99, MFL99, PGS96, RE97a, RN92, SSP96, SC99a, Smi91, SR95d, TCCT96, TWC91, WF94, WW95, WHW96, WRS^{+}97]. Stock-Specific [RE97a]. Stocked [Bea90, BBA^{+}99, HES^{+}96, KL96a, MPG93]. Stocking [DSMM91, MM92a, MC93a, WBI97, WV95]. Stockings [PBH92]. Stocks [BEC98, CS91, FSO91, FPD90, GVMP^{+}95, GNM92, PPF99, SD90, SSG93, WMWC93, RE97b, RRBA92]. Stomach [HT91, SBJ99]. Storage [BWT96, CDL^{+}99, DS90, HBMB98, TGC94]. Strains [FOS93, LGRF90, WE90]. Strait [Hea91, SM90a, SMC90]. Stranding [AKKH99]. Strategies [Goo95, LSC93]. Stratified [NLO94]. Strawberry [KOB99]. Straying [QNM91]. Stream [BVBM91, BC91, BG94, BBM91, CCD90, CW95, CST96, DR94, DV97, DPR^{+}99, ESM90, ELAD97, Fis96b, GGM95, GFD95, GHW91, HN96, HBS90, HWW^{+}97, HBB^{+}98, Hug98a, Hug98b, JMTC98, JN97, KT94, LHS95, LMC95, LAD97, LO91, LC96, LW92, LO95, MMT96, MR95a, Mef91, Mes91, OKB^{+}92, Pal94, PLL92, PPD95, RK96, RH91, RR99, RLC93, RO93, REM92, SO96, SS90b, SR95a, SJ99, Ste91, TCCT96, WP98, WDK90, WKPK93, WG94, WTP91, You98]. Stream-Dwelling [RO93]. Stream-Living [DPR^{+}99]. Stream-Resident [RLC93]. Stream-Spawning [JN97]. Streams [BMWB97, BS97a, BKM91, BLB98, BB91, CR97a, CS99, CBND97, FC93, HSCK97, HBMB98, JVH99, JMTC98, KBM^{+}98, KCSF91, KW97, LLP^{+}94, Lyo97, MW98, MRM97, ME94, Mei90, MKR97, NLO94, Pal94, PRD96, PB93, PM95, PB98, Rab97, SH97a, SR95b, SF97, SAG95, SR95c, SH98, WS99, WSB^{+}92]. Strength [Don97, FAC97, FTB90, SW95]. Stress [CBYH96, DS97, GWW91, MR93, MGM93, NWGA99, ODS92, Rei92, SC92b, WBI97, WV95]. Stressed [Gri93, GC99]. Stressor [LT92]. Stressors [Mes94]. Striped [BFOS95, BK93, BM91, BEC98, BCSD97, BK96, CHH98, CBYH96, CV99, CB90, Cou90, CRRH93, DDS96, DSL98, DLF97, DIT97, DR96, DHMS94, DWRM97, FSO91, GHR91, GNM92, Gri93, GM94, GC99, HCH96, HAHD94, HB93, HC92, HLH92, LT96, LPFA97, MHW^{+}91, MSP91, MO91, Men93a, Men93b, NWGA99, OFM91, OR92, RTA95, RE97a, RE97b, RLCB98, RC93, RD95, RT99, SJW92, SD90, SWD91, SPR94, SR95d, SC92b, TIT96, TWC91, VM94, WDM90, WDRM90, WW94, WF94, WW95, WRS^{+}97, WV95, WL94, WGGM91, WMWC93, WA91, WKSS95, ZWLB90]. Strontium [PG96]. Structure [AW94, BG90, BPL99, Bev96, BC94, BCB96, Fis96b, GSR97, HKC^{+}93, HB99, HD97, JLV96, Joh93, JCT97, KP92, McC90, MFL99, Pal94, PT94, REM92, SSP96, SS90a, SKH^{+}93, SE98]. Structured [RCR93, WP93]. Structures [Cas90]. Structuring [JO91]. Studies [BLP^{+}94, HBS90, PJ91, SK96, VHRR98, WRC90, Wil96]. Study [BMEV95, DCE97, DD98, HY97, MP94, MBD91, PH92, VSSS91]. Studying [Eil95]. Stunted [Hal91, UHG94]. Sturgeon [BRN95, CMMP98, CF99, DJTH95, FC97, HMR^{+}91, JDW^{+}97, KGJ92, KJ93, KK96, MC93c, MB96, PM97, QTBG99, RB94, RNB95, SC98, VMM99, VE99, WHW96, ZHP99]. Sturgeons [Aue96, CDK97, DP94, KHTC94, KK93, MR95b, PBM93]. Subadult [Bum91, ORT^{+}98]. Subdivision [TSM98]. Subjected [AKKH99]. Sublethal [LRW97, LRW98, LRW99]. Subpopulation [BFHB95]. Subpopulations [BSC97]. Subsampling [WM97]. Subsequent [BWD98]. Subspecies [BZBE98, WKW98]. Substrate [CDK97]. Substrates [AA90]. Subsurface [SJW92]. Subyearling [RGF90]. Success [BS99, CAB^{+}91, EHBP99, JMS^{+}98, LO95, MW98, MCB90, NAC99, VM94]. Successional [Sai92]. Sucker [CH95, HT94, JPL93, MS99, TB95, TFW^{+}98]. Suckers [HTS92, HLM97, JH99, MR95a, PM92]. Suisun [MMH94]. Suitability [AA90, BC97, CB90, PJ91]. Suitable [LLK^{+}99]. Summer [Bir98, BKM91, BBM91, CB90, ESM90, GHR91, HCH96, HC93, Ken91, KT94, LRW97, LRW99, LBKD96, MT94a, Mon93, Nak94, RSL94, RA92, SPQ95, WDB^{+}96, ZWLB90]. Sunfish [HNW97, Ret98, RN98, ZJ93]. Sunlight [MB91b]. Superior [BHSH96, BCB96, HES^{+}96, HB99, JMBK98, Lin96, OSNH95, WDK90]. Superposed [PH92]. Supersaturation [CMMP98, Lut95]. Supplemented [Cue94]. Supply [CH95, Hen90, JKO93]. Surface [BFBC92, FF92, HONCS98]. Surgically [AREK98, BW99]. Surplus [PGS96]. Surrogate [DSD99]. Survey [JROG90, RSBN98]. Surveys [BBA98, Ens94, Fis96a, HWW^{+}97, JRLK95]. Survival [BW94, BV97, CC96, DSRS93, DV99a, DV99b, FC93, FF90, FTB90, GM93, GE91a, GSS90, GM94, HES^{+}96, Hoc90, IE98, Jag95, JE96, KL96a, KOB99, KT94, KC98, KVD^{+}92, MT94a, MHW^{+}91, MSP91, Mic97a, MH94, ORM97, PM92, PW91, SHK93, SG94, SS91, SPQ95, Swi90, Swi93, Swi95, TBCK91, WV95, WPW94, WDB^{+}96]. Susceptibility [ACH^{+}98, LGRF90]. Suspended [JH99, OS93]. Suspension [SC92a]. Sustainability [JKO93]. Sustainable [Ken98, Men93b]. Sustained [HAHD94, KW99, WZ99]. Suwannee [FC97, MC93c, SC98]. Swamp [IE98]. Swarm [WE92]. Swim [FWJ96, GM93, MHC95]. Swim-up [GM93]. Swimming [ABM98, CC96, CF99, DSD99, KS97, Men93b, MO93, PMSM97, RB92, SP97, SC92b, TFH99]. Swordfish [PGS96]. Swordfish-Like [PGS96]. Symbology [VSGA94]. Symmetrical [TMWD97]. Sympatric [FS98, HB95b, LA97, MB96, VBKS97]. Sympatry [SH98]. Symposium [EN90, RMC90]. Syndrome [FABH^{+}98, Nak99]. System [BRQ93, BLP^{+}94, BS99, BCSD97, BSC97, BFM97, CBB^{+}96, Eil95, GWF99, KR96, PG96, Ton93, UHG94, Whi92]. Systematic [HB93]. Systematics [May98, SE98]. Systems [ACB^{+}99, Fle90, Fle94b, Ken95, Lem96, Mor95, Sig95, TW91, Van93].

Tabusintac [DHP99]. Tag [DHMS94, FSSH96, HSES98, IE98, WDM90, WDFY97]. Tagged [CP97, KBF99, SMC90]. Tagging [ABM98, DSD99, Hug98a]. Tags [BW99, IE98, MG91, MMSW92, WGC97]. Tailrace [IB96]. Tailraces [DJG99]. Tailwater [WH97a]. Taking [WH97a]. Taku [ENB92]. Takvatn [DHRK94]. Tambaquí [Lov98]. Tank [WKSS95]. Target [FAC97, SW95]. Tasmania [AW94]. TCDD [JHT96]. Technique [HT91, LD97]. Techniques [VHRR98, WRS^{+}97]. Telemetered [CHH98]. Telemetry [ACB^{+}99, Eil95, HDL^{+}96, LJP93, PJ91]. Temperate [HJWS99, HSH98, MS96a, SP90]. Temperature [BBM90, BM90, BJ92, BV97, Bev96, BBFW98, CC96, Cou90, DR94, DV99a, Eds99, EWE97, FDGG92, FH95, Fis99, FNAR97, GSS90, HA96, Hen90, HBMB98, JO91, JE96, KN90, Kel96, Kol91, KVD^{+}92, LCR98, LBKD96, MW98, MT94a, MK90a, MB91a, Mor92, PvdH99, RR99, RHP90, RN92, SP97, SB94, SSG94, Swi93, WBI97, WA91]. Temperature-Oxygen [Cou90]. Temperatures [BW94, DV99a, HY98, KL96b, LBKD96, MS96a, SBSN95, Swi95, WH97b]. Temporal [AD93, Bra99, EJ95, GRKM93, GLHH99, GKM^{+}94, KBSM90, MP97, PRS95, PPJ98, RA92, SO94, SRD95, ZMH97]. Ten [RM92]. Tendency [HSH98]. Tennessee [CWBC90, PB98, SH98]. Term [DE92, FSSH96, HKC^{+}93, JM96, Wat99, Wil90]. Test [RCR93, SHDB99, SHC99]. Testing [BCJ97, BS93b, CSK99, JP97, WHNW98, Wil97]. Tests [BHJ^{+}95, BFF93, SAG95, WS91]. Tetrachlorodibenzo [JHT96]. Tetracycline [Mon93, MB91b]. Texas [BJ92, CP97, Dol90, KP92, MCC^{+}91]. TFM [KKR^{+}94]. Thames [SJMM92]. Their [BB91, BBM91, DHP99, EHBP99, KL96a, LSC93, NLO94, QGWF98, SK91, Sai97, SG94, TMD96, WS99, WKW98, Arm92]. Theoretical [WP93]. Theory [Bre93a, Sai92, SB99, VRW^{+}93]. Thermal [BS96b, Eva90, HB92, KR96, KSZ99, MMH90, McC90, ME94, Mei90, PM97, SSEO99, SS90a, SF97]. Thermally [NLO94, VSGA94]. Thermoregulation [Bir98, BKN90, SG95]. Thiamin [FABH^{+}98]. Thiamine [FFCS96]. Thorny [CWK99]. Threadfin [ADHB98, BV97, DF91, DSMM91, GDT^{+}90, SRD95]. Threatened [TFH99]. Three [FSSH96, FC93, FOS93, HBS90, HSH98, SLH96, WS91, WDK90]. Threespine [LM93, RKF91]. Throw [JCT97]. Thyroid [CDAM90, Whi92]. Tidal [BG94, NBGF93, RT99]. Tidewater [SM96b]. Tiger [BM94]. Tilapia [SBSN95]. Tilefish [BS93a]. Timber [RES93, REM92]. Time [CDL^{+}99, GM93, LRCB96, LJ92, MC93a, MP93, PH92, PGS96, TMD96]. Time-Changing [PGS96]. Timing [GS98a, KV96, WPM99]. Tire [HJDI98]. Tires [HJDI98]. Tissue [FABH^{+}98, GNM92, HJLM99, Mar90]. Tissues [CSM92, HT94, HJLM99]. Today [Ney93]. Tolerance [BCWB97, BMB98, BJ92, Mor92, SH90, SI91, SF97, SB94, SSGE94]. Tolerances [LBKD96, SR95a, YC96]. Tomcod [DH94, DWRM97]. Too [WH97a]. Tool [LBBC95, LPW^{+}95]. Toronto [Ste91]. Total [DV99b, HJLM99, KT94]. Toxic [MKT92]. Toxicity [BFF93, BK96, FG93, HJDI98, HIG^{+}99, SJW92]. Toxicology [Eis95]. Trace [BHSH96]. Tracking [Car95, Eil95]. Trans [MCWN93]. Trans-2-cis-6-nonadienal [MCWN93]. Transect [BA90]. Transects [SV95]. Transfer [BLW98, UHG94]. Transferability [LAD97]. Transfers [Sca94]. Transformation [SSG94]. Transintestinal [BW99]. Transmitter [PMSM97]. Transmitters [AREK98, CF99, KL96b]. Transponder [BLP^{+}94]. Transport [MSP91, MGM93, SMDC96]. Trap [GHG98, JCT97]. Traps [HC92]. Trauma [Lut95]. Travel [Li99]. Traverse [FTB90]. Trawl [Jag99, PRR94, Sai92]. Trawl-Caught [Jag99]. Trends [BS93a, Bra99, DHR96, OB98, PF90]. Trials [TGC94]. Tributaries [BC94, KOB99, MP97, WDK90]. Tributary [BG99, KT96, RCF98, See93, TK94]. Tribute [EN90]. Tricaine [MZ97]. Trifluoromethyl [KKR^{+}94]. Triploid [LGS^{+}96, Par93b, SSS^{+}99]. Triploidy [BBM90, CNWT93]. Trophic [BLW98, DDS96, FGB^{+}95, HB95b, PS91, RO93, VHRR98]. Tropical [Ayv99, Kel99]. Trout [AW94, BC97, BVBM91, BDL97, BGR92, BFH99, BC91, Bea90, BFBC92, Bir98, BR96, BRCR93, BS96b, BBA^{+}99, BR91b, BM95, BIBV98, BCB96, BM98, CHI94, CHSM93, CCD90, CR97a, CS99, CW95, CR97b, CBFS90, CH99, CST96, CBND97, DSD99, DCE97, DR94, DW92, DV99a, DV99b, DHP99, DPR^{+}99, EO91, ESM90, Esh92, Eva90, FSSH96, FWB^{+}99, FH95, FFCS96, FC93, FOS93, FS96, GSC93, GA93, GFD95, GKM^{+}94, GHW91, Hal91, HES^{+}96, HW93, HDS91, HBS90, HJLM99, HC96b, JVH99, JMTC98, KH93, KR94a, KFTT96, KMNF94, KV96, KOB99, KT94, KT96, KHR97, KW97, Lab93, LHS95, LMC95, LA97, LLP^{+}94, LRW97, LRW98, LRW99, LCR98, LW92, Lue90, LB93, MC93a, MDS98, MO99, MMT96, MW98, MBAL94, Mat96, MSC90, MMSW92, MPG93, MR93, MGM93, MSP97, MP98]. Trout [ME94, Mei90, Mes91, NPLJ92, OB94, PKM93, PE94, PPJ98, Pow94, RR99, RM95, RLC93, SHK93, SSS^{+}99, SG94, SA92, SG95, SH98, Swa97, SB90, Swi90, Swi93, TW91, TK94, VM97, WBI97, WDK90, Wat99, WH97a, WS99, WPSD99, WSCL96, WFL^{+}91, WBD^{+}94, WGFB97, YRB97, You98, YL93]. Trouts [SS93]. trutta [DSD99]. truttaceus [AW94]. Tuna [SMPS98]. Turbid [BFHB95]. Turbidity [BGR92, GL98, MS96b, VWBM91]. Turbidity-Induced [BGR92]. Turbine [MHR94]. Turbine-Passage [MHR94]. Two [BMN91, BSS98, CP97, EA92, FF92, HC93, HMR^{+}91, HBMB98, HD94, JCP96, KBF99, KL96b, Kol91, LGRF90, LBBC95, MT94a, Mei90, MP97, OWJH98, SHDB99, SMHP92, SBJ99, TR90, WS99, WH97b, WE90, You99]. Two-Dimensional [LBBC95]. Two-Phase [SMHP92]. Two-Year [TR90]. Type [GNZO97, MG91, WM97].

U.S. [HES^{+}96]. UDP [KKR^{+}94]. UDP-Glucuronyltransferase [KKR^{+}94]. Ultrastructure [LLD96]. Umpqua [RSL94]. Undergo [Kol91]. Understanding [Wea90]. Underwater [BA90, Joh93]. Unifying [SB99]. Unimpounded [BRN95, DJTH95]. Uninflated [MHC95]. Uninhibited [Sta92]. Unique [CHI94]. Unit [LJ92]. United [FWJ96, PMV98, PM99, WGGM91, BCWB97]. Units [ADHB98]. Unlogged [CH99]. Upper [CBFS90, KBSM90, KHTC94, KMNF94, MRM97, MS99, MK91, MS97, ORM97, ORT^{+}98, OB98, Sco93, WSCL96, ZJ93]. Upstream [WTP91]. Uptake [CR97b]. Urbanization [WG94]. USA [DNR^{+}94, PKM93, PS91, WK99]. Usable [Eva90]. Use [ABM98, ACB^{+}99, BWTM90, BFBC92, BR96, BR91b, BFF93, BLP^{+}94, BM95, BB91, BBM91, CSM92, CCR98, CB95, CBND97, DCE97, DW92, DWRM97, GCM^{+}97, GNZO97, HC96a, HNW97, HBS90, HDL^{+}96, HD97, Hug98b, JMTC98, JROG90, KT96, Lab93, LHS95, LPFA97, LO91, LJP93, LB93, MR95a, MGH92, MBAL94, Mat96, MR95b, NLO94, PRD96, PBH92, PBM93, PE94, QTBG99, RE97b, RLC93, RSL94, RA92, SK91, SO94, SEDP94, SS90b, SM90a, SKR99, Swa97, SMDC96, TH91, VSGA94, WKW98, WDFY97, WGGM91, WMWC93, You98]. Used [BDL97, JRLK95, MKR97, Nak94, PJ91]. Uses [Koh97]. Using [HWW^{+}97, Ret98, MFL99]. Utah [TL96, FH95, KOB99, TL96, Tyu90, TH91]. Utility [BCJ95, SKG92]. Utilization [EWE97, Sai97].

Valdez} [MHS^{+}99, BMEV95, BSS98]. Validation [HB92, Hal91, HLM97, KBF99, LJ92, MT91, RNB95, TB95]. Validity [Mur94]. Valley [SJW92, Wat99]. Value [CB91]. Varden [BHJ^{+}95, KWE99, LA97, UMT96]. Variability [BDL97, BC91, Bro92, Cha93, DGHP99, DV97, DPR^{+}99, GKM^{+}94, LT96, PZW95, PKM93, WAW96]. Variable [BW99, BF91, BSEK99, LBCB96, Pie97]. Variables [Don97, Hou97, NW93, PPD95, SSEO99, SEDP94]. Variance [CJ95]. Variances [XC95]. Variation [BE92, BPL99, BRQ93, BEC98, CV99, CH93, Con90, EJ95, EKNP96, FGMG96, FH95, GHR91, JLV96, Kel95, MMT96, MM92b, Mes91, MK96, MIPGV97, NW93, PPJ98, PM97, RMSL92, Ret98, RE97a, SO94, SKG92, SA92, TCCT96, WE90, You99]. Variations [MSP95, OWJH98]. Varied [RM95]. Varying [BMWB97, ROI^{+}97, SPQ95]. Vateritic [BBA^{+}99]. Vegetated [JCT97]. Vegetation [BMN91, HW96, SF97]. Vehicle [DCE97]. Velocities [SS90b]. Velocity [BCJ95, CDK97, OS93]. Verification [DB95, HR95]. Vermilion [CS91, PMV98, ZMH97]. versus [KN90, SC92a]. Vertebral [GCM^{+}97]. Vertebrates [WD90]. Vertical [OW95, TF99, WK90]. Vessel [AKKH99]. Vessel-Induced [AKKH99]. Viability [SP90, Tri98]. Vibrio [ACH^{+}98]. Vibriosis [BE92]. Video [BA90, HC91]. View [Fle94a]. Virginia [Dan95, Fle94a, LO95, OFM91, SO94, SO96, SAG95]. Virus [LGRF90]. Visual [MFWP96, NW92]. Vital [DCC99]. Vitellogenin [HJWS99]. Vitellolipid [Gri93]. Vitelloprotein [Gri93]. Vitro [HJWS99]. Volitional [WKSS95]. Volume [McG91]. Vulnerability [JH99, MFWP96].

Wait [Bre93b]. Wales [BSWW98]. Walker [DV99b]. Walleye [AA90, BS95b, JCP96, JM96, MS97]. Walleyes [BR91a, BM97, Cam98, FF90, HSH91, HM92, JW98, KBF99, MTK^{+}96, MC93b, MFL99, MP97, PVE94, PS99, RBVP91, SJMM92, VWBM91, VP98, VPPH91, WEH95, WMSG90, Zim99]. Warm [GHG98, SS90a, SG95, WDB^{+}96]. Warming [BD90, HM90, HY97, JE90, KR96, LRW97, LRW98, LRW99, MMH90, McC90, Mei90, PD98]. Warmwater [BFB98, LO91, PB93, SSEO99, TZF93, TBM95]. Washington [BPL99, Bea90, BS97a, Jag95, Jag99, JMS^{+}98, LA97, MS97, SB90]. Waste [WGFB97]. Water [Arg95, CDK97, DR94, DR96, Far90, Fle90, HONCS98, Hen90, HBMB98, KH93, KVD^{+}92, KW97, LRC93, MS99, MBAL94, MS96a, MSP91, MR93, RN92, SSG94, Swi93, Van93]. Waters [BSRW90, Fis96b, HES^{+}96, Har98, ML99, MT94b, PS91, RTA95, RSV95, UMT96, WDB^{+}96, WGG92, WH95]. Watershed [CS99, DFO94, WG94]. Weakfish [GE91b, GE91a, GT96, LBCB96, TV94]. Weather [Woo94]. Web [HKC^{+}93, MP94, SKH^{+}93]. Webs [Fro97]. Wedge [TCCT96]. Wedge-Shaped [TCCT96]. Weight [BM91, HH93, HJLM99, LPW^{+}95, SM90b]. Weiss [SMD94, TMD96]. Well [DMM98]. Western [SO95, Bra99, BF91, HM92, JMBK98, LA97, MTK^{+}96, SC99b, SMPS98]. Westslope [VM97]. Where [HBB^{+}93]. White [BBM90, BRN95, CH95, CMMP98, CF99, DP94, DHR96, DJTH95, DNR^{+}94, Gra91, GDT^{+}90, GWJ94, Har98, HTS92, HT94, HA96, HMR^{+}91, JS95, JE90, JE96, KSH^{+}94, LPFA97, MB96, PBM93, Par93b, PD94, PMF90, RB94, SP97, SMD94, TB95, TMD96, TMWD97, VE99, WA91, WKSS95, CW95]. Whitefish [ABM98, ACB^{+}99, CB92, Eds99, FDGG92, FGW^{+}95, FTB90, GGGD92, LJ92, SKG92, Toh93]. Whole [HK90, Hug98b]. Whole-Lake [HK90]. Whole-Stream [Hug98b]. Wide [REM92]. Width [Ham91]. Wild [BC91, BTS^{+}99, BBA^{+}99, CAB^{+}91, CNWT93, MSP97, MP98, Mes91, MIPGV97, NCFG97, PMSM97, RE97b, UHG94, VRF94, WS99, WDB^{+}96]. Wilderness [Mat96]. Wildlife [UMT96, WSB^{+}92]. William [BSS98, FP99]. Winter [BM95, CG97, DWRM97, FP99, GH98, GSS90, HR95, HC96a, JMTC98, JE90, Kol91, LRW98, Lyo97, MM92b, MH94, Phe92, SH90, SG94, Swi95, WGFS99]. Winter-Quiescent [Kol91]. Wire [HSES98, IE98, WGC97]. Wisconsin [AA90, KOB98, Lyo97, ME94, PRS95, WMSG90, ZMDE98]. within [HCH96, LT96, PKM93, RSL94, WE90]. without [KT94]. Woody [BS97a, CST96, MKR97]. Worth [SBJ99]. Would [JVH99]. Wound [Mar90]. Wreck [HB95c]. Wyoming [HC96b, KHR97, TL96].

Year [Abr96, Bir98, BC96, CC90, CRRH93, DSMM91, DBCM92, Don97, FTB90, GDT^{+}90, JE90, MTK^{+}96, ML99, OSBS98, RLCB98, RC93, RA92, SS90a, SS90b, TR90]. Year-Class [DBCM92, Don97, FTB90]. Yearling [FDGG92]. Years [PD98]. Yellow [Cam98, DS90, ESM^{+}92, HTS92, HT94, HMPV91, HN90, LRCB96, Oli97, PT94, PvdH99, PMF90, SK96, SHDB99, SD91]. Yellow-Phase [Oli97]. Yellowstone [KHR97, SSP96, TK94]. Yellowtail [EJ95, RN92]. Yield [HN90, MBS^{+}92, PS91]. Yields [RKF91]. York [Phe92, CC90, CB91, FWLC96, PMF90, SM96a, WHW96]. Young [Abr96, AKKH99, BB98, Bir98, BKM91, BR91b, Bre93b, BC96, BBM91, BSC97, CRRH93, CBND97, DSMM91, GDT^{+}90, HBS90, HLM97, JE90, JH99, JW98, RLCB98, RC93, RA92, SS90b, Tri98, VP98]. Young-of-the-Year [BC96, CRRH93, JE90, RLCB98, RC93]. Young-of-Year [Abr96, Bir98, DSMM91, GDT^{+}90, RA92, SS90b].

Zealand [QGWF98, SA92, UJ98]. Zonation [RH91]. Zone [HIG^{+}99, Ken98, MS96a, SC99b]. Zoogeography [SP90]. Zooplankton [BV97, Bre93a, BS97b, CV99, DS92, DS96, DSB98, GDT^{+}90, JO91, LPFA97, Lin96, MCB90, VM97, WPW94, YDS96].

References

Auer:1990:CSS

[AA90] Martin T. Auer and Nancy A. Auer. Chemical suitability of substrates for walleye egg development in the Lower Fox River, Wisconsin. Transactions of the American Fisheries Society, 119 (5):871–876, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:1993:QHA

[ABG93] S. Marshall Adams, Allen M. Brown, and Ronald W. Goede. A quantitative health assessment index for rapid evaluation of fish condition in the field. Transactions of the American Fisheries Society, 122(1):63–73, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anras:1998:UAB

[ABM98] M. L. Bégout Anras, R. A. Bodaly, and R. McNicol. Use of an acoustic beam actograph to assess the effects of external tagging procedure on lake whitefish swimming activity. Transactions of the American Fisheries Society, 127(2):329–335, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Abrahams:1996:IBY

[Abr96] Mark V. Abrahams. Interaction between young-of-year fathead minnows and brook sticklebacks: Effects on growth and diet selection. Transactions of the American Fisheries Society, 125(3): 480–485, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anras:1999:MHU

[ACB^{+}99] M. L. Bégout Anras, P. M. Cooley, R. A. Bodaly, L. Anras, and R. J. P. Fudge. Movement and habitat use by lake whitefish during spawningin a boreal lake: Integrating acoustic telemetry and geographic information systems. Transactions of the American Fisheries Society, 128(5):939–952, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Arkoosh:1994:PRE

[ACC94] M. R. Arkoosh, E. Clemons, and E. Casillas. Proliferative response of English sole splenic leukocytes to mitogens. Transactions of the American Fisheries Society, 123(2):230–241, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Arkoosh:1998:ISJ

[ACH^{+}98] Mary R. Arkoosh, Ed Casillas, Paul Huffman, Ethan Clemons, Joy Evered, John E. Stein, and Usha Varanasi. Increased susceptibility of juvenile Chinook salmon from a contaminated estuary to Vibrio anguillarum. Transactions of the American Fisheries Society, 127(3):360–374, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1993:STH

[AD93] Micheal S. Allen and Dennis R. Devries. Spatial and temporal heterogeneity of larval shad in a large impoundment. Transactions of the American Fisheries Society, 122(6):1070–1079, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:1998:ESD

[ADHB98] David L. Armstrong, Jr., Dennis R. Devries, Chris Harman, and David R. Bayne. Examining similarities and differences between congeners: Do larval gizzard shad and threadfin shad act as ecologically equivalent units? Transactions of the American Fisheries Society, 127(6):1006–1020, November 1998. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:1999:SPY

[AKKH99] S. Reid Adams, Thomas M. Keevin, K. Jack Killgore, and Jan J. Hoover. Stranding potential of young fishes subjected to simulated vessel-induced drawdown. Transactions of the American Fisheries Society, 128(6):1230–1234, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:1995:MCS

[And95] Charles S. Anderson. Measuring and correcting for size selection in electrofishing mark–recapture experiments. Transactions of the American Fisheries Society, 124(5):663–676, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Angermeier:1991:EEP

[Ang91] Paul L. Angermeier. Ecological experiments: Purpose, design, and execution. Transactions of the American Fisheries Society, 120(2):270–272, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1994:E

[Ano94] Anonymous. Errata. Transactions of the American Fisheries Society, 123(3):448, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ault:1996:MSP

[AO96] Jerald S. Ault and Donald B. Olson. A multicohort stock production model. Transactions of the American Fisheries Society, 125(3):343–363, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Adams:1998:ESG

[AREK98] Noah S. Adams, Dennis W. Rondorf, Scott D. Evans, and Joseph E. Kelly. Effects of surgically and gastrically implanted radio transmitters on growth and feeding behavior of juvenile Chinook salmon. Transactions of the American Fisheries Society, 127(1):128–136, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Argent:1995:WQE

[Arg95] David G. Argent. Water quality and the early life stages of fishes. Transactions of the American Fisheries Society, 124(3):458–459, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armour:1992:IFR

[Arm92] Carl L. Armour. Influences of forest and rangeland management on salmonid fishes and their habitats. Transactions of the American Fisheries Society, 121(4):555–556, July 1992. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1992:FIC

[ASOM92] Dennis M. Allen, Susan K. Service, and M. Virginia Ogburn-Matthews. Factors influencing the collection efficiency of estuarine fishes. Transactions of the American Fisheries Society, 121 (2):234–244, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aku:1999:EHO

[AT99] Peter M. K. Aku and William M. Tonn. Effects of hypolimnetic oxygenation on the food resources and feeding ecology of cisco in Amisk Lake, Alberta. Transactions of the American Fisheries Society, 128(1):17–30, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Auer:1996:RSL

[Aue96] Nancy A. Auer. Response of spawning lake sturgeons to change in hydroelectric facility operation. Transactions of the American Fisheries Society, 125(1):66–77, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ault:1994:EHS

[AW94] T. R. Ault and R. W. G. White. Effects of habitat structure and the presence of brown trout on the population density of Galaxias truttaceus in Tasmania, Australia. Transactions of the American Fisheries Society, 123(6):939–949, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ayvazian:1999:FFT

[Ayv99] Suzanne G. Ayvazian. Fish and fisheries of tropical estuaries. Transactions of the American Fisheries Society, 128(1):183–184, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bergstedt:1990:ELT

[BA90] Roger A. Bergstedt and David R. Anderson. Evaluation of line transect sampling based on remotely sensed data from underwater video. Transactions of the American Fisheries Society, 119 (1):86–91, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bayley:1998:CCE

[Bay98] Peter B. Bayley. The catfish connection: Ecology, migration and conservation of Amazon predators. Transactions of the American Fisheries Society, 127(6):1082–1083, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bugert:1991:HUS

[BB91] R. M. Bugert and T. C. Bjornn. Habitat use by steelhead and Coho salmon and their responses to predators and cover in laboratory streams. Transactions of the American Fisheries Society, 120(4):486–493, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bain:1992:EAP

[BB92] Mark B. Bain and Suzanne E. Boltz. Effect of aquatic plant control on the microdistribution and population characteristics of largemouth bass. Transactions of the American Fisheries Society, 121(1):94–103, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bestgen:1998:EFA

[BB98] Kevin R. Bestgen and Jay M. Bundy. Environmental factors affect daily increment deposition and otolith growth in young Colorado squawfish. Transactions of the American Fisheries Society, 127(1):105–117, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bernard:1998:RHE

[BBA98] David R. Bernard, Allen E. Bingham, and Marianna Alexandersdottir. Robust harvest estimates from on-site roving-access creel surveys. Transactions of the American Fisheries Society, 127(3): 481–495, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowen:1999:VSW

[BBA^{+}99] Charles A. Bowen II, Charles R. Bronte, Ray L. Argyle, Jean V. Adams, and James E. Johnson. Vateritic sagitta in wild and stocked lake trout: Applicability to stock origin. Transactions of the American Fisheries Society, 128(5):929–938, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonner:1998:ETE

[BBFW98] T. H. Bonner, T. M. Brandt, J. N. Fries, and B. G. Whiteside. Effects of temperature on egg production and early life stages of the Fountain darter. Transactions of the American Fisheries Society, 127(6):971–978, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldwin:1990:ITW

[BBM90] Nathan W. Baldwin, Craig A. Busack, and Keith O. Meals. Induction of triploidy in white crappie by temperature shock. Transactions of the American Fisheries Society, 119(3):438–444, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bugert:1991:SHU

[BBM91] R. M. Bugert, T. C. Bjornn, and W. R. Meehan. Summer habitat use by young salmonids and their responses to cover and predators in a small Southeast Alaska stream. Transactions of the American Fisheries Society, 120(4):474–485, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beard:1991:ISO

[BC91] Thomas D. Beard, Jr. and Robert F. Carline. Influence of spawning and other stream habitat features on spatial variability of wild brown trout. Transactions of the American Fisheries Society, 120(6):711–722, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1994:AAS

[BC94] Danny J. Brown and Thomas G. Coon. Abundance and assemblage structure of fish larvae in the Lower Missouri River and its tributaries. Transactions of the American Fisheries Society, 123 (5):718–732, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buckel:1996:GER

[BC96] Jeffrey A. Buckel and David O. Conover. Gastric evacuation rates of piscivorous young-of-the-year bluefish. Transactions of the American Fisheries Society, 125(4):591–599, July 1996. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baker:1997:DEA

[BC97] Edward A. Baker and Thomas G. Coon. Development and evaluation of alternative habitat suitability criteria for brook trout. Transactions of the American Fisheries Society, 126(1):65–76, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burnham-Curtis:1996:ORD

[BCB96] Mary K. Burnham-Curtis and Charles R. Bronte. Otoliths reveal a diverse age structure for humper lake trout in Lake Superior. Transactions of the American Fisheries Society, 125(6): 844–851, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beecher:1995:NUD

[BCJ95] Hal A. Beecher, John P. Carleton, and Thom H. Johnson. Notes: Utility of depth and velocity preferences for predicting steelhead parr distribution at different flows. Transactions of the American Fisheries Society, 124(6):935–938, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beecher:1997:TIM

[BCJ97] Hal A. Beecher, John P. Carleton, and Thom H. Johnson. Testing the independence of microhabitat preferences and flow: Response to comments. Transactions of the American Fisheries Society, 126(3):541–542, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [JP97, Wil97].

Bulak:1997:RDS

[BCSD97] James S. Bulak, John S. Crane, David H. Secor, and John M. Dean. Recruitment dynamics of striped bass in the santee– Cooper system, South Carolina. Transactions of the American Fisheries Society, 126(1):133–143, January 1997. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bennett:1997:CTP

[BCWB97] Wayne A. Bennett, Rebecca J. Currie, Paul F. Wagner, and Thomas L. Beitinger. Cold tolerance and potential overwintering of the red-bellied piranha Pygocentrus nattereri in the United States. Transactions of the American Fisheries Society, 126(5): 841–849, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blumberg:1990:ECW

[BD90] Alan F. Blumberg and Dominic M. Di Toro. Effects of climate warming on dissolved oxygen concentrations in Lake Erie. Transactions of the American Fisheries Society, 119(2):210–223, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brule:1999:RGR

[BDCMSC99] Thierry Brulé, Christian Déniel, Teresa Colás-Marrufo, and Manuel Sánchez-Crespo. Red grouper reproduction in the Southern Gulf of Mexico. Transactions of the American Fisheries Society, 128(3):385–402, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baran:1997:VMU

[BDL97] P. Baran, M. Delacoste, and J. M. Lascaux. Variability of mesohabitat used by brown trout populations in the French Central Pyrenées. Transactions of the American Fisheries Society, 126 (5):747–757, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:1999:GSS

[BDZ^{+}99] Brian R. Beckman, Walton W. Dickhoff, Waldo S. Zaugg, Cameron Sharpe, Steve Hirtzel, Robin Schrock, Donald A. Larsen, Richard D. Ewing, Aldo Palmisano, Carl B. Schreck, and Conrad V. W. Mahnken. Growth, smoltification, and smolt-to-adult return of spring Chinook salmon from hatcheries on the Deschutes River, Oregon. Transactions of the American Fisheries Society, 128(6):1125–1150, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:1992:GVD

[BE92] T. D. Beacham and T. P. T. Evelyn. Genetic variation in disease resistance and growth of Chinook, Coho, and chum salmon with respect to vibriosis, furunculosis, and bacterial kidney disease. Transactions of the American Fisheries Society, 121(4):456–485, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beauchamp:1990:SDF

[Bea90] David A. Beauchamp. Seasonal and diel food habits of rainbow trout stocked as juveniles in Lake Washington. Transactions of the American Fisheries Society, 119(3):475–482, May 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1998:VLG

[BEC98] J. Jed Brown, Amir Ehtisham, and David O. Conover. Variation in larval growth rate among striped bass stocks from different latitudes. Transactions of the American Fisheries Society, 127(4): 598–610, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bevelhimer:1996:RIT

[Bev96] Mark S. Bevelhimer. Relative importance of temperature, food, and physical structure to habitat choice by smallmouth bass in laboratory experiments. Transactions of the American Fisheries Society, 125(2):274–283, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bramblett:1991:VFC

[BF91] Robert G. Bramblett and Kurt D. Fausch. Variable fish communities and the index of biotic integrity in a Western Great Plains river. Transactions of the American Fisheries Society, 120(6): 752–769, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bowen:1998:EGH

[BFB98] Zachary H. Bowen, Mary C. Freeman, and Ken D. Bovee. Evaluation of generalized habitat criteria for assessing impacts of altered flow regimes on warmwater fishes. Transactions of the American Fisheries Society, 127(3):455–468, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beck:1992:SSA

[BFBC92] J. C. Beck, K. D. Fulcher, C. F. Beck, and J. G. Cloud. Sperm surface antigen required for fertility: Identification on spermatozoa of rainbow trout by use of monoclonal antibodies. Transactions of the American Fisheries Society, 121(3):333–339, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brandt:1993:UNM

[BFF93] Oscar M. Brandt, Robert W. Fujimura, and Brian J. Finlayson. Use of Neomysis mercedis (Crustacea: Mysidacea) for estuarine toxicity tests. Transactions of the American Fisheries Society, 122(2):279–288, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baxter:1999:GLR

[BFH99] Colden V. Baxter, Christopher A. Frissell, and F. Richard Hauer. Geomorphology, logging roads, and the distribution of bull trout spawning in a forested river basin: Implications for management and conservation. Transactions of the American Fisheries Society, 128(5):854–867, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burger:1995:PSS

[BFHB95] Carl V. Burger, James E. Finn, and Leslie Holland-Bartels. Pattern of shoreline spawning by sockeye salmon in a glacially turbid lake: Evidence for subpopulation differentiation. Transactions of the American Fisheries Society, 124(1):1–15, January 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burns:1997:ASM

[BFM97] Martin D. Burns, Neil H. C. Fraser, and Neil B. Metcalfe. An automated system for monitoring fish activity patterns. Transactions of the American Fisheries Society, 126(6):1036–1040, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berlinsky:1995:AME

[BFOS95] David L. Berlinsky, Mary C. Fabrizio, John F. O’Brien, and Jennifer L. Specker. Age-at-maturity estimates for Atlantic Coast female striped bass. Transactions of the American Fisheries Society, 124(2):207–215, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartley:1990:GSG

[BG90] Devin M. Bartley and G. A. E. Gall. Genetic structure and gene flow in Chinook salmon populations of California. Transactions of the American Fisheries Society, 119(1):55–71, January 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Browder:1994:IAC

[BG94] Richard G. Browder and Greg C. Garman. Increased ammonium concentrations in a tidal freshwater stream during residence of migratory clupeid fishes. Transactions of the American Fisheries Society, 123(6):993–996, November 1994. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Braaten:1999:RBP

[BG99] Patrick J. Braaten and Christopher S. Guy. Relations between physicochemical factors and abundance of fishes in tributary confluences of the Lower Channelized Missouri River. Transactions of the American Fisheries Society, 128(6):1213–1221, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartley:1990:BGD

[BGB90] Devin M. Bartley, Graham A. E. Gall, and Boyd Bentley. Biochemical genetic detection of natural and artificial hybridization of Chinook and Coho salmon in Northern California. Transactions of the American Fisheries Society, 119(3):431–437, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barrett:1992:TIC

[BGR92] Jeffrey C. Barrett, Gary D. Grossman, and J. Rosenfeld. Turbidity-induced changes in reactive distance of rainbow trout. Transactions of the American Fisheries Society, 121(4):437–443, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brandt:1993:IAB

[BH93] Stephen B. Brandt and Kyle J. Hartman. Innovative approaches with bioenergetics models: Future applications to fish ecology and management. Transactions of the American Fisheries Society, 122(5):731–735, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bernard:1995:STM

[BHJ^{+}95] David R. Bernard, Kelly R. Hepler, J. Douglas Jones, Mary E. Whalen, and Douglas N. McBride. Some tests of the “Migration Hypothesis” for anadromous Dolly Varden (southern form). Transactions of the American Fisheries Society, 124(3):297–307, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bronte:1996:DAS

[BHSH96] Charles R. Bronte, Robert J. Hesselberg, John A. Shoesmith, and Michael H. Hoff. Discrimination among spawning concentrations of Lake Superior lake herring based on trace element profiles in sagittae. Transactions of the American Fisheries Society, 125 (6):852–859, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bunnell:1998:DMB

[BIBV98] David B. Bunnell, Jr., J. Jeffery Isely, Kyle H. Burrell, and David H. Van Lear. Diel movement of brown trout in a Southern Appalachian River. Transactions of the American Fisheries Society, 127(4):630–636, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Biro:1998:SCB

[Bir98] Peter A. Biro. Staying cool: Behavioral thermoregulation during summer by young-of-year brook trout in a lake. Transactions of the American Fisheries Society, 127(2):212–222, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bennetti:1992:FAL

[BJ92] Wayne A. Bennetti and Frank W. Judd. Factors affecting the low-temperature tolerance of Texas pinfish. Transactions of the American Fisheries Society, 121(5):659–666, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brandt:1993:SEM

[BK93] Stephen B. Brandt and Jay Kirsch. Spatially explicit models of striped bass growth potential in Chesapeake Bay. Transactions of the American Fisheries Society, 122(5):845–869, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bumguardner:1996:TOC

[BK96] Britt W. Bumguardner and Tim L. King. Toxicity of oxytetracycline and calcein to juvenile striped bass. Transactions of the American Fisheries Society, 125(1):143–145, January 1996. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bjornn:1991:RCA

[BKM91] T. C. Bjornn, S. C. Kirking, and W. R. Meehan. Relation of cover alterations to the summer standing crop of young salmonids in small Southeast Alaska streams. Transactions of the American Fisheries Society, 120(5):562–570, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bryan:1990:MPP

[BKN90] J. D. Bryan, S. W. Kelsch, and W. H. Neill. The maximum power principle in behavioral thermoregulation by fishes. Transactions of the American Fisheries Society, 119(4):611–621, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouwes:1997:FBC

[BL97] Nicolaas Bouwes and Chris Luecke. The fate of Bonneville cisco eggs in Bear Lake: Evaluating mechanisms of egg loss. Transactions of the American Fisheries Society, 126(2):240–247, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1998:IGM

[BLB98] Arthur V. Brown, Madeleine M. Lyttle, and Kristine B. Brown. Impacts of gravel mining on gravel bed streams. Transactions of the American Fisheries Society, 127(6):979–994, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buckel:1998:EDO

[BLC98] Jeffrey A. Buckel, Benjamin H. Letcher, and David O. Conover. Effects of a delayed onset of piscivory on the size of age-0 bluefish. Transactions of the American Fisheries Society, 127(4):576–587, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brannas:1994:UPI

[BLP^{+}94] Eva Brännäs, Hans Lundqvist, Earl Prentice, Monika Schmitz, Kurt Brännäs, and Bo-Sören Wiklund. Use of the Passive Integrated Transponder (PIT) in a fish identification and monitoring system for fish behavioral studies. Transactions of the American Fisheries Society, 123(3):395–401, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Budy:1998:ANE

[BLW98] Phaedra Budy, Chris Luecke, and Wayne A. Wurtsbaugh. Adding nutrients to enhance the growth of endangered sockeye salmon: Trophic transfer in an oligotrophic lake. Transactions of the American Fisheries Society, 127(1):19–34, January 1998. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:1990:TES

[BM90] Terry D. Beacham and Clyde B. Murray. Temperature, egg size, and development of embryos and alevins of five species of Pacific salmon: a comparative analysis. Transactions of the American Fisheries Society, 119(6):927–945, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1991:RRW

[BM91] Michael L. Brown and Brian R. Murphy. Relationship of relative weight (W_{r}) to proximate composition of juvenile striped bass and hybrid striped bass. Transactions of the American Fisheries Society, 120(4):509–518, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Branstetter:1994:AGE

[BM94] Steven Branstetter and John A. Musick. Age and growth estimates for the sand tiger in the Northwestern Atlantic Ocean. Transactions of the American Fisheries Society, 123(2):242–254, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1995:FWM

[BM95] Richard S. Brown and William C. Mackay. Fall and winter movements of and habitat use by cutthroat trout in the Ram River, Alberta. Transactions of the American Fisheries Society, 124(6): 873–885, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Billington:1997:GPC

[BM97] Neil Billington and Michael J. Maceina. Genetic and population characteristics of walleyes in the mobile drainage of Alabama. Transactions of the American Fisheries Society, 126(5):804–814, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Byorth:1998:CIB

[BM98] Patrick A. Byorth and James P. Magee. Competitive interactions between Arctic grayling and brook trout in the Big Hole River drainage, Montana. Transactions of the American Fisheries Society, 127(6):921–931, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bennett:1998:RGL

[BMB98] Wayne A. Bennett, Robert W. McCauley, and Thomas L. Beitinger. Rates of gain and loss of heat tolerance in channel catfish. Transactions of the American Fisheries Society, 127(6): 1051–1058, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barber:1995:EEV

[BMEV95] Willard E. Barber, Lyman L. McDonald, Wallace P. Erickson, and Mark Vallarino. Effect of the Exxon Valdez oil spill on intertidal fish: a field study. Transactions of the American Fisheries Society, 124(4):461–476, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bettoli:1991:CAT

[BMN91] Phillip W. Bettoli, Joseph E. Morris, and Richard L. Noble. Changes in the abundance of two atherinid species after aquatic vegetation removal. Transactions of the American Fisheries Society, 120(1):90–97, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barker:1997:NAD

[BMWB97] L. A. Barker, B. J. Morrison, B. J. Wicks, and F. W. H. Beamish. Notes: Age discrimination and statolith diversity in sea lamprey from streams with varying alkalinity. Transactions of the American Fisheries Society, 126(6):1021–1026, November 1997. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boreman:1997:MCI

[Bor97] John Boreman. Methods for comparing the impacts of pollution and fishing on fish populations. Transactions of the American Fisheries Society, 126(3):506–513, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beacham:1999:PSS

[BPL99] Terry D. Beacham, Susan Pollard, and Khai D. Le. Population structure and stock identification of steelhead in Southern British Columbia, Washington, and the Columbia River based on microsatellite DNA variation. Transactions of the American Fisheries Society, 128(6):1068–1084, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamesderfer:1991:ADN

[BR91a] Raymond C. Beamesderfer and Bruce E. Rieman. Abundance and distribution of northern squawfish, walleyes, and smallmouth bass in John Day Reservoir, Columbia River. Transactions of the American Fisheries Society, 120(4):439–447, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bozek:1991:AHR

[BR91b] Michael A. Bozek and Frank J. Rahel. Assessing habitat requirements of young Colorado River cutthroat trout by use of macrohabitat and microhabitat analyses. Transactions of the American Fisheries Society, 120(5):571–581, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berra:1994:RGG

[BR94a] Tim M. Berra and Victor H. Ruiz. Rediscovery of Galaxias globiceps Eigenmann from Southern Chile. Transactions of the American Fisheries Society, 123(4):595–600, July 1994. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Binkowski:1994:MDR

[BR94b] Fred P. Binkowski and Lars G. Rudstam. Maximum daily ration of Great Lakes bloater. Transactions of the American Fisheries Society, 123(3):335–343, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blanchfield:1996:USM

[BR96] Paul J. Blanchfield and Mark S. Ridgway. Use of seepage meters to measure groundwater flow at brook trout redds. Transactions of the American Fisheries Society, 125(5):813–818, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bradford:1999:TST

[Bra99] Michael J. Bradford. Temporal and spatial trends in the abundance of Coho salmon smolts from Western North America. Transactions of the American Fisheries Society, 128(5):840–846, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blinn:1993:NER

[BRCR93] Dean W. Blinn, Clay Runck, D. Andrew Clark, and John N. Rinne. Notes: Effects of rainbow trout predation on Little Colorado spinedace. Transactions of the American Fisheries Society, 122(1):139–143, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Breck:1993:FTP

[Bre93a] James E. Breck. Foraging theory and piscivorous fish: Are forage fish just big zooplankton? Transactions of the American Fisheries Society, 122(5):902–911, September 1993. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Breck:1993:HWG

[Bre93b] James E. Breck. Hurry up and wait: Growth of young bluegills in ponds and in simulations with an individual-based model. Transactions of the American Fisheries Society, 122(3):467–480, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamesderfer:1995:DDP

[BRN95] Raymond C. P. Beamesderfer, Thomas A. Rien, and Anthony A. Nigro. Differences in the dynamics and potential production of impounded and unimpounded white sturgeon populations in the Lower Columbia River. Transactions of the American Fisheries Society, 124(6):857–872, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brodeur:1992:FRV

[Bro92] Richard D. Brodeur. Factors related to variability in feeding intensity of juvenile Coho salmon and Chinook salmon. Transactions of the American Fisheries Society, 121(1):104–114, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blair:1993:VLH

[BRQ93] Gregory R. Blair, Donald E. Rogers, and Thomas P. Quinn. Variation in life history characteristics and morphology of sockeye salmon in the Kvichak River System, Bristol Bay, Alaska. Transactions of the American Fisheries Society, 122(4):550–559, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beeman:1995:NMS

[BRTV95] John W. Beeman, D. W. Rondorf, M. E. Tilson, and D. A. Venditti. A nonlethal measure of smolt status of juvenile steelhead based on body morphology. Transactions of the American Fisheries Society, 124(5):764–769, September 1995. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bryan:1992:ALM

[BRWB92] C. Frederick Bryan, D. Allen Rutherford, and Barbara Walker-Bryan. Acidification of the Lower Mississippi River. Transactions of the American Fisheries Society, 121(3):369–377, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barans:1993:TTD

[BS93a] Charles A. Barans and Bruce W. Stender. Trends in tilefish distribution and relative abundance off South Carolina and Georgia. Transactions of the American Fisheries Society, 122(2):165–178, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boisclair:1993:TAF

[BS93b] Daniel Boisclair and Pascal Sirois. Testing assumptions of fish bioenergetics models by direct estimation of growth, consumption, and activity rates. Transactions of the American Fisheries Society, 122(5):784–796, September 1993. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bergstedt:1995:ELH

[BS95a] Roger A. Bergstedt and James G. Seelye. Evidence for lack of homing by sea lampreys. Transactions of the American Fisheries Society, 124(2):235–239, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Billington:1995:NMD

[BS95b] Neil Billington and Rex Meade Strange. Notes: Mitochondrial DNA analysis confirms the existence of a genetically divergent walleye population in Northeastern Mississippi. Transactions of the American Fisheries Society, 124(5):770–776, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:1996:SMM

[BS96a] Daniel W. Beckman and Robert G. Schulz. A simple method for marking fish otoliths with alizarin compounds. Transactions of the American Fisheries Society, 125(1):146–149, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonneau:1996:DJB

[BS96b] Joseph L. Bonneau and Dennis L. Scarnecchia. Distribution of juvenile bull trout in a thermal gradient of a plunge pool in Granite Creek, Idaho. Transactions of the American Fisheries Society, 125(4):628–630, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beechie:1997:RBC

[BS97a] Timothy J. Beechie and Thomas H. Sibley. Relationships between channel characteristics, woody debris, and fish habitat in Northwestern Washington streams. Transactions of the American Fisheries Society, 126(2):217–229, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bremigan:1997:EAI

[BS97b] Mary T. Bremigan and Roy A. Stein. Experimental assessment of the influence of zooplankton size and density on gizzard shad recruitment. Transactions of the American Fisheries Society, 126(4):622–637, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bremigan:1999:LGS

[BS99] Mary T. Bremigan and Roy A. Stein. Larval gizzard shad success, juvenile effects, and reservoir productivity: Toward a framework for multi-system management. Transactions of the American Fisheries Society, 128(6):1106–1124, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burger:1997:GDS

[BSC97] Carl V. Burger, William J. Spearman, and Matthew A. Cronin. Genetic differentiation of sockeye salmon subpopulations from a geologically young Alaskan Lake system. Transactions of the American Fisheries Society, 126(6):926–938, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1999:MDM

[BSEK99] B. L. Brown, P. E. Smouse, J. M. Epifanio, and C. J. Kobak. Mitochondrial DNA mixed-stock analysis of American shad: Coastal harvests are dynamic and variable. Transactions of the American Fisheries Society, 128(6):977–994, November 1999.

CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:1990:AGB

[BSRW90] Daniel W. Beckman, A. Louise Stanley, Jeffrey H. Render, and Charles A. Wilson. Age and growth of black drum in Louisiana waters of the Gulf of Mexico. Transactions of the American Fisheries Society, 119(3):537–544, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bue:1998:EDP

[BSS98] Brian G. Bue, Samuel Sharr, and James E. Seeb. Evidence of damage to pink salmon populations inhabiting Prince William Sound, Alaska, two generations after the Exxon Valdez oil spill. Transactions of the American Fisheries Society, 127(1):35–43, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bryant:1998:CSP

[BSWW98] Mason D. Bryant, Douglas N. Swanston, Robert C. Wissmar, and Brenda E. Wright. Coho salmon populations in the karst landscape of North Prince of Wales Island, Southeast Alaska. Transactions of the American Fisheries Society, 127(3):425–433, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:1990:HIS

[BT90] F. W. H. Beamish and E. A. Trippel. Heat increment: a static or dynamic dimension in bioenergetic models? Transactions of the American Fisheries Society, 119(4):649–661, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bradford:1997:ERC

[BTA97] Michael J. Bradford, Garth C. Taylor, and J. Andrew Allan. Empirical review of Coho salmon smolt abundance and the prediction of smolt production at the regional level. Transactions of the American Fisheries Society, 126(1):49–64, January 1997. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:1992:SDP

[BTM92] Richard J. Beamish, Barbara L. Thomson, and Gordon A. Mc-Farlane. Spiny dogfish predation on Chinook and Coho salmon and the potential effects on hatchery-produced salmon. Transactions of the American Fisheries Society, 121(4):444–455, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berejikian:1999:CDB

[BTS^{+}99] Barry A. Berejikian, E. Paul Tezak, Steven L. Schroder, Thomas A. Flagg, and Curtis M. Knudsen. Competitive differences between newly emerged offspring of captive-reared and wild Coho salmon. Transactions of the American Fisheries Society, 128(5):832–839, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bumguardner:1991:MSR

[Bum91] Britt W. Bumguardner. Marking subadult red drums with oxytetracycline. Transactions of the American Fisheries Society, 120 (4):537–540, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burleson:1999:ABF

[Bur99] Mark L. Burleson. Air-breathing fishes: Evolution, diversity and adaptation. Transactions of the American Fisheries Society, 128(1):184, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Betsill:1997:ETZ

[BV97] R. K. Betsill and M. J. Van Den Avyle. Effect of temperature and zooplankton abundance on growth and survival of larval threadfin shad. Transactions of the American Fisheries Society, 126(6): 999–1011, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baltz:1991:SCM

[BVBM91] Donald M. Baltz, Bruce Vondracek, Larry R. Brown, and Peter B. Moyle. Seasonal changes in microhabitat selection by rainbow trout in a small stream. Transactions of the American Fisheries Society, 120(2):166–176, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burkhead:1991:IHN

[BW91] Noel M. Burkhead and James D. Williams. An intergeneric hybrid of a native minnow, the golden shiner, and an exotic minnow, the rudd. Transactions of the American Fisheries Society, 120 (6):781–795, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bestgen:1994:EFC

[BW94] Kevin R. Bestgen and Mark A. Williams. Effects of fluctuating and constant temperatures on early development and survival of Colorado squawfish. Transactions of the American Fisheries Society, 123(4):574–579, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baras:1999:TES

[BW99] E. Baras and L. Westerloppe. Transintestinal expulsion of surgically implanted tags by African catfish Heterobranchus longifilis of variable size and age. Transactions of the American Fisheries Society, 128(4):737–746, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bart:1998:CBC

[BWD98] Amrit N. Bart, Dwight F. Wolfe, and Rex A. Dunham. Cryopreservation of blue catfish spermatozoa and subsequent fertilization of channel catfish eggs. Transactions of the American Fisheries Society, 127(5):819–824, September 1998. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baylis:1993:ALH

[BWH93] Jeffrey R. Baylis, Daniel D. Wiegmann, and Michael H. Hoff. Alternating life histories of smallmouth bass. Transactions of the American Fisheries Society, 122(3):500–510, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bates:1996:EOP

[BWT96] Mark C. Bates, William R. Wayman, and Terrence R. Tiersch. Effect of osmotic pressure on the activation and storage of channel catfish sperm. Transactions of the American Fisheries Society, 125(5):798–802, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bain:1990:MHU

[BWTM90] Mark B. Bain, David H. Webb, Michael D. Tangedal, and Larry N. Mangum. Movements and habitat use by grass carp in a large mainstream reservoir. Transactions of the American Fisheries Society, 119(3):553–561, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bremer:1998:PCR

[BZBE98] Jaime R. Alvarado Bremer, Liang Zhang, James S. Bulak, and Bert Ely. A polymerase chain reaction–restriction fragment length polymorphism (PCR–RFLP) assay for the discrimination of mitochondrial DNA from the Florida and northern subspecies of largemouth bass. Transactions of the American Fisheries Society, 127(3):507–511, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campton:1991:RSH

[CAB^{+}91] Donald E. Campton, Fred W. Allendorf, Robert J. Behnke, Fred M. Utter, Mark W. Chilcote, Steven A. Leider, and John J. Loch. Reproductive success of hatchery and wild steelhead. Transactions of the American Fisheries Society, 120(6):816–827, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campana:1995:GSM

[CAM95] Steven E. Campana, M. Christina Annand, and James I. McMillan. Graphical and statistical methods for determining the consistency of age determinations. Transactions of the American Fisheries Society, 124(1):131–138, January 1995. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Campbell:1998:PSW

[Cam98] Elizabeth A. Campbell. Predation by small walleyes on yellow perch: Effects of prey size distribution. Transactions of the American Fisheries Society, 127(4):588–597, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:1992:GFQ

[Car92] Robert Carlson. Global freshwater quality: a first assessment. Transactions of the American Fisheries Society, 121(1):141, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carl:1995:STB

[Car95] L. M. Carl. Sonic tracking of burbot in Lake Opeongo, Ontario. Transactions of the American Fisheries Society, 124(1):77–83, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Casselman:1990:GRS

[Cas90] John M. Casselman. Growth and relative size of calcified structures of fish. Transactions of the American Fisheries Society, 119(4):673–688, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutant:1990:SHS

[CB90] Charles C. Coutant and Denise L. Benson. Summer habitat suitability for striped bass in Chesapeake Bay: Reflections on a population decline. Transactions of the American Fisheries Society, 119(4):757–778, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connelly:1991:NEV

[CB91] Nancy A. Connelly and Tommy L. Brown. Net economic value of the freshwater recreational fisheries of New York. Transactions of the American Fisheries Society, 120(6):770–775, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1992:FHW

[CB92] John H. Clark and David R. Bernard. Fecundity of humpback whitefish and least cisco in the Chatanika River, Alaska. Transactions of the American Fisheries Society, 121(2):268–273, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connelly:1995:UAD

[CB95] Nancy A. Connelly and Tommy L. Brown. Use of angler diaries to examine biases associated with 12-month recall on Mail questionnaires. Transactions of the American Fisheries Society, 124(3):413–422, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cailliet:1996:DCA

[CBB^{+}96] Gregor M. Cailliet, L. W. Botsford, John G. Brittnacher, Gary Ford, M. Matsubayashi, Aaron King, Diana L. Watters, and Robert G. Kope. Development of a computer-aided age determination system: Evaluation based on otoliths of bank rockfish off California. Transactions of the American Fisheries Society, 125 (6):874–888, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coble:1990:LTS

[CBFS90] Daniel W. Coble, Richard E. Bruesewitz, Thomas W. Fratt, and Jeffrey W. Scheirer. Lake trout, sea lampreys, and overfishing in the Upper Great Lakes: a review and reanalysis. Transactions of the American Fisheries Society, 119(6):985–995, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Curry:1997:USS

[CBND97] R. Allen Curry, Charles Brady, David L. G. Noakes, and Roy G. Danzmann. Use of small streams by young brook trout spawned in a lake. Transactions of the American Fisheries Society, 126(1): 77–83, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cech:1996:SBE

[CBYH96] Joseph J. Cech, Jr., Steven D. Bartholow, Paciencia S. Young, and Todd E. Hopkins. Striped bass exercise and handling stress in freshwater: Physiological responses to recovery environment. Transactions of the American Fisheries Society, 125(2):308–320, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chiarella:1990:SSF

[CC90] Louis A. Chiarella and David O. Conover. Spawning season and first-year growth of adult bluefish from the New York bight. Transactions of the American Fisheries Society, 119(3):455–462, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Childs:1996:TES

[CC96] Michael R. Childs and Robert W. Clarkson. Temperature effects on swimming performance of larval and juvenile Colorado squawfish: Implications for survival and species recovery. Transactions of the American Fisheries Society, 125(6):940–947, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clapp:1990:RAH

[CCD90] David F. Clapp, Richard D. Clark, Jr., and James S. Diana. Range, activity, and habitat of large, free-ranging brown trout in a Michigan stream. Transactions of the American Fisheries Society, 119(6):1022–1034, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Childs:1998:RUL

[CCR98] Michael R. Childs, Robert W. Clarkson, and Anthony T. Robinson. Resource use by larval and early juvenile native fishes in the Little Colorado River, Grand Canyon, Arizona. Transactions of the American Fisheries Society, 127(4):620–629, July 1998. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Castonguay:1990:LAC

[CDAM90] Martin Castonguay, Jean-Denis Dutil, Céline Audet, and Roberta Miller. Locomotor activity and concentration of thyroid hormones in migratory and sedentary juvenile American eels. Transactions of the American Fisheries Society, 119(6): 946–956, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chan:1997:LEW

[CDK97] Matthew D. Chan, Eric D. Dibble, and K. Jack Kilgore. A laboratory examination of water velocity and substrate preference by age-0 Gulf sturgeons. Transactions of the American Fisheries Society, 126(2):330–333, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cyr:1992:SLF

[CDL^{+}92] H. Cyr, J. A. Downing, S. Lalonde, S. B. Baines, and M. L. Pace. Sampling larval fish populations: Choice of sample number and size. Transactions of the American Fisheries Society, 121(3): 356–368, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ciereszko:1999:EET

[CDL^{+}99] A. Ciereszko, K. Dabrowski, F. Lin, S. A. Christ, and G. P. Toth. Effects of extenders and time of storage before freezing on motility and fertilization of cryopreserved muskellunge spermatozoa. Transactions of the American Fisheries Society, 128(3):542–548, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Counihan:1999:IEA

[CF99] Timothy D. Counihan and Conrad N. Frost. Influence of externally attached transmitters on the swimming performance of juvenile white sturgeon. Transactions of the American Fisheries Society, 128(5):965–970, September 1999. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cargnelli:1997:NFE

[CG97] Luca M. Cargnelli and Mart R. Gross. Notes: Fish energetics: Larger individuals Emerge from winter in better condition. Transactions of the American Fisheries Society, 126(1):153–156, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:1993:DSV

[CH93] Nicholas C. Collins and Scott G. Hinch. Diel and seasonal variation in foraging activities of pumpkinseeds in an Ontario pond. Transactions of the American Fisheries Society, 122(3):357–365, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chen:1995:GAF

[CH95] Y. Chen and H. H. Harvey. Growth, abundance, and food supply of white sucker. Transactions of the American Fisheries Society, 124(2):262–271, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Connolly:1999:BCC

[CH99] Patrick J. Connolly and James D. Hall. Biomass of coastal cutthroat trout in unlogged and previously clear-cut basins in the Central Coast Range of Oregon. Transactions of the American Fisheries Society, 128(5):890–899, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chambers:1993:PVF

[Cha93] R. Christopher Chambers. Phenotypic variability in fish populations and its representation in individual-based models. Transactions of the American Fisheries Society, 122(3):404–414, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carmichael:1998:SMT

[CHH98] John T. Carmichael, Steven L. Haeseker, and Joseph E. Hightower. Spawning migration of telemetered striped bass in the Roanoke River, North Carolina. Transactions of the American Fisheries Society, 127(2):286–297, March 1998. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carl:1994:RTA

[CHI94] L. M. Carl, C. Hunt, and P. E. Ihssen. Rainbow trout of the Athabasca River, Alberta: a unique population. Transactions of the American Fisheries Society, 123(2):129–140, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carmichael:1993:IAA

[CHSM93] Gary J. Carmichael, James N. Hanson, Maureen E. Schmidt, and Donald C. Morizot. Introgression among Apache, cutthroat, and rainbow trout in Arizona. Transactions of the American Fisheries Society, 122(1):121–130, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chen:1995:REM

[CJ95] Y. Chen and D. A. Jackson. Robust estimation of mean and variance in fisheries. Transactions of the American Fisheries Society, 124(3):401–412, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cole:1997:RAS

[CM97] Michael B. Cole and John R. Moring. Relation of adult size to movements and distribution of smallmouth bass in a Central Maine Lake. Transactions of the American Fisheries Society, 126 (5):815–821, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Counihan:1998:EDG

[CMMP98] Timothy D. Counihan, Allen I. Miller, Matthew G. Mesa, and Michael J. Parsley. The effects of dissolved gas supersaturation on white sturgeon larvae. Transactions of the American Fisheries Society, 127(2):316–322, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cormier:1993:NOT

[CNWT93] Susan M. Cormier, Timothy W. Neiheisel, Daniel E. Williams, and Terrence R. Tiersch. Natural occurrence of triploidy in a wild brown bullhead. Transactions of the American Fisheries Society, 122(3):390–392, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conover:1990:RBC

[Con90] David O. Conover. The relation between capacity for growth and length of growing season: Evidence for and implications of countergradient variation. Transactions of the American Fisheries Society, 119(3):416–430, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conquest:1998:SAI

[Con98] Loveday L. Conquest. Stock assessment in inland fisheries. Transactions of the American Fisheries Society, 127(1):153–154, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutant:1990:TOH

[Cou90] Charles C. Coutant. Temperature-oxygen habitat for freshwater and coastal striped bass in a changing climate. Transactions of the American Fisheries Society, 119(2):240–253, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coutant:1996:CEI

[Cou96] Charles C. Coutant. Comment: Effects of instream brush on juvenile Coho salmon. Transactions of the American Fisheries Society, 125(1):150–151, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [QPS96].

Cowx:1997:SAQ

[Cow97] Ian G. Cowx. Stock assessment: Quantitative methods and applications for small-scale fisheries. Transactions of the American Fisheries Society, 126(5):876–878, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chen:1994:RRA

[CP94] Y. Chen and J. E. Paloheimo. Robust regression approach to estimating fish mortality rates with a cohort-based model. Transactions of the American Fisheries Society, 123(4):508–518, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chilton:1997:DMB

[CP97] Earl W. Chilton II and Steven M. Poarch. Distribution and movement behavior of radio-tagged grass carp in two Texas reservoirs. Transactions of the American Fisheries Society, 126(3):467–476, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1997:FAC

[CR97a] Mark E. Clark and Kenneth A. Rose. Factors affecting competitive dominance of rainbow trout over brook trout in Southern Appalachian streams: Implications of an individual-based model. Transactions of the American Fisheries Society, 126(1): 1–20, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clements:1997:EHM

[CR97b] William H. Clements and David E. Rees. Effects of heavy metals on prey abundance, feeding habits, and metal uptake of brown trout in the Arkansas River, Colorado. Transactions of the American Fisheries Society, 126(5):774–785, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cowan:1993:IBM

[CRRH93] James H. Cowan, Jr., Kenneth A. Rose, Edward S. Rutherford, and Edward D. Houde. Individual-based model of young-of-the-year striped bass population dynamics. II. Factors affecting recruitment in the Potomac River, Maryland. Transactions of the American Fisheries Society, 122(3):439–458, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:1991:SVS

[CS91] Mark R. Collins and George R. Sedberry. Status of vermilion snapper and red porgy stocks off South Carolina. Transactions of the American Fisheries Society, 120(1):116–120, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clarke:1999:BTP

[CS99] K. D. Clarke and D. A. Scruton. Brook trout production dynamics in the streams of a low fertility Newfoundland watershed. Transactions of the American Fisheries Society, 128(6): 1222–1229, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cowan:1995:NPR

[CSBH95] J. H. Cowan, Jr., R. L. Shipp, H. K. Bailey IV, and D. W. Haywick. Notes: Procedure for rapid processing of large otoliths. Transactions of the American Fisheries Society, 124(2):280–282, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cochran:1999:TES

[CSK99] Philip A. Cochran, William D. Swink, and Andrew P. Kinziger. Testing and extension of a sea lamprey feeding model. Transactions of the American Fisheries Society, 128(3):403–413, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carmichael:1992:EIG

[CSM92] Gary J. Carmichael, Maureen E. Schmidt, and Donald C. Morizot. Electrophoretic identification of genetic markers in channel catfish and blue catfish by use of low-risk tissues. Transactions of the American Fisheries Society, 121(1):26–35, January 1992. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Culp:1996:SFW

[CST96] Joseph M. Culp, Garry J. Scrimgeour, and Greg D. Townsend. Simulated fine woody debris accumulations in a stream increase rainbow trout fry abundance. Transactions of the American Fisheries Society, 125(3):472–479, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cuenco:1994:MIS

[Cue94] Michael L. Cuenco. A model of an internally supplemented population. Transactions of the American Fisheries Society, 123(3): 277–288, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chick:1999:EZS

[CV99] John H. Chick and Michael J. Van Den Avyle. Effects of zooplankton spatial variation on growth of larval striped bass: an experimental approach. Transactions of the American Fisheries Society, 128(2):339–351, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clarkson:1995:TBS

[CW95] Robert W. Clarkson and Jeffrey R. Wilson. Trout biomass and stream habitat relationships in the White Mountains Area, East-Central Arizona. Transactions of the American Fisheries Society, 124(4):599–612, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clapp:1996:CFC

[CW96] David F. Clapp and David H. Wahl. Comparison of food consumption, growth, and metabolism among muskellunge: an investigation of population differentiation. Transactions of the American Fisheries Society, 125(3):402–410, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crumby:1990:CBI

[CWBC90] W. Dennis Crumby, Mark A. Webb, Frank J. Bulow, and Harold Joe Cathey. Changes in biotic integrity of a river in North-Central Tennessee. Transactions of the American Fisheries Society, 119(5):885–893, September 1990. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cox:1999:PET

[CWK99] David L. Cox, Paddy Walker, and Thomas J. Koob. Predation on eggs of the thorny skate. Transactions of the American Fisheries Society, 128(2):380–384, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conrow:1990:DAE

[CZG90] Roxanne Conrow, Alexander V. Zale, and Richard W. Gregory. Distributions and abundances of early life stages of fishes in a Florida Lake dominated by aquatic macrophytes. Transactions of the American Fisheries Society, 119(3):521–528, May 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dabrowski:1990:FN

[Dab90] Konrad Dabrowski. Fish nutrition. Transactions of the American Fisheries Society, 119(5):924–926, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Daniels:1995:FFV

[Dan95] Robert A. Daniels. Freshwater fishes of Virginia. Transactions of the American Fisheries Society, 124(2):283–284, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dicenzo:1995:VDR

[DB95] Victor J. Dicenzo and Phillip W. Bettoli. Verification of daily ring deposition in the otoliths of age-0 spotted bass. Transactions of the American Fisheries Society, 124(4):633–636, July 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donald:1992:ESO

[DBCM92] David B. Donald, John A. Babaluk, John F. Craig, and William A. Musker. Evaluation of the scale and operculum methods to determine age of adult goldeyes with special reference to a dominant year-class. Transactions of the American Fisheries Society, 121(6):792–796, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deangelis:1990:PAM

[DC90] D. L. Deangelis and R. M. Cushman. Potential application of models in forecasting the effects of climate changes on fisheries. Transactions of the American Fisheries Society, 119(2):224–239, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diamond:1999:CBQ

[DCC99] Sandra L. Diamond, Larry B. Crowder, and Lindsay G. Cowell. Catch and bycatch: The qualitative effects of fisheries on population vital rates of Atlantic croaker. Transactions of the American Fisheries Society, 128(6):1085–1105, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1997:URO

[DCE97] C. L. Davis, L. M. Carl, and D. O. Evans. Use of a remotely operated vehicle to study habitat and population density of juvenile lake trout. Transactions of the American Fisheries Society, 126 (5):871–875, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dong:1998:CCC

[DD98] Quan Dong and Donald L. Deangelis. Consequences of cannibalism and competition for food in a smallmouth bass population: an individual-based modeling study. Transactions of the American Fisheries Society, 127(2):174–191, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dettmers:1996:QRH

[DDS96] John M. Dettmers, Dennis R. Devries, and Roy A. Stein. Quantifying responses to hybrid striped bass predation across multiple trophic levels: Implications for reservoir biomanipulation. Transactions of the American Fisheries Society, 125(4):491–504, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Decker:1992:STS

[DE92] Lynn M. Decker and Don C. Erman. Short-term seasonal changes in composition and abundance of fish in Sagehen Creek, California. Transactions of the American Fisheries Society, 121(3): 297–306, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Devries:1993:FE

[Dev93] Dennis R. Devries. Fish ecology. Transactions of the American Fisheries Society, 122(2):305–308, March 1993. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1991:FBH

[DF91] Bruce M. Davis and Jeffrey W. Foltz. Food of blueback herring and threadfin shad in Jocassee Reservoir, South Carolina. Transactions of the American Fisheries Society, 120(5):605–613, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dolloff:1994:FHF

[DFO94] C. Andrew Dolloff, Patricia A. Flebbe, and Michael D. Owen. Fish habitat and fish populations in a Southern Appalachian watershed before and after Hurricane Hugo. Transactions of the American Fisheries Society, 123(4):668–678, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dolloff:1994:SFR

[DFT94] C. Andrew Dolloff, Patricia A. Flebbe, and John E. Thorpe. Salmonid flexibility: Responses to environmental extremes. Transactions of the American Fisheries Society, 123(4):606–612, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deegan:1999:IEV

[DGHP99] Linda A. Deegan, Heidi E. Golden, Chris J. Harvey, and Bruce J. Peterson. Influence of environmental variability on the growth of age-0 and adult Arctic grayling. Transactions of the American Fisheries Society, 128(6):1163–1175, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dew:1994:RGM

[DH94] C. Braxton Dew and Jack H. Hecht. Recruitment, growth, mortality, and biomass production of larval and early juvenile Atlantic tomcod in the Hudson River Estuary. Transactions of the American Fisheries Society, 123(5):681–702, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dorazio:1994:TRE

[DHMS94] Robert M. Dorazio, Kathryn A. Hattala, Carol B. McCollough, and Jorgen E. Skjeveland. Tag recovery estimates of migration of striped bass from spawning areas of the Chesapeake Bay. Transactions of the American Fisheries Society, 123(6):950–963, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doucett:1999:IAN

[DHP99] Richard R. Doucett, William Hooper, and Geoff Power. Identification of anadromous and nonanadromous adult brook trout and their progeny in the Tabusintac River, New Brunswick, by means of multiple-stable-isotope analysis. Transactions of the American Fisheries Society, 128(2):278–288, March 1999. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1996:STW

[DHR96] Gary E. Davis, Peter L. Haaker, and Daniel V. Richards. Status and trends of white abalone at the California Channel Islands. Transactions of the American Fisheries Society, 125(1):42–48, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dahl-Hansen:1994:SPP

[DHRK94] Geir A. P. Dahl-Hansen, Stein H. Rubach, and Anders Klemetsen. Selective predation by pelagic Arctic char on crustacean plankton in Takvatn, Northern Norway, before and after mass removal of Arctic char. Transactions of the American Fisheries Society, 123(3):385–394, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dibble:1991:CDR

[Dib91] Eric D. Dibble. A comparison of diving and rotenone methods for determining relative abundance of fish. Transactions of the American Fisheries Society, 120(5):663–666, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dickey:1997:SGD

[DIT97] Clinton L. Dickey, J. Jeffery Isely, and Joseph R. Tomasso. Slow growth did not decouple the otolith size–fish size relationship in striped bass. Transactions of the American Fisheries Society, 126 (6):1027–1029, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauble:1999:FCS

[DJG99] Dennis D. Dauble, Robert L. Johnson, and Aaron P. Garcia. Fall Chinook salmon spawning in the tailraces of Lower Snake River hydroelectric projects. Transactions of the American Fisheries Society, 128(4):672–679, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Devore:1995:DPP

[DJTH95] John D. Devore, Brad W. James, Charles A. Tracy, and Donna A. Hale. Dynamics and potential production of white sturgeon in the unimpounded Lower Columbia River. Transactions of the American Fisheries Society, 124(6):845–856, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diaz:1997:NDM

[DLF97] Marilyn Diaz, Gilles M. Leclerc, and Bert Ely Fishtec. Nuclear DNA markers reveal low levels of genetic divergence among Atlantic and Gulf of Mexico populations of striped bass. Transactions of the American Fisheries Society, 126(1):163–165, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Douglas:1998:MDC

[DMM98] Michael E. Douglas, Robert R. Miller, and W. L. Minckley. Multivariate discrimination of Colorado Plateau Gila spp.: The “Art of Seeing Well” revisited. Transactions of the American Fisheries Society, 127(2):163–173, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Demers:1996:APL

[DMWM96] Eric Demers, R. Scott McKinley, Allan H. Weatherley, and Donald I. McQueen. Activity patterns of largemouth and smallmouth bass determined with electromyogram biotelemetry. Transactions of the American Fisheries Society, 125(3):434–439, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunham:1994:HBG

[DNR^{+}94] Rex A. Dunham, Kimberly G. Norgren, Lynn Robison, R. Oneal Smitherman, Tom Steeger, Douglas C. Peterson, and Michael Gibson. Hybridization and biochemical genetics of black and white crappies in the Southeastern USA. Transactions of the American Fisheries Society, 123(2):141–149, March 1994. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dolman:1990:CTR

[Dol90] William B. Dolman. Classification of Texas reservoirs in relation to limnology and fish community associations. Transactions of the American Fisheries Society, 119(3):511–520, May 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donald:1997:RBY

[Don97] David B. Donald. Relationship between year-class strength for goldeyes and selected environmental variables during the first year of life. Transactions of the American Fisheries Society, 126(3):361–368, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauble:1994:RCW

[DP94] Dennis D. Dauble and Ted M. Poston. Radionuclide concentrations in white sturgeons from the Hanford reach of the Columbia River. Transactions of the American Fisheries Society, 123(4): 565–573, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunham:1999:LGV

[DPR^{+}99] J. B. Dunham, M. M. Peacock, B. E. Rieman, R. E. Schroeter, and G. L. Vinyard. Local and geographic variability in the distribution of stream-living Lahontan cutthroat trout. Transactions of the American Fisheries Society, 128(5):875–889, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeStaso:1994:IWT

[DR94] James De Staso III and Frank J. Rahel. Influence of water temperature on interactions between juvenile Colorado River cutthroat trout and brook trout in a laboratory stream. Transactions of the American Fisheries Society, 123(3):289–297, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diodati:1996:MSB

[DR96] Paul J. Diodati and R. Anne Richards. Mortality of striped bass hooked and released in salt water. Transactions of the American Fisheries Society, 125(2):300–307, March 1996. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Die:1990:EPM

[DRF90] David J. Die, Victor R. Restrepo, and William W. Fox, Jr. Equilibrium production models that incorporate fished area. Transactions of the American Fisheries Society, 119(3):445–454, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Distefano:1994:RPF

[DRWD94] Robert J. Distefano, Michael J. Roell, Bruce A. Wagner, and Jamey J. Decoske. Relative performances of four preservatives on fish and crayfish. Transactions of the American Fisheries Society, 123(5):817–823, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Diana:1990:ESG

[DS90] James S. Diana and Ronald Salz. Energy storage, growth, and maturation of yellow perch from different locations in Saginaw Bay, Michigan. Transactions of the American Fisheries Society, 119(6):976–984, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dettmers:1992:FCL

[DS92] John M. Dettmers and Roy A. Stein. Food consumption by larval gizzard shad: Zooplankton effects and implications for reservoir communities. Transactions of the American Fisheries Society, 121(4):494–507, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dettmers:1996:QLA

[DS96] John M. Dettmers and Roy A. Stein. Quantifying linkages among gizzard shad, zooplankton, and phytoplankton in reservoirs. Transactions of the American Fisheries Society, 125(1): 27–41, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1997:ERH

[DS97] Lawrence E. Davis and Carl B. Schreck. The energetic response to handling stress in juvenile Coho salmon. Transactions of the American Fisheries Society, 126(2):248–258, March 1997. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Devries:1998:PSL

[DSB98] Dennis R. Devries, Roy A. Stein, and Mary T. Bremigan. Prey selection by larval fishes as influenced by available zooplankton and gape limitation. Transactions of the American Fisheries Society, 127(6):1040–1050, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davidson:1999:ETS

[DSD99] G. W. Davidson, M. K. Sheehan, and Peter S. Davie. The effect of tagging on the swimming performance of rainbow trout as a surrogate for Kahawai Arripis trutta. Transactions of the American Fisheries Society, 128(5):971–973, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dettmers:1998:PRA

[DSL98] John M. Dettmers, Roy A. Stein, and Edward M. Lewis. Potential regulation of age-0 gizzard shad by hybrid striped bass in Ohio reservoirs. Transactions of the American Fisheries Society, 127(1):84–94, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Devries:1991:STS

[DSMM91] Dennis R. Devries, Roy A. Stein, Jeffrey G. Miner, and Gary G. Mittelbach. Stocking threadfin shad: Consequences for youngof-year fishes. Transactions of the American Fisheries Society, 120(3):368–381, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deangelis:1993:MGS

[DSRS93] D. L. Deangelis, B. J. Shuter, M. S. Ridgway, and Marten Scheffer. Modeling growth and survival in an age-0 fish cohort. Transactions of the American Fisheries Society, 122(5):927–941, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunham:1997:ISL

[DV97] Jason B. Dunham and Gary L. Vinyard. Incorporating stream level variability into analyses of site level fish habitat relationships: Some cautionary examples. Transactions of the American Fisheries Society, 126(2):323–329, March 1997. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dickerson:1999:EHC

[DV99a] Bobette R. Dickerson and Gary L. Vinyard. Effects of high chronic temperatures and diel temperature cycles on the survival and growth of Lahontan cutthroat trout. Transactions of the American Fisheries Society, 128(3):516–521, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dickerson:1999:EHL

[DV99b] Bobette R. Dickerson and Gary L. Vinyard. Effects of high levels of total dissolved solids in Walker Lake, Nevada, on survival and growth of Lahontan cutthroat trout. Transactions of the American Fisheries Society, 128(3):507–515, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dewald:1992:IBN

[DW92] Lynn Dewald and Margaret A. Wilzbach. Interactions between native brook trout and hatchery brown trout: Effects on habitat use, feeding, and growth. Transactions of the American Fisheries Society, 121(3):287–296, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunning:1997:UAT

[DWRM97] Dennis J. Dunning, John R. Waldman, Quentin E. Ross, and Mark T. Mattson. Use of Atlantic tomcod and other prey by striped bass in the Lower Hudson River Estuary during winter. Transactions of the American Fisheries Society, 126(5):857–861, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dygert:1990:SCE

[Dyg90] Peter H. Dygert. Seasonal changes in energy content and proximate composition associated with somatic growth and reproduction in a representative age-class of female English sole. Transactions of the American Fisheries Society, 119(5):791–801, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ehrhardt:1992:ATL

[EA92] Nelson M. Ehrhardt and Jerald S. Ault. Analysis of two lengthbased mortality models applied to bounded catch length frequencies. Transactions of the American Fisheries Society, 121(1):115– 122, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Escos:1995:DIP

[EAEA95] J. Escós, C. L. Alados, J. M. Emlen, and S. Alderstein. Developmental instability in the Pacific hake parasitized by myxosporeans Kudoa spp. Transactions of the American Fisheries Society, 124(6):943–945, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edsall:1999:GTR

[Eds99] Thomas A. Edsall. The growth–temperature relation of juvenile lake whitefish. Transactions of the American Fisheries Society, 128(5):962–964, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Epifanio:1999:RSA

[EHBP99] John M. Epifanio, Michael Hooe, D. Homer Buck, and David P. Philipp. Reproductive success and assortative mating among Pomoxis species and their hybrids. Transactions of the American Fisheries Society, 128(1):104–120, January 1999. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eiler:1995:RSL

[Eil95] John H. Eiler. A remote satellite-linked tracking system for studying Pacific salmon with radio telemetry. Transactions of the American Fisheries Society, 124(2):184–193, March 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eisler:1995:ATM

[Eis95] Ronald Eisler. Aquatic toxicology: Molecular, biochemical, and cellular perspectives. Transactions of the American Fisheries Society, 124(5):786–787, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eldridge:1995:TSV

[EJ95] Maxwell B. Eldridge and Brian M. Jarvis. Temporal and spatial variation in fecundity of yellowtail rockfish. Transactions of the American Fisheries Society, 124(1):16–25, January 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edsall:1993:DAR

[EKH93] Thomas A. Edsall, Gregory W. Kennedy, and William H. Horns. Distribution, abundance, and resting microhabitat of burbot on Julian’s Reef, Southwestern Lake Michigan. Transactions of the American Fisheries Society, 122(4):560–574, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Epifanio:1996:GVP

[EKNP96] John M. Epifanio, Jeffrey B. Koppelman, Michael A. Nedbal, and David P. Philipp. Geographic variation of paddlefish allozymes and mitochondrial DNA. Transactions of the American Fisheries Society, 125(4):546–561, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ensign:1997:FIS

[ELAD97] William E. Ensign, Kevin N. Leftwich, Paul L. Angermeier, and C. Andrew Dolloff. Factors influencing stream fish recovery following a large-scale disturbance. Transactions of the American Fisheries Society, 126(6):895–907, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:1990:FEJ

[EMC90] D. O. Evans, A. M. McCombie, and J. M. Casselman. Frederick Ernest Joseph Fry, 1908–1989. Transactions of the American Fisheries Society, 119(4):571–573, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:1990:IPS

[EN90] David O. Evans and William H. Neill. Introduction to the Proceedings of the Symposium “From Environment to Fish to Fisheries: a Tribute to F. E. J. Fry”. Transactions of the American Fisheries Society, 119(4):567–570, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eiler:1992:RSS

[ENB92] John H. Eiler, Bonita D. Nelson, and Robert F. Bradshaw. Riverine spawning by sockeye salmon in the Taku River, Alaska and British Columbia. Transactions of the American Fisheries Society, 121(6):701–708, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ensign:1994:SFR

[Ens94] William E. Ensign. Surveys of fisheries resources. Transactions of the American Fisheries Society, 123(4):679–680, July 1994. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elrod:1991:DJL

[EO91] Joseph H. Elrod and Robert O’Gorman. Diet of juvenile lake trout in Southern Lake Ontario in relation to abundance and size of prey fishes, 1979–1987. Transactions of the American Fisheries Society, 120(3):290–302, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eshenroder:1992:DLT

[Esh92] Randy L. Eshenroder. Decline of lake trout in Lake Huron. Transactions of the American Fisheries Society, 121(4):548–554, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ensign:1990:SFL

[ESM90] William E. Ensign, Richard J. Strange, and Stephen E. Moore. Summer food limitation reduces brook and rainbow trout biomass in a Southern Appalachian stream. Transactions of the American Fisheries Society, 119(5):894–901, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eaton:1992:FLI

[ESM^{+}92] John G. Eaton, William A. Swenson, J. Howard McCormick, Timothy D. Simonson, and Kathleen M. Jensen. A field and laboratory investigation of acid effects on largemouth bass, rock bass, black crappie, and yellow perch. Transactions of the American Fisheries Society, 121(5):644–658, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Evans:1990:MTC

[Eva90] David O. Evans. Metabolic thermal compensation by rainbow trout: Effects on standard metabolic rate and potential usable power. Transactions of the American Fisheries Society, 119(4): 585–600, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1997:TEF

[EWE97] Simon C. Ellis, Wade O. Watanabe, and Eileen P. Ellis. Temperature effects on feed utilization and growth of postsettlement stage Nassau grouper. Transactions of the American Fisheries Society, 126(2):309–315, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fynn-Aikins:1998:EDA

[FABH^{+}98] Kofi Fynn-Aikins, Paul R. Bowser, Dale C. Honeyfield, John D. Fitzsimons, and H. George Ketola. Effect of dietary amprolium on tissue thiamin and cayuga syndrome in Atlantic salmon. Transactions of the American Fisheries Society, 127(5):747–757, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fleischer:1997:SRT

[FAC97] Guy W. Fleischer, Ray L. Argyle, and Gary L. Curtis. In situ relations of target strength to fish size for Great Lakes pelagic planktivores. Transactions of the American Fisheries Society, 126(5):786–794, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farris:1990:WPB

[Far90] Jerry L. Farris. Water pollution biology. Transactions of the American Fisheries Society, 119(4):786–790, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Finn:1997:DAP

[FBHB97] James E. Finn, Carl V. Burger, and Leslie Holland-Bartels. Discrimination among populations of sockeye salmon fry with Fourier analysis of otolith banding patterns formed during incubation. Transactions of the American Fisheries Society, 126(4): 559–578, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fiss:1993:SBT

[FC93] F. C. Fiss and R. F. Carline. Survival of brook trout embryos in three episodically acidified streams. Transactions of the American Fisheries Society, 122(2):268–278, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foster:1997:SMG

[FC97] Ann M. Foster and James P. Clugston. Seasonal migration of Gulf sturgeon in the Suwannee River, Florida. Transactions of the American Fisheries Society, 126(2):302–308, March 1997. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fechhelm:1992:MST

[FDGG92] Robert G. Fechhelm, Robert E. Dillinger, Jr., Benny J. Gallaway, and William B. Griffiths. Modeling of in situ temperature and growth relationships for yearling broad whitefish in Prudhoe Bay, Alaska. Transactions of the American Fisheries Society, 121(1): 1–12, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fox:1990:EFD

[FF90] Michael G. Fox and David D. Flowers. Effect of fish density on growth, survival, and food consumption by juvenile walleyes in rearing ponds. Transactions of the American Fisheries Society, 119(1):112–121, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frierson:1992:CEA

[FF92] Edward W. Frierson and Jeffrey W. Foltz. Comparison and estimation of absorptive intestinal surface areas in two species of cichlid fish. Transactions of the American Fisheries Society, 121(4):517–523, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fisher:1996:NOT

[FFCS96] Jeffrey P. Fisher, John D. Fitzsimons, Gerald F. Combs, Jr., and Jan M. Spitsbergen. Naturally occurring thiamine deficiency causing reproductive failure in Finger Lakes Atlantic salmon and Great Lakes lake trout. Transactions of the American Fisheries Society, 125(2):167–178, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fajt:1993:OTR

[FG93] James R. Fajt and John M. Grizzle. Oral toxicity of rotenone for common carp. Transactions of the American Fisheries Society, 122(2):302–304, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fajt:1998:BRC

[FG98] James R. Fajt and John M. Grizzle. Blood respiratory changes in common carp exposed to a lethal concentration of rotenone. Transactions of the American Fisheries Society, 127(3):512–516, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fechhelm:1995:ASG

[FGB^{+}95] Robert G. Fechhelm, William B. Griffiths, James D. Bryan, Benny J. Gallaway, and William J. Wilson. Application of an in situ growth model: Inferred instance of interspecific trophic competition between anadromous fishes of Prudhoe Bay, Alaska. Transactions of the American Fisheries Society, 124(1):55–69, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fechhelm:1996:IIV

[FGMG96] Robert G. Fechhelm, William B. Griffiths, Larry R. Martin, and Benny J. Gallaway. Intra- and interannual variation in the relative condition and proximate body composition of Arctic ciscoes from the Prudhoe Bay Region of Alaska. Transactions of the American Fisheries Society, 125(4):600–612, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fechhelm:1995:IIC

[FGW^{+}95] Robert G. Fechhelm, William B. Griffiths, William J. Wilson, Benny J. Gallaway, and James D. Bryan. Intra- and interseasonal changes in the relative condition and proximate body composition of broad whitefish from the Prudhoe Bay Region of Alaska. Transactions of the American Fisheries Society, 124(4):508–519, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Filbert:1995:VCR

[FH95] Randall B. Filbert and Charles P. Hawkins. Variation in condition of rainbow trout in relation to food, temperature, and individual length in the Green River, Utah. Transactions of the American Fisheries Society, 124(6):824–835, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fisher:1996:EEN

[Fis96a] Mark R. Fisher. Estimating the effect of nonresponse bias on angler surveys. Transactions of the American Fisheries Society, 125(1):118–126, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fisher:1996:SES

[Fis96b] William L. Fisher. Stream ecology: Structure and function of running waters. Transactions of the American Fisheries Society, 125(1):154–158, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fisher:1999:ETS

[Fis99] Mark R. Fisher. Effect of temperature and salinity on size at maturity of female blue crabs. Transactions of the American Fisheries Society, 128(3):499–506, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fischer:1990:PFD

[FK90] Steven A. Fischer and William E. Kelso. Parasite fauna development in juvenile bluegills and largemouth bass. Transactions of the American Fisheries Society, 119(5):877–884, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fletcher:1990:FDF

[Fle90] R. Ian Fletcher. Flow dynamics and fish recovery experiments: Water intake systems. Transactions of the American Fisheries Society, 119(3):393–415, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fletcher:1992:NFR

[Fle92] R. Ian Fletcher. Notes: The failure and rehabilitation of a fishconserving device. Transactions of the American Fisheries Society, 121(5):678–679, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flebbe:1994:RVM

[Fle94a] Patricia A. Flebbe. A regional view of the margin: Salmonid abundance and distribution in the Southern Appalachian mountains of North Carolina and Virginia. Transactions of the American Fisheries Society, 123(4):657–667, July 1994. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fletcher:1994:FFB

[Fle94b] R. Ian Fletcher. Flows and fish behavior: Large double-entry screening systems. Transactions of the American Fisheries Society, 123(6):866–885, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fitzhugh:1997:TEO

[FNAR97] Gary R. Fitzhugh, Stephen W. Nixon, Dean W. Ahrenholz, and James A. Rice. Temperature effects on otolith microstructure and birth month estimation from otolith increment patterns in Atlantic menhaden. Transactions of the American Fisheries Society, 126(4):579–593, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foster:1993:GRH

[FOS93] Neal R. Foster, Daniel V. O’Connor, and Carl B. Schreck. Gamete ripening and hormonal correlates in three strains of lake trout. Transactions of the American Fisheries Society, 122(2): 252–267, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foy:1999:WFC

[FP99] R. J. Foy and A. J. Paul. Winter feeding and changes in somatic energy content of age-0 Pacific herring in Prince William Sound, Alaska. Transactions of the American Fisheries Society, 128(6): 1193–1200, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frank:1990:PRN

[FPD90] Kenneth T. Frank, R. Ian Perry, and Kenneth F. Drinkwater. Predicted response of Northwest Atlantic invertebrate and fish stocks to CO_{2}-induced climate change. Transactions of the American Fisheries Society, 119(2):353–365, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Frost:1997:FWI

[Fro97] Thomas M. Frost. Food Webs: Integration of patterns and dynamics. Transactions of the American Fisheries Society, 126(4): 721–724, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

France:1996:EPJ

[FS96] Robert France and Robert Steedman. Energy provenance for juvenile lake trout in small Canadian Shield lakes as shown by stable isotopes. Transactions of the American Fisheries Society, 125(4):512–518, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fukushima:1998:SHS

[FS98] Michio Fukushima and William W. Smoker. Spawning habitat segregation of sympatric sockeye and pink salmon. Transactions of the American Fisheries Society, 127(2):253–260, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fabrizio:1991:SBS

[FSO91] Mary C. Fabrizio, Ronald J. Sloan, and John F. O’Brien. Striped bass stocks and concentrations of polychlorinated biphenyls. Transactions of the American Fisheries Society, 120(5):541–551, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fabrizio:1996:CTN

[FSSH96] Mary C. Fabrizio, Bruce L. Swanson, Stephen T. Schram, and Michael H. Hoff. Comparison of three nonlinear models to describe long-term tag shedding by lake trout. Transactions of the American Fisheries Society, 125(2):261–273, March 1996. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fournie:1996:PGP

[FSW96] John W. Fournie, J. Kevin Summers, and Stephen B. Weisberg. Prevalence of Gross pathological abnormalities in estuarine fishes. Transactions of the American Fisheries Society, 125(4): 581–590, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Freeberg:1990:EEL

[FTB90] Mark H. Freeberg, William W. Taylor, and Russell W. Brown. Effect of egg and larval survival on year-class strength of lake whitefish in Grand Traverse Bay, Lake Michigan. Transactions of the American Fisheries Society, 119(1):92–100, January 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farag:1999:DEM

[FWB^{+}99] A. M. Farag, D. F. Woodward, W. Brumbaugh, J. N. Goldstein, E. MacConnell, C. Hogstrand, and F. T. Barrows. Dietary effects of metals-contaminated invertebrates from the coeur d’Alene River, Idaho, on cutthroat trout. Transactions of the American Fisheries Society, 128(4):578–592, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fries:1996:NOA

[FWJ96] Loraine T. Fries, D. Jody Williams, and S. Ken Johnson. Notes: Occurrence of Anguillicola crassus, an exotic parasitic swim bladder nematode of eels, in the Southeastern United States. Transactions of the American Fisheries Society, 125(5):794–797, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farrell:1996:NED

[FWLC96] John M. Farrell, R. G. Werner, S. R. Lapan, and K. A. Claypoole. Notes: Egg distribution and spawning habitat of northern pike and muskellunge in a St. Lawrence River marsh, New York. Transactions of the American Fisheries Society, 125(1):127–131, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Giles:1993:REE

[GA93] M. A. Giles and E. M. Attas. Rare Earth elements as internal batch marks for rainbow trout: Retention, distribution, and effects on growth of injected dysprosium, europium, and samarium. Transactions of the American Fisheries Society, 122(2):289–297, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Galat:1992:RLE

[Gal92] David L. Galat. Reservoir limnology: Ecological perspectives. Transactions of the American Fisheries Society, 121(5):696–698, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garman:1992:FPS

[Gar92] Greg C. Garman. Fate and potential significance of postspawning anadromous fish carcasses in an Atlantic coastal river. Transactions of the American Fisheries Society, 121(3):390–394, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goyke:1993:SMS

[GB93] Andrew P. Goyke and Stephen B. Brandt. Spatial models of salmonine growth rates in Lake Ontario. Transactions of the American Fisheries Society, 122(5):870–883, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gleason:1996:SSM

[GB96] Timothy R. Gleason and David A. Bengtson. Size-selective mortality of inland silversides: Evidence from otolith microstructure. Transactions of the American Fisheries Society, 125(6):

860–873, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grizzle:1999:DRS

[GC99] John M. Grizzle and Keith A. Cummins. Drinking rates of stressed one-month-old striped bass: Effects of calcium and low concentrations of sodium chloride. Transactions of the American Fisheries Society, 128(3):528–531, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gelsleichter:1997:UCF

[GCM^{+}97] James Gelsleichter, Enric Cortés, Charles A. Manire, Robert E. Hueter, and John A. Musick. Use of calcein as a fluorescent marker for elasmobranch vertebral cartilage. Transactions of the American Fisheries Society, 126(5):862–865, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gallaway:1999:DEH

[GCMR99] Benny J. Gallaway, John G. Cole, Robert Meyer, and Pasquale Roscigno. Delineation of essential habitat for juvenile red snapper in the Northwestern Gulf of Mexico. Transactions of the American Fisheries Society, 128(4):713–726, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Guest:1990:EGS

[GDT^{+}90] W. Clell Guest, Ray W. Drenner, Stephen T. Threlkeld, F. Douglas Martin, and J. Durward Smith. Effects of gizzard shad and threadfin shad on zooplankton and young-of-year white crappie production. Transactions of the American Fisheries Society, 119(3):529–536, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goshorn:1991:DSG

[GE91a] D. M. Goshorn and C. E. Epifanio. Development, survival, and growth of larval weakfish at different prey abundances. Transactions of the American Fisheries Society, 120(6):693–700, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goshorn:1991:DLW

[GE91b] D. M. Goshorn and C. E. Epifanio. Diet of larval weakfish and prey abundance in Delaware Bay. Transactions of the American Fisheries Society, 120(6):684–692, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glova:1995:BIB

[GFD95] Gordon J. Glova and Michael S. Field-Dodgson. Behavioral interaction between Chinook salmon and brown trout juveniles in a simulated stream. Transactions of the American Fisheries Society, 124(2):194–206, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffiths:1992:GCA

[GGGD92] William B. Griffiths, Benny J. Gallaway, William J. Gazey, and Robert E. Dillinger, Jr. Growth and condition of Arctic cisco and broad whitefish as indicators of causeway-induced effects in the Prudhoe Bay Region, Alaska. Transactions of the American Fisheries Society, 121(5):557–577, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gelwick:1995:IFL

[GGM95] Frances P. Gelwick, Eugene R. Gilliland, and William J. Matthews. Introgression of the Florida largemouth bass genome into stream populations of northern largemouth bass in Oklahoma. Transactions of the American Fisheries Society, 124(4): 550–562, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Giannico:1998:EFF

[GH98] Guillermo R. Giannico and Michael C. Healey. Effects of flow and food on winter movements of juvenile Coho salmon. Transactions of the American Fisheries Society, 127(4):645–651, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gadomski:1992:PNS

[GHG92] Dena M. Gadomski and Judy A. Hall-Griswold. Predation by northern squawfish on live and dead juvenile Chinook salmon. Transactions of the American Fisheries Society, 121(5):680–685, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gryska:1998:RAL

[GHG98] Andrew D. Gryska, Wayne A. Hubert, and Kenneth G. Gerow. Relative abundance and lengths of Kendall warm springs dace captured from different habitats in a specially designed trap. Transactions of the American Fisheries Society, 127(2):309–315, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gallagher:1991:VLF

[GHR91] M. L. Gallagher, M. L. Harrell, and R. A. Rulifson. Variation in lipid and fatty acid contents of Atlantic croakers, striped mullet, and summer flounder. Transactions of the American Fisheries Society, 120(5):614–619, September 1991. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grost:1991:DBT

[GHW91] Richard T. Grost, Wayne A. Hubert, and Thomas A. Wesche. Description of brown trout redds in a mountain stream. Transactions of the American Fisheries Society, 120(5):582–588, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grimes:1996:ISS

[GI96] Churchill B. Grimes and J. Jeffery Isely. Influence of size-selective mortality on growth of Gulf menhaden and king mackerel larvae. Transactions of the American Fisheries Society, 125(5):741–752, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gobalet:1995:PNA

[GJ95] Kenneth W. Gobalet and Terry L. Jones. Prehistoric native American fisheries of the Central California Coast. Transactions of the American Fisheries Society, 124(6):813–823, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gross:1994:NSD

[GKF94] Mark L. Gross, Anne R. Kapuscinski, and Anthony J. Faras. Nest-specific DNA fingerprints of smallmouth bass in Lake Opeongo, Ontario. Transactions of the American Fisheries Society, 123(4):449–459, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grewe:1994:HON

[GKM^{+}94] P. M. Grewe, C. C. Krueger, J. E. Marsden, C. F. Aquadro, and B. May. Hatchery origins of naturally produced lake trout fry captured in Lake Ontario: Temporal and spatial variability based on allozyme and mitochondrial DNA data. Transactions of the American Fisheries Society, 123(3):309–320, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gregory:1998:TRP

[GL98] Robert S. Gregory and Colin D. Levings. Turbidity reduces predation on migrating juvenile Pacific salmon. Transactions of the American Fisheries Society, 127(2):275–285, March 1998. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gorski:1999:TMD

[GLHH99] Patrick R. Gorski, Richard C. Lathrop, Susan D. Hill, and Russell T. Herrin. Temporal mercury dynamics and diet composition in the mimic shiner. Transactions of the American Fisheries Society, 128(4):701–712, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodgame:1993:EGS

[GM93] L. S. Goodgame and L. E. Miranda. Early growth and survival of age-0 largemouth bass in relation to parental size and swimup time. Transactions of the American Fisheries Society, 122 (1):131–138, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grizzle:1994:ARC

[GM94] John M. Grizzle and Alfred C. Mauldin II. Age-related changes in survival of larval and juvenile striped bass in different concentrations of calcium and sodium. Transactions of the American Fisheries Society, 123(6):1002–1005, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gulseth:1999:GBH

[GN99] Odd A. Gulseth and Kjell J. Nilssen. Growth benefit from habitat change by juvenile High-Arctic char. Transactions of the American Fisheries Society, 128(4):593–602, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grahl-Nielsen:1992:DSB

[GNM92] Otto Grahl-Nielsen and Olav Mjaavatten. Discrimination of striped bass stocks: a new method based on chemometry of the fatty acid profile in heart tissue. Transactions of the American Fisheries Society, 121(3):307–314, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Groff:1997:ICR

[GNZO97] Joseph M. Groff, Diane K. Naydan, Joseph G. Zinkl, and Bennie I. Osburn. Immunological cross-reactivity of type I–III intermediate filaments in the common carp: In situ localization with use of heterologous antibodies. Transactions of the American Fisheries Society, 126(6):948–960, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:1995:MSA

[Goo95] C. Phillip Goodyear. Mean size at age: an evaluation of sampling strategies with simulated red grouper data. Transactions of the American Fisheries Society, 124(5):746–755, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gracia:1991:SSR

[Gra91] Adolfo Gracia. Spawning stock–recruitment relationships of white shrimp in the Southwestern Gulf of Mexico. Transactions of the American Fisheries Society, 120(4):519–527, July 1991. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graham:1993:SDF

[Gra93] John H. Graham. Species diversity of fishes in naturally acidic lakes in New Jersey. Transactions of the American Fisheries Society, 122(6):1043–1057, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gray:1999:CME

[Gra99] Gerry Gray. Covariances in multiplicative estimates. Transactions of the American Fisheries Society, 128(3):475–482, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gresswell:1999:FAE

[Gre99] Robert E. Gresswell. Fire and aquatic ecosystems in forested biomes of North America. Transactions of the American Fisheries Society, 128(2):193–221, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffin:1991:CCA

[Gri91] B. R. Griffin. Characteristics of a chondroitin AC lyase produced by cytophaga columnaris. Transactions of the American Fisheries Society, 120(3):391–395, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grimes:1993:NVV

[Gri93] David V. Grimes. Notes: Vitellolipid and vitelloprotein profiles of environmentally stressed and nonstressed populations of striped bass. Transactions of the American Fisheries Society, 122(4): 636–641, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gold:1993:TSN

[GRKM93] John R. Gold, Linda R. Richardson, Timothy L. King, and Gary C. Matlock. Temporal stability of nuclear gene (allozyme) and mitochondrial DNA genotypes among red drums from the Gulf of Mexico. Transactions of the American Fisheries Society, 122(5):659–668, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grossman:1999:BR

[Gro99] Gary D. Grossman. Book reviews. Transactions of the American Fisheries Society, 128(5):974–975, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garvey:1998:CBL

[GS98a] James E. Garvey and Roy A. Stein. Competition between larval fishes in reservoirs: The role of relative timing of appearance. Transactions of the American Fisheries Society, 127(6): 1021–1039, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garvey:1998:LBG

[GS98b] James E. Garvey and Roy A. Stein. Linking bluegill and gizzard shad prey assemblages to growth of age-0 largemouth bass in reservoirs. Transactions of the American Fisheries Society, 127 (1):70–83, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gagen:1993:MBT

[GSC93] Charles J. Gagen, William E. Sharpe, and Robert F. Carline. Mortality of brook trout, mottled sculpins, and slimy sculpins during acidic episodes. Transactions of the American Fisheries Society, 122(4):616–628, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gold:1997:PSR

[GSR97] J. R. Gold, F. Sun, and L. R. Richardson. Population structure of red snapper from the Gulf of Mexico as inferred from analysis of mitochondrial DNA. Transactions of the American Fisheries Society, 126(3):386–396, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graham:1990:RBW

[GSS90] Joseph J. Graham, David K. Stevenson, and Keith M. Sherman. Relation between winter temperature and survival of larval Atlantic herring along the Maine Coast. Transactions of the American Fisheries Society, 119(4):730–740, July 1990. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grecay:1996:SPC

[GT96] Paul A. Grecay and Timothy E. Targett. Spatial patterns in condition and feeding of juvenile weakfish in Delaware Bay. Transactions of the American Fisheries Society, 125(5):803–808, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garcia-Vazquez:1995:EPN

[GVMP^{+}95] E. Garcia-Vazquez, P. Moran, A. M. Pendas, J. I. Izquierdo, and A. R. Linde. Effect of parental numbers on chromosome patterns found in artificially produced Atlantic salmon stocks. Transactions of the American Fisheries Society, 124(6):939–942, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goldstein:1999:MAC

[GWF99] J. N. Goldstein, D. F. Woodward, and A. M. Farag. Movements of adult Chinook salmon during spawning migration in a metals-contaminated system, coeur d’Alene River, Idaho. Transactions of the American Fisheries Society, 128(1):121–129, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Guy:1994:BWC

[GWJ94] Christopher S. Guy, David W. Willis, and Jeffrey J. Jackson. Biotelemetry of white crappies in a South Dakota glacial lake. Transactions of the American Fisheries Society, 123(1):63–70, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gross:1998:RAS

[GWL98] Howard P. Gross, Wayne A. Wurtsbaugh, and Chris Luecke. The role of anadromous sockeye salmon in the nutrient loading and productivity of Redfish Lake, Idaho. Transactions of the American Fisheries Society, 127(1):1–18, January 1998. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gustaveson:1991:PRL

[GWW91] A. Wayne Gustaveson, Richard S. Wydoski, and Gary A. Wedemeyer. Physiological response of largemouth bass to angling stress. Transactions of the American Fisheries Society, 120(5): 629–636, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayward:1996:TDM

[HA96] R. S. Hayward and E. Arnold. Temperature dependence of maximum daily consumption in white crappie: Implications for fisheries management. Transactions of the American Fisheries Society, 125(1):132–138, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Halbrendt:1994:EOC

[HAHD94] Catherine K. Halbrendt, Rhonda Aull-Hyde, and Dee A. Dunkers. Expert opinions on critical production factors for sustained growth of the hybrid striped bass industry. Transactions of the American Fisheries Society, 123(1):94–100, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hall:1991:AVA

[Hal91] D. L. Hall. Age validation and aging methods for stunted brook trout. Transactions of the American Fisheries Society, 120(5):

644–649, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hambright:1991:EAP

[Ham91] K. David Hambright. Experimental analysis of prey selection by largemouth bass: Role of predator mouth width and prey body depth. Transactions of the American Fisheries Society, 120(4): 500–508, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:1998:DWB

[Har98] Kyle J. Hartman. Diets of white bass in the Ohio waters of Lake Erie during June–October 1988. Transactions of the American Fisheries Society, 127(2):323–328, March 1998. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hales:1992:VOA

[HB92] L. Stanton Hales, Jr. and Mark C. Belk. Validation of otolith annuli of bluegills in a southeastern thermal reservoir. Transactions of the American Fisheries Society, 121(6):823–830, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:1993:SSB

[HB93] Kyle J. Hartman and Stephen B. Brandt. Systematic sources of bias in a bioenergetics model: Examples for age-0 striped bass. Transactions of the American Fisheries Society, 122(5):912–926, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayward:1994:GER

[HB94] Robert S. Hayward and Michael E. Bushmann. Gastric evacuation rates for juvenile largemouth bass. Transactions of the American Fisheries Society, 123(1):88–93, January 1994. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:1995:EED

[HB95a] Kyle J. Hartman and Stephen B. Brandt. Estimating energy density of fish. Transactions of the American Fisheries Society, 124(3):347–355, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:1995:TRP

[HB95b] Kyle J. Hartman and Stephen B. Brandt. Trophic resource partitioning, diets, and growth of sympatric estuarine predators. Transactions of the American Fisheries Society, 124(4):520–537, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herdendorf:1995:GSW

[HB95c] Charles E. Herdendorf and Tim M. Berra. A Greenland shark from the wreck of the SS Central America at 2,200 meters. Transactions of the American Fisheries Society, 124(6):950–953, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoff:1999:SSM

[HB99] Michael H. Hoff and Charles R. Bronte. Structure and stability of the midsummer fish communities in Chequamegon Bay, Lake Superior, 1973–1996. Transactions of the American Fisheries Society, 128(2):362–373, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1993:ABM

[HBB^{+}93] Michael J. Hansen, Daniel Boisclair, Stephen B. Brandt, Steven W. Hewett, James F. Kitchell, Martyn C. Lucas, and John J. Ney. Applications of bioenergetics models to fish ecology and management: Where do we go from here? Transactions of the American Fisheries Society, 122(5):1019–1030, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hetrick:1998:ECR

[HBB^{+}98] N. J. Hetrick, M. A. Brusven, T. C. Bjornn, R. M. Keith, and W. R. Meehan. Effects of canopy removal on invertebrates and diet of juvenile Coho salmon in a small stream in Southeast Alaska. Transactions of the American Fisheries Society, 127(6): 876–888, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hetrick:1998:CSI

[HBMB98] N. J. Hetrick, M. A. Brusven, W. R. Meehan, and T. C. Bjornn. Changes in solar input, water temperature, periphyton accumulation, and allochthonous input and storage after canopy removal along two small salmon streams in Southeast Alaska. Transactions of the American Fisheries Society, 127(6):859–875, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heggenes:1990:CTM

[HBS90] Jan Heggenes, Åge Brabrand, and Sveinjakob Saltveit. Comparison of three methods for studies of stream habitat use by young brown trout and Atlantic salmon. Transactions of the American Fisheries Society, 119(1):101–111, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hinch:1991:IDN

[HC91] Scott G. Hinch and Nicholas C. Collins. Importance of diurnal and nocturnal nest defense in the energy budget of male smallmouth bass: Insights from direct video observations. Transactions of the American Fisheries Society, 120(5):657–663, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hopkins:1992:PEC

[HC92] Todd E. Hopkins and Joseph J. Cech, Jr. Physiological effects of capturing striped bass in gill nets and fyke traps. Transactions of the American Fisheries Society, 121(6):819–822, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Headrick:1993:RSH

[HC93] Michael R. Headrick and Robert F. Carline. Restricted summer habitat and growth of northern pike in two Southern Ohio impoundments. Transactions of the American Fisheries Society, 122(2):228–236, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanson:1996:SUE

[HC96a] J. Mark Hanson and Simon C. Courtenay. Seasonal use of estuaries by winter flounder in the Southern Gulf of St. Lawrence. Transactions of the American Fisheries Society, 125(5):705–718, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubert:1996:EGA

[HC96b] Wayne A. Hubert and Charles B. Chamberlain. Environmental gradients affect rainbow trout populations among lakes and reservoirs in Wyoming. Transactions of the American Fisheries Society, 125(6):925–932, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Helser:1993:SSD

[HCG93] Thomas E. Helser, Richard E. Condrey, and James P. Geaghan. Spotted seatrout distribution in four coastal Louisiana estuaries. Transactions of the American Fisheries Society, 122(1):99–111, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haeseker:1996:SDC

[HCH96] Steven L. Haeseker, John T. Carmichael, and Joseph E. Hightower. Summer distribution and condition of striped bass within Albemarle Sound, North Carolina. Transactions of the American Fisheries Society, 125(5):690–704, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hirst:1994:APD

[HD94] Shawn C. Hirst and Dennis R. Devries. Assessing the potential for direct feeding interactions among larval black bass and larval shad in two Southeastern reservoirs. Transactions of the American Fisheries Society, 123(2):173–181, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoxmeier:1997:HUD

[HD97] R. H. John Hoxmeier and Dennis R. Devries. Habitat use, diet, and population structure of adult and juvenile paddlefish in the Lower Alabama River. Transactions of the American Fisheries Society, 126(2):288–301, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hinch:1996:UET

[HDL^{+}96] Scott G. Hinch, Ronald E. Diewert, Timothy J. Lissimore, Angela M. J. Prince, Michael C. Healey, and Michael A. Henderson. Use of electromyogram telemetry to assess difficult passage areas for river-migrating adult sockeye salmon. Transactions of the American Fisheries Society, 125(2):253–260, March 1996. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hegge:1991:ASS

[HDS91] Ola Hegge, Børre K. Dervo, and Jostein Skurdal. Age and size at sexual maturity of heavily exploited Arctic char and brown trout in Lake Atnsjø, Southeastern Norway. Transactions of the American Fisheries Society, 120(2):141–149, March 1991. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Healey:1990:ICC

[Hea90] M. C. Healey. Implications of climate change for fisheries management policy. Transactions of the American Fisheries Society, 119(2):366–373, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Healey:1991:DFR

[Hea91] M. C. Healey. Diets and feeding rates of juvenile pink, chum, and sockeye salmon in Hecate Strait, British Columbia. Transactions of the American Fisheries Society, 120(3):303–318, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hengeveld:1990:GCC

[Hen90] Henry G. Hengeveld. Global climate change: Implications for air temperature and water supply in Canada. Transactions of the American Fisheries Society, 119(2):176–182, March 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1996:CDS

[HES^{+}96] Michael J. Hansen, Mark P. Ebener, Richard G. Schorfhaar, Stephen T. Schram, Donald R. Schreiner, James H. Selgeby, and William W. Taylor. Causes of declining survival of lake trout stocked in U.S. waters of Lake Superior in 1963–1986. Transactions of the American Fisheries Society, 125(6):831–843, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heggberget:1994:CAC

[HGK^{+}94] Tor G. Heggberget, Per Grotnes, Anders Klemetsen, Roger Richardsen, and Ola Ugedal. Culture of Arctic char-possibilities in Norway. Transactions of the American Fisheries Society, 123 (3):435–444, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayward:1993:LER

[HH93] Robert S. Hayward and John F. Hiebert. Low-effort regression estimation of average daily food weight in fishes: Application to larval gizzard shad. Transactions of the American Fisheries Society, 122(5):834–844, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henry:1999:ATA

[HIG^{+}99] Theodore B. Henry, Elise R. Irwin, John M. Grizzle, Mark L. Wildhaber, and William G. Brumbaugh. Acute toxicity of an acid mine drainage mixing zone to juvenile bluegill and largemouth bass. Transactions of the American Fisheries Society, 128(5): 919–928, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartwell:1998:TST

[HJDI98] S. I. Hartwell, D. M. Jordahl, C. E. O. Dawson, and A. S. Ives. Toxicity of scrap tire leachates in estuarine salinities: Are tires acceptable for artificial reefs? Transactions of the American Fisheries Society, 127(5):796–806, September 1998. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoffman:1999:ITW

[HJLM99] Alex D. Hoffman, Correne T. Jenson, Gregory J. Lien, and James M. McKim. Individual tissue weight to total body weight relationships and total, polar, and nonpolar lipids in tissues of hatchery lake trout. Transactions of the American Fisheries Society, 128(1):178–181, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heppell:1999:ELI

[HJWS99] Scott A. Heppell, Leslie F. Jackson, Gregory M. Weber, and Craig V. Sullivan. Enzyme-Linked Immunosorbent Assay (ELISA) of vitellogenin in temperate basses (genus Morone): Plasma and in vitro analyses. Transactions of the American Fisheries Society, 128(3):532–541, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

He:1990:DIE

[HK90] Xi He and James F. Kitchell. Direct and indirect effects of predation on a fish community: a whole-lake experiment. Transactions of the American Fisheries Society, 119(5):825–835, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hewett:1993:RBG

[HK93] Steven W. Hewett and Clifford E. Kraft. The relationship between growth and consumption: Comparisons across fish populations. Transactions of the American Fisheries Society, 122(5): 814–821, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

He:1993:FWS

[HKC^{+}93] Xi He, James F. Kitchell, Stephen R. Carpenter, James R. Hodgson, Daniel E. Schindler, and Kathryn L. Cottingham. Food web structure and long-term phosphorus recycling: a simulation model evaluation. Transactions of the American Fisheries Society, 122(5):773–783, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hughes:1992:DPD

[HLH92] Steven G. Hughes, Carol A. Lemm, and Roger L. Herman. Development of a practical diet for juvenile striped bass. Transactions of the American Fisheries Society, 121(6):802–809, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoff:1997:NOM

[HLM97] Gerald R. Hoff, Daniel J. Logan, and Douglas F. Markle. Notes: Otolith morphology and increment validation in young Lost River and shortnose suckers. Transactions of the American Fisheries Society, 126(3):488–494, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hill:1990:PEG

[HM90] David K. Hill and John J. Magnuson. Potential effects of global climate warming on the growth and prey consumption of Great Lakes fish. Transactions of the American Fisheries Society, 119 (2):265–275, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:1992:EPP

[HM92] Kyle J. Hartman and F. Joseph Margraf. Effects of prey and predator abundances on prey consumption and growth of walleyes in Western Lake Erie. Transactions of the American Fisheries Society, 121(2):245–260, March 1992. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayward:1991:LCF

[HMPV91] Robert S. Hayward, F. Joseph Margraf, Jr., Donna L. Parrish, and Bruce Vondracek. Low-cost field estimation of yellow perch daily ration. Transactions of the American Fisheries Society, 120 (5):589–604, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hedrick:1991:TCL

[HMR^{+}91] Ronald P. Hedrick, Terry S. McDowell, Renee Rosemark, Diane Aronstein, and Catharine N. Lannan. Two cell lines from white sturgeon. Transactions of the American Fisheries Society, 120(4): 528–534, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henderson:1990:YYP

[HN90] Bryan A. Henderson and Stephen J. Nepszy. Yield of yellow perch in Lakes Erie and St. Clair: Community associations. Transactions of the American Fisheries Society, 119(4):741–756, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harvey:1996:ESD

[HN96] Bret C. Harvey and Rodney J. Nakamoto. Effects of steelhead density on growth of Coho salmon in a small coastal California stream. Transactions of the American Fisheries Society, 125(2): 237–243, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayward:1997:UCG

[HNW97] Robert S. Hayward, Douglas B. Noltie, and Ning Wang. Use of compensatory growth to double hybrid sunfish growth rates. Transactions of the American Fisheries Society, 126(2):316–322, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanten:1998:RBC

[HNW^{+}98] Robert P. Hanten, Jr., Robert M. Neumann, Scott M. Ward, Robert J. Carley, Christopher R. Perkins, and Robert Pirrie.

Relationships between concentrations of mercury in largemouth bass and physical and chemical characteristics of Connecticut lakes. Transactions of the American Fisheries Society, 127(5): 807–818, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hochachka:1990:SSC

[Hoc90] P. W. Hochachka. Scope for survival: a conceptual “mirror” to fry’s scope for activity. Transactions of the American Fisheries Society, 119(4):622–628, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holmes:1990:SLE

[Hol90] John A. Holmes. Sea lamprey as an early responder to climate change in the Great Lakes Basin. Transactions of the American Fisheries Society, 119(2):292–300, March 1990. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haro:1998:EWA

[HONCS98] Alex Haro, Mufeed Odeh, John Noreika, and Theodore Castro-Santos. Effect of water acceleration on downstream migratory behavior and passage of Atlantic salmon smolts and juvenile American shad at surface bypasses. Transactions of the American Fisheries Society, 127(1):118–127, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houston:1997:RCH

[Hou97] Arthur H. Houston. Review: Are the classical hematological variables acceptable indicators of fish health? Transactions of the American Fisheries Society, 126(6):879–894, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haight:1993:FED

[HPH93] Wayne R. Haight, James D. Parrish, and Therese A. Hayes. Feeding ecology of deepwater lutjanid snappers at Penguin Bank, Hawaii. Transactions of the American Fisheries Society, 122(3): 328–347, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haas:1995:AVW

[HR95] R. E. Haas and C. W. Recksiek. Age verification of winter flounder in Narragansett Bay. Transactions of the American Fisheries Society, 124(1):103–111, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heard:1997:EAC

[HSCK97] Robin M. Heard, William E. Sharpe, Robert F. Carline, and William G. Kimmel. Episodic acidification and changes in fish diversity in Pennsylvania headwater streams. Transactions of the American Fisheries Society, 126(6):977–984, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Habicht:1998:CWT

[HSES98] Christopher Habicht, Samuel Sharr, David Evans, and James E. Seeb. Coded wire tag placement affects homing ability of pink salmon. Transactions of the American Fisheries Society, 127(4): 652–657, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1991:DSF

[HSH91] Michael J. Hansen, Michael D. Staggs, and Michael H. Hoff. Derivation of safety factors for setting harvest quotas on adult walleyes from past estimates of abundance. Transactions of the American Fisheries Society, 120(5):620–628, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hodgson:1998:HTT

[HSH98] James R. Hodgson, Daniel E. Schindler, and Xi He. Homing tendency of three piscivorous fishes in a North temperate lake. Transactions of the American Fisheries Society, 127(6):1078– 1081, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayes:1991:NMR

[HT91] Daniel B. Hayes and William W. Taylor. Notes: a multipleregression technique for analyzing stomach contents in fish. Transactions of the American Fisheries Society, 120(3):396–400, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayes:1994:CCS

[HT94] Daniel B. Hayes and William W. Taylor. Changes in the composition of somatic and gonadal tissues of yellow perch following white sucker removal. Transactions of the American Fisheries Society, 123(2):204–216, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayes:1992:RYP

[HTS92] Daniel B. Hayes, William W. Taylor, and James C. Schneider. Response of yellow perch and the benthic invertebrate community to a reduction in the abundance of white suckers. Transactions of the American Fisheries Society, 121(1):36–53, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1999:ELI

[Hub99] Clark Hubbs. Effect of light intensity on brood production of livebearers Gambusia spp. Transactions of the American Fisheries Society, 128(4):747–750, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hughes:1998:RGD

[Hug98a] Nicholas F. Hughes. Reduction in growth due to electrofishing and tagging may change interannual movement behavior of stream salmonids: Evidence from Arctic grayling in an interior Alaskan stream. Transactions of the American Fisheries Society, 127(6):1072–1077, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hughes:1998:UWS

[Hug98b] Nicholas F. Hughes. Use of whole-stream patterns of age segregation to infer the interannual movements of stream salmonids: a demonstration with Arctic grayling in an interior Alaskan stream. Transactions of the American Fisheries Society, 127(6):1067– 1071, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hamilton:1990:CBM

[HW90] Steven J. Hamilton and Raymond H. Wiedmeyer. Concentrations of boron, molybdenum, and selenium in Chinook salmon. Transactions of the American Fisheries Society, 119(3):500–510, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

He:1993:EMG

[HW93] Enqiang He and Wayne A. Wurtsbaugh. An empirical model of gastric evacuation rates for fish and an analysis of digestion in piscivorous brown trout. Transactions of the American Fisheries Society, 122(5):717–730, September 1993. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayse:1996:ESD

[HW96] John W. Hayse and Thomas E. Wissing. Effects of stem density of artificial vegetation on abundance and growth of age-0 bluegills and predation by largemouth bass. Transactions of the American Fisheries Society, 125(3):422–433, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heimbuch:1997:EFA

[HWW^{+}97] Douglas G. Heimbuch, Harold T. Wilson, Stephen B. Weisberg, Jon H. Voslash;Lstad, and Paul F. Kazyak. Estimating fish abundance in stream surveys by using double-pass removal sampling. Transactions of the American Fisheries Society, 126(5):795–803, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holmes:1997:LSE

[HY97] John A. Holmes and John H. Youson. Laboratory study of the effects of spring warming and larval density on the metamorphosis of sea lampreys. Transactions of the American Fisheries Society, 126(4):647–657, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holmes:1998:EOT

[HY98] John A. Holmes and John H. Youson. Extreme and optimal temperatures for metamorphosis in sea lampreys. Transactions of the American Fisheries Society, 127(2):206–211, March 1998. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isaak:1996:MNS

[IB96] Daniel J. Isaak and Theodore C. Bjornn. Movement of northern squawfish in the tailrace of a Lower Snake River Dam relative to the migration of juvenile anadromous salmonids. Transactions of the American Fisheries Society, 125(5):780–793, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isely:1998:TRG

[IE98] J. Jeffery Isely and Arnold G. Eversole. Tag retention, growth, and survival of red swamp crayfish Procambarus clarkii marked with coded wire tags. Transactions of the American Fisheries Society, 127(4):658–660, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ihssen:1990:PMG

[IMMP90] P. E. Ihssen, L. R. McKay, I. McMillan, and R. B. Phillips. Ploidy manipulation and gynogenesis in fishes: Cytogenetic and fisheries applications. Transactions of the American Fisheries Society, 119(4):698–717, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Iwama:1996:CMA

[Iwa96] George K. Iwama. Cellular and molecular approaches to fish ionic regulation. Transactions of the American Fisheries Society, 125 (6):988–990, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jagielo:1995:ASL

[Jag95] Thomas H. Jagielo. Abundance and survival of lingcod at Cape Flattery, Washington. Transactions of the American Fisheries Society, 124(2):170–183, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jagielo:1999:MMS

[Jag99] Thomas H. Jagielo. Movement, mortality, and size selectivity of sport- and trawl-caught lingcod off Washington. Transactions of the American Fisheries Society, 128(1):31–48, January 1999. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jennings:1996:EHP

[JCP96] Martin J. Jennings, Julie E. Claussen, and David P. Philipp. Evidence for heritable preferences for spawning habitat between two walleye populations. Transactions of the American Fisheries Society, 125(6):978–982, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jordan:1997:SFV

[JCT97] Frank Jordan, Sean Coyne, and Joel C. Trexler. Sampling fishes in vegetated habitats: Effects of habitat structure on sampling characteristics of the 1-m^{2} throw trap. Transactions of the American Fisheries Society, 126(6):1012–1020, November 1997. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1997:FHA

[JDW^{+}97] James H. Johnson, David S. Dropkin, Barbara E. Warkentine, Joseph W. Rachlin, and William D. Andrews. Food habits of Atlantic sturgeon off the Central New Jersey Coast. Transactions of the American Fisheries Society, 126(1):166–170, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1990:SDW

[JE90] Timothy B. Johnson and David O. Evans. Size-dependent winter mortality of young-of-the-year white perch: Climate warming and invasion of the Laurentian Great Lakes. Transactions of the American Fisheries Society, 119(2):301–313, March 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1996:NTC

[JE96] Timothy B. Johnson and David O. Evans. Notes: Temperature constraints on overwinter survival of age-0 white perch. Transactions of the American Fisheries Society, 125(3):466–471, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1994:GAD

[Jen94] Arne J. Jensen. Growth and age distribution of a river-dwelling and a lake-dwelling population of anadromous Arctic char at the same latitude in Norway. Transactions of the American Fisheries Society, 123(3):370–376, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jessop:1993:FAA

[Jes93] B. M. Jessop. Fecundity of anadromous alewives and blueback herring in New Brunswick and Nova Scotia. Transactions of the American Fisheries Society, 122(1):85–98, January 1993. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1999:ESS

[JH99] James E. Johnson and Robert T. Hines. Effect of suspended sediment on vulnerability of young razorback suckers to predation.

Transactions of the American Fisheries Society, 128(4):648–655, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1996:DPT

[JHT96] James E. Johnson, W. Dean Heckathorn, Jr., and Andrew L. Thompson. Dispersal and persistence of 2,3,7,8tetrachlorodibenzo-p-dioxin (TCDD) in a contaminated aquatic ecosystem, Bayou Meto, Arkansas. Transactions of the American Fisheries Society, 125(3):450–457, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jonas:1996:ASC

[JKM96] Jory L. Jonas, Clifford E. Kraft, and Terry L. Margenau. Assessment of seasonal changes in energy density and condition in age-0 and age-1 muskellunge. Transactions of the American Fisheries Society, 125(2):203–210, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1993:SHD

[JKO93] Michael L. Jones, Joseph F. Koonce, and Robert O’Gorman. Sustainability of hatchery-dependent salmonine fisheries in Lake Ontario: The conflict between predator demand and prey supply. Transactions of the American Fisheries Society, 122(5): 1002–1018, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jagielo:1996:GVP

[JLV96] Thomas H. Jagielo, Larry L. Leclair, and Beth A. Vorderstrasse. Genetic variation and population structure of lingcod. Transactions of the American Fisheries Society, 125(3):372–386, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnston:1996:MFR

[JM96] Thomas A. Johnston and J. A. Mathias. Maintenance food requirements and response to short-term food deprivation of walleye larvae. Transactions of the American Fisheries Society, 125 (2):211–223, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1998:EIP

[JMBK98] Timothy B. Johnson, Doran M. Mason, Charles R. Bronte, and James F. Kitchell. Estimation of invertebrate production from patterns of fish predation in Western Lake Superior. Transactions of the American Fisheries Society, 127(3):496–506, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Juanes:1993:PAB

[JMMC93] Francis Juanes, Rick E. Marks, Kim A. McKown, and David O. Conover. Predation by age-0 bluefish on age-0 anadromous fishes in the Hudson River Estuary. Transactions of the American Fisheries Society, 122(3):348–356, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1998:CEO

[JMS^{+}98] Lyndal L. Johnson, David Misitano, Sean Y. Sol, Gregory M. Nelson, Barbara French, Gina M. Ylitalo, and Tom Hom. Contaminant effects on ovarian development and spawning success in rock sole from Puget Sound, Washington. Transactions of the American Fisheries Society, 127(3):375–392, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jakober:1998:RSI

[JMTC98] Michael J. Jakober, Thomas E. McMahon, Russell F. Thurow, and Christopher G. Clancy. Role of stream ice on fall and winter movements and habitat use by bull trout and cutthroat trout in Montana headwater streams. Transactions of the American Fisheries Society, 127(2):223–235, March 1998. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1997:DGR

[JN97] Brian L. Johnson and Douglas B. Noltie. Demography, growth, and reproductive allocation in stream-spawning longnose gar. Transactions of the American Fisheries Society, 126(3):438–466, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johannsson:1991:RPF

[JO91] Ora E. Johannsson and Robert O’Gorman. Roles of predation, food, and temperature in structuring the epilimnetic zooplankton populations in Lake Ontario, 1981–1986. Transactions of the American Fisheries Society, 120(2):193–208, March 1991. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1993:CCB

[Joh93] Sherri L. Johnson. Cover choice by bluegills: Orientation of underwater structure and light intensity. Transactions of the American Fisheries Society, 122(1):148–154, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1996:MCI

[JP96] Jeremie M. Johnson and David M. Post. Morphological constraints on intracohort cannibalism in age-0 largemouth bass. Transactions of the American Fisheries Society, 125(5):809–812, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jager:1997:CTI

[JP97] Henriette I. Jager and Edmund J. Pert. Comment: Testing the independence of microhabitat preferences and flow (part 2). Transactions of the American Fisheries Society, 126(3):537–540, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [Wil97, BCJ97].

Johnson:1993:NPR

[JPL93] James E. Johnson, Mitzi G. Pardew, and Madeleine M. Lyttle. Notes: Predator recognition and avoidance by larval razorback sucker and northern hog sucker. Transactions of the American Fisheries Society, 122(6):1139–1145, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1995:PCR

[JRLK95] C. M. Jones, D. S. Robson, H. D. Lakkis, and J. Kressel. Properties of catch rates used in analysis of angler surveys. Transactions of the American Fisheries Society, 124(6):911–928, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1990:UCS

[JROG90] Cynthia M. Jones, Douglas Robson, David Otis, and Steven Gloss. Use of a computer simulation model to determine the behavior of a new survey estimator of recreational angling. Transactions of the American Fisheries Society, 119(1):41–54, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jackson:1995:RWP

[JS95] Leslie F. Jackson and Craig V. Sullivan. Reproduction of white perch: The annual gametogenic cycle. Transactions of the American Fisheries Society, 124(4):563–577, July 1995. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jager:1999:WHC

[JVH99] H. I. Jager, W. Van Winkle, and B. D. Holcomb. Would hydrologic climate changes in Sierra Nevada streams influence trout persistence? Transactions of the American Fisheries Society, 128(2):222–240, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jonas:1998:RID

[JW98] Jory L. Jonas and David H. Wahl. Relative importance of direct and indirect effects of starvation for young walleyes. Transactions of the American Fisheries Society, 127(2):192–205, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klumb:1999:PBS

[KBF99] Robert A. Klumb, Michael A. Bozek, and Richard V. Frie. Proportionality of body to scale growth: Validation of two back-calculation models with individually tagged and recaptured smallmouth bass and walleyes. Transactions of the American Fisheries Society, 128(5):815–831, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keith:1998:RJS

[KBM^{+}98] R. M. Keith, T. C. Bjornn, W. R. Meehan, N. J. Hetrick, and M. A. Brusven. Response of juvenile salmonids to riparian and instream cover modifications in small streams flowing through second-growth forests of Southeast Alaska. Transactions of the American Fisheries Society, 127(6):889–907, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaeding:1990:TSR

[KBSM90] Lynn R. Kaeding, Bob D. Burdick, Patricia A. Schrader, and Charles W. McAda. Temporal and spatial relations between the spawning of humpback chub and roundtail chub in the Upper Colorado River. Transactions of the American Fisheries Society, 119(1):135–144, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koenig:1998:AAS

[KC98] Christopher C. Koenig and Felicia C. Coleman. Absolute abundance and survival of juvenile gags in sea grass beds of the Northeastern Gulf of Mexico. Transactions of the American Fisheries Society, 127(1):44–55, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kondolf:1991:DSP

[KCSF91] G. M. Kondolf, G. F. Cada, M. J. Sale, and T. Felando. Distribution and stability of potential salmonid spawning gravels in steep boulder-bed streams of the Eastern Sierra Nevada. Transactions of the American Fisheries Society, 120(2):177–186, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelsch:1995:PMV

[Kel95] Steven W. Kelsch. Patterns of morphometric variation in the channel and headwater catfishes. Transactions of the American Fisheries Society, 124(2):272–279, March 1995. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelsch:1996:TSP

[Kel96] Steven W. Kelsch. Temperature selection and performance by bluegills: Evidence for selection in response to available power. Transactions of the American Fisheries Society, 125(6):948–955, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kellogg:1999:BDC

[Kel99] Karen A. Kellogg. Biodiversity dynamics and conservation: The freshwater fish of tropical Africa. Transactions of the American Fisheries Society, 128(1):182–183, January 1999. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:1990:TSG

[Ken90] Robert L. Kendall. Towards a standardized genetic nomenclature for fish. Transactions of the American Fisheries Society, 119(1):

1, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendra:1991:QSH

[Ken91] Will Kendra. Quality of salmonid hatchery effluents during a summer low-flow season. Transactions of the American Fisheries Society, 120(1):43–51, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendall:1995:BCM

[Ken95] Robert L. Kendall. Bar-coded measuring systems: Priority. Transactions of the American Fisheries Society, 124(4):642, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [Mor95, Sig95].

Kennish:1998:SDS

[Ken98] Michael J. Kennish. Sustainable development in the southeastern coastal zone. Transactions of the American Fisheries Society, 127 (1):152–153, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kerr:1990:FPS

[Ker90] S. R. Kerr. The fry paradigm: Its significance for contemporary ecology. Transactions of the American Fisheries Society, 119(4): 779–785, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kieffer:1996:RBB

[KFTT96] J. D. Kieffer, R. A. Ferguson, J. E. Tompa, and B. L. Tufts. Relationship between body size and anaerobic metabolism in brook trout and largemouth bass. Transactions of the American Fisheries Society, 125(5):760–767, September 1996. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keenlyne:1992:FPS

[KGJ92] K. D. Keenlyne, E. M. Grossman, and L. G. Jenkins. Fecundity of the pallid sturgeon. Transactions of the American Fisheries Society, 121(1):139–140, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ketola:1993:IPR

[KH93] H. George Ketola and Barbara F. Harland. Influence of phosphorus in rainbow trout diets on phosphorus discharges in effluent water. Transactions of the American Fisheries Society, 122(6): 1120–1126, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kruse:1997:GID

[KHR97] Carter G. Kruse, Wayne A. Hubert, and Frank J. Rahel. Geomorphic influences on the distribution of Yellowstone cutthroat trout in the absaroka mountains, Wyoming. Transactions of the American Fisheries Society, 126(3):418–427, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keenlyne:1994:MCU

[KHTC94] K. D. Keenlyne, C. J. Henry, A. Tews, and P. Clancey. Morphometric comparisons of Upper Missouri River sturgeons. Transactions of the American Fisheries Society, 123(5):779–785, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kitchell:1990:SMC

[Kit90] James F. Kitchell. The scope for mortality caused by sea lamprey. Transactions of the American Fisheries Society, 119(4):642–648, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keenlyne:1993:ASM

[KJ93] K. D. Keenlyne and L. G. Jenkins. Age at sexual maturity of the pallid sturgeon. Transactions of the American Fisheries Society, 122(3):393–396, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kaya:1995:NRA

[KJ95] Calvin M. Kaya and Eric D. Jeanes. Notes: Retention of adaptive rheotactic behavior by F_{1} fluvial Arctic grayling. Transactions of the American Fisheries Society, 124(3):453–457, May 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kieffer:1993:AMS

[KK93] Micah C. Kieffer and Boyd Kynard. Annual movements of shortnose and Atlantic sturgeons in the Merrimack River, Massachusetts. Transactions of the American Fisheries Society, 122 (6):1088–1103, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kieffer:1996:SSS

[KK96] Micah C. Kieffer and Boyd Kynard. Spawning of the shortnose sturgeon in the Merrimack River, Massachusetts. Transactions of the American Fisheries Society, 125(2):179–186, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kieffer:1995:ECR

[KKP^{+}95] J. D. Kieffer, M. R. Kubacki, F. J. S. Phelan, D. P. Philipp, and B. L. Tufts. Effects of catch-and-release angling on nesting male smallmouth bass. Transactions of the American Fisheries Society, 124(1):70–76, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kane:1994:UGK

[KKR^{+}94] Andrew S. Kane, Myong W. Kahng, Renate Reimschuessel, Patson T. Nhamburo, and Michael M. Lipsky. UDP-Glucuronyltransferase kinetics for 3-trifluoromethyl-4-nitrophenol (TFM) in fish. Transactions of the American Fisheries Society, 123(2):217–222, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kirn:1996:GSR

[KL96a] Richard A. Kirn and George W. Labar. Growth and survival of rainbow smelt, and their role as prey for stocked salmonids in Lake Champlain. Transactions of the American Fisheries Society, 125(1):87–96, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knights:1996:EIT

[KL96b] Brent C. Knights and Becky A. Lasee. Effects of implanted transmitters on adult bluegills at two temperatures. Transactions of the American Fisheries Society, 125(3):440–449, May 1996. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kitano:1994:NSB

[KMNF94] Satoshi Kitano, Koji Maekawa, Shigeru Nakano, and Kurt D. Fausch. Notes: Spawning behavior of bull trout in the Upper Flathead Drainage, Montana, with special reference to hybridization with brook trout. Transactions of the American Fisheries Society, 123(6):988–992, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelsch:1990:TPV

[KN90] Steven W. Kelsch and William H. Neill. Temperature preference versus acclimation in fishes: Selection for changing metabolic optima. Transactions of the American Fisheries Society, 119(4): 601–610, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kempinger:1998:FKF

[KOB98] James J. Kempinger, Keith J. Otis, and Joseph R. Ball. Fish kills in the Fox River, Wisconsin, attributable to carbon monoxide from marine engines. Transactions of the American Fisheries Society, 127(4):669–672, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knight:1999:GSM

[KOB99] Curtis A. Knight, Rick W. Orme, and David A. Beauchamp. Growth, survival, and migration patterns of juvenile adfluvial Bonneville cutthroat trout in tributaries of Strawberry Reservoir, Utah. Transactions of the American Fisheries Society, 128(4): 553–563, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kynard:1996:DFO

[KOH96] Boyd Kynard, Mufeed Odeh, and Alex Haro. Design of fishways and other fish facilities, 2nd edition. Transactions of the American Fisheries Society, 125(4):631, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kohler:1997:UEC

[Koh97] Christopher C. Kohler. Uses and effects of cultured fishes in aquatic ecosystems. Transactions of the American Fisheries Society, 126(3):543–547, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kolok:1991:TCT

[Kol91] Alan S. Kolok. Temperature compensation in two centrarchid fishes: Do winter-quiescent fish undergo cellular temperature compensation? Transactions of the American Fisheries Society, 120(1):52–57, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kolok:1995:PF

[Kol95] Alan S. Kolok. The physiology of fishes. Transactions of the American Fisheries Society, 124(1):147–150, January 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kneib:1991:NGC

[KP91] R. T. Kneib and J. H. Parker. Notes: Gross conversion efficiencies of mummichog and spotfin killifish larvae from a Georgia salt marsh. Transactions of the American Fisheries Society, 120(6): 803–809, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

King:1992:PSS

[KP92] Tim L. King and Henry O. Pate. Population structure of spotted seatrout inhabiting the Texas Gulf Coast: an allozymic perspective. Transactions of the American Fisheries Society, 121(6): 746–756, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ketola:1994:RRT

[KR94a] H. George Ketola and Milo E. Richmond. Requirement of rainbow trout for dietary phosphorus and its relationship to the amount discharged in hatchery effluents. Transactions of the American Fisheries Society, 123(4):587–594, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knight:1994:FHB

[KR94b] Jeremy G. Knight and Stephen T. Ross. Feeding habits of the bayou darter. Transactions of the American Fisheries Society, 123(5):794–802, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keleher:1996:TLS

[KR96] Christopher J. Keleher and Frank J. Rahel. Thermal limits to salmonid distributions in the Rocky Mountain region and potential habitat loss due to global warming: a Geographic Information System (GIS) approach. Transactions of the American Fisheries Society, 125(1):1–13, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kraft:1993:PRL

[Kra93] Clifford E. Kraft. Phosphorus regeneration by Lake Michigan alewives in the mid-1970s. Transactions of the American Fisheries Society, 122(5):749–755, September 1993. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kolok:1997:EFA

[KS97] Alan S. Kolok and Drew Sharkey. Effect of freshwater acclimation on the swimming performance and plasma osmolarity of the euryhaline Gulf killifish. Transactions of the American Fisheries Society, 126(5):866–870, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kohler:1994:HCC

[KSH^{+}94] Christopher C. Kohler, Robert J. Sheehan, Christopher Habicht, Jeffrey A. Malison, and Terrence B. Kayes. Habituation to captivity and controlled spawning of white bass. Transactions of the American Fisheries Society, 123(6):964–974, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

King:1999:ELB

[KSZ99] J. R. King, B. J. Shuter, and A. P. Zimmerman. Empirical links between thermal habitat, fish growth, and climate change. Transactions of the American Fisheries Society, 128(4):656–665, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kocik:1994:SSG

[KT94] John F. Kocik and William W. Taylor. Summer survival and growth of brown trout with and without steelhead under equal total salmonine densities in an artificial stream. Transactions of the American Fisheries Society, 123(6):931–938, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kocik:1996:EJS

[KT96] John F. Kocik and William W. Taylor. Effect of juvenile steelhead on juvenile brown trout habitat use in a low-gradient Great Lakes tributary. Transactions of the American Fisheries Society, 125(2):244–252, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knapp:1996:SCG

[KV96] Roland A. Knapp and Vance T. Vredenburg. Spawning by California golden trout: Characteristics of spawning fish, seasonal and daily timing, redd characteristics, and microhabitat preferences. Transactions of the American Fisheries Society, 125(4): 519–531, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kroll:1992:EWT

[KVD^{+}92] Kevin J. Kroll, Joel P. Van Eenennaam, Serge I. Doroshov, Jerry E. Hamilton, and Thomas R. Russell. Effect of water temperature and formulated diets on growth and survival of larval paddlefish. Transactions of the American Fisheries Society, 121(4):538–543, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kondolf:1996:SSH

[KVR96] G. Mathias Kondolf, Jennifer C. Vick, and Timothy M. Ramirez. Salmon spawning habitat rehabilitation on the Merced River, California: an evaluation of project planning and performance. Transactions of the American Fisheries Society, 125(6):899–912, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kwak:1997:TPD

[KW97] Thomas J. Kwak and Thomas F. Waters. Trout production dynamics and water quality in Minnesota streams. Transactions of the American Fisheries Society, 126(1):35–48, January 1997. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knutsen:1999:BCN

[KW99] Christopher J. Knutsen and David L. Ward. Biological characteristics of northern pikeminnow in the Lower Columbia and Snake Rivers before and after sustained exploitation. Transactions of the American Fisheries Society, 128(6):1008–1019, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kwak:1991:ECN

[Kwa91] Thomas J. Kwak. Ecological characteristics of a northern population of the pallid shiner. Transactions of the American Fisheries Society, 120(1):106–115, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krueger:1999:SOD

[KWE99] Charles C. Krueger, Richard L. Wilmot, and Rebecca J. Everett. Stock origins of Dolly Varden collected from Beaufort Sea coastal sites of Arctic Alaska and Canada. Transactions of the American Fisheries Society, 128(1):49–57, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leary:1997:NGC

[LA97] Robb F. Leary and Fred W. Allendorf. Notes: Genetic confirmation of sympatric bull trout and Dolly Varden in Western Washington. Transactions of the American Fisheries Society, 126(4): 715–720, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Labar:1993:UBM

[Lab93] George W. Labar. Use of bioenergetics models to predict the effect of increased lake trout predation on rainbow smelt following sea lamprey control. Transactions of the American Fisheries Society, 122(5):942–950, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leftwich:1997:FIB

[LAD97] Kevin N. Leftwich, Paul L. Angermeier, and C. Andrew Dolloff. Factors influencing behavior and transferability of habitat models for a benthic stream fish. Transactions of the American Fisheries Society, 126(5):725–734, September 1997. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lavery:1992:CFB

[Lav92] Robert J. Lavery. Cichlid fishes: Behaviour, ecology and evolution. Transactions of the American Fisheries Society, 121(3): 403–404, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lockwood:1992:GSB

[LB92] Samuel F. Lockwood and John W. Bickham. Genome size in Beaufort Sea coastal assemblages of Arctic ciscoes. Transactions of the American Fisheries Society, 121(1):13–20, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luecke:1993:NEE

[LB93] Chris Luecke and Darren Brandt. Notes: Estimating the energy density of daphnid prey for use with rainbow trout bioenergetics models. Transactions of the American Fisheries Society, 122(3): 386–389, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Linner:1994:BRC

[LB94] Johan Linnér and Eva Brännäs. Behavioral response to commercial food of different sizes and self-initiated food size selection by Arctic char. Transactions of the American Fisheries Society, 123(3):416–422, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loch:1999:OGC

[LB99] John J. Loch and Scott A. Bonar. Occurrence of grass carp in the Lower Columbia and Snake Rivers. Transactions of the American Fisheries Society, 128(2):374–379, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leclerc:1995:TDH

[LBBC95] Michel Leclerc, André Boudreault, Toss A. Bechara, and Geneviève Corfa. Two-dimensional hydrodynamic modeling: a neglected tool in the instream flow incremental methodology. Transactions of the American Fisheries Society, 124(5):645–662, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lowerre-Barbieri:1996:VSA

[LBCB96] Susan K. Lowerre-Barbieri, Mark E. Chittenden, Jr., and Luiz R. Barbieri. Variable spawning activity and annual fecundity of weakfish in Chesapeake Bay. Transactions of the American Fisheries Society, 125(4):532–545, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lohr:1996:HTT

[LBKD96] S. C. Lohr, P A. Byorth, C. M. Kaya, and W. P. Dwyer. Hightemperature tolerances of fluvial Arctic grayling and comparisons with summer river temperatures of the Big Hole River, Montana. Transactions of the American Fisheries Society, 125(6): 933–939, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lobon-Cervia:1996:RSF

[LC96] Javier Lobón-Cerviá. Response of a stream fish assemblage to a severe spate in Northern Spain. Transactions of the American Fisheries Society, 125(6):913–919, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lobon-Cervia:1998:FAI

[LCR98] Javier Lobón-Cerviá and Pedro A. Rincón. Field assessment of the influence of temperature on growth rate in a brown trout population. Transactions of the American Fisheries Society, 127 (5):718–728, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lin:1996:CMSb

[LD96] Feng Lin and Konrad Dabrowski. Characteristics of muskellunge spermatozoa II: Effects of ions and osmolality on sperm motility. Transactions of the American Fisheries Society, 125(2):195–202, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lein:1997:BET

[LD97] Gregory M. Lein and Dennis R. Devries. Boat electrofishing as a technique for sampling paddlefish. Transactions of the American Fisheries Society, 126(2):334–337, March 1997. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lein:1998:PAR

[LD98] Gregory M. Lein and Dennis R. Devries. Paddlefish in the Alabama River drainage: Population characteristics and the adult spawning migration. Transactions of the American Fisheries Society, 127(3):441–454, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leibovitz:1990:DPM

[Lei90] Louis Leibovitz. Disease processes in marine bivalve molluscs. Transactions of the American Fisheries Society, 119(1):163–165, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lemke:1996:BPA

[Lem96] Michael J. Lemke. The biology of particles in aquatic systems. Transactions of the American Fisheries Society, 125(2):

341, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LaPeyre:1995:IMI

[LF95] Jerome F. La Peyre and Mohamed Faisal. Improved method for the initiation of continuous cultures of the oyster pathogen Perkinsus marinus (Apicomplexa). Transactions of the American Fisheries Society, 124(1):144–146, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lapatra:1990:SRS

[LGRF90] S. E. Lapatra, W. J. Groberg, J. S. Rohovec, and J. L. Fryer. Size-related susceptibility of salmonids to two strains of infectious hematopoietic necrosis virus. Transactions of the American Fisheries Society, 119(1):25–30, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liu:1996:IEG

[LGS^{+}96] Qinghua Liu, Cheryl A. Goudie, Bill A. Simco, Kenneth B. Davis, and Donald C. Morizot. Isozyme expression and gene–centromere distances in diploid and triploid hybrid catfish. Transactions of the American Fisheries Society, 125(1):56–65, January 1996. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lacroix:1995:SMU

[LHS95] Gilles L. Lacroix, Dale J. Hood, and James A. Smith. Stability of microhabitat use by brook trout and juvenile Atlantic salmon after stream acidification. Transactions of the American Fisheries Society, 124(4):588–598, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Li:1999:TCD

[Li99] Eric A. L. Li. A travel cost demand model for recreational snapper angling in Port Phillip Bay, Australia. Transactions of the American Fisheries Society, 128(4):639–647, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Link:1996:CPL

[Lin96] Jason Link. Capture probabilities of Lake Superior zooplankton by an obligate planktivorous fish — the lake herring. Transactions of the American Fisheries Society, 125(1):139–142, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lisle:1998:ARM

[Lis98] Thomas E. Lisle. Applied river morphology. Transactions of the American Fisheries Society, 127(5):856–858, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liu:1992:VLW

[LJ92] Kwang Ming Liu and A. L. Jensen. Validation of lake whitefish catch-per-unit-effort data with time series analysis. Transactions of the American Fisheries Society, 121(6):797–801, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lucas:1993:UPT

[LJP93] Martyn C. Lucas, Alastair D. F. Johnstone, and Imants G. Priede. Use of physiological telemetry as a method of estimating metabolism of fish in the natural environment. Transactions of the American Fisheries Society, 122(5):822–833, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Landahl:1997:ADE

[LJS^{+}97] John T. Landahl, Lyndal L. Johnson, John E. Stein, Tracy K. Collier, and Usha Varanasi. Approaches for determining effects of pollution on fish populations of Puget Sound. Transactions of the American Fisheries Society, 126(3):519–535, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lin:1996:CMSa

[LLD96] Feng Lin, Li Liu, and Konrad Dabrowski. Characteristics of muskellunge spermatozoa i: Ultrastructure of spermatozoa and biochemical composition of semen. Transactions of the American Fisheries Society, 125(2):187–194, March 1996. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liu:1999:DAF

[LLK^{+}99] Zhanjiang Liu, Ping Li, Huseyin Kucuktas, Amy Nichols, Guo Tan, Xinmin Zheng, Brad J. Argue, Rex A. Dunham, and D. Roger Yant. Development of amplified fragment length polymorphism (AFLP) markers suitable for genetic linkage mapping of catfish. Transactions of the American Fisheries Society, 128 (2):317–327, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Li:1994:CER

[LLP^{+}94] Hiram W. Li, Gary A. Lamberti, Todd N. Pearsons, Cynthia K. Tait, Judith L. Li, and John C. Buckhouse. Cumulative effects of riparian disturbances along high desert trout streams of the John Day Basin, Oregon. Transactions of the American Fisheries Society, 123(4):627–640, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luo:1991:RBB

[LM91] Jiangang Luo and John A. Musick. Reproductive biology of the bay anchovy in Chesapeake Bay. Transactions of the American Fisheries Society, 120(6):701–710, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larson:1993:FHD

[LM93] Gary L. Larson and C. David McIntire. Food habits of different phenotypes of threespine stickleback in Paxton Lake, British Columbia. Transactions of the American Fisheries Society, 122 (4):543–549, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larson:1995:EFE

[LMC95] Gary L. Larson, Stephen E. Moore, and Bart Carter. Ebb and flow of encroachment by nonnative rainbow trout in a small stream in the Southern Appalachian mountains. Transactions of the American Fisheries Society, 124(4):613–622, July 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lobb:1991:HUA

[LO91] M. Delbert Lobb III and Donald J. Orth. Habitat use by an assemblage of fish in a large warmwater stream. Transactions of the American Fisheries Society, 120(1):65–78, January 1991. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lukas:1995:FAN

[LO95] Joseph A. Lukas and Donald J. Orth. Factors affecting nesting success of smallmouth bass in a regulated Virginia stream. Transactions of the American Fisheries Society, 124(5):726–735, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lovshin:1998:FFE

[Lov98] Leonard L. Lovshin. So fruitful a fish: Ecology, conservation, and aquaculture of the Amazon’s Tambaquí. Transactions of the American Fisheries Society, 127(6):1082–1083, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1994:DJA

[LP94] V. Pernell Lewis and David S. Peters. Diet of juvenile and adult Atlantic menhaden in estuarine and coastal habitats. Transactions of the American Fisheries Society, 123(5):803–810, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Limburg:1997:CSU

[LPFA97] K. E. Limburg, M. L. Pace, D. Fischer, and K. K. Arend. Consumption, selectivity, and use of zooplankton by larval striped bass and white perch in a seasonally pulsed estuary. Transactions of the American Fisheries Society, 126(4):607–621, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Liao:1995:RWF

[LPW^{+}95] Hongsheng Liao, Clay L. Pierce, David H. Wahl, Joseph B. Rasmussen, and William C. Leggett. Relative weight (W_{r}) as a field assessment tool: Relationships with growth, prey biomass, and environmental conditions. Transactions of the American Fisheries Society, 124(3):387–400, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lange:1993:IWC

[LRC93] Ted R. Lange, Homer E. Royals, and Laurence L. Connor. Influence of water chemistry on mercury concentration in largemouth bass from Florida lakes. Transactions of the American Fisheries Society, 122(1):74–84, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Letcher:1996:SDE

[LRCB96] Benjamin H. Letcher, James A. Rice, Larry B. Crowder, and Fred P. Binkowski. Size-dependent effects of continuous and intermittent feeding on starvation time and mass loss in starving yellow perch larvae and juveniles. Transactions of the American Fisheries Society, 125(1):14–26, January 1996. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Linton:1997:MCP

[LRW97] Tyler K. Linton, Scott D. Reid, and Chris M. Wood. The metabolic costs and physiological consequences to juvenile rainbow trout of a simulated summer warming scenario in the presence and absence of sublethal ammonia. Transactions of the American Fisheries Society, 126(2):259–272, March 1997. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Linton:1998:MCP

[LRW98] Tyler K. Linton, S. D. Reid, and Chris M. Wood. The metabolic costs and physiological consequences to juvenile rainbow trout of a simulated winter warming scenario in the presence or absence of sublethal ammonia. Transactions of the American Fisheries Society, 127(4):611–619, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Linton:1999:ERR

[LRW99] Tyler K. Linton, S. D. Reid, and Chris M. Wood. Effects of a restricted ration on the growth and energetics of juvenile rainbow trout exposed to a summer of simulated warming and sublethal ammonia. Transactions of the American Fisheries Society, 128(4):758–763, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lantry:1993:EER

[LS93] Brian F. Lantry and Donald J. Stewart. Ecological energetics of rainbow smelt in the Laurentian Great Lakes: an interlake comparison. Transactions of the American Fisheries Society, 122 (5):951–976, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lipton:1997:EEP

[LS97] Douglas W. Lipton and Ivar E. Strand. Economic effects of pollution in fish habitats. Transactions of the American Fisheries Society, 126(3):514–518, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lester:1993:CES

[LSC93] Nigel P. Lester, Brian J. Shuter, and I. G. Cowx. Catch effort sampling strategies: Their application in freshwater fisheries management. Transactions of the American Fisheries Society, 122(5):1031–1033, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lorenz:1992:ECS

[LT92] Jerome J. Lorenz and Douglas H. Taylor. Effects of a chemical stressor on the parental behavior of convict cichlids with offspring in early stages of development. Transactions of the American Fisheries Society, 121(3):315–321, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laughlin:1996:HDM

[LT96] Thomas F. Laughlin and Bruce J. Turner. Hypervariable DNA markers reveal high genetic variability within striped bass populations of the Lower Chesapeake Bay. Transactions of the American Fisheries Society, 125(1):49–55, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luecke:1990:CAD

[Lue90] Chris Luecke. Changes in abundance and distribution of benthic macroinvertebrates after introduction of cutthroat trout into a previously fishless lake. Transactions of the American Fisheries Society, 119(6):1010–1021, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lutz:1995:GSG

[Lut95] Donna Schulze Lutz. Gas supersaturation and gas bubble trauma in fish downstream from a Midwestern reservoir. Transactions of the American Fisheries Society, 124(3):423–436, May 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lohr:1992:MSB

[LW92] Samuel C. Lohr and Jerry L. West. Microhabitat selection by brook and rainbow trout in a Southern Appalachian stream. Transactions of the American Fisheries Society, 121(6):729–736, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Luecke:1993:EMD

[LW93] Chris Luecke and Wayne A. Wurtsbaugh. Effects of moonlight and daylight on hydroacoustic estimates of pelagic fish abundance. Transactions of the American Fisheries Society, 122(1):

112–120, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lyons:1997:IWS

[Lyo97] John Lyons. Influence of winter starvation on the distribution of smallmouth bass among Wisconsin streams: a bioenergetics modeling assessment. Transactions of the American Fisheries Society, 126(1):157–162, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maceina:1997:PDC

[Mac97] Michael J. Maceina. Population dynamics of commercial fish in inland reservoirs. Transactions of the American Fisheries Society, 126(2):347–348, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marty:1990:WHC

[Mar90] Gary D. Marty. Wound healing in channel catfish by epithelialization and contraction of granulation tissue. Transactions of the American Fisheries Society, 119(1):145–150, January 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:1996:DMH

[Mat96] Kathleen R. Matthews. Diel movement and habitat use of California golden trout in the Golden Trout Wilderness, California. Transactions of the American Fisheries Society, 125(1):78–86, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

May:1998:MSF

[May98] Bernie May. Molecular systematics of fishes. Transactions of the American Fisheries Society, 127(5):856–858, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Medland:1991:LSB

[MB91a] T. E. Medland and F. W. H. Beamish. Lamprey statolith banding patterns in response to temperature, photoperiod, and ontogeny. Transactions of the American Fisheries Society, 120(2):255–260, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muth:1991:EST

[MB91b] Robert T. Muth and Kevin R. Bestgen. Effect of sunlight on tetracycline marks in otoliths of Colorado squawfish larvae. Transactions of the American Fisheries Society, 120(5):666–668, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1996:FRP

[MB96] Allen I. Miller and Lance G. Beckman. First record of predation on white sturgeon eggs by sympatric fishes. Transactions of the American Fisheries Society, 125(2):338–340, March 1996. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:1994:CWF

[MBAL94] Kathleen R. Matthews, Neil H. Berg, David L. Azuma, and Thomas R. Lambert. Cool water formation and trout habitat use in a deep pool in the Sierra Nevada, California. Transactions of the American Fisheries Society, 123(4):549–564, July 1994. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murdoch:1991:LSG

[MBD91] Scott P. Murdoch, F. H. William Beamish, and Margaret F. Docker. Laboratory study of growth and interspecific competition in larval lampreys. Transactions of the American Fisheries Society, 120(5):653–656, September 1991. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mallin:1992:CBM

[MBS^{+}92] Michael A. Mallin, JoAnn M. Burkholder, Michael J. Sullivan, W. E. Schaaf, and D. S. Peters. Contributions of benthic microalgae to coastal fishery yield. Transactions of the American Fisheries Society, 121(5):691–695, September 1992. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:1993:SET

[MC93a] Charles P. Madenjian and Stephen R. Carpenter. Simulation of the effects of time and size at stocking on PCB accumulation in lake trout. Transactions of the American Fisheries Society, 122(3):492–499, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madon:1993:BML

[MC93b] Sharook P. Madon and David A. Culver. Bioenergetics model for larval and juvenile walleyes: an in situ approach with experimental ponds. Transactions of the American Fisheries Society, 122 (5):797–813, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mason:1993:FGS

[MC93c] William T. Mason, Jr. and James P. Clugston. Foods of the Gulf sturgeon in the Suwannee River, Florida. Transactions of the American Fisheries Society, 122(3):378–385, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1990:ECZ

[MCB90] Thomas Miller, Larry B. Crowder, and Fred P. Binkowski. Effects of changes in the zooplankton assemblage on growth of bloater and implications for recruitment success. Transactions of the American Fisheries Society, 119(3):483–491, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCormick:1990:PCT

[McC90] Michael J. McCormick. Potential changes in thermal structure and cycle of Lake Michigan due to global warming. Transactions of the American Fisheries Society, 119(2):183–194, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morizot:1991:MHA

[MCC^{+}91] Donald C. Morizot, Stuart W. Calhoun, Lisa L. Clepper, Maureen E. Schmidt, J. Holt Williamson, and Gary J. Carmichael. Multispecies hybridization among native and introduced centrarchid basses in Central Texas. Transactions of the American Fisheries Society, 120(3):283–289, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGarey:1991:PDM

[McG91] Donald J. McGarey. Principal diseases of marine fish and shellfish, 2nd edition. volume 1: Fish. Transactions of the American Fisheries Society, 120(3):401–403, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGurk:1993:AHM

[McG93] Michael D. McGurk. Allometry of herring mortality. Transactions of the American Fisheries Society, 122(6):1035–1042, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDowall:1993:TCN

[MCWN93] R. M. McDowall, B. M. Clark, G. J. Wright, and T. G. Northcote. Trans-2-cis-6-nonadienal: The cause of cucumber odor in osmerid and retropinnid smelts. Transactions of the American Fisheries Society, 122(1):144–147, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marsh:1997:PIF

[MD97] Paul C. Marsh and Michael E. Douglas. Predation by introduced fishes on endangered humpback chub and other native species in the Little Colorado River, Arizona. Transactions of the American Fisheries Society, 126(2):343–346, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:1998:OSP

[MDS98] Charles P. Madenjian, Timothy J. Desorcie, and Ralph M. Stedman. Ontogenic and spatial patterns in diet and growth of lake trout in Lake Michigan. Transactions of the American Fisheries Society, 127(2):236–252, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McRae:1994:TCW

[ME94] Gil McRae and Clayton J. Edwards. Thermal characteristics of Wisconsin headwater streams occupied by Beaver: Implications for brook trout habitat. Transactions of the American Fisheries Society, 123(4):641–656, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meffe:1991:FIS

[Mef91] Gary K. Meffe. Failed invasion of a southeastern blackwater stream by bluegills: Implications for conservation of native communities. Transactions of the American Fisheries Society, 120(3): 333–338, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meisner:1990:PLT

[Mei90] J. Donald Meisner. Potential loss of thermal habitat for brook trout, due to climatic warming, in two Southern Ontario streams. Transactions of the American Fisheries Society, 119(2):282–291, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meng:1993:EFR

[Men93a] Lesa Meng. Estimating food requirements of striped bass larvae: an energetics approach. Transactions of the American Fisheries Society, 122(2):244–251, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meng:1993:SSS

[Men93b] Lesa Meng. Sustainable swimming speeds of striped bass larvae. Transactions of the American Fisheries Society, 122(5):702–708, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesa:1991:VFA

[Mes91] Matthew G. Mesa. Variation in feeding, aggression, and position choice between hatchery and wild cutthroat trout in an artificial stream. Transactions of the American Fisheries Society, 120(6): 723–727, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesa:1994:EMA

[Mes94] Matthew G. Mesa. Effects of multiple acute stressors on the predator avoidance ability and physiology of juvenile Chinook salmon. Transactions of the American Fisheries Society, 123(5): 786–793, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McParland:1999:GPS

[MFL99] Tara L. McParland, Moira M. Ferguson, and Arunas P. Liskauskas. Genetic population structure and mixed-stock analysis of walleyes in the Lake Erie–Lake Huron corridor using allozyme and mitochondrial DNA markers. Transactions of the American Fisheries Society, 128(6):1055–1067, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maynard:1996:VMA

[MFWP96] Desmond J. Maynard, Deborah A. Frost, F. William Waknitz, and Earl F. Prentice. Vulnerability of marked age-0 steelhead to a visual predator. Transactions of the American Fisheries Society, 125(2):330–333, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manire:1991:EMT

[MG91] Charles A. Manire and Samuel H. Gruber. Effect of M-type dart tags on field growth of juvenile lemon sharks. Transactions of the American Fisheries Society, 120(6):776–780, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matthews:1992:FHU

[MGH92] William J. Matthews, Frances P. Gelwick, and Jan Jeffrey Hoover. Food of and habitat use by juveniles of species of Micropterus and Morone in a southwestern reservoir. Transactions of the American Fisheries Society, 121(1):54–66, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDonald:1993:RHR

[MGM93] D. G. McDonald, M. D. Goldstein, and C. Mitton. Responses of hatchery-reared brook trout, lake trout, and splake to transport stress. Transactions of the American Fisheries Society, 122(6): 1127–1138, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matuszek:1990:OCO

[MGW90] John E. Matuszek, John Goodier, and Donna L. Wales. The occurrence of Cyprinidae and other small fish species in relation to pH in Ontario lakes. Transactions of the American Fisheries Society, 119(5):850–861, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miranda:1994:LDW

[MH94] L. E. Miranda and W. D. Hubbard. Length-dependent winter survival and lipid composition of age-0 largemouth bass in Bay Springs Reservoir, Mississippi. Transactions of the American Fisheries Society, 123(1):80–87, January 1994. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marty:1995:NOC

[MHC95] Gary D. Marty, David E. Hinton, and Joseph J. Cech, Jr. Notes: Oxygen consumption by larval Japanese medaka with inflated or uninflated swim bladders. Transactions of the American Fisheries Society, 124(4):623–627, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathur:1994:NTP

[MHR94] Dilip Mathur, Paul G. Heisey, and David A. Robinson. Notes: Turbine-passage mortality of juvenile American shad at a lowhead hydroelectric dam. Transactions of the American Fisheries Society, 123(1):108–111, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murphy:1999:RPS

[MHS^{+}99] Michael L. Murphy, Ronald A. Heintz, Jeffrey W. Short, Marie L. Larsen, and Stanley D. Rice. Recovery of pink salmon spawning areas after the Exxon Valdez oil spill. Transactions of the American Fisheries Society, 128(5):909–918, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1992:LHS

[MHSM92] Peter B. Moyle, Bruce Herbold, Donald E. Stevens, and Lee W. Miller. Life history and status of delta smelt in the Sacramento– San Joaquin Estuary, California. Transactions of the American Fisheries Society, 121(1):67–77, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mazik:1991:INC

[MHW^{+}91] Patricia M. Mazik, Mark L. Hinman, Douglas A. Winkelmann, Stephen J. Klaine, Bill A. Simco, and Nick C. Parker. Influence of nitrite and chloride concentrations on survival and hematological profiles of striped bass. Transactions of the American Fisheries Society, 120(2):247–254, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Michaletz:1997:FAA

[Mic97a] Paul H. Michaletz. Factors affecting abundance, growth, and survival of age-0 gizzard shad. Transactions of the American Fisheries Society, 126(1):84–100, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Michaletz:1997:IAS

[Mic97b] Paul H. Michaletz. Influence of abundance and size of age-0 gizzard shad on predator diets, diet overlap, and growth. Transactions of the American Fisheries Society, 126(1):101–111, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moran:1997:FAI

[MIPGV97] Paloma Moran, Jorge I. Izquierdo, Alberto M. Pendas, and Eva Garcia-Vazquez. Fluctuating asymmetry and isozyme variation in Atlantic salmon: Relation to age of wild and hatchery fish. Transactions of the American Fisheries Society, 126(2):194–199, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCarter:1993:MOP

[MJ93] Nigel H. McCarter and Mark R. James. A method to observe particle ingestion by fish larvae. Transactions of the American Fisheries Society, 122(4):642–644, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCauley:1990:EAT

[MK90a] Robert W. McCauley and D. Marl Kilgour. Effect of air temperature on growth of largemouth bass in North America. Transactions of the American Fisheries Society, 119(2):276–281, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meador:1990:GLB

[MK90b] Michael R. Meador and William E. Kelso. Growth of largemouth bass in low-salinity environments. Transactions of the American Fisheries Society, 119(3):545–552, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McAda:1991:MAC

[MK91] Charles W. McAda and Lynn R. Kaeding. Movements of adult Colorado squawfish during the spawning season in the Upper Colorado River. Transactions of the American Fisheries Society, 120(3):339–345, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1996:NMD

[MK96] Loren M. Miller and Anne R. Kapuscinski. Notes: Microsatellite DNA markers reveal new levels of genetic variation in northern pike. Transactions of the American Fisheries Society, 125(6): 971–977, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Monzyk:1997:CWC

[MKR97] Frederick R. Monzyk, William E. Kelso, and D. Allen Rutherford. Characteristics of woody cover used by brown madtoms and pirate perch in coastal plain streams. Transactions of the American Fisheries Society, 126(4):665–675, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Magnelia:1992:ADA

[MKT92] Stephan J. Magnelia, Christopher C. Kohler, and Donald R. Tindall. Acanthurids do not avoid consuming cultured toxic dinoflagellates yet do not become ciguatoxic. Transactions of the American Fisheries Society, 121(6):737–745, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCormick:1999:FCF

[ML99] J. Howard McCormick and Richard L. Leino. Factors contributing to first-year recruitment failure of fishes in acidified waters with some implications for environmental research. Transactions of the American Fisheries Society, 128(2):265–277, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MacEina:1992:SFL

[MM92a] Michael J. MacEina and Brian R. Murphy. Stocking Florida largemouth bass outside its native range. Transactions of the American Fisheries Society, 121(5):686–691, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mahoney:1992:NDA

[MM92b] John B. Mahoney and J. Kneeland McNulty. Notes: Diseaseassociated blood changes and normal seasonal hematological variation in winter flounder in the Hudson–Raritan Estuary. Transactions of the American Fisheries Society, 121(2):261–268, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meng:1995:SSS

[MM95] Lesa Meng and Peter B. Moyle. Status of splittail in the Sacramento–San Joaquin Estuary. Transactions of the American Fisheries Society, 124(4):538–549, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Magnuson:1990:PCT

[MMH90] John J. Magnuson, J. Donald Meisner, and David K. Hill. Potential changes in the thermal habitat of Great Lakes fish after global climate warming. Transactions of the American Fisheries Society, 119(2):254–264, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meng:1994:CAD

[MMH94] Lesa Meng, Peter B. Moyle, and Bruce Herbold. Changes in abundance and distribution of native and introduced fishes of Suisun marsh. Transactions of the American Fisheries Society, 123(4):498–507, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McAllister:1992:PNE

[MMSW92] K. W. McAllister, P. E. McAllister, R. C. Simon, and J. K. Werner. Performance of nine external tags on hatchery-reared rainbow trout. Transactions of the American Fisheries Society, 121(2):192–198, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Magee:1996:SVS

[MMT96] James P. Magee, Thomas E. McMahon, and Russell F. Thurow. Spatial variation in spawning habitat of cutthroat trout in a sediment-rich stream basin. Transactions of the American Fisheries Society, 125(5):768–779, September 1996. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meng:1991:SPL

[MO91] Lesa Meng and James J. Orsi. Selective predation by larval striped bass on native and introduced copepods. Transactions of the American Fisheries Society, 120(2):187–192, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mesa:1993:PSP

[MO93] Matthew G. Mesa and Todd M. Olson. Prolonged swimming performance of northern squawfish. Transactions of the American Fisheries Society, 122(6):1104–1110, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:1999:LEL

[MO99] Charles P. Madenjian and Daniel V. O’Connor. Laboratory evaluation of a lake trout bioenergetics model. Transactions of the American Fisheries Society, 128(5):802–814, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Monaghan:1993:NCC

[Mon93] J. P. Monaghan, Jr. Notes: Comparison of calcein and tetracycline as chemical markers in summer flounder. Transactions of the American Fisheries Society, 122(2):298–301, March 1993. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morgan:1992:LTT

[Mor92] M. Joanne Morgan. Low-temperature tolerance of American plaice in relation to declines in abundance. Transactions of the American Fisheries Society, 121(3):399–402, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morizur:1995:CBC

[Mor95] Y. Morizur. Comment: Bar-coded measuring systems. Transactions of the American Fisheries Society, 124(4):640–641, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See responses [Sig95, Ken95].

Mason:1993:MST

[MP93] Doran M. Mason and E. Vincent Patrick. A model for the space– time dependence of feeding for pelagic fish populations. Transactions of the American Fisheries Society, 122(5):884–901, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McQueen:1994:FWM

[MP94] D. J. McQueen and S. A. Popiel. Food web management: a case study of Lake Mendota. Transactions of the American Fisheries Society, 123(1):118–120, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mitro:1997:TSA

[MP97] Matthew G. Mitro and Donna L. Parrish. Temporal and spatial abundances of larval walleyes in two tributaries of Lake Champlain. Transactions of the American Fisheries Society, 126(2): 273–287, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McMichael:1998:EWJ

[MP98] Geoffrey A. McMichael and Todd N. Pearsons. Effects of wild juvenile spring Chinook salmon on growth and abundance of wild rainbow trout. Transactions of the American Fisheries Society, 127(2):261–274, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCracken:1993:GDH

[MPG93] Gary F. McCracken, Charles R. Parker, and Stanley Z. Guffey. Genetic differentiation and hybridization between stocked hatchery and native brook trout in Great Smoky Mountains National Park. Transactions of the American Fisheries Society, 122(4): 533–542, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDonald:1993:PRL

[MR93] D. G. McDonald and J. G. Robinson. Physiological responses of lake trout to stress: Effects of water hardness and genotype. Transactions of the American Fisheries Society, 122(6): 1146–1155, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matheney:1995:PMH

[MR95a] Matthew P. Matheney IV and Charles F. Rabeni. Patterns of movement and habitat use by northern hog suckers in an Ozark stream. Transactions of the American Fisheries Society, 124(6): 886–897, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moser:1995:HUM

[MR95b] Mary L. Moser and Steve W. Ross. Habitat use and movements of shortnose and Atlantic sturgeons in the Lower Cape Fear River, North Carolina. Transactions of the American Fisheries Society, 124(2):225–234, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maret:1997:FAE

[MRM97] Terry R. Maret, Christopher T. Robinson, and G. Wayne Minshall. Fish assemblages and environmental correlates in leastdisturbed streams of the Upper Snake River Basin. Transactions of the American Fisheries Society, 126(2):200–216, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matuszek:1996:NEM

[MS96a] John E. Matuszek and Brian J. Shuter. Notes: an empirical method for the prediction of daily water temperatures in the littoral zone of temperate lakes. Transactions of the American Fisheries Society, 125(4):622–627, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miner:1996:DPH

[MS96b] Jeffrey G. Miner and Roy A. Stein. Detection of predators and habitat choice by small bluegills: Effects of turbidity and alternative prey. Transactions of the American Fisheries Society, 125 (1):97–103, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Munn:1997:SHM

[MS97] Mark D. Munn and Terry M. Short. Spatial heterogeneity of mercury bioaccumulation by walleye in Franklin D. Roosevelt Lake and the Upper Columbia River, Washington. Transactions of the American Fisheries Society, 126(3):477–487, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1999:EAW

[MS99] Barbara A. Martin and Michael K. Saiki. Effects of ambient water quality on the endangered Lost River sucker in Upper Klamath Lake, Oregon. Transactions of the American Fisheries Society, 128(5):953–961, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Matuszek:1990:CLT

[MSC90] John E. Matuszek, Brian J. Shuter, and John M. Casselman. Changes in lake trout growth and abundance after introduction of cisco into Lake Opeongo, Ontario. Transactions of the American Fisheries Society, 119(4):718–729, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morizot:1994:JSA

[MSC94] Donald C. Morizot, Maureen E. Schmidt, and Gary J. Carmichael. Joint segregation of allozymes in catfish genetic crosses: Designation of Ictalurus punctatus linkage group I. Transactions of the American Fisheries Society, 123(1):22–27, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mazik:1991:IWH

[MSP91] Patricia M. Mazik, Bill A. Simco, and Nick C. Parker. Influence of water hardness and salts on survival and physiological characteristics of striped bass during and after transport. Transactions of the American Fisheries Society, 120(1):121–126, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McBride:1995:CVA

[MSP95] Richard S. McBride, Michael D. Scherer, and J. Christopher Powell. Correlated variations in abundance, size, growth, and loss rates of age-0 bluefish in a Southern New England Estuary. Transactions of the American Fisheries Society, 124(6):898– 910, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McMichael:1997:ERH

[MSP97] Geoffrey A. McMichael, Cameron S. Sharpe, and Todd N. Pearsons. Effects of residual hatchery-reared steelhead on growth of wild rainbow trout and spring Chinook salmon. Transactions of the American Fisheries Society, 126(2):230–239, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murphy:1991:NDV

[MT91] Michael D. Murphy and Ronald G. Taylor. Notes: Direct validation of ages determined for adult red drums from otolith sections. Transactions of the American Fisheries Society, 120(2):267–269, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Malloy:1994:ERL

[MT94a] Kirk D. Malloy and Timothy E. Targett. Effects of ration limitation and low temperature on growth, biochemical condition, and survival of juvenile summer flounder from two Atlantic Coast nurseries. Transactions of the American Fisheries Society, 123 (2):182–193, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murphy:1994:AGM

[MT94b] Michael D. Murphy and Ronald G. Taylor. Age, growth, and mortality of spotted seatrout in Florida waters. Transactions of the American Fisheries Society, 123(4):482–497, July 1994. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Madenjian:1996:FYG

[MTK^{+}96] Charles P. Madenjian, Jeffrey T. Tyson, Roger L. Knight, Mark W. Kershner, and Michael J. Hansen. First-year growth, recruitment, and maturity of walleyes in Western Lake Erie. Transactions of the American Fisheries Society, 125(6):821–830, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murawski:1993:CCM

[Mur93] S. A. Murawski. Climate change and marine fish distributions: Forecasting from historical analogy. Transactions of the American Fisheries Society, 122(5):647–658, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murray:1994:MPC

[Mur94] Clyde B. Murray. A method for preparing Chinook salmon otoliths for age determination, and evidence of its validity. Transactions of the American Fisheries Society, 123(3):358–367, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mooij:1999:ACF

[MVW99] W. M. Mooij, J. M. Van Rooij, and S. Wijnhoven. Analysis and comparison of fish growth from small samples of length-at-age data: Detection of sexual dimorphism in Eurasian perch as an example. Transactions of the American Fisheries Society, 128(3): 483–490, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Magoulick:1998:ETM

[MW98] Daniel D. Magoulick and Margaret A. Wilzbach. Effect of temperature and macrohabitat on interspecific aggression, foraging success, and growth of brook trout and rainbow trout pairs in laboratory streams. Transactions of the American Fisheries Society, 127(5):708–717, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGeer:1994:NNE

[MWW^{+}94] James C. McGeer, Patricia A. Wright, Chris M. Wood, Michael P. Wilkie, Carl F. Mazur, and George K. Iwama. Notes: Nitrogen excretion in four species of fish from an alkaline lake. Transactions of the American Fisheries Society, 123(5):824–829, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macavoy:1997:ETM

[MZ97] S. E. Macavoy and R. C. Zaepfel. Effects of tricaine methanesulfonate (MS-222) on hematocrit: First field measurements on blacknose dace. Transactions of the American Fisheries Society, 126(3):500–503, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nagler:1999:EPF

[NAC99] James J. Nagler, Bruce A. Adams, and Daniel G. Cyr. Egg production, fertility, and hatch success of American plaice held in captivity. Transactions of the American Fisheries Society, 128(4):727–736, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nakamoto:1994:CPU

[Nak94] Rodney J. Nakamoto. Characteristics of pools used by adult summer steelhead oversummering in the New River, California. Transactions of the American Fisheries Society, 123(5):757–765, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nakashima:1999:ELS

[Nak99] Brian S. Nakashima. Early life stage mortality syndrome in fishes of the Great Lakes and Baltic Sea. Transactions of the American Fisheries Society, 128(3):549–550, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nack:1993:SNH

[NBGF93] Steven B. Nack, Donald Bunnell, David M. Green, and John L. Forney. Spawning and nursery habitats of largemouth bass in the tidal Hudson River. Transactions of the American Fisheries Society, 122(2):208–216, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nielsen:1997:MDN

[NCFG97] Jennifer L. Nielsen, Cindy Carpanzano, Monique C. Fountain, and Christine A. Gan. Mitochondrial DNA and nuclear microsatellite diversity in hatchery and wild Oncorhynchus mykiss from freshwater habitats in Southern California. Transactions of the American Fisheries Society, 126(3):397–417, May 1997. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neves:1999:BR

[Nev99] Richard J. Neves. Book reviews. Transactions of the American Fisheries Society, 128(4):764–767, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ney:1993:BMT

[Ney93] John J. Ney. Bioenergetics modeling today: Growing pains on the cutting edge. Transactions of the American Fisheries Society, 122(5):736–748, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nielsen:1992:MSF

[Nie92] Jennifer L. Nielsen. Microhabitat-specific foraging behavior, diet, and growth of juvenile Coho salmon. Transactions of the American Fisheries Society, 121(5):617–634, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nilsson:1994:GGJ

[Nil94] J. Nilsson. Genetics of growth of juvenile Arctic char. Transactions of the American Fisheries Society, 123(3):430–434, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nielsen:1994:TSP

[NLO94] Jennifer L. Nielsen, Thomas E. Lisle, and Vicki Ozaki. Thermally stratified pools and their use by steelhead in Northern California streams. Transactions of the American Fisheries Society, 123(4):

613–626, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norcross:1991:ERM

[Nor91] Brenda L. Norcross. Estuarine recruitment mechanisms of larval Atlantic croakers. Transactions of the American Fisheries Society, 120(6):673–683, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nedbal:1994:DMD

[NP94a] Michael A. Nedbal and David P. Philipp. Differentiation of mitochondrial DNA in largemouth bass. Transactions of the American Fisheries Society, 123(4):460–468, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nichol:1994:RDR

[NP94b] Daniel G. Nichol and Ellen K. Pikitch. Reproduction of darkblotched rockfish off the Oregon Coast. Transactions of the American Fisheries Society, 123(4):469–481, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:1992:TDH

[NPLJ92] Rodger L. Nelson, William S. Platts, David P. Larsen, and Sherman E. Jensen. Trout distribution and habitat in relation to geology and geomorphology in the North Fork Humboldt River Drainage, Northeastern Nevada. Transactions of the American Fisheries Society, 121(4):405–426, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

NUrnberg:1995:AFQ

[NÜr95] Gertrud K. NÜrnberg. The anoxic factor, a quantitative measure of anoxia and fish species richness in Central Ontario lakes. Transactions of the American Fisheries Society, 124(5):677–686, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neverman:1992:VFJ

[NW92] Darcy Neverman and Wayne A. Wurtsbaugh. Visual feeding by juvenile Bear Lake sculpin. Transactions of the American Fisheries Society, 121(3):395–398, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nieland:1993:RBA

[NW93] David L. Nieland and Charles A. Wilson. Reproductive biology and annual variation of reproductive variables of black drum in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 122(3):318–327, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nieland:1995:HAB

[NW95] David L. Nieland and Charles A. Wilson. Histological abnormalities and bacterial proliferation in red drum and black drum ovaries. Transactions of the American Fisheries Society, 124 (1):139–143, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Noga:1999:CCR

[NWGA99] Edward J. Noga, Chengjie Wang, Carol B. Grindem, and Ramy Avtalion. Comparative clinicopathological responses of striped bass and palmetto bass to acute stress. Transactions of the American Fisheries Society, 128(4):680–686, July 1999. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Owens:1994:RSS

[OB94] Randall W. Owens and Roger A. Bergstedt. Response of slimy sculpins to predation by juvenile lake trout in Southern Lake Ontario. Transactions of the American Fisheries Society, 123 (1):28–36, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osmundson:1998:STE

[OB98] D. B. Osmundson and K. P. Burnham. Status and trends of the endangered Colorado squawfish in the Upper Colorado River. Transactions of the American Fisheries Society, 127(6):957–970, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olla:1992:NCP

[ODS92] Bori L. Olla, Michael W. Davis, and Carl B. Schreck. Notes: Comparison of predator avoidance capabilities with corticosteroid levels induced by stress in juvenile Coho salmon. Transactions of the American Fisheries Society, 121(4):544–547, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olney:1991:SBE

[OFM91] John E. Olney, John D. Field, and John C. McGovern. Striped bass egg mortality, production, and female biomass in Virginia Rivers, 1980–1989. Transactions of the American Fisheries Society, 120(3):354–367, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OGorman:1997:AGA

[OJS97] Robert O’Gorman, Ora E. Johannsson, and Clifford P. Schneider. Age and growth of alewives in the changing Pelagia of Lake Ontario, 1978–1992. Transactions of the American Fisheries Society, 126(1):112–126, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osborne:1992:ISL

[OKB^{+}92] Lewis L. Osborne, Steven L. Kohler, Peter B. Bayley, David M. Day, William A. Bertrand, Michael J. Wiley, and Randy Sauer. Influence of stream location in a drainage network on the index of biotic integrity. Transactions of the American Fisheries Society, 121(5):635–643, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oliveira:1997:MGR

[Oli97] Kenneth Oliveira. Movements and growth rates of yellow-phase American eels in the Annaquatucket River, Rhode Island. Transactions of the American Fisheries Society, 126(4):638–646, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olsen:1992:MFR

[OR92] Erik J. Olsen and Roger A. Rulifson. Maturation and fecundity of Roanoke River–Albemarle Sound striped bass. Transactions of the American Fisheries Society, 121(4):524–537, July 1992. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osmundson:1997:GSC

[ORM97] D. B. Osmundson, R. J. Ryel, and T. E. Mourning. Growth and survival of Colorado squawfish in the Upper Colorado River. Transactions of the American Fisheries Society, 126(4):687–698, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Orth:1996:RHH

[Ort96] Donald J. Orth. The rivers handbook: Hydrological and ecological principles. Transactions of the American Fisheries Society, 125(3):486–488, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Osmundson:1998:DPS

[ORT^{+}98] D. B. Osmundson, R. J. Ryel, M. E. Tucker, B. D. Burdick, W. R. Elmblad, and T. E. Chart. Dispersal patterns of subadult and adult Colorado squawfish in the Upper Colorado River. Transactions of the American Fisheries Society, 127(6):943–956, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OBrien:1993:ECV

[OS93] W. J. O’Brien and J. J. Showalter. Effects of current velocity and suspended debris on the drift feeding of Arctic grayling. Transactions of the American Fisheries Society, 122(4):609–615, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olsen:1998:GIB

[OSBS98] Jeffrey B. Olsen, Lisa W. Seeb, Paul Bentzen, and James E. Seeb. Genetic interpretation of broad-scale microsatellite polymorphism in odd-year pink salmon. Transactions of the American Fisheries Society, 127(4):535–550, July 1998. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ogle:1995:DFP

[OSNH95] Derek H. Ogle, James H. Selgeby, Raymond M. Newman, and Mary G. Henry. Diet and feeding periodicity of ruffe in the St. Louis River Estuary, Lake Superior. Transactions of the American Fisheries Society, 124(3):356–369, May 1995. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Orsi:1995:MVD

[OW95] Joseph A. Orsi and Alex C. Wertheimer. Marine vertical distribution of juvenile Chinook and Coho salmon in Southeastern Alaska. Transactions of the American Fisheries Society, 124(2): 159–169, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ososkov:1996:DSB

[OW96] Inna Ososkov and Judith S. Weis. Development of social behavior in larval mummichogs after embryonic exposure to methylmercury. Transactions of the American Fisheries Society, 125(6): 983–987, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OBeirn:1998:MVG

[OWJH98] Francis X. O’Beirn, Randal L. Walker, Michelle L. Jansen, and Peter B. Heffernan. Microgeographical variations in gametogenesis and sex ratios in the eastern oyster at two marsh sites in Georgia. Transactions of the American Fisheries Society, 127 (2):298–308, March 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paller:1994:RBF

[Pal94] Michael H. Paller. Relationships between fish assemblage structure and stream order in South Carolina coastal plain streams. Transactions of the American Fisheries Society, 123(2):150–161, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parrish:1993:EFC

[Par93a] James D. Parrish. The ecology of fishes on coral reefs. Transactions of the American Fisheries Society, 122(3):511–514, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1993:CTD

[Par93b] Glenn R. Parsons. Comparisons of triploid and diploid white crappies. Transactions of the American Fisheries Society, 122 (2):237–243, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:1993:CRF

[PB93] James T. Peterson and Peter B. Bayley. Colonization rates of fishes in experimentally defaunated warmwater streams. Transactions of the American Fisheries Society, 122(2):199–207, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pipas:1998:HBR

[PB98] James C. Pipas and Frank J. Bulow. Hybridization between redeye bass and smallmouth bass in Tennessee streams. Transactions of the American Fisheries Society, 127(1):141–146, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paragamian:1992:UDG

[PBH92] Vaughn L. Paragamian, Edward C. Bowles, and Brian Hoelscher. Use of daily growth increments on otoliths to assess stockings of hatchery-reared Kokanees. Transactions of the American Fisheries Society, 121(6):785–791, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsley:1993:SRH

[PBM93] Michael J. Parsley, Lance G. Beckman, and George T. McCabe, Jr. Spawning and rearing habitat use by white sturgeons in the Columbia River downstream from McNary Dam. Transactions of the American Fisheries Society, 122(2):217–227, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Popper:1998:ASO

[PC98] Arthur N. Popper and Thomas J. Carlson. Application of sound and other stimuli to control fish behavior. Transactions of the American Fisheries Society, 127(5):673–707, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powell:1990:NCS

[PCGW90] Allyn B. Powell, Alexander J. Chester, John J. Govoni, and Stanley M. Warlen. Nutritional condition of spot larvae associated with the Mississippi River plume. Transactions of the American Fisheries Society, 119(6):957–965, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pope:1994:IBL

[PD94] Kevin L. Pope and Dennis R. Devries. Interactions between larval white crappie and gizzard shad: Quantifying mechanisms in small ponds. Transactions of the American Fisheries Society, 123(6): 975–987, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:1998:CNC

[PD98] R. O. Parker, Jr. and R. L. Dixon. Changes in a North Carolina reef fish community after 15 years of intense fishing — global warming implications. Transactions of the American Fisheries Society, 127(6):908–920, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Partridge:1999:RGM

[PD99] David G. Partridge and Dennis R. DeVries. Regulation of growth and mortality in larval bluegills: Implications for juvenile recruitment. Transactions of the American Fisheries Society, 128(4): 625–638, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pert:1994:HUA

[PE94] Edmund J. Pert and Don C. Erman. Habitat use by adult rainbow trout under moderate artificial fluctuations in flow. Transactions of the American Fisheries Society, 123(6):913–923, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petersen:1994:ISP

[Pet94] James H. Petersen. Importance of spatial pattern in estimating predation on juvenile salmonids in the Columbia River. Transactions of the American Fisheries Society, 123(6):924–930, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Porch:1990:SDT

[PF90] Clay E. Porch and William W. Fox, Jr. Simulating the dynamic trends of fisheries regulated by small daily bag limits. Transactions of the American Fisheries Society, 119(5):836–849, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pender:1996:HHB

[PG96] Peter J. Pender and Roland K. Griffin. Habitat history of barramundi Lates calcarifer in a North Australian River system based on barium and strontium levels in scales. Transactions of the American Fisheries Society, 125(5):679–689, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Purcell:1990:MAA

[PGG90] Jennifer E. Purcell, Daniel Grosse, and Jill J. Grover. Mass abundances of abnormal Pacific herring larvae at a spawning ground in British Columbia. Transactions of the American Fisheries Society, 119(3):463–469, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prager:1996:ASP

[PGS96] Michael H. Prager, C. Phillip Goodyear, and Gerald P. Scott. Application of a surplus production model to a swordfish-like simulated stock with time-changing gear selectivity. Transactions of the American Fisheries Society, 125(5):729–740, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prager:1992:CWD

[PH92] Michael H. Prager and John M. Hoenig. Can we determine the significance of key-event effects on a recruitment time series? — a power study of superposed epoch analysis. Transactions of the American Fisheries Society, 121(1):123–131, January 1992. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phelan:1992:WFM

[Phe92] Beth A. Phelan. Winter flounder movements in the Inner New York Bight. Transactions of the American Fisheries Society, 121 (6):777–784, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Poe:1991:FPF

[PHV^{+}91] Thomas P. Poe, Hal C. Hansel, Steven Vigg, Douglas E. Palmer, and Linda A. Prendergast. Feeding of predaceous fishes on outmigrating juvenile salmonids in John Day Reservoir, Columbia River. Transactions of the American Fisheries Society, 120(4): 405–420, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1997:VCB

[Pie97] Rodney B. Pierce. Variable catchability and bias in population estimates for northern pike. Transactions of the American Fisheries Society, 126(4):658–664, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petering:1991:NSC

[PJ91] Raymond W. Petering and David L. Johnson. Notes: Suitability of a cyanoacrylate adhesive to close incisions in black crappies used in telemetry studies. Transactions of the American Fisheries Society, 120(4):535–537, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1995:HDI

[PJN95] J. M. Phillips, J. R. Jackson, and R. L. Noble. Hatching date influence on age-specific diet and growth of age-0 largemouth bass. Transactions of the American Fisheries Society, 124(3): 370–379, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perkins:1993:HBT

[PKM93] David L. Perkins, Charles C. Krueger, and Bernie May. Heritage brook trout in Northeastern USA: Genetic variability within and among populations. Transactions of the American Fisheries Society, 122(4):515–532, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearsons:1992:IHC

[PLL92] Todd N. Pearsons, Hiram W. Li, and Gary A. Lamberti. Influence of habitat complexity on resistance to flooding and resilience of stream fish assemblages. Transactions of the American Fisheries Society, 121(4):427–436, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Plumb:1990:MME

[Plu90] John A. Plumb. Methods for the microbiological examination of fish and shellfish. Transactions of the American Fisheries Society, 119(3):562–566, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Papoulias:1992:EFA

[PM92] Diana Papoulias and W. L. Minckley. Effects of food availability on survival and growth of larval razorback suckers in ponds. Transactions of the American Fisheries Society, 121(3):340–355, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pinder:1995:IPH

[PM95] Michael J. Pinder and Raymond P. Morgan II. Interactions of pH and habitat on cyprinid distributions in Appalachian streams of Maryland. Transactions of the American Fisheries Society, 124 (1):94–102, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Power:1997:LVL

[PM97] M. Power and R. S. McKinley. Latitudinal variation in lake sturgeon size as related to the thermal opportunity for growth. Transactions of the American Fisheries Society, 126(4):549–558, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Potts:1999:OAG

[PM99] Jennifer C. Potts and Charles S. Manooch III. Observations on the age and growth of graysby and coney from the Southeastern United States. Transactions of the American Fisheries Society, 128(4):751–757, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prout:1990:DGP

[PMF90] Mark W. Prout, Edward L. Mills, and John L. Forney. Diet, growth, and potential competitive interactions between age-0 white perch and yellow perch in Oneida Lake, New York. Transactions of the American Fisheries Society, 119(6):966–975, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peake:1997:ITA

[PMSM97] S. Peake, R. S. McKinley, D. A. Scruton, and R. Moccia. Influence of transmitter attachment procedures on swimming performance of wild and hatchery-reared Atlantic salmon smolts. Transactions of the American Fisheries Society, 126(4):707–714, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Potts:1998:AGV

[PMV98] Jennifer C. Potts, Charles S. Manooch III, and Douglas S. Vaughan. Age and growth of vermilion snapper from the Southeastern United States. Transactions of the American Fisheries Society, 127(5):787–795, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:1999:GSP

[PNSF99] Mark S. Peterson, Larry C. Nicholson, Douglas J. Snyder, and Gregory L. Fulling. Growth, spawning preparedness, and diet of Cycleptus meridionalis (Catostomidae). Transactions of the American Fisheries Society, 128(5):900–908, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Power:1994:QEB

[Pow94] M. Power. Quantitative ecology and the brown trout. Transactions of the American Fisheries Society, 123(6):1006–1008, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Power:1994:MDS

[PP94] M. Power and G. Power. Modeling the dynamics of smolt production in Atlantic salmon. Transactions of the American Fisheries Society, 123(4):535–548, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Putman:1995:RBE

[PPD95] John H. Putman, Clay L. Pierce, and David M. Day. Relationships between environmental variables and size-specific growth rates of Illinois stream fishes. Transactions of the American Fisheries Society, 124(2):252–261, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paragamian:1999:MDA

[PPF99] Vaughn L. Paragamian, Madison S. Powell, and Joyce C. Faler. Mitochondrial DNA analysis of burbot stocks in the Kootenai River Basin of British Columbia, Montana, and Idaho. Transactions of the American Fisheries Society, 128(5):868–874, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Post:1998:STV

[PPJ98] John R. Post, E. A. Parkinson, and N. T. Johnston. Spatial and temporal variation in risk to piscivory of age-0 rainbow trout: Patterns and population level consequences. Transactions of the American Fisheries Society, 127(6):932–942, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pascual:1995:FAH

[PQF95] Miguel A. Pascual, Thomas P. Quinn, and Howard Fuss. Factors affecting the homing of fall Chinook salmon from Columbia River hatcheries. Transactions of the American Fisheries Society, 124(3):308–320, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paller:1996:UFC

[PRD96] Michael H. Paller, Marcel J. M. Reichert, and John M. Dean. Use of fish communities to assess environmental impacts in South Carolina coastal plain streams. Transactions of the American Fisheries Society, 125(5):633–644, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1996:BCF

[PRL96] Clay L. Pierce, Joseph B. Rasmussen, and William C. Leggett. Back-calculation of fish length from scales: Empirical comparison of proportional methods. Transactions of the American Fisheries Society, 125(6):889–898, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parkinson:1994:CAT

[PRR94] Eric A. Parkinson, Bruce E. Rieman, and Lars G. Rudstam. Comparison of acoustic and trawl methods for estimating density and age composition of Kokanee. Transactions of the American Fisheries Society, 123(6):841–854, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Post:1995:TSD

[PRS95] John R. Post, Lars G. Rudstam, and Denise M. Schael. Temporal and spatial distribution of pelagic age-0 fish in Lake Mendota, Wisconsin. Transactions of the American Fisheries Society, 124 (1):84–93, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peters:1991:EMT

[PS91] D. S. Peters and W. E. Schaaf. Empirical model of the trophic basis for fishery yield in coastal waters of the Eastern USA. Transactions of the American Fisheries Society, 120(4):459–473, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1999:GDL

[PS99] Todd A. Phillips and Robert C. Summerfelt. Gill development of larval walleyes. Transactions of the American Fisheries Society, 128(1):162–168, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paszkowski:1994:EPS

[PT94] Cynthia A. Paszkowski and William M. Tonn. Effects of prey size, abundance, and population structure on piscivory by yellow perch. Transactions of the American Fisheries Society, 123(6): 855–865, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paszkowski:1999:BR

[PT99] Cynthia A. Paszkowski and William M. Tonn. Book reviews. Transactions of the American Fisheries Society, 128(4):764–767, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1997:HBI

[PV97] Patrick C. Pierce and Michael J. Van Den Avyle. Hybridization between introduced spotted bass and smallmouth bass in reservoirs. Transactions of the American Fisheries Society, 126(6): 939–947, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Power:1999:AYP

[PvdH99] M. Power and M. R. van den Heuvel. Age-0 yellow perch growth and its relationship to temperature. Transactions of the American Fisheries Society, 128(4):687–700, July 1999. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parrish:1994:NAP

[PVE94] Donna L. Parrish, Bruce Vondracek, and William J. Eckmayer. Notes: Accuracy and precision of daily age estimates for walleyes from Ohio hatcheries and Lake Erie. Transactions of the American Fisheries Society, 123(4):601–605, July 1994. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Philipp:1991:SGN

[PW91] David P. Philipp and Gregory S. Whitt. Survival and growth of northern, Florida, and reciprocal Fl hybrid largemouth bass in Central Illinois. Transactions of the American Fisheries Society, 120(1):58–64, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Petersen:1999:DCB

[PW99] James H. Petersen and David L. Ward. Development and corroboration of a bioenergetics model for northern pikeminnow feeding on juvenile salmonids in the Columbia River. Transactions of the American Fisheries Society, 128(5):784–801, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:1995:VBC

[PZW95] Robert M. Parker, Mark P. Zimmerman, and David L. Ward. Variability in biological characteristics of northern squawfish in the Lower Columbia and Snake Rivers. Transactions of the American Fisheries Society, 124(3):335–346, May 1995. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:1998:GPD

[QGWF98] T. P. Quinn, E. Graynoth, C. C. Wood, and C. J. Foote. Genotypic and phenotypic divergence of sockeye salmon in New Zealand from their ancestral British Columbia populations. Transactions of the American Fisheries Society, 127(4):517–534, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:1991:HSP

[QNM91] Thomas P. Quinn, Richard S. Nemeth, and Donald O. McIsaac. Homing and straying patterns of fall Chinook salmon in the Lower Columbia River. Transactions of the American Fisheries Society, 120(2):150–156, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:1996:EIB

[QPS96] Thomas P. Quinn, N. Phil Peterson, and S. Spalding. Effects of instream brush on juvenile Coho salmon: Response to comment. Transactions of the American Fisheries Society, 125(1):151–153, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [Cou96].

Quist:1999:OHU

[QTBG99] Michael C. Quist, Jeff S. Tillma, Matthew N. Burlingame, and Christopher S. Guy. Overwinter habitat use of shovelnose sturgeon in the Kansas River. Transactions of the American Fisheries Society, 128(3):522–527, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:1990:FPD

[Qui90] Terrance J. Quinn II. Fish population dynamics: Implications for management. Transactions of the American Fisheries Society, 119(2):390–392, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rountree:1992:FHG

[RA92] Rodney A. Rountree and Kenneth W. Able. Foraging habits, growth, and temporal patterns of salt-marsh creek habitat use by young-of-year summer flounder in New Jersey. Transactions of the American Fisheries Society, 121(6):765–776, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rabeni:1997:SSS

[Rab97] Charles F. Rabeni. Sediment in streams: Sources, biological effects, and control. Transactions of the American Fisheries Society, 126(6):1048–1051, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1992:LAS

[RB92] Robert M. Ross and Thomas W. H. Backman. Larval American shad: Effects of age and group size on swimming and feeding behavior. Transactions of the American Fisheries Society, 121(4): 508–516, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rien:1994:APW

[RB94] Thomas A. Rien and Raymond C. Beamesderfer. Accuracy and precision of white sturgeon age estimates from pectoral fin rays. Transactions of the American Fisheries Society, 123(2):255–265, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1992:NGS

[RBL92] Robert M. Ross, Thomas W. H. Backman, and Karin E. Limburg. Notes: Group-size-mediated metabolic rate reduction in American shad. Transactions of the American Fisheries Society, 121(3):385–390, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rudstam:1994:BMA

[RBM94] Lars G. Rudstam, Fred P. Binkowski, and Michael A. Miller. A bioenergetics model for analysis of food consumption patterns of bloater in Lake Michigan. Transactions of the American Fisheries Society, 123(3):344–357, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rieman:1991:ELJ

[RBVP91] Bruce E. Rieman, Raymond C. Beamesderfer, Steven Vigg, and Thomas P. Poe. Estimated loss of juvenile salmonids to predation by northern squawfish, walleyes, and smallmouth bass in John Day Reservoir, Columbia River. Transactions of the American Fisheries Society, 120(4):448–458, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:1993:IBM

[RC93] Kenneth A. Rose and James H. Cowan, Jr. Individual-based model of young-of-the-year striped bass population dynamics. I. Model description and baseline simulations. Transactions of the American Fisheries Society, 122(3):415–438, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robinson:1998:DLF

[RCF98] Anthony T. Robinson, Robert W. Clarkson, and Robert E. Forrest. Dispersal of larval fishes in a regulated river tributary. Transactions of the American Fisheries Society, 127(5):772–786, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rice:1993:IBS

[RCR93] James A. Rice, Larry B. Crowder, and Kenneth A. Rose. Interactions between size-structured predator and prey populations: Experimental test and model comparison. Transactions of the American Fisheries Society, 122(3):481–491, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rulifson:1995:LHP

[RD95] Roger A. Rulifson and Michael J. Dadswell. Life history and population characteristics of striped bass in Atlantic Canada. Transactions of the American Fisheries Society, 124(4):477–507, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1997:SSV

[RE97a] R. Anne Richards and Christine Esteves. Stock-specific variation in scale morphology of Atlantic striped bass. Transactions of the American Fisheries Society, 126(6):908–918, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1997:USM

[RE97b] R. Anne Richards and Christine Esteves. Use of scale morphology for discriminating wild Stocks of Atlantic striped bass. Transactions of the American Fisheries Society, 126(6):919–925, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reinert:1992:BIS

[Rei92] Robert E. Reinert. Biological indicators of stress in fish. Transactions of the American Fisheries Society, 121(2):274–276, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rutherford:1992:DWE

[REM92] D. Allen Rutherford, Anthony A. Echelle, and O. Eugene Maughan. Drainage-wide effects of timber harvesting on the structure of stream fish assemblages in Southeastern Oklahoma. Transactions of the American Fisheries Society, 121(6):716–728, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reeves:1993:DJA

[RES93] Gordon H. Reeves, Fred H. Everest, and James R Sedell. Diversity of juvenile anadromous salmonid assemblages in coastal Oregon Basins with different levels of timber harvest. Transactions of the American Fisheries Society, 122(3):309–317, May 1993. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rettig:1998:VSC

[Ret98] Jessica E. Rettig. Variation in species composition of the larval assemblage in four Southwest Michigan lakes: Using allozyme analysis to identify larval sunfish. Transactions of the American Fisheries Society, 127(4):661–668, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reynolds:1994:NCS

[Rey94] James B. Reynolds. Northern chars: Scientific challenges and management opportunities. Transactions of the American Fisheries Society, 123(3):368–369, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rondorf:1990:FES

[RGF90] Dennis W. Rondorf, Gerard A. Gray, and Robert B. Fairley. Feeding ecology of subyearling Chinook salmon in riverine and reservoir habitats of the Columbia River. Transactions of the American Fisheries Society, 119(1):16–24, January 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rahel:1991:FAH

[RH91] Frank J. Rahel and Wayne A. Hubert. Fish assemblages and habitat gradients in a Rocky Mountain–Great Plains stream: Biotic zonation and additive patterns of community change. Transactions of the American Fisheries Society, 120(3):319–332, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Regier:1990:ITC

[RHP90] Henry A. Regier, John A. Holmes, and Daniel Pauly. Influence of temperature changes on aquatic ecosystems: an interpretation of empirical data. Transactions of the American Fisheries Society, 119(2):374–389, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roberts:1999:ARC

[RJK^{+}99] Steven B. Roberts, Leslie F. Jackson, William King V, Ronald G. Taylor, Harry J. Grier, and Craig V. Sullivan. Annual reproductive cycle of the common snook: Endocrine correlates of maturation. Transactions of the American Fisheries Society, 128(3): 436–445, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radtke:1996:EML

[RK96] Richard L. Radtke and Robert A. Kinzie III. Evidence of a marine larval stage in endemic Hawaiian stream gobies from isolated high-elevation locations. Transactions of the American Fisheries Society, 125(4):613–621, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rutherford:1995:IPC

[RKBC95] D. Allen Rutherford, William E. Kelso, C. Frederick Bryan, and Glenn C. Constant. Influence of physicochemical characteristics on annual growth increments of four fishes from the Lower Mississippi River. Transactions of the American Fisheries Society, 124(5):687–697, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rico:1991:DPY

[RKF91] Ciro Rico, Urs Kuhnlein, and Gerald J. Fitzgerald. A DNA probe that yields highly informative DNA fingerprints for the threespine stickleback. Transactions of the American Fisheries Society, 120 (6):809–815, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reed:1992:GFM

[RKR92] Bobby C. Reed, William E. Kelso, and D. Allen Rutherford. Growth, fecundity, and mortality of paddlefish in Louisiana. Transactions of the American Fisheries Society, 121(3):378–384, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rinoon:1993:MUS

[RLC93] Pedro A. Rinóon and Javier Lobón-Cerviá. Microhabitat use by stream-resident brown trout: Bioenergetic consequences. Transactions of the American Fisheries Society, 122(4):575–587, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robichaud-Leblanc:1998:DGY

[RLCB98] Kimberly A. Robichaud-Leblanc, Simon C. Courtenay, and Tillmann J. Benfey. Distribution and growth of young-of-the-year striped bass in the Miramichi River Estuary, Gulf of St. Lawrence. Transactions of the American Fisheries Society, 127 (1):56–69, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rand:1994:EDS

[RLO^{+}94] Peter S. Rand, Brian F. Lantry, Robert O’Gorman, Randall W. Owens, and Donald J. Stewart. Energy density and size of pelagic prey fishes in Lake Ontario, 1978–1990: Implications for salmonine energetics. Transactions of the American Fisheries Society, 123(4):519–534, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rieman:1992:IFD

[RM92] Bruce E. Rieman and Deborah L. Myers. Influence of fish density and relative productivity on growth of Kokanee in ten oligotrophic lakes and reservoirs in Idaho. Transactions of the American Fisheries Society, 121(2):178–191, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rieman:1995:OBT

[RM95] Bruce E. Rieman and John D. McIntyre. Occurrence of bull trout in naturally fragmented habitat patches of varied size. Transactions of the American Fisheries Society, 124(3):285–296, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Regier:1990:IPS

[RMC90] Henry A. Regier, John J. Magnuson, and Charles C. Coutant. Introduction to Proceedings: Symposium on Effects of Climate Change on Fish. Transactions of the American Fisheries Society, 119(2):173–175, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reisenbichler:1992:GVS

[RMSL92] R. R. Reisenbichler, J. D. McIntyre, M. F. Solazzi, and S. W. Landino. Genetic variation in steelhead of Oregon and Northern California. Transactions of the American Fisheries Society, 121 (2):158–169, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1992:ISA

[RN92] Michael R. Ross and Gary A. Nelson. Influences of stock abundance and bottom-water temperature on growth dynamics of haddock and yellowtail flounder on Georges Bank. Transactions of the American Fisheries Society, 121(5):578–587, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richardson:1998:GEM

[RN98] Dennis J. Richardson and Brent B. Nickol. Gastric evacuation and movement of materials through the alimentary canal of green sunfish. Transactions of the American Fisheries Society, 127(6): 1059–1066, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rossiter:1995:VAE

[RNB95] A. Rossiter, D. L. G. Noakes, and E. W. H. Beamish. Validation of age estimation for the lake sturgeon. Transactions of the American Fisheries Society, 124(5):777–781, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roell:1993:TBP

[RO93] Michael J. Roell and Donald J. Orth. Trophic basis of production of stream-dwelling smallmouth bass, rock bass, and flathead catfish in relation to invertebrate bait harvest. Transactions of the American Fisheries Society, 122(1):46–62, January 1993. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raibley:1997:NLB

[ROI^{+}97] Paul T. Raibley, Timothy M. O’Hara, Kevin S. Irons, K. Douglas Blodgett, and Richard E. Sparks. Notes: Largemouth bass size distributions under varying annual hydrological regimes in the Illinois River. Transactions of the American Fisheries Society, 126(5):850–856, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Railsback:1999:BMS

[RR99] Steven F. Railsback and Kenneth A. Rose. Bioenergetics modeling of stream trout growth: Temperature and food consumption effects. Transactions of the American Fisheries Society, 128(2): 241–256, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rottiers:1992:NDS

[RRBA92] D. V. Rottiers, L. A. Redell, H. E. Booke, and S. Amaral. Notes: Differences in Stocks of American shad from the Columbia and Delaware Rivers. Transactions of the American Fisheries Society, 121(1):132–136, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rmstrong:1994:RBS

[RS94] Michael P. Rmstrong and Bethany A. Starr. Reproductive biology of the smooth flounder in Great Bay Estuary, New Hampshire. Transactions of the American Fisheries Society, 123(1): 112–114, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rasmussen:1998:BCI

[RSBN98] Paul W. Rasmussen, Michael D. Staggs, T. Douglas Beard, Jr., and Steven P. Newman. Bias and confidence interval coverage of creel survey estimators evaluated by simulation. Transactions of the American Fisheries Society, 127(3):469–480, May 1998. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roper:1994:SDH

[RSL94] Brett B. Roper, Dennis L. Scarnecchia, and Tim J. La Marr. Summer distribution of and habitat use by Chinook salmon and steelhead within a major basin of the South Umpqua River, Oregon. Transactions of the American Fisheries Society, 123(3): 298–308, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rand:1993:MSP

[RSS^{+}93] Peter S. Rand, Donald J. Stewart, Paul W. Seelbach, Michael L. Jones, and Leslie R. Wedge. Modeling steelhead population energetics in Lakes Michigan and Ontario. Transactions of the American Fisheries Society, 122(5):977–1001, September 1993. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1995:AGM

[RSV95] Jeffrey L. Ross, Tonya M. Stevens, and Douglas S. Vaughan. Age, growth, mortality, and reproductive biology of red drums in North Carolina waters. Transactions of the American Fisheries Society, 124(1):37–54, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rulifson:1999:SBS

[RT99] Roger A. Rulifson and Katherine A. Tull. Striped bass spawning in a tidal bore river: The Shubenacadie Estuary, Atlantic Canada. Transactions of the American Fisheries Society, 128(4): 613–624, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Render:1995:RDS

[RTA95] Jeffrey H. Render, Bruce A. Thompson, and Robert L. Allen. Reproductive development of striped mullet in Louisiana estuarine waters with notes on the applicability of reproductive assessment methods for isochronal species. Transactions of the American Fisheries Society, 124(1):26–36, January 1995. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Render:1992:RBS

[RW92] Jeffrey H. Render and Charles A. Wilson. Reproductive biology of sheepshead in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 121(6):757–764, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruzycki:1999:OHS

[RW99] James R. Ruzycki and Wayne A. Wurtsbaugh. Ontogenetic habitat shifts of juvenile Bear Lake sculpin. Transactions of the American Fisheries Society, 128(6):1201–1212, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snowdon:1992:NAV

[SA92] R. J. Snowdon and K. D. Adam. Notes: Allozyme variation in rainbow trout from Central North Island, New Zealand. Transactions of the American Fisheries Society, 121(6):810–814, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smogor:1995:DAA

[SAG95] Roy A. Smogor, Paul L. Angermeier, and Clark K. Gaylord. Distribution and abundance of American eels in Virginia streams: Tests of null models across spatial scales. Transactions of the American Fisheries Society, 124(6):789–803, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1992:AFG

[Sai92] Saul B. Saila. Application of fuzzy graph theory to successional analysis of a multispecies trawl fishery. Transactions of the American Fisheries Society, 121(2):211–233, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1997:LMR

[Sai97] Saul B. Saila. Living marine resources: Their utilization and management. Transactions of the American Fisheries Society, 126(1):171–172, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shaklee:1990:GNP

[SAMW90] James B. Shaklee, Fred W. Allendorf, Donald C. Morizot, and Gregory S. Whitt. Gene nomenclature for protein-coding loci in fish. Transactions of the American Fisheries Society, 119(1):2– 15, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saunders:1991:EMA

[Sau91] R. L. Saunders. Ecology and management of Atlantic salmon. Transactions of the American Fisheries Society, 120(6):828–833, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swartzman:1990:SER

[SB90] Gordon L. Swartzman and David A. Beauchamp. Simulation of the effect of rainbow trout introduction in Lake Washington. Transactions of the American Fisheries Society, 119(1):122–134, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stauffer:1994:EST

[SB94] Jay R. Stauffer, Jr. and Suzanne E. Boltz. Effect of salinity on the temperature preference and tolerance of age-0 Mayan cichlids. Transactions of the American Fisheries Society, 123(1):101–107, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sharber:1999:EUP

[SB99] Norman G. Sharber and Jane Sharber Black. Epilepsy as a unifying principle in electrofishing theory: a proposal. Transactions of the American Fisheries Society, 128(4):666–671, July 1999. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sommer:1997:RSS

[SBH97] Ted Sommer, Randall Baxter, and Bruce Herbold. Resilience of splittail in the Sacramento–San Joaquin Estuary. Transactions of the American Fisheries Society, 126(6):961–976, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stockwell:1999:KFD

[SBJ99] Jason D. Stockwell, Krista L. Bonfantine, and Brett M. Johnson. Kokanee foraging: a Daphnia in the stomach is worth two in the lake. Transactions of the American Fisheries Society, 128(1):169– 174, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schrock:1994:MGS

[SBRH94] Robin M. Schrock, John W. Beeman, Dennis W. Rondorf, and Philip V. Haner. A microassay for gill sodium, potassiumactivated ATPase in juvenile Pacific salmonids. Transactions of the American Fisheries Society, 123(2):223–229, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Starling:1995:NPL

[SBSN95] Scott M. Starling, Rita M. Bruckler, R. Kirk Strawn, and William H. Neill. Notes: Predicting the lethality of fluctuating low temperatures to blue tilapia. Transactions of the American Fisheries Society, 124(1):112–117, January 1995. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sanderson:1992:ECS

[SC92a] S. Laurie Sanderson and Joseph J. Cech, Jr. Energetic cost of suspension feeding versus particulate feeding by juvenile Sacramento blackfish. Transactions of the American Fisheries Society, 121(2):149–157, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strange:1992:RSP

[SC92b] Richard J. Strange and Joseph J. Cech, Jr. Reduced swimming performance of striped bass after confinement stress. Transactions of the American Fisheries Society, 121(2):206–210, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sulak:1998:ELH

[SC98] Kenneth J. Sulak and James P. Clugston. Early life history stages of Gulf sturgeon in the Suwannee River, Florida. Transactions of the American Fisheries Society, 127(5):758–771, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seeb:1999:AMD

[SC99a] Lisa W. Seeb and Penelope A. Crane. Allozymes and mitochondrial DNA discriminate Asian and North American populations of chum salmon in mixed-stock fisheries along the South Coast of the Alaska Peninsula. Transactions of the American Fisheries Society, 128(1):88–103, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seeb:1999:HGH

[SC99b] Lisa W. Seeb and Penelope A. Crane. High genetic heterogeneity in chum salmon in Western Alaska, the contact zone between northern and southern lineages. Transactions of the American Fisheries Society, 128(1):58–87, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scarratt:1994:ITM

[Sca94] David J. Scarratt. Introductions and transfers of marine species: Achieving a balance between economic development and resource protection. Transactions of the American Fisheries Society, 123 (2):273–276, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Samis:1993:PRF

[SCJ93] Andrew J. W. Samis, Patrick W. Colgan, and Peter H. Johansen. Pentachlorophenol and reduced food intake of bluegill. Transactions of the American Fisheries Society, 122(6):1156–1160, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scoppettone:1993:IBN

[Sco93] G. Gary Scoppettone. Interactions between native and nonnative fishes of the Upper Muddy River, Nevada. Transactions of the American Fisheries Society, 122(4):599–608, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schill:1990:IDA

[SD90] W. Bane Schill and Robert M. Dorazio. Immunological discrimination of Atlantic striped bass stocks. Transactions of the American Fisheries Society, 119(1):77–85, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Szalai:1991:RPP

[SD91] Alexander J. Szalai and Terry A. Dick. Role of predation and parasitism in growth and mortality of yellow perch in Dauphin Lake, Manitoba. Transactions of the American Fisheries Society, 120(6):739–751, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stark:1998:GSS

[SE98] William J. Stark and Anthony A. Echelle. Genetic structure and systematics of smallmouth bass, with emphasis on interior highlands populations. Transactions of the American Fisheries Society, 127(3):393–416, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scholz:1994:UQR

[SEDP94] Paul M. Scholz, Don Edwards, John M. Dean, and Roger C. Palm, Jr. Use of a quadratic response model to estimate optimum combinations of nutritional variables. Transactions of the American Fisheries Society, 123(2):266–272, March 1994. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seelbach:1993:PBS

[See93] Paul W. Seelbach. Population biology of steelhead in a stableflow, low-gradient tributary of Lake Michigan. Transactions of the American Fisheries Society, 122(2):179–198, March 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1997:TTV

[SF97] Ryan K. Smith and Kurt D. Fausch. Thermal tolerance and vegetation preference of Arkansas darter and Johnny darter from Colorado plains streams. Transactions of the American Fisheries Society, 126(4):676–686, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1994:SRT

[SG94] R. W. Smith and J. S. Griffith. Survival of rainbow trout during their first winter in the Henry’s Fork of the Snake River, Idaho. Transactions of the American Fisheries Society, 123(5):747–756, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snucins:1995:CWE

[SG95] E. J. Snucins and J. M. Gunn. Coping with a warm environment: Behavioral thermoregulation by lake trout. Transactions of the American Fisheries Society, 124(1):118–123, January 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saunders:1990:IPG

[SH90] Richard L. Saunders and Paul R. Harmon. Influence of photoperiod on growth of juvenile Atlantic salmon and development of salinity tolerance during winter–spring. Transactions of the American Fisheries Society, 119(4):689–697, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strand:1994:GSD

[SH94] Rita Strand and Tor G. Heggberget. Growth and sex distribution in an anadromous population of Arctic char in Northern Norway. Transactions of the American Fisheries Society, 123(3):377–384, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scott:1997:FAI

[SH97a] Mark C. Scott and Lenwood W. Hall, Jr. Fish assemblages as indicators of environmental degradation in Maryland coastal plain streams. Transactions of the American Fisheries Society, 126(3): 349–360, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stafford:1997:MCM

[SH97b] Craig P. Stafford and Terry A. Haines. Mercury concentrations in Maine sport fishes. Transactions of the American Fisheries Society, 126(1):144–152, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strange:1998:NNL

[SH98] Richard J. Strange and James W. Habera. No net loss of brook trout distribution in areas of sympatry with rainbow trout in Tennessee streams. Transactions of the American Fisheries Society, 127(3):434–440, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schulz:1999:IBI

[SHC99] Eric J. Schulz, Mark V. Hoyer, and Daniel E. Canfield, Jr. An index of biotic integrity: a test with limnological and fish data from sixty Florida lakes. Transactions of the American Fisheries Society, 128(4):564–577, July 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaeffer:1999:FTT

[SHDB99] Jeffrey S. Schaeffer, Robert C. Haas, James S. Diana, and James E. Breck. Field test of two energetic models for yellow perch. Transactions of the American Fisheries Society, 128(3): 414–435, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savino:1993:FAF

[SHK93] Jacqueline F. Savino, Mary G. Henry, and Harold L. Kincaid. Factors affecting feeding behavior and survival of juvenile lake trout in the Great Lakes. Transactions of the American Fisheries Society, 122(3):366–377, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shuter:1991:CBF

[SI91] Brian J. Shuter and Peter E. Ihssen. Chemical and biological factors affecting acid tolerance of smallmouth bass. Transactions of the American Fisheries Society, 120(1):23–33, January 1991. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siegwarth:1994:IHR

[Sie94] Gary L. Siegwarth. Identification of hatchery-reared channel catfish by means of pectoral spine cross sections. Transactions of the American Fisheries Society, 123(5):830–834, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sigler:1994:EMB

[Sig94] Michael F. Sigler. An electronic measuring board with bar codes. Transactions of the American Fisheries Society, 123(1):115–117, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sigler:1995:BCM

[Sig95] Michael F. Sigler. Bar-coded measuring systems: Response to comment. Transactions of the American Fisheries Society, 124 (4):641–642, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [Mor95, Ken95].

Smith:1991:NPC

[SJ91] Barry D. Smith and Glen S. Jamieson. Notes: Possible consequences of intensive fishing for males on the mating opportunities of Dungeness crabs. Transactions of the American Fisheries Society, 120(5):650–653, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smithson:1999:MPS

[SJ99] Elizabeth B. Smithson and Carol E. Johnston. Movement patterns of stream fishes in a Ouachita Highlands stream: an examination of the restricted movement paradigm. Transactions of the American Fisheries Society, 128(5):847–853, September 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1992:CCL

[SJMM92] Ian R. Smith, Allan F. Johnson, Donald MacLennan, and Harold Manson. Chemical contaminants, lymphocystis, and dermal sarcoma in walleyes spawning in the Thames River, Ontario. Transactions of the American Fisheries Society, 121(5):608–616, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saiki:1992:TAS

[SJW92] Michael K. Saiki, Mark R. Jennings, and Raymond H. Wiedmeyer. Toxicity of agricultural subsurface drainwater from the San Joaquin Valley, California, to juvenile Chinook salmon and striped bass. Transactions of the American Fisheries Society, 121(1):78–93, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sabo:1991:RBM

[SK91] Matthew J. Sabo and William E. Kelso. Relationship between morphometry of excavated floodplain ponds along the Mississippi River and their use as fish nurseries. Transactions of the American Fisheries Society, 120(5):552–561, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Savino:1996:CBN

[SK96] Jacqueline F. Savino and Cynthia S. Kolar. Competition between nonindigenous ruffe and native yellow perch in laboratory studies. Transactions of the American Fisheries Society, 125(4):562–571, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shields:1992:MDV

[SKG92] Barbara A. Shields, Anne R. Kapuscinski, and Kevin S. Guise. Mitochondrial DNA variation in four Minnesota populations of lake whitefish: Utility as species and population markers. Transactions of the American Fisheries Society, 121(1):21–25, January 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schindler:1993:FWS

[SKH^{+}93] Daniel E. Schindler, James F. Kitchell, Xi He, Stephen R. Carpenter, James R. Hodgson, and Kathryn L. Cottingham. Food web structure and phosphorus cycling in lakes. Transactions of the American Fisheries Society, 122(5):756–772, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snedden:1999:DSP

[SKR99] Gregg A. Snedden, William E. Kelso, and D. Allen Rutherford. Diel and seasonal patterns of spotted gar movement and habitat use in the Lower Atchafalaya River Basin, Louisiana. Transactions of the American Fisheries Society, 128(1):144–154, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stauffer:1996:CFR

[SLH96] Jay R. Stauffer, Jr., Thomas J. Lovullo, and Ho Yeon Han. Commensalistic feeding relationships of three Lake Malaŵi fish species. Transactions of the American Fisheries Society, 125(2):

224–229, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1990:GAS

[SM90a] Barry D. Smith and Gordon A. McFarlane. Growth analysis of Strait of Georgia lingcod by use of length-frequency and lengthincrement data in combination. Transactions of the American Fisheries Society, 119(5):802–812, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Springer:1990:PRW

[SM90b] Timothy A. Springer and Brian R. Murphy. Properties of relative weight and other condition indices. Transactions of the American Fisheries Society, 119(6):1048–1058, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1991:FMN

[SM91] Douglas G. Smith and Paula J. S. Martin. Freshwater macroinvertebrates of Northeastern North America. Transactions of the American Fisheries Society, 120(1):130–132, January 1991. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shepherd:1996:DFD

[SM96a] William C. Shepherd and Edward L. Mills. Diel feeding, daily food intake, and Daphnia consumption by age-0 gizzard shad in Oneida Lake, New York. Transactions of the American Fisheries Society, 125(3):411–421, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swenson:1996:FET

[SM96b] Ramona O. Swenson and Aria T. McCray. Feeding ecology of the tidewater goby. Transactions of the American Fisheries Society, 125(6):956–970, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1990:MMT

[SMC90] Barry D. Smith, Gordon A. McFarlane, and Alan J. Cass. Movements and mortality of tagged male and female lingeod in the Strait of Georgia, British Columbia. Transactions of the American Fisheries Society, 119(5):813–824, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1994:NHB

[SMD94] Stephen M. Smith, Michael J. Maceina, and Rex A. Dunham. Natural hybridization between black crappie and white crappie in Weiss Lake, Alabama. Transactions of the American Fisheries Society, 123(1):71–79, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swanson:1996:USA

[SMDC96] Christina Swanson, Randall C. Mager, Serge I. Doroshov, and Joseph J. Cech, Jr. Use of salts, anesthetics, and polymers to minimize handling and transport mortality in delta smelt. Transactions of the American Fisheries Society, 125(2):326–329, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Soriano:1992:NFA

[SMHP92] Mina Soriano, Jacques Moreau, John M. Hoenig, and Daniel Pauly. New functions for the analysis of two-phase growth of juvenile and adult fishes, with application to Nile perch. Transactions of the American Fisheries Society, 121(4):486–493, July 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1991:MAF

[Smi91] Stephen J. Smith. Mathematical analysis of fish stock dynamics. Transactions of the American Fisheries Society, 120(5):669–672, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sissenwine:1998:CWA

[SMPS98] Michael P. Sissenwine, Pamela M. Mace, Joseph E. Powers, and Gerald P. Scott. A commentary on Western Atlantic bluefin tuna assessments. Transactions of the American Fisheries Society, 127 (5):838–855, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:1996:PEP

[Sny96] Randal J. Snyder. Physiological ecology of Pacific salmon. Transactions of the American Fisheries Society, 125(5):819–820, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sabo:1994:TVM

[SO94] Matthew J. Sabo and Donald J. Orth. Temporal variation in microhabitat use by age-0 smallmouth bass in the North Anna River, Virginia. Transactions of the American Fisheries Society, 123(5):733–746, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saiki:1995:EIR

[SO95] Michael K. Saiki and R. Scott Ogle. Evidence of impaired reproduction by western mosquitofish inhabiting seleniferous agricultural drainwater. Transactions of the American Fisheries Society, 124(4):578–587, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sabo:1996:ASS

[SO96] Matthew J. Sabo and Donald J. Orth. Absence of size-selective mortality among larval smallmouth bass in a Virginia stream. Transactions of the American Fisheries Society, 125(6):920–924, November 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shuter:1990:CPV

[SP90] B. J. Shuter and J. R. Post. Climate, population viability, and the zoogeography of temperate fishes. Transactions of the American Fisheries Society, 119(2):314–336, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smiley:1997:EPT

[SP97] Peter C. Smiley, Jr. and Glenn R. Parsons. Effects of photoperiod and temperature on swimming performance of white crappie. Transactions of the American Fisheries Society, 126(3):495–499, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spangler:1990:COM

[Spa90] George R. Spangler. Co-operative management of local fisheries. Transactions of the American Fisheries Society, 119(6): 1059–1062, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spalding:1995:SDS

[SPQ95] S. Spalding, N. P. Peterson, and T. P. Quinn. Summer distribution, survival, and growth of juvenile Coho salmon under varying experimental conditions of brushy instream cover. Transactions of the American Fisheries Society, 124(1):124–130, January 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stellwag:1994:MDD

[SPR94] Edmund J. Stellwag, E. Sturgis Payne, and Roger A. Rulifson. Mitochondrial DNA diversity of Roanoke River striped bass. Transactions of the American Fisheries Society, 123(3):321–334, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [WW95, SR95d].

Shively:1996:FRN

[SPS96] Rip S. Shively, Thomas P. Poe, and Sally T. Sauter. Feeding response by northern squawfish to a hatchery release of juvenile salmonids in the Clearwater River, Idaho. Transactions of the American Fisheries Society, 125(2):230–236, March 1996. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smale:1995:HHT

[SR95a] Martin A. Smale and Charles F. Rabeni. Hypoxia and hyperthermia tolerances of headwater stream fishes. Transactions of the American Fisheries Society, 124(5):698–710, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smale:1995:IHH

[SR95b] Martin A. Smale and Charles F. Rabeni. Influences of hypoxia and hyperthermia on fish species composition in headwater streams. Transactions of the American Fisheries Society, 124(5): 711–725, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sowa:1995:RER

[SR95c] Scott P. Sowa and Charles F. Rabeni. Regional evaluation of the relation of habitat to distribution and abundance of smallmouth bass and largemouth bass in Missouri streams. Transactions of the American Fisheries Society, 124(2):240–251, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stellwag:1995:MDS

[SR95d] E. J. Stellwag and Roger A. Rulifson. Mitochondrial DNA stability and striped bass stock identification: Response to comment. Transactions of the American Fisheries Society, 124(6): 956–959, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [SPR94, WW95].

Schael:1995:STD

[SRD95] Denise M. Schael, James A. Rice, and Donald J. Degan. Spatial and temporal distribution of threadfin shad in a southeastern reservoir. Transactions of the American Fisheries Society, 124 (6):804–812, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schertzer:1990:TSL

[SS90a] W. M. Schertzer and A. M. Sawchuk. Thermal structure of the Lower Great Lakes in a warm year: Implications for the occurrence of hypolimnion anoxia. Transactions of the American Fisheries Society, 119(2):195–209, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simonson:1990:CSV

[SS90b] Timothy D. Simonson and William A. Swenson. Critical stream velocities for young-of-year smallmouth bass in relation to habitat use. Transactions of the American Fisheries Society, 119(5): 902–909, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snucins:1991:SIS

[SS91] E. J. Snucins and B. J. Shuter. Survival of introduced smallmouth bass in low-pH lakes. Transactions of the American Fisheries Society, 120(2):209–216, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stearley:1993:PPT

[SS93] R. F. Stearley and G. R. Smith. Phylogeny of the Pacific trouts and salmons (Oncorhynchus) and genera of the family Salmonidae. Transactions of the American Fisheries Society, 122(1):1–33, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scheller:1999:AIA

[SSEO99] Robert M. Scheller, Virginia M. Snarski, John G. Eaton, and Gary W. Oehlert. An analysis of the influence of annual thermal variables on the occurrence of fifteen warmwater fishes. Transactions of the American Fisheries Society, 128(2):257–264, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shaklee:1993:ECM

[SSG93] James B. Shaklee, John Salini, and Rod N. Garrett. Electrophoretic characterization of multiple genetic stocks of barramundi perch in Queensland, Australia. Transactions of the American Fisheries Society, 122(5):685–701, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Staurnes:1994:ESC

[SSG94] Magne Staurnes, Trygve Sigholt, and Odd A. Gulseth. Effects of seasonal changes in water temperature on the parr–smolt transformation of Atlantic salmon and anadromous Arctic char. Transactions of the American Fisheries Society, 123(3):408–415, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Staurnes:1994:EMS

[SSGE94] Magne Staurnes, Trygve Sigholt, Odd A. Gulseth, and Robert Eliassen. Effects of maturation on seawater tolerance of anadromous Arctic char. Transactions of the American Fisheries Society, 123(3):402–407, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Steyermark:1999:BIH

[SSGI99] Anthony C. Steyermark, James R. Spotila, Deborah Gillette, and Hadar Isseroff. Biomarkers indicate health problems in brown bullheads from the industrialized Schuylkill River, Philadelphia. Transactions of the American Fisheries Society, 128(2):328–338, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scarnecchia:1996:ASY

[SSP96] Dennis L. Scarnecchia, Phillip A. Stewart, and Greg J. Power. Age structure of the Yellowstone–Sakakawea paddlefish stock, 1963–1993, in relation to reservoir history. Transactions of the American Fisheries Society, 125(2):291–299, March 1996. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sheehan:1999:BGA

[SSS^{+}99] Robert J. Sheehan, Scott P. Shasteen, Arul V. Suresh, Anne R. Kapuscinski, and James E. Seeb. Better growth in all-female diploid and triploid rainbow trout. Transactions of the American Fisheries Society, 128(3):491–498, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stansley:1992:PCU

[Sta92] William Stansley. Postmortem changes in uninhibited brain cholinesterase activity in the fathead minnow. Transactions of the American Fisheries Society, 121(1):137–139, January 1992. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Steedman:1991:OEC

[Ste91] Robert J. Steedman. Occurrence and environmental correlates of black spot disease in stream fishes near Toronto, Ontario. Transactions of the American Fisheries Society, 120(4):494–499, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Spyker:1995:DAP

[SV95] Kay A. Spyker and Eric P. Van Den Berghe. Diurnal abundance patterns of Mediterranean fishes assessed on fixed transects by scuba divers. Transactions of the American Fisheries Society, 124(2):216–224, March 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stanley:1995:ESD

[SW95] David R. Stanley and Charles A. Wilson. Effect of scuba divers on fish density and target strength estimates from stationary dual-beam hydroacoustics. Transactions of the American Fisheries Society, 124(6):946–949, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swanberg:1997:MHU

[Swa97] Tim R. Swanberg. Movements of and habitat use by fluvial bull trout in the Blackfoot River, Montana. Transactions of the American Fisheries Society, 126(5):735–746, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Secor:1991:IML

[SWD91] David H. Secor, Miller G. White, and John Mark Dean. Immersion marking of larval and juvenile hatchery-produced striped bass with oxytetracycline. Transactions of the American Fisheries Society, 120(2):261–266, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swink:1990:ELT

[Swi90] William D. Swink. Effect of lake trout size on survival after a single sea lamprey attack. Transactions of the American Fisheries Society, 119(6):996–1002, November 1990. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swink:1991:HSS

[Swi91] William D. Swink. Host-size selection by parasitic sea lampreys. Transactions of the American Fisheries Society, 120(5):637–643, September 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swink:1993:EWT

[Swi93] William D. Swink. Effect of water temperature on sea lamprey growth and lake trout survival. Transactions of the American Fisheries Society, 122(6):1161–1166, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swink:1995:GSN

[Swi95] William D. Swink. Growth and survival of newly parasitic sea lampreys at representative winter temperatures. Transactions of the American Fisheries Society, 124(3):380–386, May 1995. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1995:VAE

[TB95] Kip R. Thompson and Daniel W. Beckman. Validation of age estimates from white sucker otoliths. Transactions of the American Fisheries Society, 124(4):637–639, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1991:RSC

[TBCK91] Jay M. Thompson, Eric P. Bergersen, Clarence A. Carlson, and Lynn R. Kaeding. Role of size, condition, and lipid content in the overwinter survival of age −0 Colorado squawfish. Transactions of the American Fisheries Society, 120(3):346–353, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Travnichek:1995:RWF

[TBM95] Vincent H. Travnichek, Mark B. Bain, and Michael J. Maceina. Recovery of a warmwater fish assemblage after the initiation of a minimum-flow release downstream from a hydroelectric dam. Transactions of the American Fisheries Society, 124(6):836–844, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tringali:1999:GIC

[TBS99] Michael D. Tringali, Theresa M. Bert, and Seifu Seyoum. Genetic identification of centropomine fishes. Transactions of the American Fisheries Society, 128(3):446–458, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Terrell:1996:MSF

[TCCT96] James W. Terrell, Brian S. Cade, Jeanette Carpenter, and Jay M. Thompson. Modeling stream fish habitat limitations from wedgeshaped patterns of variation in standing stock. Transactions of the American Fisheries Society, 125(1):104–117, January 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

TeWinkel:1999:VMN

[TF99] Leslie M. TeWinkel and Guy W. Fleischer. Vertical migration and nighttime distribution of adult bloaters in Lake Michigan. Transactions of the American Fisheries Society, 128(3):459–474, May 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Toepfer:1999:SPT

[TFH99] Conrad S. Toepfer, William L. Fisher, and Jason A. Haubelt. Swimming performance of the threatened leopard darter in relation to road culverts. Transactions of the American Fisheries Society, 128(1):155–161, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Terceiro:1992:CPM

[TFS92] Mark Terceiro, David A. Fournier, and John R. Sibert. Comparative performance of MULTIFAN and Shepherd’s Length Composition Analysis (SRLCA) on simulated length-frequency distributions. Transactions of the American Fisheries Society, 121(5): 667–677, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiersch:1998:CSE

[TFW^{+}98] Terrence R. Tiersch, Chester R. Figiel, Jr., William R. Wayman, J. Holt Williamson, Gary J. Carmichael, and Owen T. Gorman. Cryopreservation of sperm of the endangered razorback sucker. Transactions of the American Fisheries Society, 127(1):95–104, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tiersch:1994:CCC

[TGC94] Terrence R. Tiersch, Cheryl A. Goudie, and Gary J. Carmichael. Cryopreservation of channel catfish sperm: Storage in cryoprotectants, fertilization trials, and growth of channel catfish produced with cryopreserved sperm. Transactions of the American Fisheries Society, 123(4):580–586, July 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tyus:1991:DHU

[TH91] Harold M. Tyus and G. Bruce Haines. Distribution, habitat use, and growth of age-0 Colorado squawfish in the Green River Basin, Colorado and Utah. Transactions of the American Fisheries Society, 120(1):79–89, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tomasso:1996:NPR

[TIT96] Anne Osborn Tomasso, J. Jeffery Isely, and Joseph R. Tomasso, Jr. Notes: Physiological responses and mortality of striped bass angled in freshwater. Transactions of the American Fisheries Society, 125(2):321–325, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thedinga:1995:RJI

[TJ95] John F. Thedinga and Scott W. Johnson. Retention of jetinjected marks on juvenile Coho and sockeye salmon. Transactions of the American Fisheries Society, 124(5):782–785, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thurow:1994:AYC

[TK94] Russell F. Thurow and John G. King. Attributes of Yellowstone cutthroat trout redds in a tributary of the Snake River, Idaho. Transactions of the American Fisheries Society, 123(1):37–50, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Teshima:1992:KPB

[TKOS92] Chisato Teshima, Shigeharu Kudo, Yoshimi Ohtani, and Asako Saito. Kidney pathology from the bacterium Hafnia alvei: Experimental evidence. Transactions of the American Fisheries Society, 121(5):599–607, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Teuscher:1996:CBK

[TL96] David Teuscher and Chris Luecke. Competition between Kokanees and Utah chub in Flaming Gorge Reservoir, Utah– Wyoming. Transactions of the American Fisheries Society, 125 (4):505–511, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Travnichek:1996:HTE

[TMD96] Vincent H. Travnichek, Michael J. Maceina, and Rex A. Dunham. Hatching time and early growth of age-0 black crappies, white crappies, and their naturally produced F_{1} hybrids in Weiss Lake, Alabama. Transactions of the American Fisheries Society, 125 (2):334–337, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Travnichek:1997:SHB

[TMWD97] Vincent H. Travnichek, Michael J. Maceina, Michael C. Wooten, and Rex A. Dunham. Symmetrical hybridization between black crappie and white crappie in an Alabama reservoir based on analysis of the cytochrome-b gene. Transactions of the American Fisheries Society, 126(1):127–132, January 1997. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tisa:1991:CAG

[TN91] Mark S. Tisa and John J. Ney. Compatibility of alewives and gizzard shad as reservoir forage fish. Transactions of the American Fisheries Society, 120(2):157–165, March 1991. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tohtz:1993:LWD

[Toh93] Joel Tohtz. Lake whitefish diet and growth after the introduction of Mysis relicta to Flathead Lake, Montana. Transactions of the American Fisheries Society, 122(4):629–635, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tonn:1990:CCF

[Ton90] William M. Tonn. Climate change and fish communities: a conceptual framework. Transactions of the American Fisheries Society, 119(2):337–352, March 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tonn:1993:DAH

[Ton93] William M. Tonn. The development of an aquatic habitat classification system for lakes. Transactions of the American Fisheries Society, 122(6):1167–1170, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tidwell:1990:CPF

[TR90] James H. Tidwell and H. Randall Robinette. Changes in proximate and fatty acid composition of fillets from channel catfish during a two-year growth period. Transactions of the American Fisheries Society, 119(1):31–40, January 1990. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trippel:1998:ESV

[Tri98] Edward A. Trippel. Egg size and viability and seasonal offspring production of young Atlantic cod. Transactions of the American Fisheries Society, 127(3):339–359, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Trenham:1998:BIA

[TSM98] Peter C. Trenham, H. Bradley Shaffer, and Peter B. Moyle. Biochemical identification and assessment of population subdivision in morphologically similar native and invading smelt species (Hypomesus) in the Sacramento–San Joaquin Estuary, California. Transactions of the American Fisheries Society, 127(3):417–424, May 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:1994:DOS

[TV94] Malcolm H. Taylor and Enrico P. Villoso. Daily ovarian and spawning cycles in weakfish. Transactions of the American Fisheries Society, 123(1):9–14, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tabor:1991:PRI

[TW91] Roger A. Tabor and Wayne A. Wurtsbaugh. Predation risk and the importance of cover for juvenile rainbow trout in lentic systems. Transactions of the American Fisheries Society, 120(6): 728–738, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tsai:1991:RFP

[TWC91] Chu-Fa Tsai, Martin Wiley, and Ai-Ling Chai. Rise and fall of the Potomac River striped bass stock: a hypothesis of the role of sewage. Transactions of the American Fisheries Society, 120(1): 1–22, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tyus:1990:PRC

[Tyu90] Harold M. Tyus. Potamodromy and reproduction of Colorado squawfish in the Green River Basin, Colorado and Utah. Transactions of the American Fisheries Society, 119(6):1035–1047, November 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Travnichek:1993:ELW

[TZF93] Vincent H. Travnichek, Alexander V. Zale, and William L. Fisher. Entrainment of lchthyoplankton by a warmwater hydroelectric facility. Transactions of the American Fisheries Society, 122(5):709–716, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ugedal:1994:GWS

[UHG94] Ola Ugedal, Tor G. Heggberget, and Gjert E. Grande. Growth of wild stunted Arctic char after transfer to a commercial rearing system. Transactions of the American Fisheries Society, 123(3): 423–429, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Unwin:1998:ODA

[UJ98] M. J. Unwin and G. D. James. Occurrence and distribution of adult Chinook salmon in the New Zealand commercial fishery. Transactions of the American Fisheries Society, 127(4):560–575, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Underwood:1996:RAL

[UMT96] Tevis J. Underwood, Michael J. Millard, and Laura A. Thorpe. Relative abundance, length frequency, age, and maturity of Dolly Varden in nearshore waters of the Arctic National Wildlife Refuge, Alaska. Transactions of the American Fisheries Society, 125(5):719–728, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanHassel:1993:WQN

[Van93] John H. Van Hassel. Water quality in North American River systems. Transactions of the American Fisheries Society, 122 (4):645, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vaughan:1994:EBG

[VB94] Douglas S. Vaughan and Michael L. Burton. Estimation of von Bertalanffy growth parameters in the presence of size-selective mortality: a simulated example with red grouper. Transactions of the American Fisheries Society, 123(1):1–8, January 1994. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanSnikGray:1997:FRP

[VBKS97] Ellen Van Snik Gray, Jeff M. Boltz, Karen A. Kellogg, and Jay R. Stauffer, Jr. Food resource partitioning by nine sympatric darter species. Transactions of the American Fisheries Society, 126(5): 822–840, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Veinott:1999:EES

[VE99] Geoffrey I. Veinott and R. Douglas Evans. An examination of elemental stability in the fin ray of the white sturgeon with laser ablation sampling–inductively coupled plasma–mass spectrometry (LAS-ICP-MS). Transactions of the American Fisheries Society, 128(2):352–361, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanderZanden:1998:ASI

[VHRR98] M. Jake Vander Zanden, Menno Hulshof, Mark S. Ridgway, and Joseph B. Rasmussen. Application of stable isotope techniques to trophic studies of age-0 smallmouth bass. Transactions of the American Fisheries Society, 127(5):729–739, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDenAvyle:1994:ESI

[VM94] Michael J. Van Den Avyle and Monica A. Maynard. Effects of saltwater intrusion and flow diversion on reproductive success of striped bass in the Savannah River Estuary. Transactions of the American Fisheries Society, 123(6):886–903, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Verschuren:1997:FZH

[VM97] Dirk Verschuren and Leo F. Marnell. Fossil zooplankton and the historical status of westslope cutthroat trout in a headwater lake of Glacier National Park, Montana. Transactions of the American Fisheries Society, 126(1):21–34, January 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanEenennaam:1999:KAG

[VMM99] A. L. Van Eenennaam, J. D. Murray, and J. F. Medrano. Karyotype of the American green sturgeon. Transactions of the American Fisheries Society, 128(1):175–177, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vanderkooy:1998:CCS

[VP98] Steven J. Vanderkooy and Mark S. Peterson. Critical current speeds for young Gulf Coast walleyes. Transactions of the American Fisheries Society, 127(1):137–140, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vigg:1991:RCJ

[VPPH91] Steven Vigg, Thomas P. Poe, Linda A. Prendergast, and Hal C. Hansel. Rates of consumption of juvenile salmonids and alternative prey fish by northern squawfish, walleyes, smallmouth bass, and channel catfish in John Day Reservoir, Columbia River. Transactions of the American Fisheries Society, 120(4):421–438, July 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanWinkle:1993:IBA

[VRC93] Webster Van Winkle, Kenneth A. Rose, and R. Christopher Chambers. Individual-based approach to fish population dynamics: an overview. Transactions of the American Fisheries Society, 122(3):397–403, May 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Villarreal:1994:FAP

[VRF94] Beverly Winkler Villarreal, Paul M. Rosenblum, and Loraine T. Fries. Fatty acid profiles in red drum muscle: Comparison between wild and cultured fish. Transactions of the American Fisheries Society, 123(2):194–203, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanWinkle:1993:LLH

[VRW^{+}93] Webster Van Winkle, Kenneth A. Rose, Kirk O. Winemiller, Donald L. Deangelis, Sigurd W. Christensen, Robert G. Otto, and Brian J. Shuter. Linking life history theory, environmental setting, and individual-based modeling to compare responses of different fish species to environmental change. Transactions of the American Fisheries Society, 122(3):459–466, May 1993. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Volk:1994:UBC

[VSGA94] Eric C. Volk, Steven L. Schroder, Jeffrey J. Grimm, and H. Sprague Ackley. Use of a bar code symbology to produce multiple thermally induced otolith marks. Transactions of the American Fisheries Society, 123(5):811–816, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Valcarce:1991:DSF

[VSSS91] Ron Valcarce, Douglas Stevenson, Grant Gill Smith, and John W. Sigler. Discriminant separation of fishes: a preliminary study of fish eggs by Curie-point pyrolysis–mass spectrometry– chemometric analysis. Transactions of the American Fisheries Society, 120(6):796–802, November 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vignier:1992:GFC

[VVF92] Veronique Vignier, John H. Vandermeulen, and Angus J. Fraser. Growth and food conversion by Atlantic salmon parr during 40 days’ exposure to crude oil. Transactions of the American Fisheries Society, 121(3):322–332, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanDoornik:1998:IAC

[VW98] Donald M. Van Doornik and Gary A. Winans. Inheritance of allozymes in Coho salmon. Transactions of the American Fisheries Society, 127(5):833–837, September 1998. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vandenbyllaardt:1991:RBT

[VWBM91] Lenore Vandenbyllaardt, F. J. Ward, C. R. Braekevelt, and D. B. McIntyre. Relationships between turbidity, piscivory, and development of the retina in juvenile walleyes. Transactions of the American Fisheries Society, 120(3):382–390, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woiwode:1991:ETP

[WA91] John G. Woiwode and Ira R. Adelman. Effects of temperature, photoperiod, and ration size on growth of hybrid striped bass × white bass. Transactions of the American Fisheries Society, 120 (2):217–229, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:1995:SBB

[Wal95] John R. Waldman. Sea bass: Biology, exploitation, and conservation. Transactions of the American Fisheries Society, 124(4): 643–644, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warlen:1992:AGS

[War92] Stanley M. Warlen. Age, growth, and size distribution of larval Atlantic menhaden off North Carolina. Transactions of the American Fisheries Society, 121(5):588–598, September 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warren:1993:RCM

[War93] Melvin L. Warren, Jr. River conservation and management. Transactions of the American Fisheries Society, 122(1):155–156, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1999:LTT

[Wat99] Thomas F. Waters. Long-term trout production dynamics in Valley Creek, Minnesota. Transactions of the American Fisheries Society, 128(6):1151–1162, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winans:1996:AVO

[WAW96] Gary A. Winans, Paul B. Aebersold, and Robin S. Waples. Allozyme variability of Oncorhynchus nerka in the Pacific Northwest, with special consideration to populations of Redfish Lake, Idaho. Transactions of the American Fisheries Society, 125(5): 645–663, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodward:1994:ERT

[WBD^{+}94] Daniel F. Woodward, William G. Brumbaugh, Aaron J. Delonay, Edward E. Little, and Charlie E. Smith. Effects on rainbow trout fry of a metals-contaminated diet of benthic invertebrates from the Clark Fork River, Montana. Transactions of the American Fisheries Society, 123(1):51–62, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1997:CET

[WBI97] Eric J. Wagner, Thomas Bosakowski, and Steven Intelmann. Combined effects of temperature and high pH on mortality and the stress response of rainbow trout after stocking. Transactions of the American Fisheries Society, 126(6):985–998, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webb:1990:LBL

[WD90] Paul W. Webb and Vivian De Buffrénil. Locomotion in the biology of large aquatic vertebrates. Transactions of the American Fisheries Society, 119(4):629–641, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilkie:1996:PSW

[WDB^{+}96] Michael P. Wilkie, Kevin Davidson, Mark A. Brobbel, James D. Kieffer, Richard K. Booth, Alex T. Bielak, and Bruce L. Tufts.

Physiology and survival of wild Atlantic salmon following angling in warm summer waters. Transactions of the American Fisheries Society, 125(4):572–580, July 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilson:1997:MDC

[WDFY97] Raymond R. Wilson, Jr., Kim A. Donaldson, Marc E. Frischer, and Thomas B. Young. Mitochondrial DNA control region of common snook and its prospect for use as a genetic tag. Transactions of the American Fisheries Society, 126(4):594–606, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1990:APP

[WDK90] Thomas F. Waters, Michael T. Doherty, and Charles C. Krueger. Annual production and production: Biomass ratios for three species of stream trout in Lake Superior tributaries. Transactions of the American Fisheries Society, 119(3):470–474, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:1990:NME

[WDM90] John R. Waldman, Dennis J. Dunning, and Mark T. Mattson. Notes: a morphological explanation for size-dependent anchor tag loss from striped bass. Transactions of the American Fisheries Society, 119(5):920–923, September 1990. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:1990:RDH

[WDRM90] John R. Waldman, Dennis J. Dunning, Quentin E. Ross, and Mark T. Mattson. Range dynamics of Hudson River striped bass along the Atlantic Coast. Transactions of the American Fisheries Society, 119(5):910–919, September 1990. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Withler:1990:GVR

[WE90] R. E. Withler and T. P. T. Evelyn. Genetic variation in resistance to bacterial kidney disease within and between two strains of Coho salmon from British Columbia. Transactions of the American Fisheries Society, 119(6):1003–1009, November 1990. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilde:1992:GSP

[WE92] Gene R. Wilde and Anthony A. Echelle. Genetic status of pecos pupfish populations after establishment of a hybrid swarm involving an introduced congener. Transactions of the American Fisheries Society, 121(3):277–286, May 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weatherley:1990:AUF

[Wea90] A. H. Weatherley. Approaches to understanding fish growth. Transactions of the American Fisheries Society, 119(4):662–672, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wahl:1995:EEP

[WEH95] David H. Wahl, Lisa M. Einfalt, and Michael L. Hooe. Effect of experience with piscivory on foraging behavior and growth of walleyes. Transactions of the American Fisheries Society, 124 (5):756–763, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:1994:PSD

[WF94] John R. Waldman and Mary C. Fabrizio. Problems of stock definition in estimating relative contributions of Atlantic striped bass to the coastal fishery. Transactions of the American Fisheries Society, 123(5):766–778, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodward:1991:SGC

[WFL^{+}91] D. F. Woodward, A. M. Farag, E. E. Little, B. Steadman, and R. Yancik. Sensitivity of greenback cutthroat trout to acidic pH and elevated aluminum. Transactions of the American Fisheries Society, 120(1):34–42, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weaver:1994:UWH

[WG94] L. Alan Weaver and Greg C. Garman. Urbanization of a watershed and historical changes in a stream fish assemblage. Transactions of the American Fisheries Society, 123(2):162–172, March 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1997:RPC

[WGC97] D. Scott Waters, Christopher S. Guy, and Christopher P. Clouse. Relative position of coded wire tags in paddlefish rostrums. Transactions of the American Fisheries Society, 126(2):338–342, March 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woodward:1997:CTA

[WGFB97] Daniel F. Woodward, Jack N. Goldstein, Ayda M. Farag, and William G. Brumbaugh. Cutthroat trout avoidance of metals and conditions characteristic of a mining waste site: Coeur d’Alene River, Idaho. Transactions of the American Fisheries Society, 126(4):699–706, July 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wright:1999:PHW

[WGFS99] Russell A. Wright, James E. Garvey, Aimee H. Fullerton, and Roy A. Stein. Predicting how winter affects energetics of age-0 largemouth bass: How do current models fare? Transactions of the American Fisheries Society, 128(4):603–612, July 1999. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wing:1992:ASM

[WGG92] Bruce L. Wing, Charles M. Guthrie III, and Anthony J. Gharrett. Atlantic salmon in marine waters of Southeastern Alaska. Transactions of the American Fisheries Society, 121(6):814–818, November 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wirgin:1991:UDF

[WGGM91] Isaac I. Wirgin, Cheryl Grunwald, S. J. Garte, and Charles Mesing. Use of DNA fingerprinting in the identification and management of a striped bass population in the Southeastern United States. Transactions of the American Fisheries Society, 120(3): 273–282, May 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walker:1995:SRN

[WH95] Randal L. Walker and Peter B. Heffernan. Sex ratio of the northern quahog according to age, size, and habitat in coastal Waters of Georgia. Transactions of the American Fisheries Society, 124 (6):929–934, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weiland:1997:CDL

[WH97a] Mark A. Weiland and Robert S. Hayward. Cause for the decline of large rainbow trout in a tailwater fishery: Too much putting or too much taking? Transactions of the American Fisheries Society, 126(5):758–773, September 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitledge:1997:LEB

[WH97b] Gregory W. Whitledge and Robert S. Hayward. Laboratory evaluation of a bioenergetics model for largemouth bass at two temperatures and feeding levels. Transactions of the American Fisheries Society, 126(6):1030–1035, November 1997. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitesel:1992:PTH

[Whi92] Timothy A. Whitesel. Plasma thyroid hormone levels in migratory and lake-resident Coho salmon juveniles from the Karluk River System, Alaska. Transactions of the American Fisheries Society, 121(2):199–205, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitledge:1998:TBM

[WHNW98] Gregory W. Whitledge, Robert S. Hayward, Douglas B. Noltie, and Ning Wang. Testing bioenergetics models under feeding regimes that elicit compensatory growth. Transactions of the American Fisheries Society, 127(5):740–746, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:1996:SCN

[WHW96] John R. Waldman, John T. Hart, and Isaac I. Wirgin. Stock composition of the New York bight Atlantic sturgeon fishery based on analysis of mitochondrial DNA. Transactions of the American Fisheries Society, 125(3):364–371, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilimovsky:1990:MTJ

[Wil90] Norman J. Wilimovsky. Misuses of the term “Julian day”. Transactions of the American Fisheries Society, 119(1):162, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:1996:LSM

[Wil96] John G. Williams. Lost in space: Minimum confidence intervals for idealized PHABSIM studies. Transactions of the American Fisheries Society, 125(3):458–465, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:1997:CTI

[Wil97] John G. Williams. Comment: Testing the independence of microhabitat preferences and flow (part 1). Transactions of the American Fisheries Society, 126(3):536–537, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [JP97, BCJ97].

Winemiller:1994:ELH

[Win94] Kirk O. Winemiller. The evolution of life histories. Transactions of the American Fisheries Society, 123(5):835–839, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wissing:1995:FEF

[Wis95] Thomas E. Wissing. Feeding ecology of fish. Transactions of the American Fisheries Society, 124(6):960–962, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Witherell:1990:NVD

[WK90] David B. Witherell and Boyd Kynard. Notes: Vertical distribution of adult American shad in the Connecticut River. Transactions of the American Fisheries Society, 119(1):151–155, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weis:1991:NRP

[WK91] Judith S. Weis and Anwar A. Khan. Notes: Reduction in prey capture ability and condition of mummichogs from a polluted habitat. Transactions of the American Fisheries Society, 120 (1):127–129, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whittier:1999:IFN

[WK99] Thomas R. Whittier and Thomas M. Kincaid. Introduced fish in Northeastern USA lakes: Regional extent, dominance, and effect on native species richness. Transactions of the American Fisheries Society, 128(5):769–783, September 1999. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1993:PDS

[WKPK93] Thomas F. Waters, John P. Kaehler, Taylor J. Polomis, and Thomas J. Kwak. Production dynamics of smallmouth bass in a small Minnesota stream. Transactions of the American Fisheries Society, 122(4):588–598, July 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woods:1995:VTS

[WKSS95] L. Curry Woods III, Christopher C. Kohler, Robert J. Sheehan, and Craig V. Sullivan. Volitional tank spawning of female striped bass with male white bass produces hybrid offspring. Transactions of the American Fisheries Society, 124(4):628–632, July 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:1998:URA

[WKW98] D. Jody Williams, Steven Kazianis, and Ronald B. Walter. Use of random amplified polymorphic DNA (RAPD) for identification of largemouth bass subspecies and their intergrades. Transactions of the American Fisheries Society, 127(5):825–832, September 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winger:1994:ESS

[WL94] Parley V. Winger and Peter J. Lasier. Effects of salinity on striped bass eggs and larvae from the Savannah River, Georgia. Transactions of the American Fisheries Society, 123(6):904– 912, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winner:1997:ENT

[WM97] Brent L. Winner and Robert H. McMichael, Jr. Evaluation of a new type of Box splitter designed for subsampling estuarine ichthyofauna. Transactions of the American Fisheries Society, 126(6):1041–1047, November 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weaver:1993:ADL

[WMC93] Melissa J. Weaver, John J. Magnuson, and Murray K. Clayton. Analyses for differentiating littoral fish assemblages with catch data from multiple sampling gears. Transactions of the American Fisheries Society, 122(6):1111–1119, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wiener:1990:FIM

[WMSG90] James G. Wiener, Robert E. Martini, Thomas B. Sheffy, and Gary E. Glass. Factors influencing mercury concentrations in walleyes in Northern Wisconsin lakes. Transactions of the American Fisheries Society, 119(5):862–870, September 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wirgin:1993:UMD

[WMWC93] Isaac Wirgin, Lorraine Maceda, John R. Waldman, and Robert N. Crittenden. Use of mitochondrial DNA polymorphisms to estimate the relative contributions of the Hudson River and Chesapeake Bay striped bass stocks to the mixed fishery on the Atlantic Coast. Transactions of the American Fisheries Society, 122(5):669–684, September 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wooster:1994:MCW

[Woo94] Warren S. Wooster. Marine climate, weather and fisheries. Transactions of the American Fisheries Society, 123(3):445–447, May 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walters:1993:DDG

[WP93] Carl J. Walters and John R. Post. Density-dependent growth and competitive asymmetries in size-structured fish populations: a theoretical model and recommendations for field experiments. Transactions of the American Fisheries Society, 122(1):34–45, January 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warren:1998:RCB

[WP98] Melvin L. Warren, Jr. and Mitzi G. Pardew. Road crossings as barriers to small-stream fish movement. Transactions of the American Fisheries Society, 127(4):637–644, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1995:IPJ

[WPL95] David L. Ward, James H. Petersen, and John J. Loch. Index of predation on juvenile salmonids by northern squawfish in the Lower and Middle Columbia River and in the Lower Snake River. Transactions of the American Fisheries Society, 124(3):321–334, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whalen:1999:MTA

[WPM99] Kevin G. Whalen, Donna L. Parrish, and Stephen D. McCormick. Migration timing of Atlantic salmon smolts relative to environmental and physiological factors. Transactions of the American Fisheries Society, 128(2):289–301, March 1999. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilhelm:1999:SFH

[WPSD99] Frank M. Wilhelm, Brian R. Parker, David W. Schindler, and David B. Donald. Seasonal food habits of bull trout from a small alpine lake in the Canadian Rocky Mountains. Transactions of the American Fisheries Society, 128(6):1176–1192, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Welker:1994:GSL

[WPW94] Michael T. Welker, Clay L. Pierce, and David H. Wahl. Growth and survival of larval fishes: Roles of competition and zooplankton abundance. Transactions of the American Fisheries Society, 123(5):703–717, September 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whillans:1990:FJF

[WRC90] Thomas H. Whillans, Henry A. Regier, and W. Jack Christie. F. E. J. Fry’s field studies: Good field data provoke new questions. Transactions of the American Fisheries Society, 119(4):574–584, July 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:1997:ECS

[WRS^{+}97] John R. Waldman, R. Anne Richards, W. Bane Schill, Isaac Wirgin, and Mary C. Fabrizio. An empirical comparison of stock identification techniques applied to striped bass. Transactions of the American Fisheries Society, 126(3):369–385, May 1997. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1990:RPR

[WS90] Bruce R. Ward and Pat A. Slaney. Returns of pen-reared steelhead from riverine, estuarine, and marine releases. Transactions of the American Fisheries Society, 119(3):492–499, May 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wahl:1991:FCG

[WS91] David H. Wahl and Roy A. Stein. Food consumption and growth of three esocids: Field tests of a bioenergetic model. Transactions of the American Fisheries Society, 120(2):230–246, March 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weiss:1999:PHR

[WS99] Steven Weiss and Stefan Schmutz. Performance of hatcheryreared brown trout and their effects on wild fish in two small Austrian streams. Transactions of the American Fisheries Society, 128(2):302–316, March 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

West:1992:AMO

[WSB^{+}92] Robin L. West, Michael W. Smith, Willard E. Barber, James B. Reynolds, and Haakon Hop. Autumn migration and overwintering of Arctic grayling in coastal streams of the Arctic National Wildlife Refuge, Alaska. Transactions of the American Fisheries Society, 121(6):709–715, November 1992. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:1996:GDP

[WSCL96] R. N. Williams, D. K. Shiozawa, J. E. Carter, and R. F. Leary. Genetic detection of putative hybridization between native and introduced rainbow trout populations of the Upper Snake River. Transactions of the American Fisheries Society, 125(3):387–401, May 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitmore:1992:GAP

[WTC92] Donald H. Whitmore, To H. Thai, and Cheryl M. Craft. Gene amplification permits minimally invasive analysis of fish mitochondrial DNA. Transactions of the American Fisheries Society, 121(2):170–177, March 1992. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winston:1991:UEF

[WTP91] Matthew R. Winston, Christopher M. Taylor, and Jimmie Pigg. Upstream extirpation of four minnow species due to damming of a prairie stream. Transactions of the American Fisheries Society, 120(1):98–105, January 1991. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallin:1995:IES

[WV95] Julie E. Wallin and Michael J. Van Den Avyle. Interactive effects of stocking site salinity and handling stress on survival of striped bass fingerlings. Transactions of the American Fisheries Society, 124(5):736–745, September 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:1994:OPD

[WW94] John R. Waldman and Isaac I. Wirgin. Origin of the present Delaware River striped bass population as shown by analysis of mitochondrial DNA. Transactions of the American Fisheries Society, 123(1):15–21, January 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waldman:1995:CMD

[WW95] John R. Waldman and Isaac I. Wirgin. Comment: Mitochondrial DNA stability and striped bass stock identification. Transactions of the American Fisheries Society, 124(6):954–956, November 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [SPR94] and response [SR95d].

Walters:1996:IHS

[WW96] Jody P. Walters and Jeffrey R. Wilson. Intraspecific habitat segregation by smallmouth bass in the Buffalo River, Arkansas. Transactions of the American Fisheries Society, 125(2):284–290, March 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1999:RSB

[WZ99] David L. Ward and Mark P. Zimmerman. Response of smallmouth bass to sustained removals of northern pikeminnow in the Lower Columbia and Snake Rivers. Transactions of the American Fisheries Society, 128(6):1020–1035, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Xu:1995:PEV

[XC95] Xucai Xu and S. Y. Chiu. Population estimates with variances of Spanish sardine along the Southern Fujian Coast, China, from length-based cohort analysis. Transactions of the American Fisheries Society, 124(3):437–452, May 1995. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:1996:ETR

[YC96] Paciencia S. Young and Joseph J. Cech, Jr. Environmental tolerances and requirements of splittail. Transactions of the American Fisheries Society, 125(5):664–678, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yako:1996:FPO

[YDS96] Lisa A. Yako, John M. Dettmers, and Roy A. Stein. Feeding preferences of omnivorous gizzard shad as influenced by fish size and zooplankton density. Transactions of the American Fisheries Society, 125(5):753–759, September 1996. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:1990:FRL

[YHW90] Michael K. Young, Wayne A. Hubert, and Thomas A. Wesche. Fines in redds of large salmonids. Transactions of the American Fisheries Society, 119(1):156–162, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yule:1993:LTC

[YL93] Daniel L. Yule and Chris Luecke. Lake trout consumption and recent changes in the fish assemblage of Flaming Gorge Reservoir. Transactions of the American Fisheries Society, 122(6): 1058–1069, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:1998:AAC

[You98] Michael K. Young. Absence of autumnal changes in habitat use and location of adult Colorado River cutthroat trout in a small stream. Transactions of the American Fisheries Society, 127 (1):147–151, January 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:1999:ECM

[You99] Kyle A. Young. Environmental correlates of male life history variation among Coho salmon populations from two Oregon coastal basins. Transactions of the American Fisheries Society, 128(1): 1–16, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Young:1997:IMD

[YRB97] Michael K. Young, Russell B. Rader, and Timberley A. Belish. Influence of macroinvertebrate drift and light on the activity and movement of Colorado River cutthroat trout. Transactions of the American Fisheries Society, 126(3):428–437, May 1997. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zehfuss:1999:AGS

[ZHP99] Katherine Potak Zehfuss, Joseph E. Hightower, and Kenneth H. Pollock. Abundance of Gulf sturgeon in the Apalachicola River, Florida. Transactions of the American Fisheries Society, 128 (1):130–143, January 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:1999:FHS

[Zim99] Mark P. Zimmerman. Food habits of smallmouth bass, walleyes, and northern pikeminnow in the Lower Columbia River Basin during outmigration of juvenile anadromous salmonids. Transactions of the American Fisheries Society, 128(6):1036–1054, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zigler:1993:GML

[ZJ93] Steven J. Zigler and Cecil A. Jennings. Growth and mortality of larval sunfish in backwaters of the Upper Mississippi River. Transactions of the American Fisheries Society, 122(6):1080– 1087, November 1993. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zorn:1998:ISH

[ZMDE98] Sarah A. Zorn, Terry L. Margenau, James S. Diana, and Clayton J. Edwards. The influence of spawning habitat on natural reproduction of muskellunge in Wisconsin. Transactions of the American Fisheries Society, 127(6):995–1005, November 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhao:1997:AGT

[ZMH97] Boxian Zhao, John C. McGovern, and Patrick J. Harris. Age, growth, and temporal change in size at age of the vermilion snapper from the South Atlantic bight. Transactions of the American Fisheries Society, 126(2):181–193, March 1997. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhang:1998:SCC

[ZT98] Quiyang Zhang and Terrence R. Tiersch. Standardization of the channel catfish karyotype with localization of constitutive heterochromatin and restriction enzyme banding. Transactions of the American Fisheries Society, 127(4):551–559, July 1998. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zhang:1994:RID

[ZTC94] Quiyang Zhang, Terrence R. Tiersch, and Richard K. Cooper. Rapid isolation of DNA for genetic screening of catfishes by polymerase chain reaction. Transactions of the American Fisheries Society, 123(6):997–1001, November 1994. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zimmerman:1999:IPJ

[ZW99] Mark P. Zimmerman and David L. Ward. Index of predation on juvenile salmonids by northern pikeminnow in the Lower Columbia River Basin, 1994–1996. Transactions of the American Fisheries Society, 128(6):995–1007, November 1999. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zale:1990:LCA

[ZWLB90] Alexander V. Zale, Jerry D. Wiechman, Robert L. Lochmiller, and Jim Burroughs. Limnological conditions associated with summer mortality of striped bass in Keystone Reservoir, Oklahoma. Transactions of the American Fisheries Society, 119(1):

REFERENCES 213 72–76, January 1990. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).